A Complete Bibliography of Publications in North American Journal of Fisheries

Management: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

21 August 2024 Version 1.05

Title word cross-reference

13 [HTPS10]. ^{15} [HTPS10]. δ [HTPS10]. × [NLK17, PVST11]. W_{s} [BBC10, BBC11, BBS19].

0.1-ha [BK16].

11-Year [SC18]. 12-mm [BRB^{+}17, TPC^{+}15].

2011 [FRL16]. 20E [CGM13, FGS13]. 222 [CGM13, WWS^{+}14]. 28-Year [MCS^{+}17]. 29 [MFRS11].

30 [BBC11]. 305-mm [SSP^{+}19].

8- [TPC^{+}15]. 8-mm [OL17].

9- [TPC^{+}15].

Abandonment [DLSW12]. Abdominally [TGKI10]. Ability [KMK^{+}15, MCB^{+}19]. Abiotic [FQ19, RHG13, SPT^{+}12]. Above [CFBB15, DB13, WBB10]. Above-Water [CFBB15]. Absence [SMP^{+}14]. Abundance [AJ13, BS12, BK16, BHN11, BM14, BS10, BGP10, CWH^{+}14, DCL19, FAW14, GAWC10, GBHO11, GHL12, HKMH10, HDKB10, HHCH15, HFCH15, HHMQ15, HWFN14, HK17, HH15, KTN14, KDMH10, KSM16, KSL^{+}19, LS13, MKRH14, MJCM15, MR16, MJ11, MH11b, MGS14, MLEK15, NCSR17, NKL18, OHL10, OHSP17, PKKR10, PSC15, PTH^{+}10, PNHA12, RRH^{+}11, RPBM10, RHMH15, SFW11, SBL^{+}18a, SCG11a, TCB18, VKH^{+}19, WWB^{+}17, WEC^{+}19, ZC15]. Abundances [SBL11, WS11b]. Abundant [KFW12]. Acadian [SDPB17]. Acanthopagrus [GMB10]. Acceptable [PS10]. Access [APP13, DBK11, HWB11]. Accessibility [GMG^{+}15]. Acclimated [CFW^{+}10, WALL17]. Acclimating [CCC16]. Acclimation [RCP^{+}13]. Accommodate [APD^{+}19]. Account [CA11]. Accounting [CJ14, GGCF14, SJM15, Vin14]. Accumulation [LJR11]. Accuracy [BGW^{+}11, BMW^{+}19, CWH^{+}14, DIK19, HKMH10, HDB^{+}12, MS16, MH11b, PZC^{+}12]. Accurate [SFW11]. Accurately [GHK19]. Achieve [HTF^{+}13, Rey15, SL15b]. Achievement [LBB16]. Acid [PHFK10]. Acknowledgments [Ano18h, DDD^{+}19]. Acoustic [BHC^{+}10, DHLD11, FBL^{+}19, JPWR10, MCB^{+}19, SAYJ17, TMC10, WRNJ^{+}19, WHG^{+}16, WEC^{+}19]. Acoustic-Tagged [DHLD11]. across [ACRBB10, HFH12, HFH18, HEL^{+}15, Kae13, RMTW16, WPJ^{+}18]. Act [SC14]. Actions [GGC17, JBRWH18]. Active [HVW^{+}17]. Actively [SMS13]. Activity [MSK19b, PRRD15]. Adaptation [ACMB^{+}14]. Adaptive [DCS11]. Added [DR19]. Addition [GCLC18, HW10, MODC10]. Additional [Bea17, Dag17]. Additions [Kra18]. Adequacy [CVD^{+}14]. Adfluvial [BSM17, MHDD16, RM14, SRSS12]. Adjacent [FWWTJD16, PVH^{+}13]. Adjusting [Der14]. Adult [BRS^{+}12, BHN11, BG12, BRA^{+}18b, CSH^{+}18, CBHB16, CKC^{+}12, CMD12, DHLD11, DR10, EMK^{+}17, GDP14, GKS13, GZ13, HFH12, HFH18, HO13, HLM^{+}18, ITY13, JM12a, JHK11, KDP^{+}10, KJN^{+}17, KAZ15, KRJT13, KPP^{+}18, MS18, MHR^{+}18, MCK^{+}15, MW19, MGWC13, MWS15, NTJ^{+}14, POWF11, RWPH11, SG13, SSE19, SL15a, SZHC15, WBG^{+}17, WMMB19, WHG^{+}16]. Advanced [LSK^{+}14, PTD10, TPL18]. Advanced-Fingerling [PTD10]. Aegean [GPLT12]. Aerial [AWG^{+}18, HWB11, SPW^{+}12, SLR10]. Aerial-Access [HWB11]. Affect [Bal15, LN12c, Pul17, RW11, SC17, SBD^{+}11, TCG^{+}17]. Affected [BKP^{+}14]. Affecting [EDN10, KPL^{+}16, KBH^{+}19, OHL10, RFQ14, SS11a]. Affects [HLM^{+}18]. Afraid [SCB^{+}19]. African [NEH10]. AFS [PBA^{+}17]. After [SHMJ19a, CBC^{+}17, EBS18, ES14, FMA^{+}12, GLH^{+}18, HIS16, HMS11, HWZ19, Mea12, NWZ14, RIMR19, SWH^{+}16, SHMJ19b, TGKI10, WHGC16]. Age [AJR^{+}19, AKP15, BCZ^{+}18, BK16, BCZ17, BBW18, BGW^{+}17, BMW^{+}19, BE11, CWB16, CF18, CGG^{+}18, CCDF18, Cle15, CPM^{+}10, CHDW18, DM12, DSB19, DWB^{+}15, DIK19, DRW^{+}18, DLC11, DJS^{+}16, EDN10, FC17, FB14, GMG^{+}15, GWB^{+}10, hGgXfC^{+}13, HNH11, HM17, HSC^{+}16, HFN11, HFH12, HFH18, HLM^{+}18, HB16, HUB^{+}18, HB14b, HT11, IWB10, JJ10, JF13, KSM^{+}18, KNS18, KMPP12, LOV17, MH13, MSK18, MSK19b, OEZ17, OLOI13, OI17, PC12, PTD10, PPO18, RFLP14, RZ10, SBS14, SPT^{+}12, SPHK16, SPL18, SWW14, SDL^{+}15, SKS^{+}15, SM14, TCG^{+}17, TLC^{+}19, TCM19, TPC^{+}15, TP17, WS16, WRNJ^{+}19, WSS13, WKTQ15, WF17, WSA18, WD10, ZWH^{+}15]. Age-0 [AKP15, BK16, BBW18, BGW^{+}17, BE11, CWB16, CGG^{+}18, DM12, DWB^{+}15, DRW^{+}18, EDN10, GMG^{+}15, HNH11, HFH12, HFH18, HB16, HB14b, JJ10, MSK18, PTD10, SBS14, SPT^{+}12, SWW14, SKS^{+}15, SM14, TPC^{+}15, WS16, WRNJ^{+}19, WF17]. Age-1 [MSK19b, SM14, WSS13]. Age-Specific [CCDF18]. Age-Structured [FC17, GWB^{+}10, RZ10]. Agent [GKMC15]. Ages [BKM16, BKM19, BSM^{+}15, TP17, WI17, WHV15]. Aggregation [DHA12]. Aging [KBK13a, OLM15, SP13, SS13b]. Agonistic [HB15]. Agricultural [CHJ^{+}16]. Aid [HCH13, PDD^{+}18, PPO18, SLR^{+}16]. Aimed [APC^{+}14]. Ain’t [SCB^{+}19]. Air [CMWM18, CDQG12, GMG16b, LM16a, RHC^{+}14, RSQH18, RSQ^{+}19, WMMB19]. al [SCPC14]. Alabama [CF18, HWDA12, SG13]. Alarm [BWHM16]. Alaska [CAS16, CSWW18, FHL18, HTF^{+}13, JBHM19, LS10, LL12, MS17a, MBM13, POWF11, RIMR19, SMW16, Sta12, SCA11, SC14, SF16, WHGC16]. Alaskan [SS13b]. Alberta [MSFS12]. Alcohol [GF18]. Aleutian [TS13]. Alevins [MKRH14]. Alewife [CBC^{+}17, GBFR18, KSL^{+}19, RRH^{+}11, SBB19]. Alewives [CSV13, SOSL10]. Alizarin [WMD^{+}19]. Allegheny [BCMB14]. Allele [SH11]. Alligator [BDSB15, BKDB17, DPK^{+}17, DBS19, MAH^{+}13, SBS14, SBDB16, WHKB17]. Allocation [MSBP10, RMTW16]. Allowing [WH11]. Alone [LMC^{+}16]. along [DV17, McB14]. Alosa [WBZ19]. Alosines [HH10]. Alpha [DBKG14]. Alphanumeric [MLHS13]. Alpine [CMK^{+}19, FBB18, KKMS11, KMK^{+}15]. Alsea [JPWR10]. Altamaha [BFP12]. Alterations [SHMS16]. Altered [SPL18]. Altering [GH17]. Alternative [BF13, DC13, FBM^{+}19, NFQ16, SBD^{+}11, STSH10, SLB10, WHMS14]. Alternatives [IWB10]. Ambient [SHS15]. Ambiguity [HSC^{+}16]. America [KFW12, KCCW19]. American [Ano16b, Ano19h, ABBG13, BFP12, BZ13, BBO16, BC18, CBHS16, CVD^{+}14, GHK19, HH11a, HH11b, LCC^{+}15, MHR^{+}18, MSK^{+}16, MS17b, PBA^{+}17, PT11, SRFW14, SR16, SBW^{+}17b, SC14, SEM14, TGS^{+}15, TO10, TCB18, UBR^{+}17, WHB^{+}14, WBZ19, WA14, ZC12]. americanus [BC18]. Ammocoete [AM17]. Ammocoetes [RM12]. Ammonia [WMSH13]. Ammonium [SAF16, WMSH13]. Among [CGL18, DIK19, AW10, Aka10, ASL^{+}13, BSTC13, CSF12, DWW17, GKS13, GSCR11, HB15, HNPB15, JM15, Kae13, KTHM13, LSQ12, OLM15, RS11, SFH^{+}13, SPT^{+}12, SQGK10, SEP^{+}19, VWB^{+}11]. Among-Reader [DIK19].

Amphibian [FBB18]. Amphidromous [CK10, SS13b, SMCS14]. Amphipods [FRH10]. Amplified [SDF12]. Anadromous [BSM17, BBC^{+}13, CJ14, HMBT16, HHP^{+}12b, HH15, LPY16, OEZ17, WWB^{+}17]. Anal [BKM19]. Analyses [ACRBB10, DPK^{+}17]. Analysis [AHVJ11, BBD19, CAS16, DCL19, FB14, FJB16, FBC^{+}10, HTPS10, HH12, HM12, HWZ19, JHS15b, LMC^{+}16, LS10, LL12, MBBS10, PHS^{+}13, RR17, STW^{+}12, SS13b, SMCS14, SBB19, TBIJ19, VMA^{+}16, WHB^{+}14, WTB^{+}15]. Anatolia [GPLT12]. Anatomical [DBKG14, ZWH^{+}15]. Anchor [BF14, BR12, HHP12a, KBK13b, RWPH11]. Anchored [UWL13]. Anchoring [BSH10]. Androscoggin [WBZ19]. Anesthesia [CGM13, VMC^{+}11]. Anesthetize [DBW13]. Anesthetizing [HJK11]. Angled [KB11, LLW^{+}18, TS11]. Angler [AMQSK15, APP13, AJ13, AWG^{+}18, BvPCH17, BHHA11, BRWQ18, BW17, Car13, Cle15, DHH17, GBL^{+}13, HV18, HHSB13, HMD16, LM16b, MP14, McC15, McC17, MS14, PZC^{+}12, QA16, RD10, RPJ18, See11, SSE19, SQS^{+}17, SLR10, THI^{+}19, WQP13, WAW^{+}19]. Angler-Caught [BRWQ18]. Anglers [BCBW18, BHHA11, BS11b, CSF12, DA12, FBD^{+}19, GSG18, HAW^{+}18, HAS^{+}19, HM18, LM16a, LS10, LN12b, MEL^{+}12, PZC^{+}12, PLUH19, SF14, SFH^{+}13, VPV18]. Angling [AJ10, BRN16, BLSS19, BLM^{+}11, BGHL10, DLSW12, DA12, EBS18, FSW18, HBS^{+}19, KDA^{+}18, MSE^{+}18, PB13, RSQ^{+}19, SFH12, SMP^{+}14, SK18b, SBL^{+}18b, SMS13, TLL^{+}18]. Anguilla [DA12]. Animal [Hen15]. Annual [BACRB14, CMC^{+}19, HPA^{+}11, SCS17a, SCWP17]. Annuli [RHPH14]. Annulus [BC14, BK12]. Antenna [BRB^{+}17, BSTC13, FAW14, KCS^{+}17]. Antennas [KZ12, OHL10, SBL11]. Anthropogenic [ACMB^{+}14]. Antimycin [BJZ11, BHG^{+}13, FSS^{+}10, GKMC15, HMW^{+}10]. Antimycin-A [BHG^{+}13]. antipodarum [SAF16]. Apache [BB19]. Apex [SEB^{+}19]. Appalachia [HAS^{+}19]. Appalachian [DP13, HH17, KKM16]. Appalachians [HMP10]. Apparatus [SE12]. Appear [Bal15]. Appendaged [BDH^{+}14]. Applicability [MB17, OH15]. Application [ASD^{+}18, BLM^{+}11, CHJ^{+}16, FBC^{+}10, HD14, HO13, McC17, NBP15, PLCG13, PW13, SSK19, ZM17, WHC14b]. Applications [BKSH^{+}12, GMG^{+}16a, SOS^{+}15]. Applied [DGS^{+}11]. Applying [HWB11, JDH16]. Approach [BGCS11, CDD^{+}11, DBK11, DZ17, DRC17, FJB16, FCH^{+}13, GWCA10, HV18, HWWT15, LC14, MAM^{+}12, MMHM17, SSK19, SBVB12, SL17a, XM14]. Approaches [RÅW18, STSH10]. Approaching [WBZ19]. Aquacalm [CGM13]. Aquaculture [HH12, HL15]. Aquatic [CJA^{+}13, DFMP18, GKMC15, HSM^{+}18, MI10, MKWM18, RWAJ16, SPQ12, TL15, TKH^{+}16]. AQUI [CGM13, FGS13]. AQUI-S [CGM13, FGS13]. Arapaima [HL15]. Arctic [ASP^{+}17, GD10, GAH^{+}18, HMBT16, SSS15, SZKK10, SS13b]. Area [HW16, MCP15, Sam19, HM17, HDM15]. Areas [SM19, TFP^{+}11]. Argentine [RGOR10]. Argyrosomus [BFLT17]. Arizona [AKP15, GDP14, SK18b]. Arkansas [EJL10, HGW10, JHB10, LSQ12, Qui10, QA16]. Arm [BFS^{+}19]. Arrays [BGM17, BSTC13]. Artificial [APD16, BOM16, Blo13, CKBP19, DDAS14, HQ10, JHS15a, KPL12, SDKS16, SBL^{+}12]. Artisanal [NEH10]. Ascending [MSK19a]. Asian [CBDW15, LMC^{+}16, LKM^{+}17, PW13, SGCB19, ZS16]. Asp [FG13]. Aspius [FG13]. Assay [DSJS16, MEJ10]. Assemblage [BPQ13, FQ19, SHL16, TKH^{+}16]. Assemblages [DWW12, DV17, HMCT16, MSFS12, MCTM14, NQ11, NW12, PP11, SS17, SHMJ19a, SHMJ19b]. Assess [APP13, GHL12, HBS^{+}19, KD10, LSC^{+}16, LS11, MSW^{+}14, NKC16, SHM^{+}10, SK18b, WKS^{+}10]. Assessed [SOS^{+}15]. Assessing [BvPCH17, CEH^{+}15, CC15, DBJD12, DGS^{+}11, FOH^{+}12, FCH^{+}13, FOH^{+}15, GMG^{+}16a, Ger11, GTB^{+}19, HFCH15, HNPB15, MS16, MAM^{+}12, MCB^{+}19, NW12, PSCL13, RH14, RCGF18, RFW^{+}10, RFW^{+}11, RV12, RTMJ10, RZ10, RJC11, RPJ18, Sin18, THZ^{+}17, BZ13]. Assessment [Ask16, BBG19, BCD^{+}10, BKM16, BEK^{+}18, BCMB14, CBC^{+}16, CVD^{+}14, CST^{+}18, DBS19, Der14, DC13, FC17, FHH16, GWB^{+}10, IWB10, KKSH12, KHH13, LBB16, MFLS18, MR16, NEH10, PLCG13, PLUH19, PS10, RWAJ16, SSK19, SEP^{+}14, SMW12, SPK10, TKH^{+}16, WWB^{+}17, ZWS^{+}19, ZKM13]. Assessments [PCP^{+}19]. Assigning [CBHB16]. Assignment [WWP18]. Assignments [CHDW18]. Assisted [MR16]. Associated [BDC^{+}13, Bur11, PFD17, SVFG10, ZDC14]. Associations [ACRBB10, AJ10, BPQ13, GAH^{+}18, PSB^{+}10a, SQP12, WHC12]. Assumption [WWB^{+}17]. Assumptions [HMD16]. Atchafalaya [PKKR10]. Atlantic [AES11, CS17, GSCR11, RKB15, AZE^{+}18, BMM^{+}12, BRS^{+}12, Bal15, BSS^{+}14, BC18, BG12, BG19, CC10, CHL13, DHA12, DJS^{+}16, FDM^{+}13, GBHO11, GBL15, GSG18, GL16, JM15, KSM^{+}18, LBL11, McB14, MFRS11, MSHC15, MHC10, NMG^{+}17, SW17, SHL16, SZHC15, SMW12, SDL^{+}15, VPV18, VKH^{+}19, WS11a, WHB^{+}14, WBF^{+}15, ZDC14, ZC15, ZOC^{+}14]. Attaching [BF13]. Attitudes [DA12, HHSB13]. Attractant [BDS^{+}11]. Attractants [CWB16]. Attributable [CC10]. Attributes [MPL^{+}15]. Australasian [HCL^{+}12]. Australia [FWG15, FWR^{+}15b, GMB10, HADG10]. Authors [Ano11, Ano12c, Ano13, Ano15, Ano19h, Car14, MS17b]. Automated [KPL^{+}16, LKB19, NTJ^{+}14, WA14]. Autonomous [TMC10]. Autumn [WMB19]. Auxiliary [FB14]. Availability [RSI16, RKB15, SG13]. Available [McC17, SAF16]. Avian [ERC^{+}11, FOH^{+}12, FCH^{+}13]. Avoid [MQJK19, WMB^{+}15]. Avoidance [WLL^{+}15]. Axial [ABČ^{+}15]. Axial-Flow [ABČ^{+}15].

Back [MH13]. Back-Calculate [MH13]. Backpack [DT11, DTS19, HBDA13, SLP13]. Backwater [CBDW15]. Backwaters [DM12]. Bacterial [MEJ10]. Bad [JM12b]. Bag [Ask16, BGCF11, MM18b, ORL^{+}16, Ryp15]. Baited [SMS13]. Balanced [MMHM17]. Balancing [SCA11]. Banff [PT15]. Bank [SBG^{+}19]. Bar [BF14, BR12, HHP12a, RWPH11]. Barbed [Blo13, RS11]. Barbless [Blo13, RS11]. Barcoding [KIST16]. Bargaining [Lee12]. Barge [MSD11]. Barotrauma [CGL18, ES17]. Barramundi [HH12]. Barrenlands [CEH^{+}15]. Barrier [FWT14, MJH17, PMP^{+}14]. Barriers [TWL^{+}17]. Based [AM17, AWG^{+}18, BKHS14, BP15, CJ14, CHDW18, DRC17, EWS^{+}17, HV18, HWNR14, HCH13, HKP^{+}14, JDH16, LS13, MRC12, MKHC17, MEJ10, MG10, NEH10, PCT^{+}14, PHS^{+}13, PLUH19, PCP^{+}19, PDH^{+}19, RGH^{+}13, RBP16, SLB10, SPW^{+}12, SLS16, WHMS14, WPJ^{+}18, WGS16, ND15]. Baseline [WBB10]. Basin [ADMC10, FSS^{+}10, HDKB10, HMW^{+}10, HER^{+}12, HADG10, KEC11, MSM^{+}14, PKKR10, TFH^{+}16, ZEM11, ZLB10, ACRBB10, AM17, BSS^{+}17, BGD^{+}19, BWT^{+}13, BSG10, Cle15, DCS^{+}16, EHP^{+}16, FDHWP16, HLM^{+}18, ILBJ12, KACS18, MCK^{+}15, MRA14, RS18, SBW^{+}17a, WPJ^{+}18, WBH13]. Basins [RBB10a, UACZ^{+}16]. Bass [ABVS14, ADD14, ALCD17, AZE^{+}18, BDH^{+}14, BLSS19, BKME17, BKM19, BBC10, BBC11, BBS19, BRA^{+}18b, BRA18a, BE11, Bre13, BEH^{+}19, BRSA15, CGDB14, CI10, CF18, CS17, DVWS10, DWB^{+}15, DLSW12, DDEW18, DAKP13, DLC11, DN11, FC17, FSW18, FHE^{+}12, GKS13, HGW10, HWDA12, HBS^{+}19, HAW^{+}18, HRKB19, HBS14a, HVW^{+}17, HB14b, ITY13, IMM13, JHU^{+}13, JHB10, KB10, KIST16, KDA^{+}18, KADH15, KHDA14, LGC11, LAA^{+}18, LHS^{+}14, LN12c, LSK^{+}14, MMS^{+}17, MASN16, MSK18, MMB19, MCS^{+}17, MB17, NTJ15, NSW16, NBFL17, NKL18, PSB^{+}10a, PB13, PLCN17, PTD10, RD10, SHMS16, SG13, Sam16, Sam19, SPP^{+}18, SBW^{+}17a, SBL^{+}18a, SLK15, SHK^{+}17, SL15a, SB18, SOBW10, SKBK12, SQS^{+}17, SVD^{+}18, SS16, SSP^{+}19, SHS15, TTS^{+}18, TLSB18, TGA15, TPB^{+}16, THI^{+}19, TPL18, THZ^{+}17, TP17, WBG^{+}17, WI17]. Bass [WS11b, WMB19, WWD13, WCWA10]. Batch [GL16, WGM^{+}11]. Bathythermal [BATK16]. Battery [HBDA13]. Bay [BGCF11, BSB^{+}15, MKWM18, SEB^{+}19, WHB^{+}14, BLM10, CKCF12, ERC^{+}11, FB14, FJB16, FBM^{+}19, HWWT15, JHU^{+}13, KTN14, LN12c, LBL11, OHSP17, SBD^{+}11, TO10, UML^{+}11]. Bayesian [BGCS11, BC17, GRSF^{+}16, SBL^{+}18a, SBVB12, VMA^{+}16]. Bayou [HGW10]. Be [BFLT17, HTF^{+}13, SH18]. Beach [JJ10]. Beam [HH15]. Beaver [JBRWH18]. Bed [AD13, AHAC15, PRR^{+}13, SCB^{+}19, WBK^{+}10, WBK^{+}13]. Bedding [HAW^{+}18]. Before [SHMJ19a, HWZ19, SHMJ19b, WBB10]. Behavior [ABČ^{+}15, DDEW18, HB15, HW16, HHP^{+}12b, ITY13, JPWR10, LGC11, MCTM14, RHC^{+}14, SDKS16, STP19, SOBW10, SQS^{+}17, VJK13, WRNJ^{+}19]. Behavioral [BLM^{+}11, SWG^{+}16]. Behaviors [CGL18, HAW^{+}18, MCK^{+}15]. Being [GLH^{+}18]. Believed [BHHA11]. below [MH11a]. Benchmarks [HB11]. Benefit [AHVJ11]. Benefits [CSWW18, ES17, GGC17, SPW^{+}12]. Benthic [RWT17, SWHA15, TGKZ19]. BENZOAK [CGM13]. Bering [SGF^{+}13, Sta12, SC14]. Bertalanffy [OI17]. Better [LP17]. Between [APD^{+}18, KIST16, AJ10, BMM^{+}12, BHN11, BK12, CEP15, CBC^{+}17, CAWB17, DA12, FBB19, FOH^{+}15, HPD^{+}19, HDB^{+}12, HFH12, HFH18, JBZ^{+}10, KPL12, KTN14, KHDA14, KEC11, LDP^{+}10, LLP10, LDW10, LM16b, MHBR19, PDM12, PDL^{+}18, RW11, RWAJ16, SLR^{+}16, SPWC15, UWL13, WCWA10]. Beverton [CBS13]. Bias [HBS14a, MQJK19, MQS13b, OLM15, OM10, PC12, RFW^{+}10, SR16, SLR10, SCPC14, WHC14a, WHC14b, YC11]. Biased [TP17]. Biases [Ger10, Ger11, RFW^{+}11, RV12, SCPC14]. Big [DC13]. Bigger [LP17]. Bighead [CSWP19, LDP^{+}10, RJP^{+}15]. Bight [DJS^{+}16, CS17]. Billy [RSHS15]. Bingham [BCD^{+}10]. Bioacoustic [FWT14]. Bioassessment [OH15]. Bioeconomic [BLM10, HWWT15]. Bioelectrical [SMCS14]. Bioenergetic [RBP16, SOSL10]. Bioenergetic-Based [RBP16]. Bioenergetics [FCH^{+}13, GHK19, SHK^{+}17]. Biological [AW10, JHU^{+}13, LSL^{+}13, PS10, SCA11, WBT17, ZKM13]. Biology [Gre11, TS13]. Biomass [Der14, ESL^{+}15, HW10, JWC10, KW13, Kra18]. Biota [DFMP18]. Biotic [MSK19b, OH15, RHG13]. Bitterroot [NEC^{+}13]. Black [ABVS14, BK16, BDH^{+}14, CS17, FBB19, ITT10, IWB10, ISC10, Ise10, IMM13, KCCW19, LSMD15, MMS^{+}17, MIH15, PDM12, Sam19, SHS15, TSNH10, JWC10]. Blackfish [FHL18]. Blackfoot [TGS^{+}15]. Blacknose [CC10]. Block [MJH17]. Blood [DP13]. Blue [AOS^{+}17, ASD^{+}18, BS10, BF13, BF14, BLM10, GFM^{+}18, HBLL12, HWWT15, NLK17, RH16, SS13a, ZC10]. Blueback [DR10, DG12, Gur12, GBTL14]. Bluefin [GSG18, MG14]. Bluegill [BW18, BLSS19, HWDA12, Ryp15, SEP^{+}19, THI^{+}19]. Bluegills [KWMB11, KMPP12, NKL18, RLC12, WWD13]. Bluehead [HUST17]. Boat [ADMC10, BS12, BS11b, HPD^{+}19, MQJK19, MK13, MBPG12, NQ11, PA16, THZ^{+}17]. Boats [MK13]. Bodied [AD13, KCS^{+}17, RPC^{+}19, WBK^{+}10, WBK^{+}13, BGW^{+}11]. Bodies [BEH^{+}19]. Body [NKL18, Ran18, SKBB15]. Boise [ND11]. Bone [WSA18]. Bones [PC12]. Bonneville [BWR12]. Bonytail [BSS^{+}17, DHK^{+}19, SCG11b]. Boot [SM13]. Boreal [McG14]. Both [FMA^{+}12, ZS16]. Bowfishing [Qui10]. Branch [Kra18]. Brand [NLK17]. Branding [Bea17, DK16, Dag17]. BRD [ASD^{+}18, GFM^{+}18]. BRDs [GFM^{+}18]. Breton [GBL15]. Brief [Bal15]. Brink [GBHO11]. British [BWLB12, BCRM^{+}14a, BCRM^{+}14b, PRTS14, BRBP10, MMP12, VMA^{+}16]. Broad [MSC13, PLCG13]. Broad-Scale [MSC13, PLCG13]. Broodstock [BKHS14, BEK^{+}18, MW19, MEJ10]. Brook [AW10, BTD^{+}12, BHG^{+}13, CCMA^{+}10, DKWG14, GFP^{+}12, HNH11, HH17, HAS^{+}19, HK17, HDM15, HBA^{+}12, KKM16, KHH13, KSH^{+}16, KMK^{+}15, KW13, Kra18, LKB19, LSL^{+}13, LLW^{+}18, MJ11, MSC13, OL17, PFRM12, PVST11, RZ10, SMH17, SRSS12, SNTZ14, SQGK10, TCM19, WDD^{+}14, WWP18]. Brown [ACS19, DWCK16, DTA18, DSD^{+}13, HM17, JWC10, Mat10, MSC13, MB12, NWZ14, QK11, SHSCA16, WMSV18, JM15]. Brunswick [LAA^{+}18, SDL^{+}15, ALCD17, WBZ19]. Bubble [FWT14, GLCD13, MJH17, SWW14, ZS16]. Budd [TPBF15]. Budgetary [GTB^{+}19]. Built [DTS19]. Bulkley [TEB^{+}19]. Bull [ACRBB10, ASP^{+}17, BPB13, BBD19, BB12, BPT^{+}19, BHG^{+}13, DB13, DGS^{+}11, ES14, GMG^{+}16a, HS12, MAM^{+}12, MHDD16, MGS14, NEC^{+}13, PVST11, RSHS15, WGJ10]. Burbot [AAYC14, AJR^{+}19, BQHR17, BRWQ18, BQR^{+}19, GM12, KQRS16, NCJ^{+}11]. Burden [BHC^{+}10, LFD^{+}17]. butcheri [GMB10]. Bycatch [ASD^{+}18, BFP12, BATK16, BSS^{+}14, BM14, BC18, GGC17, GFM^{+}18, MCH12, RGC^{+}12, RJMG13, SMW12, ZC15]. Bypass [APPR14].

C [HTPS10]. Cage [SC18]. calcarifer [HH12]. Calcein [BFLT17, CMF^{+}14, EKM10, HQ10, LP12, MKRH14, WMD^{+}19]. Calcified [BCZ^{+}18, BKM16, BMW^{+}19, OLM15, WMD^{+}19]. Calculate [MH13]. Calculating [CHBY13]. Calibration [WKP14]. California [BS11a, BBS18, DK16, DBE11, FOH^{+}15, GWCA10, GRR^{+}17, HDB^{+}12, HW10, HPM^{+}18, MGB^{+}14, PRM15, RR17, SHMS16]. californicus [HDB^{+}12]. Callinectes [AOS^{+}17, ASD^{+}18]. Calling [HBS^{+}19]. Camera [AWG^{+}18, MI10]. Cameras [PA16, SPW^{+}12, SK18b]. Can [FBL^{+}19, HTF^{+}13, Hen15, HVW^{+}17, MQJK19, MSK19b, RW11, SCW13]. Canada [CJH16, GD10, GBHO11, HMBT16, Hax15, LAA^{+}18, PT15, SPHK16, SDL^{+}15, VMA^{+}16, WHB^{+}14]. Candidate [WMSH13]. Cannot [LMC^{+}16]. Canyon [GDP14, PHS^{+}13]. Capacity [GCA^{+}13]. Cape [GBL15]. Captive [BDC^{+}13, VJK13]. Captive-Reared [VJK13]. Capture [Bal15, Blo13, BS10, CCDF18, GHL12, HMP10, JHK11, KD10, PP10, SWHA15, TRB^{+}13, TEB^{+}19, YS12]. Captured [KHBH19, McC16, SS11b]. Captures [KCCW19]. Capturing [CDBW17, HPD^{+}19]. Carbon [CST^{+}18, KAZ15, VMC^{+}11]. Carcass [DPS^{+}17, HW10, MPM10]. Carcasses [SHM^{+}10]. Caribbean [CK10]. Carolina [DBJD12, GRCR10, RH14, AS10, BR14, FMP19, HH11a, HH11b, HH15, ORA17, WHC12]. Carp [BS12, BGBG17, BFS^{+}19, BDB^{+}19, CSWP19, CWB16, CBHS16, CBDW15, CDBW17, CPSG12, DHLD11, HPD^{+}19, HB14a, KML14, KCCW19, LDP^{+}10, LKV^{+}17, LMC^{+}16, LKM^{+}17, MKBB13, PW13, RSLG10, RHC^{+}14, RJP^{+}15, SP13, SDF^{+}13, SJM15, SC18, SPWC15, SCWP17, WSH^{+}14, WALL17, WKTQ15, WHB11, ZGW11, ZS16]. Case [ACMB^{+}14, BACRB14, BT16, CWH^{+}14, DASB16, FJB16, GHK19, HVW^{+}17, MKBB13, MWS15, MCS^{+}17, MSK19a, MHC10, RWAJ16, RGOR10, SBL^{+}18a, SDF12]. Caspian [ERC^{+}11]. Cataraft [HHMQ15]. Cataraft-Mounted [HHMQ15]. Catch [AS10, Aka10, BHHA11, BGCH17, BHN11, BDH^{+}14, BDSB15, BGHL10, BMMMLML15, CvPW15, CWB16, CSF12, CI10, CBDW15, CKCF12, DKWG14, DLSW12, ES17, EBS18, FF15, FC17, FSW18, FB14, FWR^{+}15b, GBL^{+}13, HHSB13, KD10, Kro19, LM16a, LFD^{+}17, LJR11, MQJK19, MMS^{+}17, MQS12, McC15, McC16, MH11a, MSK18, PZC^{+}12, PDM12, PB13, PA16, PS10, Pul17, RR17, RZ10, RGOR10, RPJ18, RSQ^{+}19, Sam16, Sam19, SCS17a, SBD^{+}11, SS11a, SLK15, SMP^{+}14, SBC^{+}11, See11, STW^{+}12, SCW^{+}17, SL12, SC14, SBL^{+}18b, SEP^{+}19, WHC14a, WHC14b]. Catch-And-Release [LFD^{+}17, BGHL10, CvPW15, DKWG14, ES17, EBS18, FSW18, GBL^{+}13, MMS^{+}17, McC16, MH11a, PB13, RZ10, RSQ^{+}19, Sam19, SMP^{+}14, SBL^{+}18b]. Catch-at-Age [FB14]. Catch-Curve [STW^{+}12]. Catch-Free [FC17]. Catch-Oriented [BHHA11]. Catch-per-Unit-Effort [WHC14a, WHC14b]. Catch-Related [CvPW15, HHSB13]. Catchability [KSH^{+}16, THZ^{+}17, WEC^{+}19]. Catchable [KM11, MSK^{+}16]. Catchable-Sized [MSK^{+}16]. Catches [BDS^{+}11, CS17, Mat10, OMT^{+}15, SE12, SEP11, TSNH10]. Catching [CL12]. Categories [BBC10, BBC11, BBS19, CAWB17, DN11]. Catfish [BSL18, BS10, BF13, BF14, BSO^{+}16, CPM^{+}10, EPM16, FMP19, HGW16, HBLL12, HHSB13, LDW10, MCSF^{+}17, NLK17, PDD^{+}18, QA16, SS13a, SEB^{+}19, SPHK16, Sin18, SL12, ZC10]. Catostomid [SRFS15]. Catostomus [UMC14]. Catskill [GBFR18]. Caught [BEK^{+}18, BRWQ18, CMWM18, GMG16b, GMB10, Kae13, RHC^{+}14, RS11, RSQH18, SS13a, SVFG10]. Caught-and-Released [RHC^{+}14, RSQH18]. Causes [MHC10]. Ceded [HS19, MSE^{+}18]. Central [HAS^{+}19, MGB^{+}14, BPQ13, BS11a, BG13, CBM13, FWWTJD16, FOH^{+}15, HWDA12, HMBT16, HMP10, MP14, PLCG13, SHMS16, SSK18, WFWSV14]. Centuries [Hen15]. Ceratomyxa [HHB12]. cerebralis [ZLB10]. Certain [HW12]. Certification [ZGW11]. Cessation [SDF12]. Chain [BCZ^{+}18, HSHH17]. Challenges [CKC^{+}12, KDP^{+}10]. Champlain [HMS11, MMS^{+}17, PPO18]. Change [BB19, DN11, KLR^{+}12, PM14, RFHP17]. Changes [BRBP10, CBC^{+}16, CBC^{+}17, ES14, GBFR18, HH17, HMS11, NWZ14, RPBM10, SBC^{+}11, SHMJ19a, VEW^{+}13, WH12]. Changing [HGK15]. Channel [BSL18, Bre13, CPM^{+}10, DHB^{+}10, EPM16, HWZ19, LDW10, MRC12, NLK17, PDD^{+}18, QA16, SB15, SPHK16, Sin18, SD11, SL12, TRB11]. Channels [BSH10]. Char [GD10, HMBT16, Nak13]. Characteristics [AJ10, BQR^{+}19, DHH17, DGF11, GD10, HHR^{+}18, Hax15, HPM^{+}18, HEL^{+}15, HDM15, JHB10, KRB^{+}16, KQRS16, MHDD16, MKHC17, NKC16, PVST11, PDH^{+}19, UACZ^{+}16, WLL^{+}15, WQP13, WKP14]. Characterization [BSS^{+}14, HS19, HPA^{+}11, HM18]. Characterize [BLM^{+}11, Mea12]. Characterizing [GSG18, MSE^{+}18, PBA^{+}17, SS17, THI^{+}19]. Checks [CHDW18]. Chemical [FMA^{+}12, GPP^{+}10, JTB16]. Chemistry [DP13, RCGF18, WPJ^{+}18]. Chesapeake [OHSP17, BLM10, HWWT15, JHU^{+}13, KTN14, LN12c, LBL11, SBD^{+}11, SEB^{+}19, TO10, UML^{+}11]. Chewing [LKV^{+}17]. Chinook [RSHS15, AT18, BSM17, BHC^{+}10, BC17, BSG10, BSBF13, BBS18, CBS^{+}12, CWH^{+}14, CSH^{+}17, CBC^{+}16, CBC^{+}17, CCC16, CMD12, FHC16, FBL^{+}19, FCE^{+}12, GLH^{+}18, HD14, HLM^{+}18, HJB14, HPM^{+}18, HWFN14, HQ10, HRH10, JKN^{+}10, JDG10, JF13, JF14, JNC^{+}16, KJN^{+}17, KAZ15, KRB^{+}16, KTHM13, KPP^{+}18, KSL^{+}19, MS18, MNYV12, MFLS18, MQS12, MQS13a, MQS13b, MSD11, MSS^{+}12, MGB^{+}14, MEJ10, MPM10, NKC^{+}18, NB10, PRM15, PSB^{+}10b, PFD17, RM14, RCP^{+}13, SHMS16, SEP^{+}14, SM19, SLG^{+}15, TFP^{+}11, TBH^{+}14, TPC^{+}15, VWB^{+}11, VEW^{+}13, VJK13, WWS^{+}14, WS16, WHMS14, ZWS^{+}19]. Chloride [MKK^{+}12, SCG11b]. Choice [CSF12, DC13, MQJK19]. Choke [LKM^{+}17]. Chub [DASB16, GDP14, GPLT12, GFP^{+}12, HB16, HW19, PHS^{+}13, UMC14, WPY^{+}15]. Chum [CJ14, FBC^{+}10, GLCD13, KSMN18, POWF11, SOS^{+}15]. Chutes [GMG^{+}15]. Cichla [NBFL17]. Cichlid [BFN16]. Circle [BDH^{+}14, BMMMLML15, KSH^{+}16, SKH^{+}16, SMS13]. Cisco [FBM^{+}19, KMPV13, MRB^{+}11, RHG12, RHG13]. Clackamas [MAM^{+}12]. Clair [KYD18, BSB^{+}15]. Clam [FWG15]. Clark [DB13]. clarkii [LPY16, LPY16]. Class [FC17, HFH12, HFH18, LOV17, LO18, NTJ15]. Classes [DS12]. Classification [CSWP19]. Claw [DKSW15]. Clear [BE11]. Clearwater [DCS^{+}16, FBD^{+}19]. Climate [ACMB^{+}14, BB19, KTN14, KLR^{+}12, PM14, RFHP17]. Closure [ES14, HMS11, Mat10, WWB^{+}17]. Cluster [DCS11]. Coarse [LGC11, WWW15, WS11b]. Coast [DPK^{+}17, CDD^{+}11, DC13, HR11a, LS17, McB14, WHB^{+}14, WBF^{+}15]. Coastal [BR14, BS11a, BA18, BRA^{+}18b, BRA18a, BRSA15, BKP^{+}14, CBWB14, JBZ^{+}10, KML14, LPY16, MS17a, PSCL13, Pto13, RGC^{+}12, RMTW16, SPM15, SBTW17, VBBK16, ZC15, GWCA10]. Coaster [LSL^{+}13]. Cod [BC18, DHA12, FWR^{+}15b, FWR^{+}15a, MFRS11, ZDC14, ZC15]. Coded [AAYC14, BSL18, BR12, ERC^{+}11, GL16, JSDJ16, KPL^{+}16, KBH^{+}19, LKB19, LXM^{+}12]. Coded-Wire-Tagged [KBH^{+}19]. Coded-Wire-Tagging [JSDJ16]. Coeur [SD11]. Coexistence [PJLB14]. Cogradient [APD^{+}18]. Coho [BCD^{+}10, BKP^{+}14, CBWB14, GDLM17, HLS^{+}17, JM12b, JCB^{+}18, KAZ15, KLR^{+}12, RFQ14, RPBM10]. Cohort [BSM^{+}15, PTD10]. Cohorts [WH11]. Coincide [RRH^{+}11]. Coldwater [FDHWP16, GH17, WMB19]. Collapse [CBC^{+}17]. Collected [NTJ^{+}14, UWL13, WBF^{+}15]. Collecting [BGBG17]. Collection [MW19, NSW16]. Collector [WBGM13, WBF13]. Colonial [EHP^{+}16]. Colony [ERC^{+}11, SCMP13]. Colorado [BSS^{+}17, BWT^{+}13, AKP15, BRN16, BFS^{+}19, CCDF18, GDP14, HHRB15, HFRH17, LN12b, WBH13]. Columbia [CJH16, HLM^{+}18, JBZ^{+}10, KRJT13, NTJ^{+}14, SHMJ19a, SHMJ19b, VMA^{+}16, ACRBB10, AM17, BWLB12, BCRM^{+}14a, BCRM^{+}14b, BKSH^{+}12, BRBP10, HWFN14, JKN^{+}10, JKSK13, MMP12, MRA14, PRTS14, RS18, SEP^{+}14, SWH^{+}16, SM19, TBH^{+}14]. Combinations [WGM^{+}11]. Combined [RFHP17]. Combining [HWNR14, HH15, WMB^{+}15]. Comment [ABP^{+}18, AD13, BCRM^{+}14b, Bea17, Can17, Dag17, Erm12, FSS^{+}10, FSV12, GGCF14, Ger11, HMW^{+}10, HBS14b, PRTS14, RFW^{+}11, RV12, Rey15, SCPC14, SL15b, Vin14, WBK^{+}13, WCY^{+}18]. Commercial [ASS17, BFP12, BATK16, CPSG12, CH17, DCP^{+}19, GRCR10, GMB10, HMBT16, HMS11, JBHM19, MBM13, OMT^{+}15, RB15, SS11b, SC18, WHB11].

Commercially [SAF16]. Commitment [DA12]. Common [ASL^{+}16, BS12, BA11, CWB16, CBHB16, CBHS16, CPM^{+}10, CPSG12, HB14a, HMP10, RHC^{+}14, RM12, SQS^{+}17, SLG^{+}15, SC18, WKTQ15, WHB11, ZS16]. Common-Garden [SLG^{+}15]. Commonwealth [ASS17]. Communicating [FHH16]. Communities [AHAC15, BSBC14, Mea12, MOJ12, PBA^{+}17]. Community [BBB^{+}14, CJA^{+}13, KMB^{+}12, MKWM18, PKKR10, PRRD15, SVPD13, SSPP19, WKS^{+}10]. Comparative [BTD^{+}12, FSS^{+}10, HMW^{+}10, UMC14]. Compared [DLSW12, NSW16, PP11, TPL18]. Comparing [GGCF14, KRB^{+}16, LS13, LN12b, MHAO11, ORC^{+}16, SF14, TP17, UBR^{+}17, Vin14]. Comparison [APF18, AHAC15, BA18, BGCF11, BCZ17, BFN16, BKME17, BA11, BGW^{+}17, BGM17, BSTC13, BBC^{+}13, CSV13, CBDW15, CDBW17, CPM^{+}10, DSJS16, DCHC13, EJL10, FSW16, FCE^{+}12, hGgXfC^{+}13, HPD^{+}19, HHMQ15, JPWR10, KPL12, KSH^{+}16, KAZ15, KSG^{+}18, KDMH10, KYD18, Kro19, MHBR19, NQ11, NCMT12, PDM12, PBA^{+}17, PN19, PVH^{+}13, PTD10, PDL^{+}18, RBB10a, RBB^{+}10b, RMTW16, SKH^{+}16, STSH10, SBL11, SS17, SGC13, TSFS19, WKS^{+}10, WI17, WD10, WMD^{+}19, ZWH^{+}15]. Comparisons [Ano12b, ITT10, JY11, OLM15, PC12, TSNH10]. Compatibility [RBB^{+}10b]. Compensatory [CA11, SB11]. Competition [MSK19b]. Complete [RH14]. Complex [BSVB19, FBC^{+}10, HCH13, PT15, RBB10a, RPJ18, SSS15, SHMJ19a, SHMJ19b]. Complexes [CDD^{+}11, DSB19, HCH13]. Complexity [HKMH10, SS16, WWW15, ZEM11]. Compliance [Car13, SWH^{+}16]. Composition [DLC11, JKN^{+}10, PLCN17, PTD10, PHFK10, SCKZ17, SHL16]. Compounds [DBW13, SAF16]. Compression [LMS^{+}18]. Computed [MHBR19]. Concealment [HW16]. Concentration [LLP10]. Concentrations [HPA^{+}11, RRH^{+}11]. Concern [CI10]. Concurrently [BvPCH17]. Condition [CF18, CGG^{+}18, HH17, KJN^{+}17, KSG^{+}18, LS11, NKL18, SCG11a]. Condition-Dependent [KJN^{+}17]. Conditional [SWHA15]. Conditioned [TPL18]. Conditions [CSF12, HHR^{+}18, HKP^{+}14, JWC10, LMS^{+}18, McG14, PPFC18, RHC^{+}14, SCB^{+}19, UML^{+}11, VJK13]. Conductivity [DT11, HKMH10, LCR19, MKK^{+}12]. Configurations [SBW^{+}16, VKBL11]. Conflicts [BKHS14]. Congruent [OHSP17]. Conjoint [MBBS10]. Connecticut [BBB^{+}14]. Connectivity [AW10, MC10, RH14, WFWSV14]. Consensus [DRC17]. Consequences [Der14, DKSW15, GGC17, HBA^{+}12, LSL^{+}13, PM14, RHC^{+}14, SOSL10, SC14, TEB^{+}19, TP17]. Conservation [ACMB^{+}14, BACRB14, BKHS14, CCMA^{+}10, CDQG12, DA12, GMG^{+}16a, HS19, MAG^{+}15, PT11, SPQ12, SQP12, SHM^{+}10, TPB^{+}16, WQF15, WHMS14]. Conservative [RD10]. Conserving [CGL18]. Consideration [WCWA10]. Consistency [KTHM13]. Consistently [DZ17, FBB19]. Consisting [MJH17]. Conspecifics [SM14]. Constant [HB14b]. Constraints [Bre13, GTB^{+}19]. Constructed [GMG^{+}15, PQB10].

Construction [HWZ19, PFRM12]. Consumed [SHK^{+}17, TGA15]. Consumers [WHC12]. Consumption [CSWW18, FH16, GHK19, LRHR17, SCMP13]. Consumption-Oriented [FH16, LRHR17]. Consumptive [HGW16]. Contact [SBW^{+}17b]. Content [NSW16, SMCS14]. Context [BHHA11, DHH17, HCL^{+}12, PCP^{+}19, Rey15, SL15b]. Context-Dependent [HCL^{+}12, PCP^{+}19]. Continuous [TPBF15, TKH^{+}16]. Contrast [RGH^{+}13]. Contrasting [DWCK16, HMCT16, Sta12, SPWC15]. Contribute [HVW^{+}17]. Contribution [LAM16]. Contributions [BSB^{+}15, DSB19, OMT^{+}15]. Control [ACMB^{+}14, BT16, CPSG12, ILBJ12, KML14, RWAJ16, SDF^{+}13, SC18, WWM13, ZC12]. Controlling [SAF16]. Controls [NKC^{+}18, VPV18, WBT17]. Conundrum [Can17]. Conventional [HPD^{+}19, RGOR10]. Convergence [SBG^{+}19]. Conversion [DC13]. Conversions [PSCS11]. Conveyance [MGWC13]. Cool [BB19, CMC^{+}19]. Cool-Season [CMC^{+}19]. Cooling [MMP12]. Cormorant [SCMP13]. Cormorants [MSK^{+}16]. Correcting [BHW13, SLR10]. Correction [SR16]. Correlations [Kae13]. Corrosion [LMS^{+}18]. Corrugated [PSCL13]. Cortisol [FGS13]. Cost [AHVJ11, CBWB14, CBDW15, SPW^{+}12, SGC13]. Cost-Effective [SGC13]. Costs [SMK13]. Count [HWB11, MSW^{+}14]. Counters [SB15]. Counts [AWG^{+}18, HS12, PFRM12, SLB10, SPW^{+}12, SLR10, VBBK16, WKP14, XM14]. County [GWCA10]. Cover [CBWB14, CKBP19, DR19, HW16]. Coverage [NTJ15]. Covered [RJMG13]. Cowlitz [KTHM13]. Crab [AHVJ11, ASD^{+}18, BLM10, GCLC18, GFM^{+}18, HWWT15, MBM13, RJMG13, RH16, SGF^{+}13, SCW^{+}17, TDM19, WHC12]. Crabs [ADMC10, AOS^{+}17, DKSW15]. Crank [FMP19]. Crappie [ABVS14, BK16, ITT10, IWB10, MDM12, MIH15, PCP^{+}19, TSNH10]. Crappies [BDS^{+}11, DM12, FBB19, ISC10, Ise10]. Crayfish [BA18, ETD16, RB15, RD10, WS13]. Created [SCPC14]. Creek [BCD^{+}10, Bri17, CWH^{+}14, DK16, DBE11, RM14, SHK^{+}17]. Creeks [GRCR10]. Creel [AMQSK15, KD10, LSLC15, MQS13a, MQS13b, MM17, MIH15]. Crested [MSK^{+}16, SCMP13]. Criteria [GSCCP18, VMA^{+}16]. Critical [CMMR19]. Croaker [NEH10]. Cross [NEH10]. Crossing [PQB10]. Cryptic [WAW^{+}19]. Cucumber [HDB^{+}12]. Cue [BWHM16]. Cui [SRFW14, SRFS15]. Cui-ui [SRFW14, SRFS15]. Cull [GCLC18, RH16]. Culter [LXM^{+}12, LXM^{+}12]. Cultured [CMF^{+}14, CSH^{+}18]. Culvert [SHMJ19a, SHMJ19b, WWP18]. Culverts [BG13, MSFS12, PSCL13, PN19, PQB10]. Current [DT11, DDD^{+}18, GD10, GFC^{+}15, SW17, SPQ12]. Curtain [ZS16]. Curtains [CSH^{+}17, SWW14]. Curve [STW^{+}12]. Curves [RBP16, WMB^{+}15]. Cusk [ZC15]. Cutthroat [ACS19, APAC12, ASL^{+}13, ASL^{+}16, BACRB14, BBO16, BWR12, BKP^{+}14, CMK^{+}19, DZS10, DCS^{+}16, EMK^{+}17, Gre11, HHRB15, HFRH17, HGK15, JBZ^{+}10, Kae12, Kae13, KEC11, LPY16, MKHC17, ND11, PSCL13, Pto13, RFHP17, RSQ^{+}19, SRSS12, SBVB12, TGS^{+}15, UACZ^{+}16, VBBK16, WQF15, ZRR^{+}19]. Cycle [HPA^{+}11, Pul17]. Cycles [DT11]. Cyprinid [HW19, LMC^{+}16, NASG12, PLCG13]. Cyprinidae [NBP15]. Cyprinids [NBP15].

D [UWL13]. D-Frame [UWL13]. Dace [CC10, PFWG16]. Daily [BR14, GHL12, HB16, HB14b, MIH15, Ryp15, SBS14]. Dakota [BKME17, CWB16, DWB^{+}15, HB14a, JWC10, SDDC16]. d’Alene [SD11]. Dalles [JKSK13, KRJT13]. Dam [APPR14, AZE^{+}18, BPT^{+}19, GLH^{+}18, HHB12, KRJT13, KLR^{+}12, MS18, MPL^{+}15, NMG^{+}17, RH14, SZHC15, STSH10, SM19, SWW14, TCB18, WA14, WBB10, WDW13, FRL16, JKSK13, MH11a, MSK19a, WBZ19]. Damage [LLW^{+}18]. Dams [CKC^{+}12, HH11b, HMCT16, JM12a, DB13]. Dark [SF14]. Darling [HADG10]. Darter [DCS11]. Darters [RTJJ16]. Data [AWG^{+}18, BRN16, BGCH17, BG12, BRWQ18, BC17, CAS16, DPS^{+}17, DCL19, GHL12, GCA^{+}13, GTB^{+}19, HWNR14, LS13, MM18b, MPM10, RFW^{+}10, RKB15, SRCC^{+}18, VEIW^{+}16, WHC14a, WHC14b, WWM13, ZSW11]. Data-Analysis [DCL19]. Data-Limited [SRCC^{+}18]. Data-Poor [BC17, CAS16, WWM13]. Day [BKME17]. Days [Lee12, PTD10]. Days-at-Sea [Lee12]. DC [CC15, DP13, HRH10]. Dead [SLB10]. Death [SJM15]. Decade [MG10]. Decades [Kae13]. Decision [BBD19, Ols19, SLR^{+}16, VMA^{+}16]. Decisions [BLM^{+}11, JM12b]. Decline [KMPV13]. Declines [DRG10]. Decreased [CCC16]. Dee [HH11a]. Deep [BDH^{+}14, CMWM18, RTMJ10, SKH^{+}16, SMS13, TS11]. Deep-Hooking [CMWM18]. Deeply [DVWS10, RLC12]. Deepwater [HR11b, SM14]. Defensible [VEIW^{+}16]. Defining [CDD^{+}11, ILBJ12]. Deflection [ZS16]. Deflector [SMW12]. Degraded [RGH^{+}13]. Delaware [WHB^{+}14, WBF^{+}15]. Delayed [SS13a, WH11]. Deleterious [CCDF18, SH18]. Delta [BBS18, CMF^{+}14, CSH^{+}18, SSS15, WHG^{+}16, BSBF13, NB10, PSB^{+}10b]. deltoides [FWG15]. Demand [HGW16, MBBS10]. Demersal [MFRS11]. Demographic [MHDD16, SMK13]. Demographics [CPM^{+}10, KCF15, MCSF^{+}17, SSP^{+}19, SPWC15]. Demography [DFPS10]. Demonstration [PW13]. Densities [BKP^{+}14, DM12, WS11a]. Density [BW18, BG12, CA11, CF18, DLP16, FHL18, GGCF14, HFH12, HFH18, KZ12, MPL^{+}15, PLCG13, SLK15, SFH12, SW17, SB18, SS16, TWL^{+}17, TRB11, THZ^{+}17, Vin14, WDD^{+}14]. Density-Dependent [GGCF14, SW17, Vin14]. Density-Independent [GGCF14, Vin14]. Dependence [BW18, CA11]. Dependent [APP13, GGCF14, GZ13, Han19, HCL^{+}12, KJN^{+}17, LXM^{+}12, PCP^{+}19, SW17, Vin14, WHC14a, WHC14b]. Deplete [CWB16]. Depleted [OHSP17]. Depletion [DP13, HKMH10, PD13]. Deployment [AG15, TMC10]. Deposition [BMM^{+}12, BGCF11, HB14b, SBS14]. Depredation [DCP^{+}19]. Depth [JHS15a, TFP^{+}11]. Derby [TSFS19]. Derived [BS10, KMPP12, PZC^{+}12]. Deriving [PS10]. Descending [ES17]. Deschutes [ZLB10]. Describe [McG14]. Describing [BSVB19, CJH16, WS11a]. Description [GGC17, KCCW19]. Desert [PSC15, RW11, SLMM10]. Design [ES17, HWB11, JJ10, MQS12, SMW12, SKS^{+}15]. Designed [APD^{+}19]. Designs [BOM16, BFL^{+}10, DRG10, Kro19, MQS13a, MJCM15, SC17]. Detect [DRG10, DCHC13, RFKB17]. Detecting [FAW14, HF13, NCMT12]. Detection [AD13, BHW13, BE11, BSTC13, GRSF^{+}16, HRH10, HEL^{+}15, KCS^{+}17, KDMH10, NSW16, NTJ^{+}14, OHL10, SLSS16, SAYJ17, WBK^{+}10, WBK^{+}13, ZLB10]. Detections [WPS^{+}19]. Detector [BDT^{+}16]. Detectors [HHRB15]. Detects [HBS14a]. Deter [SWW14]. Determinants [BW17, SF16]. Determination [GSCCP18, HT11, Ise10, MAP14, THQ15]. Determine [DRC17, KHH13, NKC16, SLP17]. Determined [MS16]. Determining [BHHA11, BDT^{+}16, BE11, FCE^{+}12, PK18, WGM^{+}11]. Detroit [KYD18]. Develop [BB19]. Developing [LWJS16, MB17, RBP16, ZSW11]. Development [Ano12b, BFLT17, CK10, DV17, ETD16, JY11, LGC11, PW13, RFW^{+}10, RTMJ10, TFP^{+}11, UML^{+}11, ZC15, ZKM13, ZM17]. Device [ASD^{+}18, MRC12, MGWC13]. Devices [GFM^{+}18, RGC^{+}12]. Devils [CFBB15, KPL12]. Diagnose [FDHWP16]. Diameter [Ano12b, JY11]. Diamondback [ASD^{+}18]. Diary [GBL^{+}13]. Did [Mat10]. DIDSON [PNHA12]. Diel [BATK16, HLS^{+}17, JM15, MMB19, RMTW16]. Diet [AZE^{+}18, CGG^{+}18, CJH16, KIST16, PTD10, SGCB19, SKBSA11]. Diets [EF10, KFM12, LMD^{+}19]. Differ [SHS15]. Difference [BMM^{+}12]. Differences [DWW17, LSQ12, PHFK10, SLG^{+}15]. Different [ASL^{+}13, BLM10, CBHS16, CHPP11, DSB19, hGgXfC^{+}13, PPFC18, RS17, SH11, See11, SWG^{+}16]. Differentiating [Pto13, UWL13]. Differentiation [GKS13]. Differing [SBW^{+}16, SCB^{+}19]. Digestive [LMB^{+}18]. Digital [MI10, SBB19]. Diked [SBTW17]. Dimensional [BBG19]. Dimensions [HSG16, LKvR^{+}18]. Dimorphism [MAH^{+}13]. Diodes [SWG^{+}16]. Dioxide [CST^{+}18, KAZ15, VMC^{+}11]. Diploid [BTD^{+}12, CMK^{+}19, KSG^{+}18, KKMS11, KM11]. Direct [DT11, DDD^{+}18, GFC^{+}15, TPBF15]. Directed [DHA12]. Direction [GBTL14]. Directly [CFW^{+}10]. Disaster [AG11]. Discard [BGCH17, CZH^{+}19, FWWTJD16, RGOR10, RPMB19, SBG^{+}19]. Discarded [HR11b, MFRS11]. Discards [GSCCP18, SOB18]. Discharge [AKP15, GKT11]. Discharges [PK18]. Discoveries [AM17]. Discovery [BFS^{+}19]. Discrimination [MGB^{+}14, WPJ^{+}18]. Discussion [HCH13]. Disease [CEP15, GLCD13, MEJ10, SBG^{+}19]. Disinfection [SM13]. Disk [BKD^{+}10]. Dispelling [HMD16]. Dispersal [BRSA15, KB10, MMS^{+}17, ORA17, SKBSA11]. Displacement [BRSA15]. Disposal [PPR11]. Disruption [DHA12, SOSL10, SBL^{+}18b]. Dissolved [GLCD13, KDA^{+}18, SM14, UML^{+}11, WLL^{+}15, WALL17]. Distinct [APD^{+}18, DWW17, LAA^{+}18, MGB^{+}14]. Distinguish [LMC^{+}16, MPL^{+}15, THQ15]. Distinguishes [DSJS16]. Distinguishing [GPP^{+}10, MKHC17]. Distribution [AOS^{+}17, BPB13, BQHR17, BFL^{+}10, BRA^{+}18b, BE11, DCS^{+}16, DN11, DGF11, FBD^{+}19, HB14a, HDM15, JHS15a, JHB10, KTHM13, MNYV12, MFLS18, MCK^{+}15, MJ11, RH14, SEB^{+}19, SPQ12, TWL^{+}17, UML^{+}11]. Distributions [BATK16, BHW13, CMMR19, CDD^{+}11, GAH^{+}18, NP12, SEP^{+}19, TBH^{+}14]. Disturbance [BEH^{+}19]. Divergence [GSCR11, SSK18]. Diversion [HWNR14, MC10, PMP^{+}14]. Diversions [GRR^{+}17]. Diversity [AT18, BACRB14, CAWB17, FSS^{+}10, HMW^{+}10, HGK15, JF14, SOS^{+}15, SQGK10]. Diverts [ZS16]. Dividing [SE12]. Diving [HHMQ15]. DNA [ABP^{+}18, AES11, KIST16, PBA^{+}17, RPC^{+}19, SDF12, SLP17, UBR^{+}17, WCY^{+}18]. Do [Bre13, BWHM16, Can17, HBD^{+}17, LSM^{+}16, ORL^{+}16, PZC^{+}12, SBD^{+}11, SKBB15, SHS15, WBG^{+}17, WP11]. Does [Bal15, CEP15, DLC11, GHK19, HW10, PA16, Pul17]. Dogfish [BR14]. Dolly [SS13b, SMCS14]. Dolphinfish [MHBR19, RPMB19]. Dominated [HTH^{+}19, SEP^{+}19]. Donax [FWG15]. Dorsal [BKM19, IWB10, WKTQ15, WI17, WD10]. Double [MSK^{+}16, SCMP13]. Double-Crested [MSK^{+}16, SCMP13]. Downriver [DG12]. Downstream [AZE^{+}18, DBE11, KRJT13, MAP14, NMG^{+}17, SEM14]. Dozer [HPD^{+}19]. Drainage [ETD16, HB10, NCJ^{+}11, NEC^{+}13, TGS^{+}15]. Drawdowns [NTJ15]. Dredge [PPR11, SHM^{+}10, SMW12]. Dredging [ADMC10]. Drift [HHMQ15]. Drifting [BFL^{+}10]. Driftless [HM17, HDM15]. Drive [BBO16, HBD^{+}17]. Driven [KMPV13]. Drivers [AOS^{+}17]. Drought [JWC10]. Drum [DSB19, DBJD12, LSMD15, WFWSV14]. Dual [HHMQ15, HB12, HH15, MLEK15, PNHA12]. Dual-Frequency [HHMQ15, HH15, MLEK15, PNHA12]. Dual-Grid [HB12]. Dude [WPS^{+}19]. Due [AHVJ11, Bea17, Dag17, SL14, DK16]. Dungeness [AHVJ11, MBM13, SCW^{+}17]. Duplex [BRB^{+}17, BSTC13, SLSS16]. Duration [MQS13a, SM10]. During [Bal15, HHR^{+}18, RSQ^{+}19, SHS15, BPB13, BKSH^{+}12, BBL^{+}11, BKP^{+}14, CMD12, CHW12, DBW13, DSD^{+}13, FRL16, FCH^{+}13, GDP14, HFN11, HPA^{+}11, ISC10, Ise10, JM15, KLR^{+}12, LM16a, MSD11, RHG12, RHG13, SM14, TSFS19, WBF^{+}15]. Duty [DT11]. Dwelling [ETD16, JSDJ16, KSH^{+}16, MHG^{+}11, MB17, PFRM12, SMS13]. Dynamic [AJ13, Ask16, DBS19, SBL^{+}18a, SCS17b, SFH12]. Dynamics [APP13, BGD^{+}19, BBL^{+}11, DS12, GBL15, HSHH17, HHH14, MD10, MSBP10, NFQ16, OHSP17, RHG12, RHG13, SPM15, SCMP13, SRFS15, SEP11, SCB^{+}19, WDD^{+}14, ZDC14].

Eagle [Bri17, CCMA^{+}10]. Early [BWR12, BKP^{+}14, CSED10, DDD^{+}18, DSD^{+}13, PTD10, SLG^{+}15]. Early-Cohort [PTD10]. Earning [WBT17]. Earthen [BK16]. East [BSO^{+}16, Kra18, NEC^{+}13, ZEM11]. Eastern [BKME17, GBL15, KSMN18, KTN14, LBF17, PKKR10, PDL^{+}18, SGF^{+}13, Sta12, SC14, SKS^{+}15, WEC^{+}19]. EBFM [JDH16]. Ecological [CDD^{+}11, DKSW15, JM12b, KPL12, SMK13]. Ecology [BWR12, SEB^{+}19, SKS^{+}15, WQF15]. Economic [ASS17, BRN16, BM14, CL12, HHS^{+}13, ILBJ12, KHDA14, LS17, LS10, LL12, LN12b, LHS^{+}14, OMT^{+}15, SMW16, SL17a, SCA11, SBG^{+}19, TPB^{+}16, WHGC16]. Ecopath [McG14]. Ecosystem [ESL^{+}15, HMS11, JDH16, McG14, WHMS14]. Ecosystem-Based [JDH16, WHMS14]. Ecosystems [CDBW17, DCL19, FQ19, GFC^{+}15, KFM12]. Ecotype [LR12]. eDNA [MFLS18]. Education [CH17]. Eel [DA12, PT11, SGCB19, SEM14, TCB18]. Eels [CVD^{+}14, WA14]. Effect [AS10, BJZ11, DDEW18, DR10, FHE^{+}12, GF18, HFCH15, KWMB11, KDA^{+}18, MQS12, MP14, PZC^{+}12, SRSS12, SPS16, TCB18, THZ^{+}17, WLK^{+}16, WQF15]. Effective [GSCR11, SGC13, WBG^{+}17, WMSV18]. Effectiveness [APD^{+}19, AHS15, ACMB^{+}14, BOM16, BWB18, CBDW15, DBW13, DRW^{+}18, ES17, FHH16, HDKB10, HR11b, HFRH17, JHS15b, KCRR19, LCC^{+}15, MC10, Mea12, ORC^{+}16, PMP^{+}14, PQB10, RB15, RÅW18, RM12, Sam19, SPQ12, SPW^{+}12, TKH^{+}16, ZC12]. Effects [ADMC10, ADD14, AJR^{+}19, AKP15, Bea17, BDH^{+}14, BW18, BKSH^{+}12, BS10, BSO^{+}16, BDC^{+}13, BGHL10, BRA18a, BHR^{+}15, CJA^{+}13, CF18, CCDF18, CHW12, DK16, Dag17, DDAS14, DBKG14, DVWS10, DP13, DLSW12, DMC^{+}19, EPM16, FH16, FSS^{+}10, FQ19, GM12, GLCD13, GKMC15, GRSF^{+}16, GRCR10, GMG16b, HMW^{+}10, HWNR14, HB16, HBLL12, HRH10, HHSB13, JHS15a, JSDJ16, KAZ15, LJR11, LXM^{+}12, MSFS12, MAM^{+}12, MCB^{+}19, MW19, MCTM14, NTJ15, Nak13, NKC^{+}18, NKL18, NBP15, NASG12, OL17, PNWC11, PCP^{+}13, PKKR10, RGC^{+}12, RSLG10, Rey15, RFLP14, RZ10, RD10, RIMR19, RSQH18, RSQ^{+}19, Ryp15, Sam19, SPP^{+}18, SMH17, SBDB16, SBVB12, SSE19, SL15a, SB18, SOBW10, SQS^{+}17, SCG11a, SSP18, SSPP19, SS16, SL15b, SLS16, SEM14, SCG11b, TFP^{+}11, TRB11, TPF^{+}13, WF17, WGJ10, WMMB19, ZW12]. Efficacy [BRB^{+}17, BDS^{+}11, BRWQ18, FGS13, HHRB15, MMP12, SAYJ17, SK18b, SAF16]. Efficiencies [HMP10, PDM12]. Efficiency [BBW18, Blo13, BS10, BSTC13, FK16, GRR^{+}17, GH17, HNH11, KCS^{+}17, PSC15, PP10, SLSS16, SCW^{+}17, UBR^{+}17, WBGM13, WBF13, YS12]. Effort [AJ13, AWG^{+}18, FWR^{+}15b, KHBH19, LJR11, MQJK19, MQS13a, MSBP10, NKC16, PA16, RPJ18, RMTW16, SCS17a, SFH12, SLR10, SLP13, TSFS19, WHC14a, WHC14b]. Egg [BGCF11, CF13, FDM^{+}13, LBF17, PPFC18, SEM14, WBGM13, WBF13]. Egg-Per-Recruit [SEM14]. Egg-to-Larvae [LBF17]. Eggs [BWB18, CSWP19, KKSH12, LMC^{+}16, PCP^{+}13, RSLG10, RRH^{+}11]. Eight [HR11a]. Elastomer [LN12a, MM16, YHP10, ZC10]. Electric [GFC^{+}15, HKMH10, WBG^{+}17]. Electrical [FMA^{+}12, JTB16, NASG12]. Electricity [FVH^{+}17]. Electrified [HPD^{+}19]. Electroanesthesia [VMC^{+}11]. Electrode [BGM17, MK13]. Electrofishers [DT11, DTS19, MK13]. Electrofishing [BS12, BHN11, BFN16, BKME17, BS10, BGBG17, BDB^{+}19, BHG^{+}13, CC10, CC15, DP13, DCHC13, FK16, FMP19, HPD^{+}19, HMP10, HBDA13, JJ10, KDMH10, MQJK19, MK13, Mea12, MH11b, MB17, NQ11, PD13, PP10, Rey15, RMTW16, SFW11, SLK15, SHL16, SNTZ14, SOBW10, SBL11, SLP13, SL15b, TKH^{+}16, THZ^{+}17, YS12]. Electrofishing-Derived [BS10]. Electronarcosis [Bal15, HJK11, KAZ15]. Electronic [FVH^{+}17, SB15]. Electroshock [Bea17, DK16, Dag17, Rey15, SL14, SL15b]. Electroshocked [HRH10, SPS16]. Electroshocking [AM17]. Electrotrawling [SVD^{+}18]. Elements [ASP^{+}17]. Elevated [BGHL10, KLR^{+}12]. Elevation [BPB13, BHR^{+}15, JM12a]. Elimination [HIS16]. Elk [SHK^{+}17]. elongatus [NEH10]. Elwha [BPT^{+}19, MPL^{+}15]. Embryo [NASG12]. Embryos [ASL^{+}13, ASL^{+}16, BGM17, CDQG12, GFC^{+}15, SPS16, TGKZ19]. Emergent [RWCP11]. Emerging [JDG10, TDM19]. Emigration [HHP^{+}12b, TRB11]. Emitting [SWG^{+}16]. Empirical [BCD^{+}10, GPLT12, GFP^{+}12, HWNR14, LS10, RHMH15, XM14, JHS15b]. Enclosed [RJP^{+}15]. Endangered [BWT^{+}13, BBC^{+}13, CMMR19, GBL15, GKT11, HPA^{+}11, PDD^{+}18, WBH13, WHG^{+}16]. Endemic [GPLT12]. Endoscopic [TGKI10]. Energy [SPT^{+}12]. England [AG11, TDM19]. Enhance [APPR14, HW10, MC10]. Enhancement [ADD14, BFLT17, CEH^{+}15, CKBP19, LBF17, Ols19]. Enhances [RCP^{+}13]. Enlightened [EF10]. Enough [KSM16]. Enriched [WGM^{+}11]. Entering [DMRJ10]. Entosphenus [DSJS16]. Entrained [DMRJ10]. Entrainment [FRL16, MHR^{+}18]. Entrapment [CBDW15]. Entry [ASD^{+}18]. Enumerate [CKC^{+}12]. Environment [HGK15, NCJ^{+}11, Ols19, OOWW12, UWL13]. Environment-Related [UWL13]. Environmental [ABP^{+}18, AOS^{+}17, BW18, GRSF^{+}16, MOJ12, PBA^{+}17, PFD17, RPC^{+}19, SBVB12, SLP17, UBR^{+}17, WCY^{+}18]. Environments [HQ10]. Enzyme [MEJ10]. Enzyme-Linked [MEJ10]. Epinephelus [SNT^{+}16]. Epoxies [IH11]. Equal [SCPC14]. Equation [BBC10, BBC11, BBS19, GPLT12, GFP^{+}12]. Equations [CK10, Ger11, RFW^{+}10, RFW^{+}11, RV12]. Equilibrium [CBS13, SFH12]. Equivocal [ES17]. Eradicate [GFC^{+}15, KMK^{+}15]. Eradicating [SMH17]. Eradication [BHG^{+}13, KKM16, PT15, RB15, SNTZ14]. Erie [BSB^{+}15, LJR11, MKWM18, RMTW16, SBTW17, SSK18, VKBL11, VBT^{+}12, WEC^{+}19, ZEM11, ZKM13]. Erosion [MB12]. Errata [Ano10a]. Erratum [Ano10b, Ano12b, Ano14a, Ano14b, Ano14c, Ano16a, Ano17b, BBC11, HFH18, MFRS11, Ano12a, Ano17a, Ano18a, Ano19a]. Error [AMQSK15, APF18, HSM^{+}18, KBK13a]. Errors [WPS^{+}19]. Escambia [WHKB17]. Escanaba [EDN10, HIS16, HFN11, SSE19]. Escape [DR10, RJMG13, SCW^{+}17]. Escapement [CKC^{+}12, FWT14, MPM10, PNHA12, SBW^{+}16, VEIW^{+}16, WFS13, WDW13]. Escapes [SWW14]. Establishing [GSCCP18, MP14]. Estimate [BRN16, BS12, BRWQ18, CZH^{+}19, FC17, HWB11, HSM^{+}18, HH15, LKV^{+}17, LSLC15, MG14, MFLS18, McC16, McC17, MEL^{+}12, MS14, MWS15, OHL10, POWF11, PVH^{+}13, SC17, SLK15, SHK^{+}17, SHL16, TP17, VKH^{+}19, WKP14, WHV15, WD10]. Estimated [OMT^{+}15]. Estimates [AWG^{+}18, BGCF11, BGCH17, BM14, BGW^{+}11, BB12, BMW^{+}19, CBS^{+}12, CKC^{+}12, CCDF18, CPM^{+}10, DIK19, Der14, FWWTJD16, GF18, HT11, KMPP12, MQS12, MM18a, MH11b, MB17, PZC^{+}12, Ran18, SL12, TP17, WWB^{+}17, WKTQ15, WGS16, WEC^{+}19]. Estimating [BGCS11, BCZ^{+}18, BKM16, BKM19, BSO^{+}16, BSS18, Bou10, BSM^{+}15, DPS^{+}17, DLC11, FAW14, GRSF^{+}16, HKMH10, HDKB10, HV18, HHMQ15, IWB10, KADH15, KSM16, MBBS10, MKRH14, MQS13a, MJCM15, McC15, MLEK15, MPM10, OI17, PSB^{+}10b, PSC15, PNHA12, PA16, PPO18, RPBM10, RPMB19, RHMH15, SCS17a, SL17a, STSH10, SBL11, SPL18, VEW^{+}13, WH12, WFS13, WI17, XM14, ZM17, ZWH^{+}15]. Estimation [AD13, BC18, BGP10, FHL18, hGgXfC^{+}13, HD14, HSC^{+}16, HCL^{+}12, HUB^{+}18, KNS18, KDMH10, MFLS18, MQS12, MR16, MM17, MM18a, PP10, SFW11, SLR10, SJM15, SBB19, TCM19, VBT^{+}12, WSA18, WBK^{+}10, WBK^{+}13, YC11]. Estimators [AWG^{+}18, CWH^{+}14, MS16, SLB10]. Estuaries [FTSF13, TMC10]. Estuarine [Ols19, SCS17b]. Estuary [GH17, JPWR10, NEH10, ADMC10, DBJD12, FWWTJD16, SSS15, TBH^{+}14]. Eulachon [HO13]. European [DA12, McG14, SVD^{+}18]. Evacuation [TGA15]. Evaluate [ERC^{+}11, FDM^{+}13, HSG16, PRR^{+}13, RFKB17, STP19, SEM14, WHMS14]. Evaluating [BOM16, BBW18, BE11, BWB18, BSG10, CvPW15, CA14, CWH^{+}14, CSH^{+}17, Der14, DLP16, FBM^{+}19, GCA^{+}13, GTB^{+}19, HB11, KMK^{+}15, MMHM17, PMP^{+}14, SKBK12, SDF12, SK18b, TFH^{+}16, TCB18, WFS13, FBL^{+}19]. Evaluation [ADMC10, ABČ^{+}15, BSL18, BP14, BGD^{+}19, BKM19, BF13, BP15, BT16, BHC^{+}10, BC17, BTD^{+}12, CWB16, CMF^{+}14, CJ14, CKBP19, CMD12, DKWG14, DMC^{+}19, EKM10, FMP19, FWT14, GZ13, HP14, HWDA12, HO13, HRKB19, HH10, HWNR14, HW19, HS12, HWWT15, HADG10, HPTH18, IH11, KNC14, KYD18, LKB19, LOS^{+}11, LCC^{+}15, LP12, LSLC15, LSK^{+}14, MKRH14, MJCM15, McC15, McC17, MASN16, MIH15, NBFL17, NFQ16, ORC^{+}16, PSC15, PT11, RS17, SFW11, SBW^{+}17a, SP13, Sin18, SSP^{+}19, TGKZ19, TGKI10, TSNH10, VMC^{+}11, VKBL11, WS11a, WBP13, WRNJ^{+}19, WSS13, WMSH13, WPY^{+}15, WSA18, WWM13, WDW13, YHP10, ZC10]. Evaluations [MRC12, Pie10, SRCC^{+}18]. Everglades [SK18a]. Evidence [APD^{+}18, BSB^{+}15, CCC16, EWS^{+}17, HHT^{+}13, JBZ^{+}10, LKM^{+}17, LAA^{+}18, PRM15, SB11, TPBF15, WFWSV14]. Evolution [SMW12]. Evolutionarily [EF10]. Examination [CHL13, CS17, SHL16]. Examining [Lee12, MHBR19, SDKS16]. Example [BGCH17, FC17, KADH15, MGB^{+}14, SC17, SHSCA16, SK18b]. Exceeding [SKBB15]. Excision [WSA18]. Exclosures [DFMP18]. Excluder [RGC^{+}12, SHM^{+}10]. Exclusion [ASD^{+}18]. Exclusive [KHDA14]. Exhibits [KSL^{+}19]. Existing [DSJS16, GTB^{+}19]. Expanding [SPW^{+}12]. Expansion [LN12c, NTJ15]. Expenditures [LM16b]. Experience [DDEW18, HCL^{+}12]. Experienced [BEH^{+}19]. Experiences [MSHC15]. Experiment [CVD^{+}14, NKC^{+}18, PT11, SLG^{+}15, ZM17]. Experimental [BTD^{+}12, CHJ^{+}16, CMD12, CHL13, GDP14, GZ13, MJH17, Rey15, SL15b, TRB11, THQ15, VKBL11]. Experimentally [MGWC13, MSK19b]. Expert [MAM^{+}12, MSHC15]. Explicit [CBC^{+}16]. Exploitation [BSO^{+}16, BSS18, CI10, CAS16, HAS^{+}19, MM18a, MDM12, MS14, MASN16, SC17, Sam16, WHB11]. Exploited [DZ14, DZ17, FJB16, HVW^{+}17, SPM15]. Exploratory [ACRBB10]. Explore [SLS16]. Exploring [CA11, FHH16, LKvR^{+}18, SLP17]. Exports [NB10]. Exposed [ABČ^{+}15, CBS^{+}12, GLH^{+}18, LMS^{+}18, SBL^{+}18b]. Exposure [CMWM18, GMG16b, LM16a, RHC^{+}14, RSQH18, RSQ^{+}19, WWS^{+}14, WMMB19]. Exposures [NASG12]. Extended [TSFS19]. Extension [CBS13]. External [BF13, DRC17, HKP^{+}14, SNT^{+}16].

Facilities [CBS^{+}12]. Facility [MHR^{+}18]. Factors [BW18, DDAS14, EDN10, FQ19, KSV^{+}10, KPL^{+}16, KBH^{+}19, KW13, LFD^{+}17, LSK^{+}14, OHL10, PFD17, RFQ14, RHG13, SS11a, SVFG10, SPT^{+}12, SNTZ14]. Fail [MCS^{+}17]. Failed [SRFS15]. Fails [RGOR10]. Failure [AG11, WP11]. Failures [SMW16]. Fall [ABBG13, BBS18, CJ14, CCC16, JKN^{+}10, PFD17, RCP^{+}13]. Fall-Run [BBS18, PFD17]. Fallacy [Ger10]. Fallback [RS18]. Fallfish [JNC^{+}16]. Falls [BBO16]. False [HSC^{+}16]. Family [NBP15]. Farmers [MRC12]. Fat [CVD^{+}14]. Fatty [PHFK10]. FCRPS [SWH^{+}16]. Feasibility [DGS^{+}11, GMG^{+}16a, McC15, WHG^{+}16]. Features [DRC17]. Fecundity [BMM^{+}12, BCZ17, SLMM10, SBB19]. Federal [PA16, SMW16, SEP^{+}14, SWH^{+}16, TSFS19]. Federally [MAM^{+}12]. Feeding [BR14, DVWS10, FMA^{+}12, KFM12, MCTM14, SEB^{+}19, SOBW10, SQS^{+}17, WWD13]. fellowesii [GPLT12]. Female [BMM^{+}12, BR14, BEK^{+}18, LS11, TO10]. Ferry [AKP15, BRN16]. Fetch [SCKZ17]. Fidelity [HR11a, JHK11]. Field [AM17, BKSH^{+}12, BWB18, DMC^{+}19, DRC17, DG12, GBTL14, HKP^{+}14, MJ11, MRC12, MKHC17, NKC^{+}18, NASG12, PVST11, WDW13]. Field-Based [DRC17, MRC12, MKHC17]. Fields [GFC^{+}15, HKMH10]. Fight [CMWM18, WMMB19]. Fillet [LRHR17]. Filling [WHC14a, WHC14b]. Fin [CHW12, HTPS10, HUB^{+}18, MB12, RHPH14, SH18]. Fine [BSVB19, MAP14, SCS17b]. Fine-Scale [BSVB19, MAP14, SCS17b]. Finespotted [HGK15]. Fingerling [LSK^{+}14, PTD10, TPL18, ZC10]. Fingerlings [LC14, RS17]. Fire [CJA^{+}13]. First [BC14, BSS^{+}17, CVD^{+}14, KFW12]. Fish [ABČ^{+}15, AHAC15, APP13, AJ13, BRB^{+}17, BPQ13, BOM16, BBB^{+}14, BKD^{+}10, BP14, BK12, BSBC14, BGW^{+}11, BG13, CEH^{+}15, CL12, CMMR19, CMC^{+}19, CC15, CKCF12, DDAS14, DBKG14, DWW12, DZ17, Der14, DCL19, DHK^{+}19, DGS^{+}11, DV17, ESL^{+}15, FK16, FAW14, FJB16, FQ19, FMA^{+}12, GMG^{+}16a, GPLT12, GAH^{+}18, GHK19, GTB^{+}19, HWNR14, HMCT16, HMP10, HVW^{+}17, HEL^{+}15, HJK11, HADG10, HBD^{+}17, HHB12, JJ10, JM12b, KMB^{+}12, KSM16, LJR11, LCR19, MSFS12, MS16, MJCM15, McC16, Mea12, MSS^{+}12, MRC12, MGWC13, MEL^{+}12, MOJ12, MKWM18, MCTM14, NQ11, NKC16, OHL10, ORC^{+}16, OH15, OOWW12, PMP^{+}14, PBA^{+}17, PKKR10, PC12, PLUH19, PQB10, PP11, PRRD15, Pul17, Rey15, RFKB17, RD10, RGOR10, RM12, RCP^{+}13, SFW11, SBD^{+}11, SSK19, SLSS16, SCMP13, SFH^{+}13, SVPD13, SHL16, SPQ12, SQP12]. Fish [SCG11a, SBW^{+}17b, SS17, SMW12, SHMJ19a, SE12, SBTW17, SL15b, SEP^{+}19, TRB11, TKH^{+}16, UML^{+}11, WKS^{+}10, WBG^{+}17, WMSV18, WPS^{+}19, WKP14, XM14, ZSW11]. Fished [RH16]. Fisher [TDM19, RJC11]. Fisheries [APF18, Ano16b, Ano19h, APP13, Ask16, BCBW18, BATK16, BSS^{+}14, BW17, CvPW15, Can17, CJ14, CMWM18, CBC^{+}16, CHBY13, CGL18, CH17, DC13, DCP^{+}19, EF10, FF15, FH16, FC17, FHH16, FWR^{+}15a, GSCCP18, HP14, HSHH17, HPTH18, HHS^{+}13, HMD16, ISC10, JBHM19, LKvR^{+}18, LTG^{+}11, LRHR17, MBBS10, MCP15, MQS12, MQS13a, MQS13b, McC16, MM18a, MDM12, MS14, MS17b, MSE^{+}18, MG10, Ols19, PM14, PBA^{+}17, PLUH19, PG14, RGOR10, SRCC^{+}18, SBL^{+}18a, SCW13, SCA11, SLS16, SC14, THI^{+}19, UBR^{+}17, UML^{+}11, VMA^{+}16, WHB^{+}14, WLK^{+}16, WWM13, YC11]. Fishers [Aka10, CGL18, SPW^{+}12]. Fishery [ASS17, AS10, AJ13, BFP12, BDH^{+}14, BM14, BC18, BSB^{+}15, BLM10, CZH^{+}19, CJ14, DS12, DHA12, DMC^{+}19, DKSW15, FWG15, FBM^{+}19, FBC^{+}10, GMG16b, GFM^{+}18, HIS16, HBS^{+}19, HMBT16, HWB11, HSG16, HF13, HMS11, HVW^{+}17, HCH13, HWWT15, JDH16, LFD^{+}17, LS17, LBB16, LAM16, LHS^{+}14, LN12c, LBL11, MG14, MBM13, McB14, MP14, MFRS11, MCH12, MBPG12, MMHM17, NEH10, PA16, PDL^{+}18, RGC^{+}12, RR17, RGOR10, RPJ18, RPMB19, SOB18, SCS17a, SCS17b, SPM15, SMW16, SOS^{+}15, SS11b, SMW12, SBG^{+}19, SF16, TDM19, TEB^{+}19, VPV18, WHC14a, WHC14b, WQP13, WAW^{+}19, WHGC16, WMMB19, ZC12, ZC15, ZKM13]. Fishery-Dependent [WHC14a, WHC14b]. Fishery-Independent [FWG15]. Fishery-Related [WMMB19]. Fishery-Scale [CZH^{+}19]. Fishes [ADMC10, APD16, AJ10, AD13, BWT^{+}13, BWB18, CK10, DS12, DP13, DR19, EJL10, FDHWP16, GBL^{+}13, GRSF^{+}16, HGW16, HH15, KCS^{+}17, LDW10, LCR19, MKK^{+}12, PD13, PDD^{+}18, PP10, RH14, RCGF18, RWT17, SK18a, SBW^{+}16, WBH13, WBK^{+}10, WBK^{+}13, WBT17]. Fishing [APD16, Aka10, BSO^{+}16, DHH17, DZ14, DZ17, FWR^{+}15b, GGCF14, HFN11, HBD^{+}17, HHSB13, KHBH19, KADH15, LMS^{+}18, LMB^{+}18, LS17, LL12, MBM13, MP14, ML13, MSBP10, SPP^{+}18, SCS17b, SF14, SM10, SMS13, SF16, TSFS19, Vin14, WHGC16]. Fishway [APD^{+}19, AR12, BJ19, HFRH17, WBZ19]. Fishways [BWHM16, KDP^{+}10, MJCM15, NTJ^{+}14]. Fishy [WMB^{+}15]. Fitness [RSQ^{+}19]. Five [ASL^{+}16, MS16, UMC14]. Fixed [BSTC13, LCC^{+}15, PCP^{+}13]. Fixed-Station [LCC^{+}15]. Flannelmouth [HUST17]. Flat [MRC12]. Flat-Plate [MRC12]. Flathead [BSO^{+}16, HB16, HGW16, MCSF^{+}17, SEB^{+}19, Sin18]. Flats [LFD^{+}17]. Fleet [MSBP10]. Flies [Blo13]. Flint [GKS13, ITY13, Sam16]. Floating [FAW14]. Flood [FRL16]. Floodplain [CDBW17, CHJ^{+}16, EJL10]. Florida [FWWTJD16, BSBC14, CGL18, HBS^{+}19, HH12, HRKB19, HL15, KADH15, MASN16, MG10, PTD10, SGCB19, SK18a, SQS^{+}17, TPB^{+}16, TPL18, WHKB17, WFWSV14]. Flounder [AS10, AK12, FTSF13, HF13, MBPG12, MMHM17, SOB18, SCS17b, SS11b]. Flow [ABČ^{+}15, GDP14, HGW10, LTG^{+}11, MMP12, Nak13, RBP16, SCB^{+}19]. Flowing [SPWC15]. Fluctuating [HB14b, PCP^{+}13]. Fluctuations [LSM^{+}16, SC14]. Flume [PRR^{+}13]. Fluorescent [SDF12]. Fluvial [DZS10, DCS^{+}16, TWL^{+}17]. Following [BPT^{+}19, GBFR18, JCB^{+}18, KKM16, MMS^{+}17, PDD^{+}18, PB13, TLL^{+}18, WRNJ^{+}19, BRSB^{+}16, CFWC11, LBF17, MHAP13, RH14, SDF12, WWD13]. Food [HTH^{+}19, MSK19b, SPT^{+}12, SOSL10]. Foothills [MSFS12]. Forage [SSPP19, WBT17]. Foraging [RCMB18, SL15a, SSP18, SS16]. Forced [CFWC11]. Forecast [HP14]. Forecasting [FF15]. Forestlands [MS17a]. Forests [CSWP19, JBHM19]. Forestwide [DRG10]. Fork [DB13, DCS^{+}16, JF13, MNYV12, NEC^{+}13, PSCS11, SD11]. Formalin [GF18, SCG11b]. Format [MLHS13]. Formation [BC14, BK12, LO18, RFLP14, RHPH14, WS16]. Forms [JBZ^{+}10, VMC^{+}11]. Formulations [RB15]. Four [ASD^{+}18, DVWS10, DWW12, KBH^{+}19, Kro19, MK13, OLM15, PLCG13, PN19, SNTZ14, SBW^{+}17b, SPL18, SEP^{+}19]. Fraction [BGCS11, DZ17]. Fractions [DZ14]. Fragmented [HNPB15]. Frame [UWL13]. Framework [BGP10, DGS^{+}11, GMG^{+}16a, NCSR17, SRCC^{+}18]. Francisco [ERC^{+}11]. Fraser [CHPP11]. Free [FC17, SPWC15]. Free-Flowing [SPWC15]. Freeze [NLK17]. Frequency [AWG^{+}18, BHW13, DS12, FAW14, HHMQ15, HH15, KD10, LBB16, MLEK15, PD13, PNHA12, RFKB17]. Frequent [BWT^{+}13]. Freshwater [AT18, BS11a, DSB19, FRH10, FCH^{+}13, GPLT12, GFC^{+}15, KADH15, MS16, RCGF18, SK18a, See11, TBH^{+}14]. Friendly [ZSW11]. Frogs [BKSH^{+}12]. Fry [CMK^{+}19, GLCD13, HQ10, HLS^{+}17, MPL^{+}15, SKBSA11]. Function [NB10, OI17]. Functions [BRWQ18]. Fundy [WHB^{+}14]. Funnel [BGCF11]. Future [AHS15, ALCD17, CB10, DBK11, HW12, LSC^{+}16, MD10, TSFS19, TSNH10, ZRR^{+}19]. Fuzzy [MSHC15]. Fyke [BSBC14, CBDW15, GHL12, Kro19, SBW^{+}16, SS17]. Fyke-Net [GHL12, Kro19].

Gags [SKS^{+}15]. Galvanized [PDM12]. Game [DR19, ESL^{+}15]. Gap [SPQ12]. Gaps [RJMG13, WHC14a, WHC14b]. Gar [BKDB17, DPK^{+}17, DBS19, MAH^{+}13, SBS14, SBDB16, SL17b]. Garden [SLG^{+}15]. Gars [BDSB15, WHKB17]. Gas [GLCD13, RCMB18, WLL^{+}15, WALL17]. Gastric [BBW18, CSV13, HNH11]. Gear [BGW^{+}17, CKCF12, GRSF^{+}16, MSK18, MSBP10, NCMT12, TEB^{+}19, KKSH12]. Gear- [MSK18]. Gear-Type [CKCF12]. Gears [CBDW15, CDBW17, DWW12, EJL10, KQDC19, SEP^{+}19, WMB^{+}15]. Gender [SB11]. Gender- [SB11]. Genera [DSJS16]. Generating [MHAP13, MAP14, PMP^{+}14]. Genetic [ADD14, AES11, BKHS14, BSVB19, BSB^{+}15, DPK^{+}17, DASB16, DBJD12, DSJS16, DCP^{+}19, DLC11, GSCR11, HD14, HS19, HRKB19, HMBT16, HHP^{+}12b, HWFN14, HBS14a, HVW^{+}17, JF14, LMC^{+}16, LSMD15, MNYV12, MWS15, NEC^{+}13, ORD19, PLCN17, SBW^{+}17a, SOS^{+}15, SSK18, SQGK10, TTS^{+}18, TLSB18, TBH^{+}14, VWB^{+}11, VEW^{+}13, WHB^{+}14, WAW^{+}19, WBB10, WWP18]. Genetically [HVW^{+}17, LAA^{+}18, MGB^{+}14]. Genetics [AK12, HDM15, OHSP17]. Genotype [TPBF15]. Geographic [WWB^{+}17]. Geographically [APD^{+}18, DWW17]. Geography [LR12]. Georges [SBG^{+}19]. Georgia [Sam16, BFP12, GKS13, ITY13, RPJ18, TBIJ19]. Georgian [CKCF12]. Geostatistical [WHC14a, WHC14b]. German [DA12]. Ghost [MBM13, SCB^{+}19]. Ghosts [SCB^{+}19]. Giant [RJMG13]. Gibbon [SZKK10]. Gila [HGW16, RW11]. Gill [AS10, BFN16, BM14, BDSB15, DHA12, GMB10, Han19, KHBH19, LJR11, SBDB16, SS11a, SR16, SS11b, SBW^{+}17b, VKBL11, WSH^{+}14, WKS^{+}10, ZM17]. Gill-Net [AS10, BM14, BDSB15, DHA12, KHBH19, LJR11, SS11a, SS11b, VKBL11, ZM17]. Gippsland [GMB10]. GIS [ND15]. GIS-Based [ND15]. Gizzard [CA11, DSB19, FSHR17, NKL18, WSH^{+}14]. Glacial [DWW12, DWB^{+}15, HB14a]. Global [HAW^{+}18]. Globally [SC18]. Gloves [WBG^{+}17]. Go [FBL^{+}19, HHP^{+}12b]. Goals [BKHS14, HTF^{+}13]. Goby [NCMT12]. Golden [FWR^{+}15b, FWR^{+}15a]. Good [JM12b]. GPS [RFKB17]. Grab [PP11]. Graded [Ask16]. Gradient [DV17, SPS16]. Gradients [PDH^{+}19]. Grand [BGCF11, GDP14, PHS^{+}13]. Grande [AR12, GHL12, ZRR^{+}19]. Grange [LKV^{+}17]. Granite [APPR14, MSK19a]. Grass [BFS^{+}19, BDB^{+}19, CSWP19, KML14, MKBB13, SDF^{+}13, SJM15, ZGW11]. Grassland [CJA^{+}13]. Gravel [MODC10, PRR^{+}13]. Gray [AES11, FWWTJD16]. Grayling [ASP^{+}17, GAH^{+}18, SZKK10]. Grazing [DFMP18]. Great [BT16, JBRWH18, LSC^{+}16, BJ19, Car13, DR19, FDHWP16, FSS^{+}10, GKMC15, HMW^{+}10, HSG16, ILBJ12, LTG^{+}11, ML13, MSK18]. Greatest [SPQ12, SQP12]. Green [MLEK15, TPF^{+}13, WKS^{+}10, WBB10, AT18, BQR^{+}19]. Grid [HB12]. Grinnell [GD10]. Gross [MHBR19]. Ground [BP15]. Ground-Based [BP15]. Groundfish [AG11, LS17]. Groundfishes [CDD^{+}11]. Grounds [KHBH19]. Groundwater [DSB19, HSF18]. Grouper [BMMMLML15, SNT^{+}16]. Groups [OOWW12, RBB^{+}10b, SLSS16, SSK18]. Growing [ZWH^{+}15]. Growth [APD^{+}18, BK16, BSL18, BRSB^{+}16, BKP^{+}14, CGDB14, DSB19, DVWS10, DDEW18, DJS^{+}16, FILH15, FBL^{+}19, FWR^{+}15a, GLCD13, GBFR18, GF18, GKT11, HNH11, HDB^{+}12, HLM^{+}18, HW10, HB16, HF13, ITT10, KWMB11, KSM^{+}18, KSG^{+}18, KSL^{+}19, LLP10, MCB^{+}19, MSK19b, NKL18, OL17, OLOI13, OI17, PPO18, SHSCA16, SW17, SSE19, SPHK16, SKBSA11, SGF^{+}13, SLG^{+}15, Sta12, SDL^{+}15, SDF^{+}13, SDPB17, TRB11, TPC^{+}15, WS16, WRNJ^{+}19, WF17, WS13, WMB^{+}15, WWD13, ZOC^{+}14]. Growth-Increment [WS16]. Grunt [ORD19]. GSI [HWFN14]. Guadalupe [PSB^{+}10a]. Guide [Ano11, Ano12c, Ano13, Ano15, Ano19h, BG19, BWHM16, Car14, EF10, MS17b, PRRD15]. Guiding [GBTL14]. Gulf [AOS^{+}17, BMMMLML15, JHS15a, RCMB18, SC17, SFH^{+}13, SKS^{+}15, SM14, WGS16, AES11, CZH^{+}19, CK10, CAS16, DPK^{+}17, FWWTJD16, GGC17, GGCF14, HB12, MCH12, PA16, RGC^{+}12, RR17, RKB15, RGOR10, SDKS16, Sta12, SDPB17, TSFS19, Vin14, ZC12, ZC15]. Gulls [FOH^{+}15]. Gun [CDQG12, RJP^{+}15].

Habitat [ACRBB10, APAC12, ABVS14, ASL^{+}13, ASL^{+}16, BPQ13, BOM16, BS10, Bre13, BC17, CEH^{+}15, CMMR19, CJMD10, CGG^{+}18, CHJ^{+}16, CKBP19, CMC^{+}19, DDAS14, DFMP18, DTA18, DV17, FDM^{+}13, GDP14, GAH^{+}18, GKS13, GDLM17, GCA^{+}13, HKMH10, HGW10, HH10, HH11a, HH11c, HPM^{+}18, KPL12, KCRR19, KYD18, LGC11, LBF17, LS13, MRB^{+}11, MHAO11, MS17a, MNYV12, MJ11, MODC10, MMB19, MSHC15, ND15, ORC^{+}16, ORA17, Pto13, RGH^{+}13, RFKB17, RSI16, RÅW18, RABK18, RBB^{+}10b, RBP16, RMTW16, RIMR19, SHMS16, SG13, SPM15, SPT^{+}12, SQP12, SBL11, SCB^{+}19, TFP^{+}11, UML^{+}11, WWW15, WA16, WHKB17, WS11b, ZWS^{+}19, ZSW11]. Habitat-Specific [SBL11]. Habitats [AM17, BKSH^{+}12, JJ10, KCCW19, MFLS18, TBH^{+}14, TKH^{+}16]. Habits [Bal15, BR14, MSK19b]. Haddock [CZH^{+}19, PSCS11]. Hake [RR17, RGOR10]. Half [BRB^{+}17, BSTC13, SLSS16]. Half-Duplex [BRB^{+}17, BSTC13, SLSS16]. Halibut [LS11, SF16]. Hampshire [SBB19]. Hand [ASD^{+}18, BSO^{+}16, FMP19]. Hand-Crank [FMP19]. Handling [BLSS19, FGS13, KDA^{+}18, LLW^{+}18, SMP^{+}14, SQS^{+}17, SCG11b, TEB^{+}19, WBG^{+}17]. Handling-Induced [SMP^{+}14]. Hang [AS10]. Hang-in [AS10]. Hanson [WBB10]. Harbor [BM14]. Hardness [RSLG10]. Harry [STP19]. Harvest [CPSG12, DAKP13, ES14, FBM^{+}19, FWR^{+}15b, GD10, GRCR10, HIS16, HHCH15, HV18, HWB11, HCL^{+}12, Hen15, HSHH17, HCH13, JKN^{+}10, LWJS16, LS10, LSLC15, McC17, MMHM17, RPJ18, RIMR19, Sam16, SCW13, SL17a, SSE19, SM10, SLS16, SGC13, TSFS19, VPV18, WAW^{+}19, WWM13, ZC12]. Harvested [BCZ17, ISC10]. Hatch [PCP^{+}13]. Hatcheries [MEJ10]. Hatchery [AR12, APP13, BZ13, BGCS11, BKHS14, Bri17, BSG10, CGDB14, CFW^{+}10, CFWC11, DBKG14, DDEW18, EKM10, FHC16, FDM^{+}13, GPP^{+}10, HTF^{+}13, HM12, JPWR10, JF13, JNC^{+}16, JCB^{+}18, KPP^{+}18, KSL^{+}19, MNYV12, MB12, MW19, MSK^{+}16, MPM10, NCJ^{+}11, ND11, OMT^{+}15, PNWC11, PHFK10, RS18, RCP^{+}13, SKH^{+}16, SMK13, SSP18, SLG^{+}15, SPK10, SPP12, TRB11, TCG^{+}17, TLC^{+}19, VEW^{+}13, WBP13, WMSV18, WPJ^{+}18, WMMB19]. Hatchery- [KPP^{+}18, MPM10]. Hatchery-Dependent [APP13]. Hatchery-Origin [WPJ^{+}18]. Hatchery-Produced [GPP^{+}10]. Hatchery-Reared [AR12, CGDB14, DBKG14, DDEW18, JPWR10, KSL^{+}19, MB12, PHFK10, SKH^{+}16, SSP18, SPK10, SPP12, WBP13, WMSV18]. Hatching [RSLG10]. Haul [KPP^{+}18, TSNH10]. Head [CSH^{+}17, MSS^{+}12, WBZ19]. Head-of-Reservoir [CSH^{+}17]. Head-of-Tide [WBZ19]. Headwater [HAS^{+}19, HBA^{+}12, PT15, UACZ^{+}16, WHMS14, WWP18]. Healing [HRH10, WRNJ^{+}19]. Health [HHB12]. Heavily [SPM15]. Height [Ano12b, JY11]. Helps [WMB^{+}15]. Hematological [JTB16]. Hemorrhagic [PCT^{+}14, TPBF15]. Herbicides [MKBB13]. Herring [BSS^{+}14, DRW^{+}18, DR10, DG12, Gur12, GBTL14, OHSP17, SB15]. Hessenauer [SCPC14]. Heterogeneity [SC17]. Hickory [HH11c]. Hidden [DPS^{+}17, HTF^{+}13]. Hidden-Process [DPS^{+}17]. Hierarchical [NB10]. High [AWG^{+}18, DM12, FBB19, GSCCP18, GMB10, HHR^{+}18, HR11a, KZ12, LFD^{+}17, McC15, McC16, SCS17a, SK18b, WAW^{+}19]. High-Density [KZ12]. High-Effort [SCS17a]. High-Frequency [AWG^{+}18]. High-Profile [WAW^{+}19]. High-Relief [HR11a]. High-Use [SK18b]. High-Water [HHR^{+}18]. Higher [ORL^{+}16]. Highlands [TLSB18]. Highway [PFRM12]. Hills [JWC10]. Hilsa [NCSR17]. Hired [SF16]. Historical [GD10, KBH^{+}19, McG14]. History [AT18, BACRB14, BSM17, BPT^{+}19, BWR12, CAWB17, DS12, DASB16, DSD^{+}13, EMK^{+}17, HGK15, JBRWH18, JBZ^{+}10, KRB^{+}16, MCS^{+}17, RM14, TRB^{+}13, UACZ^{+}16]. Hokkaido [KSMN18]. Holders [TBIJ19]. Hole [CFBB15, KPL12]. Holt [CBS13]. Homarus [BC18]. Homing [HTF^{+}13]. Hook [BDH^{+}14, BMMMLML15, DVWS10, FSW18, RLC12, RPMB19, SOB18, SKH^{+}16, SEP11]. Hook-and-Line [RPMB19, SOB18]. Hooked [DVWS10, RLC12]. Hooking [BDH^{+}14, CMWM18, MH11a, MHBR19, RS11, SS13a, SKH^{+}16, SMS13]. Hooks [BDT^{+}16, KSH^{+}16, LMS^{+}18, RS11, SMS13]. Hoop [CBDW15, Sin18, SL12]. Hopper [PPR11]. Horizon [SM14]. Horizontal [MRC12]. Hormone [HPA^{+}11]. Hornyhead [HW19]. Hourly [XM14]. House [SF14]. Howard [WBB10]. HSIs [LS13]. Hudson [KSM^{+}18]. Humpback [GDP14, PHS^{+}13, WPY^{+}15]. Huron [CBC^{+}17, RRH^{+}11, BATK16, BSB^{+}15, CBC^{+}16, EWS^{+}17, FB14, HER^{+}12, JDG10, KSL^{+}19, PM13, SBHJ17, ZM17]. Hurricanes [PKKR10]. Hybrid [BK16]. Hybridization [ACMB^{+}14, BKDB17, CEP15, KEC11, KACS18, LDP^{+}10, ND11, PNWC11, TLSB18]. Hybrids [DZS10, GWB^{+}10, MKHC17, NLK17, PVST11, SDF12]. Hydrilla [KML14, MKBB13, NTJ15, SDF^{+}13]. Hydroacoustic [DCL19, GH17, XM14]. Hydrochloride [WMD^{+}19]. Hydroelectric [AZE^{+}18, CBS^{+}12, CKC^{+}12, GBTL14, HFCH15, MHAP13, MAP14]. Hydrokinetic [ABČ^{+}15]. Hydrology [NCSR17]. Hydropower [SEP^{+}14]. Hydrosystem [RS18]. Hydroturbine [CBS^{+}12]. Hydroxide [WMSH13]. Hypotheses [WS11a].

IBI [OH15]. Ice [BLSS19, FHL18, TLL^{+}18]. Ice-Angling [BLSS19, TLL^{+}18]. Ichthyoplankton [AG15]. Ictalurus [SLS16]. Idaho [CMK^{+}19, CWH^{+}14, CMWM18, DB13, ES14, FBD^{+}19, GKT11, HHH14, KKMS11, KM11, KMK^{+}15, KD10, MNYV12, MQS13b, MM18a, MHDD16, MEL^{+}12, MGS14, MSM^{+}14, MS14, MSK^{+}16, ND11, NFQ16, SLMM10, SD11, TGS^{+}15, VWB^{+}11]. Identification [BS11b, BSB^{+}15, CSWP19, DBJD12, DSJS16, DCP^{+}19, FG13, FAW14, GL16, HH11a, HJB14, HHMQ15, HWFN14, HH15, McB14, MLEK15, PNHA12, PVST11, RFKB17, SM19, TBH^{+}14, WAW^{+}19, PZC^{+}12]. Identify [AM17, CSH^{+}18, DBK11, FDM^{+}13, LS11, MB12, MM18b, SDF12]. Identifying [HH10, RABK18]. IFQ [WHGC16]. II [AHS15, MRC12, WHC14b]. ilisha [NCSR17]. Ill [SCG11a]. Illinois [LDP^{+}10, CBDW15, DWW17, FHE^{+}12, HGW10, KCF15, LKV^{+}17, MCSF^{+}17, PVH^{+}13]. Illustration [MBBS10]. Image [SBB19]. Imagery [BOM16, HSF18]. Imaging [BP15, BWB18]. Immersion [SL17b]. Immobilization [BDB^{+}19, FMA^{+}12, LCR19, WBG^{+}17]. Immunosorbent [MEJ10]. Imnaha [FHC16]. Impact [BR14, BPB13, BLSS19, BRA18a, BLM10, CHPP11, DDD^{+}18, LS10, LHS^{+}14, MSK19b, SL17a]. Impacted [VBBK16]. Impacts [ACS19, ERC^{+}11, FCH^{+}13, HMCT16, HKP^{+}14, HHS^{+}13, KBK13a, KBK13b, LSC^{+}16, LS17, LBL11, NEH10, PFRM12, SDDC16, SW17, SMW16, SMW12, TPB^{+}16, YC11]. Impair [SKBB15]. Impairment [FDHWP16, SBL^{+}18b]. Impedance [SMCS14]. Impede [WBG^{+}17]. Impediment [HCH13, MRB^{+}11]. Impediments [JM12a]. Imperfect [BHW13]. Imperiled [BWHM16, DCS11, HW19, RTJJ16]. Impervious [UML^{+}11]. Impinged [BKD^{+}10, BP14]. Impingement [KSV^{+}10]. Implant [DBKG14, LN12a, MM16, MLHS13, NBP15, YHP10, ZC10]. Implantation [BRSB^{+}16, BHC^{+}10, DBW13, FVH^{+}17, KWMB11, WWS^{+}14, WRNJ^{+}19, WF17]. Implantations [AAYC14]. Implanted [FBL^{+}19, GM12, GLH^{+}18, HBLL12, TPC^{+}15]. Implanting [CHL13, OL17]. Implants [CSV13]. Implement [JDH16]. Implementation [VMA^{+}16, WHGC16]. Implementing [RABK18]. Implications [AK12, BACRB14, BATK16, BKDB17, BGBG17, CBS^{+}12, CGG^{+}18, DS12, DBS19, DC13, DKSW15, DGF11, FILH15, GAH^{+}18, GRR^{+}17, HS19, HMBT16, HB14a, HK17, HPTH18, HHB12, HT11, KLR^{+}12, MD10, MCP15, MQS13b, MP14, MCSF^{+}17, PS10, PDH^{+}19, Sam16, SCS17b, SMP^{+}14, SKBSA11, SSK18, TPBF15, TSFS19, WDD^{+}14, WLK^{+}16, WQP13, WHC12, ZDC14, BWHM16]. Importance [ABP^{+}18, BHHA11, BEH^{+}19, GBTL14, SM10, TBIJ19, WCY^{+}18]. Impounded [SPWC15]. Impoundment [NKL18, SSPP19]. Impoundments [FSW16, FHE^{+}12, HWDA12, HBA^{+}12, KSG^{+}18, LDW10, LSQ12, MOJ12, WWD13]. Improve [BB12, CKC^{+}12, KIST16, MB17, ORL^{+}16, WEC^{+}19]. Improved [GH17, HHCH15]. Improves [JCB^{+}18]. Improving [DIK19, HB12, HR11b]. In-Person [CFBB15]. In-Season [CJ14]. In-Situ [TGKZ19]. In-Water [Rey15, SL15b]. Inbreeding [SH11]. Incidence [HRH10, LMB^{+}18].

Incidental [CAS16]. Include [HBS14b, SPW^{+}12]. Inconnu [SSS15]. Incorporating [BB12, RFKB17, SKS^{+}15]. Increase [CWB16, KLR^{+}12, Mat10, RPBM10, SHMS16, TFH^{+}16]. Increased [AKP15, BM14, CCC16, LBF17, PM13, SBG^{+}19]. Increasing [BDS^{+}11, BDSB15, BGM17, BKP^{+}14, CBWB14, GCLC18, RRH^{+}11]. Increment [HB16, HB14b, WS16]. Incrementally [SDF^{+}13]. Incubators [PPFC18]. Independent [FWG15, GGCF14, Vin14]. Index [Ano12b, CKCF12, JY11, PW13]. Indian [NCSR17]. Indiana [MCSF^{+}17]. Indicate [CCDF18]. Indicator [FSW18, WP11]. Indicators [HSG16, MG10]. Indices [BS10, LS13, OH15, PLCG13, THZ^{+}17, ZC15]. Indirect [BSS18, TPBF15]. Individual [CSH^{+}18, HWNR14, HEL^{+}15, LS17, MSS^{+}12, MWS15, PHS^{+}13, RGH^{+}13, SF16, WHGC16]. Individual-Based [HWNR14, PHS^{+}13, RGH^{+}13]. Indoor [TPL18]. Induce [BDB^{+}19]. Induced [SMP^{+}14, SOSL10]. Induction [JTB16, LP12, ZGW11]. Industry [AG11]. Inequality [SBC^{+}11, SCW13]. Inert [HSF18]. Infection [SMP^{+}14]. Inferred [AES11, BFL^{+}10]. Inferring [BG12]. Infested [SM13]. Influence [BRN16, BMMMLML15, CSF12, CBS^{+}12, CF13, DHH17, DASB16, DWB^{+}15, FHC16, HNH11, HHP^{+}12b, HSHH17, HEL^{+}15, JWC10, KMPP12, LSM^{+}16, LBB16, LLW^{+}18, MCP15, MOJ12, NCSR17, PTH^{+}10, RS18, RLC12, SHSCA16, SM14, TLC^{+}19, TRB^{+}13, WKTQ15, WP11, WCWA10]. Influences [CBM13, CKCF12, DCS^{+}16, ES17, FWR^{+}15a, HW16, MNYV12, MSHC15, WWW15]. Influencing [HPM^{+}18, KSV^{+}10, LFD^{+}17, LSK^{+}14, RHG13, SNTZ14]. Inform [AWG^{+}18]. Information [Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, FB14, FJB16, FHH16, LTG^{+}11]. Informative [Kae13]. Infrared [HSF18]. Infusion [SNT^{+}16]. Ingestion [LMB^{+}18, SPP^{+}18]. Initial [LS17]. Injected [PFWG16]. Injuries [CC10, KHBH19, MHBR19, SVD^{+}18]. Injury [CC15, DMRJ10, HRH10, ILBJ12, SOB18]. Inland [FC17, JHK11]. Inner [WHB^{+}14]. Insects [Erm12, FSV10, FSV12]. Inshore [AS10, WFWSV14]. Insight [OLOI13]. Insights [DPK^{+}17, RIMR19]. Install [IH11]. Installation [PPFC18]. Installed [ABVS14]. Instantaneous [AWG^{+}18, WH11]. Instantaneous-Rates [WH11]. Instream [DFMP18, GDLM17, KCRR19, MRA14, PPFC18, RBP16]. Intake [SPT^{+}12]. Integrated [AAYC14, BRB^{+}17, BR12, Bur11, BSTC13, CHJ^{+}16, FOH^{+}12, GM12, HHP12a, HUST17, HWWT15, KWMB11, KZ12, MM16, MHG^{+}11, NKC^{+}18, OL17, SNT^{+}16, SLSS16, SKBK12, WF17, WHG^{+}16, YHP10]. Integrating [CH17, WEC^{+}19]. Integration [BKHS14, FB14]. Integrity [BKHS14, OH15]. Intensify [MDM12]. Intensity [SLP13]. Interact [PM14]. Interaction [RCP^{+}13]. Interactions [BRS^{+}12, LDW10, MOJ12, RW11, SPT^{+}12, SCB^{+}19, WCWA10]. Interannual [KTN14]. Interior [ACRBB10, MRA14, MAG^{+}15]. Intermixed [HSHH17]. Internal [WBP13]. Internally [WS13].

International [WHGC16]. Internationally [DZ14, DZ17]. Interpretation [DSD^{+}13, OH15, PC12]. Interstate [MCSF^{+}17]. Intertidal [WHC12]. Intra [SCWP17]. Intra-Annual [SCWP17]. Intracoelomic [BDC^{+}13, FVH^{+}17]. Introduced [ASL^{+}13, DASB16, FMP19, HTH^{+}19, HGW16, HHH14, MSC13, PT15, SLR10, WGJ10]. Introducing [NWZ14]. Introduction [DASB16, KMB^{+}12, NBFL17]. Introductions [GBFR18]. Introgressed [DLC11]. Introgression [BSVB19, HRKB19, VEW^{+}13]. Introgressive [LDP^{+}10]. Invaded [LC14]. Invasions [RFHP17]. Invasive [BS12, Bal15, BT16, BWHM16, CBHS16, CSWW18, FBB19, LSC^{+}16, NCMT12, RWAJ16, SGCB19, SEB^{+}19, SOSL10, SC18, SPWC15, TGKZ19, TGKI10]. Invasive-Induced [SOSL10]. Invertivorous [SS16]. Investigating [LDW10, TCM19]. Investigation [AZE^{+}18]. Investigations [SWW14]. Iowa [BPQ13, FSW16, NQ11, SPQ12, SQP12, WFS13]. Irish [HT11]. Islands [CH17, TS13]. Isle [SQGK10]. Isolated [RFHP17]. Isolation [BBO16]. Isotope [HTPS10]. Isotopes [FDM^{+}13, THQ15, WGM^{+}11]. Isotopic [EWS^{+}17]. Issue [Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g]. Issues [WLK^{+}16]. Italy [PLCG13]. Iteroparity [Cle15]. Iteroparous [TFH^{+}16]. IVb [TPBF15].

J [KSH^{+}16, SMS13]. James [ABBG13, BRS^{+}12]. Japan [KSMN18, OMT^{+}15]. Japanese [Nak13]. japonicus [BFLT17]. Jaw [VBT^{+}12]. Jaw-Tagged [VBT^{+}12]. Jersey [ADMC10, MBPG12, SOB18]. Jet [AG15, DMRJ10]. Jet-Propelled [AG15]. Jigs [RS11]. Joaquin [BGD^{+}19, BSBF13, BBS18, NB10, PSB^{+}10b]. Job [HBS14b]. John [ALCD17, LAA^{+}18, SDL^{+}15]. Johnson [CWH^{+}14]. Joint [BGP10, LTG^{+}11]. Jonah [TDM19]. Jordan [PPFC18]. Journal [Ano16b, Ano19h, MS17b]. Juan [CMMR19]. Juglines [SS13a]. Juvenile [APPR14, AT18, ASL^{+}13, ASL^{+}16, AAYC14, BQHR17, BSL18, BFLT17, BKHS14, BB12, BG12, BRA^{+}18b, BDB^{+}19, BHC^{+}10, BSBF13, BBS18, BKP^{+}14, Bur11, CBS^{+}12, CGM13, CSH^{+}17, CCDF18, CC10, CDBW17, CHJ^{+}16, CHL13, CJH16, CHW12, DMRJ10, DBJD12, DWW17, DBE11, DG12, ERC^{+}11, FBL^{+}19, FOH^{+}12, FCH^{+}13, FTSF13, GLH^{+}18, GPP^{+}10, GDLM17, Gur12, GBTL14, HHT^{+}13, HW10, HJB14, HHP^{+}12b, HPM^{+}18, HRH10, JKSK13, JM15, KWMB11, KSM^{+}18, KTN14, KLR^{+}12, KDMH10, KSM16, LKV^{+}17, LP12, MHR^{+}18, MFLS18, MCB^{+}19, MSD11, MMB19, MCTM14, NSW16, NB10, OHSP17, ORA17, PSB^{+}10b, RGOR10, SMP^{+}14, SB18, SBL11, SMK13, SL17b, SSP18, SHMJ19b, SS16, TFP^{+}11, WBP13, WWS^{+}14, WWW15, WPY^{+}15, WMSV18, WHMS14, ZWS^{+}19]. Juvenile-Based [BKHS14]. Juveniles [Bou10, TBH^{+}14].

Kansas [KNC14, KSG^{+}18, NKC16, NKL18]. Katrina [PKKR10]. Kayaks [AG15]. Keeping [CL12]. Kelt [HFB^{+}13, TFH^{+}16]. Kelts [BS11a]. Kettle [DFPS10]. Key [LKM^{+}17]. Kidney [MEJ10]. Kiyis [LOV17]. Klamath [EHP^{+}16]. Know [PZC^{+}12]. Knowledge [TDM19]. Known [BMW^{+}19, CHDW18]. Known-Age [BMW^{+}19]. Koi [DHLD11]. Kokanee [AJ13, Ask16, BMW^{+}19, Han19, KQDC19, LR12, PJLB14, WAW^{+}19, WGJ10]. Kootenai [BQHR17]. Kootenay [LR12].

Labeling [WMD^{+}19]. Laboratory [BP14, FWT14, OL17, RHC^{+}14, SWW14, WPY^{+}15, WDW13]. Lacs [KMPV13]. Lacustrine [AW10]. Laden [DHB^{+}10]. Lagos [Aka10]. Lahontan [APAC12, SRSS12, SBVB12]. Lake [AHS15, ABVS14, BBG19, BGCF11, BATK16, BRBP10, BBL^{+}11, BSB^{+}15, BGM17, BHR^{+}15, CSF12, CSED10, Car13, CJMD10, CBHB16, CSH^{+}17, CSWW18, CDQG12, DWW12, DFPS10, DGF11, EWS^{+}17, EDN10, FVH^{+}17, GZ13, HTF^{+}13, HIS16, HBS^{+}19, HFN11, HHT^{+}13, HFCH15, HER^{+}12, HB14a, HMS11, HUB^{+}18, HHH14, JHS15b, Kae12, KPL^{+}16, KBH^{+}19, KYD18, KMPV13, LOS^{+}11, LLP10, LBF17, LWJS16, LMD^{+}19, MD10, MHAO11, MKRH14, MCB^{+}19, MHAP13, MAP14, McG14, MKWM18, NCJ^{+}11, NFQ16, PJLB14, PCP^{+}13, PSR16, PM13, PDL^{+}18, RBB10a, RSHS15, RTMJ10, RRH^{+}11, RPJ18, RMTW16, SCKZ17, SS11a, SH11, SRFW14, SRFS15, SBVB12, SSE19, SBHJ17, SKBSA11, SKBB15, SHMJ19b, SSK18, SQGK10, SGC13, TS11, TGKZ19, TPBF15, WA16, WFS13, WHV15, WSA18, WEC^{+}19, WMD^{+}19, ZEM11, ZM17, ZWH^{+}15, AJ13, BATK16, BS11b, BFS^{+}19]. Lake [BBL^{+}11, BSB^{+}15, CCMA^{+}10, CB10, DRC17, EWS^{+}17, FB14, FDM^{+}13, HER^{+}12, HMS11, HHH14, JDG10, Kae12, Kae13, KRB^{+}16, KBH^{+}19, LOS^{+}11, LR12, LSL^{+}13, LOV17, LJR11, MD10, MRB^{+}11, MMS^{+}17, MHDD16, PNWC11, PM13, PTD10, PPO18, RHG12, RHG13, SG13, SS17, SHMJ19a, SBTW17, SOSL10, SQGK10, TFP^{+}11, TGKZ19, TLL^{+}18, VKBL11, VBT^{+}12, WSS13, WPJ^{+}18, ZKM13, ZM17]. Lake-Origin [NCJ^{+}11]. Lakes [BS12, BKME17, BSBC14, CWB16, CMK^{+}19, CBDW15, DWB^{+}15, DRW^{+}18, DWW17, EJL10, FBB18, GKT11, GMB10, HHCH15, HSHH17, HADG10, ILBJ12, ISC10, JHK11, KIST16, KMB^{+}12, KKMS11, KM11, KMK^{+}15, LSM^{+}16, LLP10, LC14, LAM16, LP17, McC15, ND15, ORL^{+}16, PHFK10, QA16, RWCP11, RS17, RSI16, Ryp15, SEP^{+}19, TPB^{+}16, WS11b, BT16, BJ19, Car13, CBC^{+}16, CBC^{+}17, HSG16, JBRWH18, KSL^{+}19, LSC^{+}16, LTG^{+}11, ML13, RRH^{+}11, SBHJ17]. Lakeshore [LGC11, LSL^{+}13, TFP^{+}11]. Lampetra [DSJS16]. Lamprey [AM17, BT16, BWHM16, CKC^{+}12, DSJS16, GRR^{+}17, ILBJ12, JM12a, JSDJ16, KDP^{+}10, MSW^{+}14, RWAJ16, RM12]. Lampreys [APD^{+}19, CGM13, DCHC13, GDLM17, MCK^{+}15, MJH17, NTJ^{+}14, SHK^{+}17, SBL^{+}12]. Land [BG19, Bre13, MKK^{+}12]. Land-Use [Bre13]. Landed [MBPG12]. Landing [LLW^{+}18]. Landings [MG10, SCS17a]. Landings-Based [MG10]. Landlocked [OEZ17, PRM15]. Landscape [ESL^{+}15, MJ11, Pto13]. Lane [GSCR11]. Laparoscopy [FG13]. Lapse [AWG^{+}18, SK18b]. Large [AZE^{+}18, BEH^{+}19, CF18, CVD^{+}14, DTA18, FSHR17, GRSF^{+}16, GDLM17, HDKB10, HNPB15, HPTH18, KFM12, LSM^{+}16, McG14, MLHS13, NCJ^{+}11, PBA^{+}17, PT11, SG13, SSK19, SS11a, SPT^{+}12, SE12, SEP11, TPB^{+}16, TMC10, WKS^{+}10, WQP13]. Large-Format [MLHS13]. Large-Reservoir [SSK19]. Large-River [GRSF^{+}16]. Large-Scale [CVD^{+}14, PT11]. Largemouth [BLSS19, BKM19, BRSA15, CI10, CF18, DVWS10, DLSW12, DDEW18, DAKP13, DLC11, FC17, FSW18, FHE^{+}12, GKS13, HWDA12, HBS14a, HVW^{+}17, KIST16, KDA^{+}18, KADH15, LGC11, LN12c, LSK^{+}14, MASN16, MCS^{+}17, NTJ15, NSW16, NKL18, PLCN17, PTD10, SPP^{+}18, SLK15, SL15a, SB18, SOBW10, SKBK12, SQS^{+}17, SS16, TGA15, THI^{+}19, THZ^{+}17, TP17, WBG^{+}17, WI17, WS11b, WWD13]. Larger [SL12]. Largest [GRR^{+}17, HMBT16, HWB11]. Larus [FOH^{+}15]. Larvae [BFL^{+}10, BFS^{+}19, JSDJ16, LBF17, PRR^{+}13, SPS16, UWL13]. Larval [DCHC13, FRH10, GDLM17, GZ13, GF18, KYD18, PRRD15, RFLP14, SHK^{+}17, WGM^{+}11]. Laser [RWT17]. Late [CMD12, JKN^{+}10]. Lates [HH12]. Latitudinal [KMPP12]. Laurentian [BT16, LSC^{+}16]. Lavage [BBW18, HNH11]. Lead [MSK19b]. Leaders [SBD^{+}11]. Leading [SBG^{+}19]. Learned [KCRR19]. Learning [CSWP19]. Lees [AKP15, BRN16]. Legacies [FQ19]. Legal [HIS16]. Length [BMM^{+}12, BS10, BBC10, BBC11, BBS19, BGW^{+}11, BHW13, CI10, Car13, CF18, DBKG14, DMC^{+}19, DC13, DN11, EPM16, FRH10, Ger11, ITT10, JF13, LKV^{+}17, MSW^{+}14, MH13, MCS^{+}17, MBPG12, MIH15, NEH10, OI17, Pie10, PSCS11, PA16, PCP^{+}19, PDH^{+}19, RFW^{+}10, RFW^{+}11, Ran18, RV12, RWT17, RD10, SHSCA16, SLMM10, SR16, SSP^{+}19, SLS16, SEP^{+}19]. Length-Based [NEH10, PCP^{+}19, PDH^{+}19, SLS16]. Length-Related [Ger11, RFW^{+}10, RFW^{+}11, RV12]. Lengths [BK12]. Lentic [BKSH^{+}12, BQR^{+}19, FQ19, MSK^{+}16, TCM19]. Lepisosteids [BKDB17, BR12]. Less [HW12, VPV18]. Less-Certain [HW12]. Less-Restrictive [VPV18]. Lessons [KCRR19]. Lethal [CJH16, NSW16, SK18a]. Level [DBK11, LSM^{+}16, TPF^{+}13]. Levels [Aka10, DT11, GAWC10, ILBJ12]. License [HBD^{+}17, TBIJ19]. Life [AT18, BACRB14, BSM17, BPT^{+}19, BWR12, CSED10, CAWB17, DS12, DDD^{+}18, DSD^{+}13, EMK^{+}17, HGK15, JBZ^{+}10, KRB^{+}16, POWF11, RM14, SJM15, UACZ^{+}16]. Life-History [DSD^{+}13]. Light [BDS^{+}11, FWT14, MJH17, SWG^{+}16]. Light-Emitting [SWG^{+}16]. Lighting [SBL^{+}12]. Lights [SWW14]. Lightweight [HBDA13]. Like [BJ19]. Likelihood [BGP10]. Limit [Car13, EPM16, Pie10, PG14, PS10, RD10, Ryp15, SHSCA16, SCS17a, SSP^{+}19]. Limitations [MSK19b, SPW^{+}12]. Limited [HKMH10, KSL^{+}19, MM18b, SRCC^{+}18, SOBW10]. Limiting [KW13]. Limits [Ask16, CI10, DMC^{+}19, MCS^{+}17, MM18b, MIH15, ORL^{+}16, THQ15]. linbin [WTB^{+}15]. Lindgren [SEP11]. Line [Ano12b, Ger10, JY11, RPMB19, SOB18]. Linear [SBL^{+}18a]. Linked [KSL^{+}19, MEJ10]. Linking [APP13, BG19, UML^{+}11]. Lipid [SMCS14]. Liquid [WMSH13]. Listed [MAM^{+}12]. Litigation [SBG^{+}19]. Littoral [DWW12, EJL10, RSI16, WWW15, WS11b]. Littoral-Zone [EJL10]. Live [BEH^{+}19, Hen15, SVFG10, SLB10, SHS15]. Live-Release [SVFG10]. Live-Well [BEH^{+}19, SHS15]. Lived [PC12]. Lobster [BC18, LCC^{+}15, ZC12, ZC15]. Local [BBO16, EWS^{+}17]. Location [BS10, RPC^{+}19, Sam19, SCW^{+}17, WKTQ15]. Locational [WPS^{+}19]. Locations [AHAC15, BFL^{+}10, BWB18, CCC16, KBH^{+}19, Ols19, WSS13]. Loci [LR12]. Lock [MH11a]. Logic [MSHC15]. Long [APC^{+}14, BCD^{+}10, BDC^{+}13, HWZ19, HHRB15, KML14, KCRR19, PC12, Pie10, PRRD15, RWPH11, SBC^{+}11, YHP10]. Long-Lived [PC12]. Long-Term [APC^{+}14, BCD^{+}10, BDC^{+}13, HWZ19, HHRB15, KML14, KCRR19, Pie10, PRRD15, RWPH11, SBC^{+}11, YHP10]. Longevity [FSW18]. Longitudinal [MC10, SVPD13]. Longline [MFRS11]. Longnose [DC13, PC12]. Lord [HT11]. Loss [BKHS14, Bur11, KBK13b, KPL^{+}16, KD10, LKB19, MM18a, SC17, SBG^{+}19, WPY^{+}15]. Losses [GRR^{+}17]. Lost [AHVJ11, Bur11, BHR^{+}15, EHP^{+}16]. Lotic [MJ11]. Louisiana [PKKR10]. Low [CBWB14, CCDF18, GAWC10, GBHO11, HKMH10, HGW10, HHR^{+}18, JM12a, PNHA12, RRH^{+}11]. Low- [HHR^{+}18]. Low-Abundance [PNHA12]. Low-Conductivity [HKMH10]. Low-Cost [CBWB14]. Low-Elevation [JM12a]. Lower [SK18a, ZOC^{+}14, APPR14, BBD19, EJL10, GMG^{+}15, HHR^{+}18, JBZ^{+}10, KCF15, KSM^{+}18, KB11, LO18, MCSF^{+}17, MSK19a, RHPH14, SHMJ19a, SHMJ19b, SPK10, SPP12, ZLB10]. Lower-Mode [ZOC^{+}14]. Lowland [TWL^{+}17]. lucumonis [GFP^{+}12]. Lunar [Pul17]. Lures [SPP^{+}18]. luteiventris [BKSH^{+}12].

Machine [CSWP19]. Mackerel [SPM15]. Macroinvertebrate [FSS^{+}10, HMW^{+}10, HMCT16]. Macroinvertebrates [GKMC15]. Macrophthalmia [GRR^{+}17]. Macrophyte [SS16]. Madison [BGW^{+}17]. Madtoms [CKBP19]. Magnitude [HFCH15, MHC10]. Mail [OM10]. Main [DB13, HER^{+}12, KJN^{+}17, SEP^{+}14, SM19]. Main-Stem [DB13, SEP^{+}14, SM19]. Maine [BC18, CZH^{+}19, HB12, MCH12, SZHC15, SDPB17, WBZ19, ZC12, ZC15, ZOC^{+}14]. Majority [ISC10]. Make [JM12b]. Making [SLR^{+}16]. Malaclemys [ASD^{+}18]. Male [Bal15, BT16, SLG^{+}15, SMH17]. Males [ISC10]. Mammal [BM14, EF10]. Managed [MS17a, PDD^{+}18, SCS17a, SCMP13]. Management [AG11, AW10, AK12, Ano16b, Ano19h, BB19, BGD^{+}19, BGBG17, BRA18a, BQR^{+}19, BW17, BLM10, BBC^{+}13, CvPW15, Can17, CBS^{+}12, CA14, CJ14, CBHS16, CAWB17, DBS19, DLP16, DA12, DAKP13, EF10, FF15, FILH15, FJB16, FQ19, FWR^{+}15a, GGC17, Gre11, HW12, HTF^{+}13, HP14, HWDA12, HMBT16, HB14a, HSHH17, HVW^{+}17, HCH13, HWWT15, HPTH18, HT11, JDH16, JBRWH18, KBK13a, KBK13b, LTG^{+}11, LBB16, LRHR17, MBBS10, MKBB13, MCP15, McB14, MP14, MS17b, MCSF^{+}17, MEJ10, NFQ16, PM14, PCT^{+}14, PPO18, RCGF18, RWAJ16, RGOR10, RIMR19, RJMG13, SCS17b, SW17, SDKS16, SCW13, SKBSA11, SOS^{+}15, Sta12, SSK18, TPBF15, TSFS19, UML^{+}11, WLK^{+}16, WQP13, ZDC14]. Managers [MBBS10]. Managing [Ask16, BKDB17, HGK15, McB14, PJLB14, PLUH19, ZC12].

Mandatory [CI10]. Manipulation [BRA18a]. Manistee [GAH^{+}18]. Manitoba [SPHK16]. Manual [BSTC13, SMH17, TL15]. Mapping [ESL^{+}15, RWCP11, WA16]. Margin [PPO18]. Marine [BM14, EF10, McB14, MG10]. Mark [BGP10, CHBY13, EKM10, FOH^{+}12, HD14, HV18, HUB^{+}18, HQ10, KDMH10, MWS15, RFLP14, RHMH15, SC17, VEIW^{+}16, WWB^{+}17, WEC^{+}19, WGM^{+}11, YC11, ZEM11]. Mark-Selective [CHBY13, YC11]. Marked [SFW11, SLSS16]. Marker [MSS^{+}12]. Markers [MWS15, SDF12]. Market [Lee12, SC14, WHGC16]. Marking [BFLT17, CMF^{+}14, CHW12, JDG10, LN12a, LP12, NBP15, SL17b, WGM^{+}11, ZC10]. Markings [OMT^{+}15]. Marks [CSH^{+}18, MM16]. Marlin [GMG16b]. Martin [SG13]. Maryland [MKK^{+}12]. Marys [RWAJ16]. Mass [BMM^{+}12, EKM10, JDG10, OMT^{+}15]. Mass-Marking [JDG10]. Massachusetts [KHDA14]. Matching [AW10]. Material [KZ12]. Materials [RM12]. Maternal [AJR^{+}19]. Matter [Bre13, DLC11]. Maturation [FWR^{+}15a, HLM^{+}18, MSK19b, SLG^{+}15, Sta12, SDPB17, TLC^{+}19]. Mature [BR14]. Maturity [BWB18, JF13, LS11, SLMM10, TO10]. Maumee [MKWM18]. Maxillae [WHV15]. Maxillary [MM16, WSA18]. Maximize [ASD^{+}18, CvPW15]. Maximum [BHC^{+}10, CI10, DT11, HWB11, HT11]. Maximum-Count [HWB11]. May [MDM12, SH18]. McNeil [POWF11]. Mean [CF18]. Measurably [RPBM10]. Measure [HS12, RBB^{+}10b]. Measured [MK13]. Measurement [AMQSK15, DC13, HDB^{+}12, RWT17]. Measurements [BGW^{+}11, CBHB16, Gur12, Rey15, SL15b]. Measures [ACMB^{+}14, CHPP11]. Measuring [BM14, LM16b]. Mechanisms [WFS13]. Mechanistic [BB19]. Media [HSF18]. Mediate [CEP15]. Medium [KDMH10]. Medium-Sized [KDMH10]. Meet [DZ14]. Melanophore [MSS^{+}12]. Mendocino [GWCA10]. Menhaden [LBL11, SW17]. Menippe [DKSW15, GCLC18]. mercenaria [GCLC18]. Mercury [LLP10]. Mesh [LLW^{+}18, WKS^{+}10]. Meta [HM12]. Meta-Analysis [HM12]. Metacommunity [HMCT16]. Metal [BDT^{+}16, PSCL13]. Metallic [LMB^{+}18]. Method [BSS18, DCS11, DZ17, ES17, Ger10, HJB14, MFLS18, OOWW12, PLUH19, RSI16, RWT17, TCG^{+}17, WA14, ZM17]. Methodologies [VEIW^{+}16]. Methodology [ABP^{+}18, PK18, WCY^{+}18, WHC14a]. Methods [BA18, CHBY13, CJH16, DSJS16, ETD16, FSW16, FOH^{+}12, hGgXfC^{+}13, HTPS10, HH10, HCL^{+}12, Kae13, KD10, KCCW19, LKB19, LSLC15, McC17, MM18a, NSW16, PN19, PVH^{+}13, RHMH15, SCS17a, TGKZ19, WI17]. Methow [SMK13]. Metolius [RSHS15]. Metric [MQJK19]. Metrics [HDB^{+}12, Kae13, LRHR17, SBDB16]. Metropolitan [SK18b]. Mexico [AOS^{+}17, AES11, BMMMLML15, CK10, FWWTJD16, GGC17, GGCF14, HGW16, JHS15a, PA16, RGC^{+}12, RCMB18, RKB15, SC17, SDKS16, SFH^{+}13, SKS^{+}15, SM14, TSFS19, Vin14, WGS16]. Michigan [CBC^{+}16, KSL^{+}19, WPJ^{+}18, BGCF11, BG13, CBC^{+}17, DRC17, ESL^{+}15, GAH^{+}18, HHT^{+}13, HWZ19, KRB^{+}16, KBH^{+}19, LMD^{+}19, MD10, MRB^{+}11, NWZ14, PSR16, PDL^{+}18, RRH^{+}11, SS17, TPBF15, ZW12]. Microacoustic [GLH^{+}18]. Microchemistry [CFG16, DPK^{+}17, MPL^{+}15]. Microhabitat [HLS^{+}17, PSB^{+}10a, WMB19]. Microsatellites [FCE^{+}12]. Microstructure [DSD^{+}13]. Microtag [FBL^{+}19]. Mid [SHL16]. Mid-Atlantic [SHL16]. Middle [CS17, RPJ18]. Midwater [BSS^{+}14]. Midwest [CBHS16, FSHR17, OH15]. Midwestern [BS12, SP13, WSH^{+}14]. Migrant [CVD^{+}14, DBE11, SPM15, WA14]. Migrants [Bou10]. Migration [BBO16, BSM^{+}15, BJ19, CMD12, JM12a, JKN^{+}10, KJN^{+}17, MMP12, MS18, MHAP13, MCK^{+}15, MSK19a, NMG^{+}17, PM14, PSB^{+}10b, PFD17, RFQ14, TRB^{+}13, VKH^{+}19]. Migrations [EMK^{+}17, FK16]. Migratory [ABBG13, BGP10, CBHS16, ES14, JPWR10, MJCM15, NCSR17, RH14, SBL^{+}12, TWL^{+}17, TCG^{+}17]. Mille [KMPV13]. Mini [BSBC14]. Mini-Fyke [BSBC14]. Minimal [DIK19]. Minimizing [BDSB15]. Minimum [MIH15, NKC16, SHSCA16, SSP^{+}19]. Minimum-Length [SHSCA16]. Minnesota [GHK17, BCBW18, CSF12, CI10, DFPS10, DWCK16, GHK19, HDM15, ITT10, ISC10, IMM13, KMB^{+}12, KMPV13, LAM16, PNWC11, PCT^{+}14, Pie10, RS17, SS11a, SF14, SCMP13, Sin18, SQGK10, TS11, WSS13]. Minnow [AR12, GHL12, NBP15, PDM12, PSC15]. Minnows [PSC15, RPC^{+}19]. Mirror [SHMJ19b, SHMJ19a]. Mississippi [AJ10, DLP16, HBS^{+}19, HHS^{+}13, LDP^{+}10, LKM^{+}17, MH11a, PTH^{+}10, PVH^{+}13]. Missouri [FRL16, GMG^{+}15, SPK10, SPP12, ABVS14, BFL^{+}10, CFG16, GWB^{+}10, HPD^{+}19, HB10, HPTH18, MOJ12, PVH^{+}13, RD10, STP19, SSP^{+}19, WCWA10]. Misunderstood [ALCD17]. Mitigation [CGL18]. Mitochondrial [AES11]. Mixed [FBC^{+}10, HJB14, WHB^{+}14, WLK^{+}16, ZEM11]. Mixed-Mode [WLK^{+}16]. Mixed-Origin [ZEM11]. Mixed-Race [HJB14]. Mixed-Stock [FBC^{+}10, WHB^{+}14]. Mixing [WH11]. Mixture [GRSF^{+}16, MB17, VKH^{+}19]. mm [BRB^{+}17, OL17, SSP^{+}19, TPC^{+}15]. Mobile [BGM17, RFKB17, SBL11]. Mode [BvPCH17, WLK^{+}16, ZOC^{+}14]. Model [AWG^{+}18, BB19, CBS13, CHPP11, DPS^{+}17, FHH16, GHK19, HF13, KYD18, MR16, ND15, PSCL13, PHS^{+}13, RGH^{+}13, RZ10, See11, SCG11a, SC14, SEM14, WHGC16, WHMS14, WEC^{+}19, ZKM13]. Model-Assisted [MR16]. Model-Based [AWG^{+}18]. Modeled [CSH^{+}17]. Modeling [AG11, ASD^{+}18, BLM^{+}11, BSO^{+}16, BRA18a, CB10, DMC^{+}19, EPM16, McG14, MSHC15, NCSR17, NB10, PP10, SBDB16, SCMP13, SLS16, SMCS14, VKH^{+}19]. Modelling [BBG19]. Models [BBC^{+}13, CWH^{+}14, CHBY13, DLP16, FTSF13, GRSF^{+}16, HP14, HHCH15, HWNR14, McC16, MJ11, MB17, NW12, NEH10, PS10, Pto13, RBP16, SBL^{+}18a, SLMM10, SHK^{+}17, SPQ12, WH11, WH12]. Modernizing [VEIW^{+}16]. Modification [SHMJ19a, SHMJ19b]. Modified [BP14, GMG^{+}15, HB12, JJ10, Kro19, SBW^{+}16, SS17]. Mohawk [DG12, Gur12]. Mohn [Der14]. Moines [SCWP17]. Moisture [SMCS14]. Molt [SGF^{+}13]. Monetary [JBHM19]. mongolicus [LXM^{+}12].

Monitoring [BRB^{+}17, BCD^{+}10, Bou10, CFBB15, DRG10, DCS11, GWCA10, GTB^{+}19, HWFN14, HK17, KCRR19, LCC^{+}15, PRRD15, RÅW18, RBB^{+}10b, SB15, SKS^{+}15, VWB^{+}11, VEW^{+}13, WDD^{+}14, WA14]. Montana [BKSH^{+}12, BGW^{+}17, DB13, FBB18, NEC^{+}13, SGC13]. Montane [CJA^{+}13, HHRB15, PJLB14]. Month [APF18]. Months [WBF^{+}15]. Moore [WBGM13, WBF13]. Mooring [TMC10]. Morphological [CSWP19, CBHB16, DSJS16, MGB^{+}14, PVST11]. Morphology [DHB^{+}10, HWZ19, LMC^{+}16, LPY16]. Morphometric [LC14]. Mortality [APD16, AHVJ11, BSL18, BEK^{+}18, BDSB15, BRA^{+}18b, BGM17, Bur11, CGDB14, CZH^{+}19, CBS^{+}12, CSED10, CHBY13, DPS^{+}17, DMRJ10, FWWTJD16, GGCF14, GMG16b, HFN11, KSH^{+}16, KJN^{+}17, KADH15, LXM^{+}12, MS16, MG14, MH11a, MS14, MHC10, NKC^{+}18, ORA17, PB13, PTH^{+}10, PVH^{+}13, PTD10, RS11, RPMB19, SHMS16, SC17, SCS17b, SS13a, SB11, SVFG10, SKH^{+}16, SPHK16, SBHJ17, SBG^{+}19, SEM14, TS11, UWL13, Vin14, WPY^{+}15, WS13, WGS16]. Motivations [BHHA11, HAW^{+}18]. Mottled [RHMH15]. Mountain [BGW^{+}17, Erm12, FSV10, FSV12, IH11, McC15, Nak13, GBFR18, MH11b, PN19, SNTZ14]. Mountains [GKMC15, KKM16, RFHP17]. Mounted [HHMQ15]. Mouth [SBD^{+}11]. Movement [APAC12, ASL^{+}13, BQHR17, BB12, BKP^{+}14, DCS^{+}16, EBS18, FBD^{+}19, GDP14, GKS13, HGW10, HR11a, HH11b, HNPB15, HB14a, HSHH17, HHRB15, MHAO11, MJH17, MMB19, NCJ^{+}11, PSB^{+}10a, QK11, RBB10a, SSS15, SKBB15, SCB^{+}19, WHKB17, YS12, ZS16]. Movements [ABBG13, BRB^{+}17, CBC^{+}17, DHLD11, KHDA14, MAP14, TPF^{+}13]. MS [CGM13, WWS^{+}14]. MS-222 [CGM13, WWS^{+}14]. Much [KSM16, RPBM10]. Mud [RJMG13]. Mudsnails [SM13, SAF16]. Mulloway [BFLT17]. Multi [RPJ18, VMA^{+}16]. Multi-Criteria [VMA^{+}16]. Multi-Lake [RPJ18]. Multicompartment [PRR^{+}13]. Multidecadal [WFWSV14]. Multifilament [SBDB16]. Multimesh [Han19]. Multinomial [MB17]. Multiple [BA11, BFL^{+}10, CWH^{+}14, MJCM15, NASG12, NMG^{+}17, PD13, SHL16, SQP12, WMB^{+}15]. Multiple-Pass [PD13, SHL16]. Multiregional [SL17a]. Multiscale [WTB^{+}15]. Multiseason [GRSF^{+}16]. Multispecies [Lee12, MSBP10, NW12]. Multistock [WQP13]. Multiyear [MEJ10]. Mummichog [BRSB^{+}16]. Mummichogs [GRCR10]. Murray [HADG10, FWR^{+}15b, FWR^{+}15a]. Murrumbidgee [FWR^{+}15b]. Muscle [HTPS10]. Muskellunge [APD^{+}18, AZE^{+}18, BCBW18, DWW17, DMC^{+}19, EDN10, FILH15, FH16, JHK11, KFM12, KMB^{+}12, KMK^{+}15, MW19, MWS15, ND15, ORA17, RWPH11, SSE19, SSP18, SWW14, WWW15, WRNJ^{+}19, WF17, WDW13, YHP10]. my [WPS^{+}19]. mykiss [HFB^{+}13, HHP^{+}12b, TRB^{+}13]. Myxobolus [ZLB10].

N [HTPS10, MB17, VKH^{+}19]. N-Mixture [MB17, VKH^{+}19]. Nahanni [ASP^{+}17]. Naive [OOWW12]. Natal [CFG16]. National [FSS^{+}10, GKMC15, HMW^{+}10, JBHM19, LSL^{+}13, PT15, SK18a, SZKK10]. Native [BWHM16, CEP15, CBHS16, CSWW18, DGS^{+}11, FBB19, GMG^{+}16a, HW12, HGW16, HB10, HBA^{+}12, Kae12, KKM16, ND11, SHMS16, TTS^{+}18, WQF15, WMSV18, WGJ10]. Natural [BGCS11, BKME17, BKDB17, BWHM16, CSH^{+}18, DR19, HFN11, HQ10, KPL12, KPP^{+}18, MS16, MSS^{+}12, MPM10, OOWW12, RHC^{+}14, RCP^{+}13, SKBSA11, TRB11, VBBK16, VJK13]. Natural-Origin [KPP^{+}18, MPM10]. Naturalized [KRB^{+}16, PNWC11]. Naturally [GWB^{+}10, JPWR10, PTH^{+}10]. Nature [BJ19]. Nature-Like [BJ19]. near [BBO16, DG12, KHBH19, MRA14]. Nearby [SSK18]. Nearshore [DV17, FWWTJD16, HPM^{+}18, SCKZ17, SS17, TFP^{+}11, WFWSV14]. Nebraska [HHR^{+}18, RHPH14]. Nechako [MMP12]. Necropsy [MHBR19]. Need [RFW^{+}10, SPQ12, SQP12]. Needed [NKC16, RPBM10]. Needs [DBK11, SCA11]. Negatively [LN12c]. Neosho [MMB19, TLSB18]. Nest [DLSW12, HBS14a, WS11b]. Nesting [LGC11]. Net [AS10, Ano12b, BDS^{+}11, BM14, BDSB15, BFL^{+}10, CSWW18, DHA12, GHL12, JY11, KHBH19, KSMN18, Kro19, LJR11, LLW^{+}18, RJMG13, SS11a, SMW16, SR16, SS11b, VKBL11, ZM17, PMP^{+}14]. Net-Covered [RJMG13]. Net-Pens [KSMN18]. Nets [AG15, BFN16, BSBC14, CBDW15, GMB10, Han19, SBD^{+}11, SBDB16, Sin18, SBW^{+}16, SBW^{+}17b, SS17, SL12, UWL13, WKS^{+}10, ZM17]. Netting [Ano12b, JY11]. Network [LTG^{+}11, VMA^{+}16]. Networks [BC17, WTB^{+}15]. Neutral [LR12]. Neutralizing [GKMC15]. Nevada [FSS^{+}10, HMW^{+}10, SRFW14, SRFS15, SBVB12]. Newly [WH11]. Nigeria [Aka10, NEH10]. Night [BKME17, CKC^{+}12, HDKB10]. Nighttime [SFW11]. Nipigon [SQGK10]. Nipissing [TLL^{+}18]. No [CCC16, SVD^{+}18, SCB^{+}19]. Nonchemical [PT15]. Noncompliance [HF13]. Nongame [DR19]. Nonintrogressive [PNWC11]. Nonlethal [ASP^{+}17, BKM19, CJH16, GL16, HTPS10, IWB10, Ise10, NSW16, TCM19, WSA18]. Nonmigratory [SMK13]. Nonnatal [CGDB14]. Nonnative [ACS19, BQR^{+}19, BHG^{+}13, CEP15, CDQG12, DGF11, Kae12, KKM16, KQRS16, ND11, NFQ16, PDD^{+}18, RFHP17, SHMS16, SMH17, SK18a, SNTZ14, VEW^{+}13, WQF15]. Nonphysical [MJH17]. Nonrandom [AHAC15]. Nonresident [LS10]. Nonselective [YC11]. Nonspawning [LS11]. Nontrout [LN12b]. Nonwadeable [NQ11]. Norms [CGL18]. North [FMP19, HH11a, HH11b, HH15, ORA17, AS10, Ano16b, Ano19h, BR14, KFW12, KCCW19, MS17b, MASN16, RH14, SCKZ17, SR16, SBW^{+}17b, SD11, WS11b]. Northeast [BGCH17, HCH13, Lee12, BM14, SM14, WS11a]. Northeastern [KW13, LMD^{+}19, SOS^{+}15, TCB18]. Northern [BKM16, CBHS16, CSWW18, GF18, hGgXfC^{+}13, HS19, HKP^{+}14, LSM^{+}16, LLP10, LN12c, MD10, MCH12, OLOI13, OLM15, ORL^{+}16, PC12, Pie10, RFLP14, SDDC16, SF14, SKBSA11, Sta12, TS13, WQF15, FILH15, FH16, GBL^{+}13, GHK17, HFH12, HHCH15, HFH18, JHS15a, ND15, RCMB18, RSI16, SC17, SFH^{+}13, WGS16]. Northwest [ASP^{+}17, SMW12, BSS^{+}14, CEH^{+}15, DCHC13, FBB18, GDLM17, KLR^{+}12, MFRS11, RBB^{+}10b]. Northwestern [HW10]. Nothing [Can17]. Nova [GBHO11]. Novel [SKBK12]. Nuclear [PMP^{+}14]. Nucleotide [SOS^{+}15]. Nulhegan [Kra18]. Number [McC16]. Numbers [SHK^{+}17]. Numerical [BBG19, OMT^{+}15]. Nunavut [GD10]. Nurseries [WFWSV14]. Nursery [PHFK10]. Nutritional [JHU^{+}13].

Oahe [FRL16]. Objectives [LBB16, MP14]. Observational [BE11]. Observations [AR12, BWT^{+}13, BSM17, Rey15, SL15b, WBZ19]. Observed [DK16, PDD^{+}18, SCPC14]. Obtained [BMW^{+}19]. Occupancy [GRSF^{+}16, MFLS18]. Occupying [DSB19]. Occurrence [LS13]. Occurring [PTH^{+}10]. Ocean [CHBY13, YC11, AES11, GSCR11, RKB15, SOS^{+}15, TS13]. Octopus [RS11]. Off [BvPCH17, DC13, MRC12, Mat10, WFWSV14, WBF^{+}15]. Off-Site [BvPCH17]. Off-Stream [MRC12]. Offs [SLR^{+}16]. Offshore [GMG16b, LOS^{+}11, SDKS16]. Offspring [THQ15]. Ohio [KSV^{+}10, PCP^{+}19, PDH^{+}19, RMTW16, SVPD13, SDF12]. Oil [RCMB18, SM14]. Okaloosa [RTJJ16]. Okanagan [AJ13]. Oklahoma [SSPP19, SL12]. On-Site [MM17, SM10, McC15]. Oncorhynchus [DBE11, HFB^{+}13, HHP^{+}12b, LPY16, TRB^{+}13]. One [APF18]. One- [APF18]. ongus [SNT^{+}16]. Onset [MSS^{+}12]. Ontario [BS11b, BBL^{+}11, CB10, FBM^{+}19, FDM^{+}13, HFCH15, LOS^{+}11, LBF17, SOSL10, TLL^{+}18]. Ontogenetic [SGCB19]. Ontogeny [SB18]. Open [APP13]. Open-Access [APP13]. Opening [DHA12]. Openings [CvPW15]. Operated [DT11, SPW^{+}12]. Operation [RJP^{+}15]. Operations [JKSK13, KRJT13, PPR11, RS18]. Opercular [PC12]. Opinions [DA12]. Opportunities [CKC^{+}12]. Opportunity [CBHS16]. Optimal [ASD^{+}18, MM18b, SLP13]. Optimization [SOB18]. Optimizing [DZ14, LP17]. Optimum [BDB^{+}19]. Options [FJB16, FWR^{+}15a]. Orconectes [RB15]. Order [Ano12b, JY11]. Oregon [HHB12, Bri17, CFW^{+}10, CCC16, Cle15, DASB16, HR11a, JM12a, JPWR10, JF14, MAM^{+}12, MP14, RSHS15, RM14, SHK^{+}17, ZLB10]. Oreille [BBD19, HHH14, MHDD16]. Orientation [Ano12b, CSF12, JY11]. Oriented [BHHA11, FH16, Han19, LRHR17]. Origin [FCE^{+}12, HHP^{+}12b, KPP^{+}18, MPM10, NCJ^{+}11, WPJ^{+}18, WBF^{+}15, ZEM11]. Origins [CFG16, SDF12]. Osmotic [LP12]. Other [DZ17, HJK11, LMC^{+}16, LRHR17, MSK19b, NBP15, WBT17, LDW10]. Otolith [BK12, CFG16, DPK^{+}17, DSD^{+}13, GPP^{+}10, HB14b, MPL^{+}15, OMT^{+}15, RCGF18, SL17b, SPL18, WPJ^{+}18]. Otolith-Chemistry-Based [WPJ^{+}18]. Otoliths [BC14, BKM19, CHW12, HM17, IWB10, KMPP12, SBS14, TP17, WS16, WI17, WGM^{+}11]. Ottawa [Hax15]. Otter [TSNH10]. Outcome [WP11]. Outcomes [FWG15]. Outdoor [EKM10]. Outperforms [CFBB15]. Output [DT11, DTS19]. Ova [BMM^{+}12]. Overfishing [MM18b, PS10, SHSCA16]. Overlap [FBB19, KIST16]. Overshoot [RS18]. Overwintering [PHFK10]. Owasco [CJMD10]. Owned [HWDA12]. Oxygen [KDA^{+}18, SM14, UML^{+}11]. Oxytetracycline [CHW12, RFLP14, SL17b, WMD^{+}19]. Oyster [KTN14, WHC12]. Ozark [JHB10, MMB19, QK11, RD10, SSP^{+}19, TLSB18].

Pacific [SOS^{+}15, APD^{+}19, APC^{+}14, BBC^{+}13, BWHM16, CGM13, CKC^{+}12, CAS16, DZ14, DZ17, DC13, DBE11, DCHC13, FOH^{+}15, GDLM17, HR11a, HB11, JM12a, JBHM19, KDP^{+}10, KLR^{+}12, LS17, LS11, MSW^{+}14, MCK^{+}15, NTJ^{+}14, PG14, RR17, RABK18, RBB^{+}10b, TS13]. Package [WTB^{+}15]. Pactola [SDDC16]. Paddlefish [BCMB14, CBS13, HPTH18, LSQ12, STP19]. Pads [LKV^{+}17]. Paired [AG15, RWT17, SS17]. Paired-Laser [RWT17]. Pallid [BFL^{+}10, FGS13, JTB16, SPT^{+}12, SPK10, SPP12, SWHA15, SH18]. Palmetto [DN11, PB13]. Panel [Ano12b, JY11, MAM^{+}12]. Panfish [FSW16, Kro19, LRHR17]. Panmixia [JBZ^{+}10]. Panther [HBS^{+}19]. Parameters [BDB^{+}19, DP13, NW12, WEC^{+}19]. Parastichopus [HDB^{+}12]. Park [FSS^{+}10, HMW^{+}10, PT15, SK18a, GKMC15, SZKK10]. Parr [GKT11, HDKB10, JNC^{+}16, TRB11, WS11a]. Partial [LOV17, RH14, RFQ14, WH11, WH12, WSA18]. Partial-Year [WH12]. Participation [CB10, RD10, SFH^{+}13]. Partitioning [JM15]. Party [MBPG12]. Pass [DP13, HKMH10, JKSK13, KZ12, PD13, SHL16]. Pass-By [KZ12]. Passage [APD^{+}19, APPR14, BG13, CBS^{+}12, CKC^{+}12, DG12, GLH^{+}18, HFRH17, KRJT13, KLR^{+}12, MAP14, MSK19a, NMG^{+}17, PSCL13, PN19, PQB10, SZHC15, STSH10, SBTW17, SEM14, WA14]. Passing [MS18]. Passive [AAYC14, BRB^{+}17, BR12, Bur11, BSTC13, CHJ^{+}16, FOH^{+}12, GM12, HHP12a, HUST17, KWMB11, KZ12, MM16, MHG^{+}11, NKC^{+}18, OL17, SNT^{+}16, SLSS16, SKBK12, WF17, WHG^{+}16, YHP10]. Passively [SMS13]. Past [ALCD17, CB10, GBTL14, JHS15b]. Pathologic [BDC^{+}13]. Patterns [ACRBB10, APAC12, BSVB19, BB12, BRSA15, Der14, HAW^{+}18, HH11b, HK17, MS18, MMS^{+}17, MSC13, MSS^{+}12, ND11, OHSP17, RBB10a, RCMB18, SSS15]. Peacock [NBFL17]. Peak [DT11]. Pectoral [HUB^{+}18, RHPH14]. Pedigree [HBS14a]. Pee [HH11a]. Pelagic [Han19]. Pelagic-Oriented [Han19]. Pelican [SRFW14]. Pelicans [GHK19, MSK^{+}16, TGS^{+}15]. Pellet [PTD10, TPL18]. Pellet-Reared [PTD10, TPL18]. Penaeid [RGC^{+}12]. Pend [BBD19, HHH14, MHDD16]. Pennsylvania [DKWG14, SBW^{+}17a]. Penobscot [SZHC15, ZOC^{+}14]. Pens [KSMN18]. Per-Recruit [DLP16]. Percentage [AS10]. Percentile [Ger10]. Perceptions [SFH^{+}13, TDM19]. Perch [BK12, DWB^{+}15, DRC17, FBB19, FRH10, FWR^{+}15b, FWR^{+}15a, GBFR18, HTH^{+}19, HH12, KWMB11, LSM^{+}16, LJR11, LSK^{+}14, MIH15, MSK19b, RBB10a, SB11, SPL18, TS13]. Performance [APPR14, AHS15, ASL^{+}16, BLSS19, BSG10, BTD^{+}12, CA14, CMK^{+}19, CFW^{+}10, CHPP11, DTS19, FHC16, GAWC10, JHS15b, KDP^{+}10, KKMS11, KM11, LAM16, MK13, NW12, PSCL13, PDL^{+}18, PPFC18, RPC^{+}19, RBB^{+}10b, RCP^{+}13, SC17, SLB10, TCG^{+}17, UMC14, WBP13, WALL17, YHP10]. Periods [APF18, RMTW16, WMB19]. Permanently [IH11]. Permanganate [GKMC15]. Permits [McC17]. Persistence [BBO16, HSC^{+}16, HGK15, MGS14, RFHP17, ZRR^{+}19]. Person [CFBB15]. Personal [DA12, MSHC15]. Perspectives [Bea17, BW17, Dag17]. Peshtigo [CSED10]. PHABSIM [BCD^{+}10]. Pharyngeal [LKV^{+}17]. Phase [TCB18, WA14]. Phenetic [WBB10]. Phenology [PM14]. Phenotypic [MKHC17]. Phocoena [BM14]. Photogrammetric [BGW^{+}11, RWT17]. Photonic [CMF^{+}14]. Physical [CC10, HPM^{+}18, KHBH19]. Physiologic [BDC^{+}13]. Physiological [BEH^{+}19, HBLL12, MGWC13, SBL^{+}18b, SCG11b, WWS^{+}14]. Physiology [HHP^{+}12b, RHC^{+}14]. Pickerel [BCZ^{+}18]. Pickering [PMP^{+}14]. Pictured [LSL^{+}13]. Piedmont [KCS^{+}17, KML14, MKBB13]. Piers [MCTM14]. Pike [BKM16, CBHS16, CSWW18, GF18, HKP^{+}14, LLP10, OLM15, ORL^{+}16, Pie10, RFLP14, SDDC16, SF14, SKBSA11, WQF15]. Pikeminnow [CCDF18]. Pilot [GWCA10]. Pink [OMT^{+}15]. Pipe [KZ12]. Piscicide [CST^{+}18, WMSH13]. Piscivory [PM13, SB18]. PIT [BF14, BRSB^{+}16, CSV13, DBW13, FAW14, HHRB15, KCS^{+}17, OHL10, PFWG16, RFKB17, RWPH11, SLSS16, SBL11, TPC^{+}15, WS13]. PIT-Tagged [OHL10, RFKB17, SBL11, FAW14]. Pitfall [SLSS16]. Pitman [SEP11]. Placement [DBKG14, GH17]. Plain [BA18, KML14]. Plains [DR19, MSK18]. Plan [LTG^{+}11]. Planning [APC^{+}14, BG19, RABK18]. Plant [RWCP11]. Plants [CMC^{+}19, KSV^{+}10]. Plasma [HPA^{+}11]. Plastic [SPP^{+}18, SNT^{+}16]. Plastic-Infusion-Process [SNT^{+}16]. Plate [MRC12]. Platform [AM17]. Platforms [RCMB18]. Platte [HHR^{+}18, RHPH14]. Plausible [CAS16]. Point [LKM^{+}17, TKH^{+}16]. Points [JHU^{+}13, PG14, ZKM13]. Pole [PP11]. Policies [FBM^{+}19]. Pollock [PSCS11, SC14]. Polyethylene [KZ12]. Polymorphic [SDF12]. Polymorphisms [SOS^{+}15]. Pond [RJP^{+}15, TPL18]. Pond-Reared [TPL18]. Ponds [BK16]. Pool [APD^{+}19]. Pool-and-Weir [APD^{+}19]. Poor [BC17, CAS16, WWM13]. Population [AK12, AD13, BGCS11, BWLB12, BCRM^{+}14a, BCRM^{+}14b, BGD^{+}19, BSO^{+}16, BRA^{+}18b, BRA18a, BRWQ18, BQR^{+}19, BBL^{+}11, BHG^{+}13, CWH^{+}14, CBC^{+}17, CPM^{+}10, CPSG12, DPK^{+}17, DFPS10, DGF11, ES14, FH16, FWG15, FJB16, FHL18, GD10, HIS16, HSC^{+}16, HHR^{+}18, HWNR14, HVW^{+}17, HB11, HHH14, HS12, HDM15, JKN^{+}10, JHB10, JDG10, JF14, JCB^{+}18, KML14, KQRS16, KMPV13, LAA^{+}18, LSMD15, LR12, LWJS16, MD10, MS18, MHDD16, MWS15, MCS^{+}17, MB17, Nak13, NFQ16, NWZ14, NEC^{+}13, ORD19, OHSP17, PLCG13, PDD^{+}18, PHS^{+}13, PNHA12, PRTS14, PDH^{+}19, PPO18, RZ10, SBW^{+}17a, SBDB16, SFH12, SCMP13, SRSS12, SRFS15, SBVB12, SPHK16, SZKK10, SPP12, SSP^{+}19, SDF^{+}13, SPWC15, TTS^{+}18, TCM19, TPF^{+}13, WDD^{+}14, WHB11, WBK^{+}10, WBK^{+}13, ZEM11, ZWH^{+}15]. Population-Level [TPF^{+}13]. Population-Specific [MS18]. Populations [AW10, ASL^{+}13, ASL^{+}16, BCZ17, BW18, BQR^{+}19, BSB^{+}15, Bri17, BSM^{+}15, CAWB17, DRG10, DASB16, DLC11, EWS^{+}17, EPM16, GBHO11, GBL15, HVW^{+}17, HUST17, HBA^{+}12, ITT10, KKM16, KFW12, KSM16, LSQ12, MS16, MW19, MGB^{+}14, NKC16, OEZ17, OHSP17, PNWC11, PRM15, PCP^{+}19, RFHP17, RD10, SBL^{+}18a, SMH17, SH11, SB15, SCWP17, VWB^{+}11, VEW^{+}13, WWB^{+}17, WKP14, ZRR^{+}19, Sin18]. Porpoise [BM14]. Portable [BDT^{+}16, HHRB15, HJK11, KCS^{+}17, OHL10, SBL^{+}12]. Portfolio [JDH16, McB14]. Portrait [MG10]. Possible [HBS14b, LSC^{+}16]. Post [GZ13, ITY13, TPF^{+}13]. Post-Release [ITY13]. Post-Rescue [TPF^{+}13]. Post-Spawned [GZ13]. Postdrift [KYD18]. Posthandling [CSV13]. Postrelease [Bal15, CFW^{+}10, GMG16b, KBH^{+}19, LFD^{+}17, MG14, RCP^{+}13, SKH^{+}16, SBHJ17, SS11b, TLL^{+}18]. Poststocking [PTD10, TPL18]. Potamodromous [BACRB14, FK16, SRFS15]. Potamopyrgus [SAF16]. Potassium [GKMC15, SCG11b]. Potential [APD16, BZ13, BRS^{+}12, BR14, BATK16, BKM19, CCDF18, CMC^{+}19, DIK19, FILH15, FCH^{+}13, GFC^{+}15, HMBT16, HW19, MAM^{+}12, MMHM17, RCP^{+}13, SHM^{+}10, TPF^{+}13, WBT17, ZWS^{+}19]. Potomac [LN12c, OLOI13]. Pound [CBDW15, SBD^{+}11]. Powell [BFS^{+}19]. Power [CGL18, DRG10, DT11, KSV^{+}10, Lee12, SWH^{+}16]. Practical [BBC^{+}13]. Practices [HM12, SOBW10, TEB^{+}19]. Prairie [DBE11]. Pre [MHAP13]. Pre-Spawn [MHAP13]. Precision [BCZ^{+}18, BM14, BGW^{+}11, BMW^{+}19, CWH^{+}14, DIK19, DRW^{+}18, KNC14, KNS18, OLM15, PC12, SPL18, SL12, SS13b, XM14]. Predation [AZE^{+}18, DWB^{+}15, DBE11, ERC^{+}11, EHP^{+}16, FOH^{+}12, FCH^{+}13, FHE^{+}12, GZ13, JNC^{+}16, KIST16, LFD^{+}17, LSK^{+}14, MSK^{+}16, SAYJ17, SRFW14, TRB11, TGS^{+}15]. Predator [LFD^{+}17, OOWW12, SHMS16, SEB^{+}19]. Predator-Naive [OOWW12]. Predators [OOWW12, WMSV18]. Predatory [BR14, HTH^{+}19]. Predict [BRA18a, GHK19, SCKZ17]. Predicted [PDD^{+}18]. Predicting [BC17, CBS13, CB10, HHCH15, PP11, SLMM10, SCG11a, ZRR^{+}19]. Prediction [ESL^{+}15, WHC14a, WHC14b]. Predictive [FTSF13, ND15, Pto13]. Predicts [KHBH19]. Preference [CSF12, LL12]. Preferences [DHH17, GSG18, HHSB13, LM16b, SDKS16, THI^{+}19]. Preferred [PK18]. Preliminary [AZE^{+}18, MCH12, PT11]. Prenant [WLL^{+}15, WALL17]. prenanti [WLL^{+}15, WALL17]. Prepositioned [FK16]. Prescribed [CJA^{+}13]. Prescribed-Fire [CJA^{+}13]. Presence [BDT^{+}16, BWB18, CBS^{+}12, MPL^{+}15, TPBF15]. Present [ALCD17]. Preservation [GF18]. Prespawn [DPS^{+}17, NKC^{+}18, WMMB19]. Pressure [FSW18]. Prestocking [KPL^{+}16]. Prevent [MM18b]. Previously [BHHA11]. Prey [DCL19, NSW16, RD10, SL15a, SSP18, WCWA10]. Prey-Fish [DCL19]. Priest [NFQ16]. Primary [BHHA11]. Prior [ZWS^{+}19, BSM17, McG14]. Prioritization [FDHWP16]. Private [TSFS19]. Privately [HWDA12]. Proactive [KACS18]. Probabilistic [MJCM15, McC15, PSCL13]. Probabilities [MGS14, PSB^{+}10b]. Probability [BE11, KDMH10, OHL10, RH16, SWHA15, WPS^{+}19]. Probing [PP11]. Procedures [Bal15, CHL13, HBS14a, KDA^{+}18]. Process [DPS^{+}17, SNT^{+}16]. Procypris [WALL17]. Produce [DM12]. Produced [BKHS14, GPP^{+}10, GWB^{+}10]. Production [BCD^{+}10, BKHS14, BG19, HFCH15, LSM^{+}16, NW12, RPBM10]. Productivity [AT18, AJ13, BRBP10]. Profile [ASS17, WAW^{+}19]. Profiles [HLM^{+}18]. Profitability [RR17]. Program [Cle15, GRSF^{+}16, HADG10, KPP^{+}18, LSL^{+}13, LS17, SPQ12, WHGC16]. Programmatic [ORC^{+}16]. Programs [BFLT17, BRA18a, RJC11, RÅW18, SKBSA11]. Progress [PSR16]. Project [BBG19, ORC^{+}16, MPL^{+}15]. Projects [KCRR19, SEP^{+}14]. Prolonged [WWS^{+}14]. Propelled [AG15]. Proportional [DN11]. Proportionally [PA16]. Proportions [WAW^{+}19]. Proposed [BBC10, BBS19, DN11, GFP^{+}12, HL15, NBFL17, BBC11]. Protect [RM12]. Protocol [Ano12b, JY11, SHL16]. Protocols [RBB^{+}10b]. Prototype [SE12]. Provide [ES17, WBG^{+}17]. Proximate [PHFK10]. Pseudotolithus [NEH10]. Psychological [LKvR^{+}18]. Public [LN12b]. Published [MS16]. Puerto [ASS17, NBFL17, PLCN17]. Puget [LPY16, LSLC15, PQB10, RFQ14, TWL^{+}17]. Pulse [DDD^{+}18, SVD^{+}18]. Pulsed [BBW18, CPSG12, DT11, DP13, DDD^{+}18, CC15, HRH10]. Pulsed-DC [CC15, HRH10]. Pumped [MHR^{+}18]. Pupfish [CFBB15, KPL12, RW11]. Purpose [Sta12]. Purse [DRW^{+}18]. Push [AG15]. Putative [FDM^{+}13]. Pyramid [SRFW14, SRFS15].

Quality [BC17, DMC^{+}19, LBL11]. Quantification [GHK17, LSK^{+}14]. Quantify [BMM^{+}12, DCL19, NCSR17]. Quantifying [ACMB^{+}14, ASP^{+}17, CSWW18, HJB14, HSM^{+}18, JBHM19, RSI16, SBL^{+}18a, TL15]. Quantile [Ran18]. Quantitative [HPTH18, Kae13]. Quantity [Bre13]. Quartzville [RM14]. Quasi [BC18]. Quasi-Stationary [BC18]. Quaternary [SAF16]. Quebec [LLP10]. Quota [LS17, SF16, WHGC16]. Quotas [DZ14, HHCH15, RR17].

R [WTB^{+}15]. rabaudi [WALL17]. Race [HJB14]. Raceway [MJH17]. Raceways [EKM10]. Radio [CMD12, DHLD11, DR10, FAW14, NKC^{+}18, RFKB17, TGA15, WPS^{+}19, YS12]. Radio- [DHLD11]. Radio-Tagged [CMD12, YS12]. Radio-Tagging [DR10]. Radiotelemetry [FCE^{+}12, FOH^{+}15, POWF11]. Rafael [BWT^{+}13]. Raft [FAW14]. Rafts [MK13]. Rainbow [AKP15, BJZ11, BJ19, CCMA^{+}10, DBW13, DBKG14, DZS10, EKM10, Erm12, FRL16, FSV10, FSV12, GBFR18, GZ13, GFC^{+}15, HW16, JM15, KKM16, KKMS11, KM11, KEC11, KACS18, LC14, LN12a, MP14, MPL^{+}15, MGWC13, MHG^{+}11, MSK^{+}16, MKHC17, ND11, OEZ17, PNWC11, QK11, SDDC16, SMP^{+}14, SKH^{+}16, SSPP19, THQ15, WMSV18]. Rainy [TS11]. Raised [EKM10]. Rakers [WSH^{+}14]. Ramp [PA16]. Rana [BKSH^{+}12]. Random [AHAC15, CSWP19, SSK19, See11]. Randomly [SDF12]. Range [HB10, KCCW19, PPO18, TTS^{+}18]. Rapid [BPT^{+}19, KHH13, MFLS18, MHAP13, WWP18]. Rapid-Assessment [MFLS18]. Rapidly [CBWB14, HGK15]. Rare [DCS11]. Rate [BHN11, BRA18a, BRWQ18, BMMMLML15, CZH^{+}19, GF18, GKT11, HDB^{+}12, MQS12, SKH^{+}16, SSP18, TRB^{+}13, WLK^{+}16]. Rates [AS10, BCBW18, Bea17, BDH^{+}14, BDSB15, CWB16, CSED10, CMWM18, Cle15, CBDW15, CAS16, CC15, DK16, Dag17, DBS19, GBL^{+}13, GM12, GMG16b, GMB10, IMM13, LMS^{+}18, LLP10, MS16, MM16, McC16, MEL^{+}12, MS14, MSK18, Pul17, RS17, RS18, RWPH11, SNT^{+}16, SLK15, SFH^{+}13, SM10, SEP^{+}19, UACZ^{+}16, VBT^{+}12, WH11, WH12]. Rathbun [WFS13]. Ratio [DHK^{+}19, JF13, SLMM10]. Ration [BR14]. Ratios [ISC10]. Ray [BDT^{+}16, HUB^{+}18, SH18]. Rays [RHPH14]. Razorback [HPA^{+}11]. Reach [LKV^{+}17, Pto13]. Reaches [DM12, HNPB15, MCK^{+}15]. Reactions [DR10]. Reader [DIK19]. Reared [AR12, CGDB14, DBKG14, DDEW18, HB14b, JPWR10, KSL^{+}19, MHAO11, MB12, NCJ^{+}11, PTD10, PHFK10, SKH^{+}16, SSP18, SPK10, SPP12, TPL18, VJK13, WBP13, WMSV18]. Rearing [BRBP10, DDEW18, FCH^{+}13, HLM^{+}18, HQ10, PK18, ZWS^{+}19]. Recall [APF18, OM10]. Recapture [BCBW18, BGP10, FOH^{+}12, HD14, HV18, HF13, HUB^{+}18, KDMH10, MWS15, RHMH15, SC17, VEIW^{+}16, WWB^{+}17, WEC^{+}19, ZEM11]. Recapture-Based [HV18]. Recaptures [WH11, WH12]. Received [HSM^{+}18]. Receiver [TMC10]. Receivers [BSH10, GH17]. Recently [PQB10]. Recognize [OOWW12]. Recolonization [BPT^{+}19, BKP^{+}14, WBB10]. Recommendations [DCL19, FSHR17, MSK18, STW^{+}12]. Reconditioned [HFB^{+}13]. Reconnaissance [ZW12]. Reconstruction [CA14, HBS14a]. Recordings [MJCM15]. Recoveries [FOH^{+}12]. Recovery [APC^{+}14, BG19, ERC^{+}11, FWG15, JTB16, KKM16, LAM16, MEL^{+}12, PNWC11, PDD^{+}18, WBG^{+}17, WFWSV14]. Recreational [APF18, APP13, AJ13, Ask16, BDH^{+}14, BSB^{+}15, CZH^{+}19, DBK11, DCP^{+}19, FF15, FC17, FMP19, FWR^{+}15b, GSG18, GMG16b, HWB11, HCL^{+}12, HSG16, HF13, HBD^{+}17, HHS^{+}13, LKvR^{+}18, LWJS16, LHS^{+}14, LSLC15, LRHR17, MG14, ML13, MBPG12, MMHM17, PLUH19, PA16, RPMB19, SCS17a, SFH12, SFH^{+}13, SBC^{+}11, See11, SCW13, SBHJ17, SPW^{+}12, SM10, TSFS19, VMA^{+}16, VPV18, WLK^{+}16, WQP13]. Recreationally [FJB16]. Recruit [CBS13, DLP16, SLR^{+}16, SEM14]. Recruitment [AKP15, BBO16, DWB^{+}15, EDN10, GSCCP18, HHT^{+}13, KSG^{+}18, RSHS15, RHG12, RHG13, SHSCA16, SKS^{+}15, SM14, ZKM13]. Red [BMMMLML15, DBJD12, GGC17, GGCF14, GPP^{+}10, JHS15a, PA16, RKB15, SC17, SCS17a, SL14, SM14, TSFS19, Vin14, WGS16, WFWSV14, WMD^{+}19, LAM16]. Redband [BSVB19, MSM^{+}14, MAG^{+}15, ND11, SLMM10]. Redbelly [PFWG16]. Redd [GCA^{+}13, HS12, KTHM13, LPY16, MSW^{+}14, MR16, MPL^{+}15, MPM10, PFRM12, VBBK16, WGJ10]. Redds [BPB13, FDM^{+}13, SLP17]. Redesigned [MC10]. Redfish [SDPB17]. Redtail [LXM^{+}12]. Reduce [CDQG12, NSW16, PA16, RGOR10, SOB18, SCW13]. Reduced [ORL^{+}16, Ryp15, SLSS16]. Reduces [KACS18]. Reducing [FWT14, MSK19b, RCP^{+}13, RJMG13, WHC14a, WHC14b]. Reduction [ASD^{+}18, FRH10, GFM^{+}18, LBL11, Nak13, NKL18, PA16]. Reductions [MRA14]. Reef [APD16, CGL18, HR11a, JHS15a, MKRH14, PSR16, Pul17, SDKS16, SFH^{+}13]. Reefs [APD16, WHC12]. Reference [APF18, BCZ17, JHU^{+}13, PG14, WP11, ZKM13]. Refined [BGBG17]. Refining [HWB11]. Reflex [SBL^{+}18b, WBG^{+}17]. Refuge [AHS15, JHS15b, MD10]. Refuges [APD16, KPL12]. Refugia [SD11]. Regimes [DLP16]. Region [JBRWH18, MSS^{+}12, PQB10, SHL16]. Regional [BW17, EWS^{+}17, GWCA10, HHS^{+}13, IMM13, KADH15, KTN14, MASN16, RÅW18, ZSW11]. Registration [CH17]. Regression [Ger10, Ran18]. Regulated [MI10, PFD17, Sam19]. Regulation [AJ13, JDG10, MMP12, PCP^{+}19, PDH^{+}19, SHSCA16]. Regulations [Ask16, BLM10, CSF12, Car13, DKWG14, DAKP13, EPM16, LWJS16, Pie10, RZ10, SLS16]. Rehabilitation [DTA18, MD10, MRB^{+}11, PSR16, WP11]. Reidentification [MSS^{+}12]. Reintroduction [BBD19, BSM17, BCMB14, KPP^{+}18, MAM^{+}12, RW11, ZWS^{+}19]. Reintroductions [APC^{+}14, DGS^{+}11]. Related [CvPW15, Ger11, HHSB13, RFW^{+}10, RFW^{+}11, RV12, SCW^{+}17, UWL13, WPY^{+}15, WFS13, WMMB19]. Relation [CS17, HFN11, TWL^{+}17]. Relationship [BHN11, BK12, DA12, FOH^{+}15, HFH12, HFH18, KTN14, LLP10, LM16b]. Relationships [GKT11, HDB^{+}12, ITT10, JBRWH18, MKK^{+}12]. Relative [BW18, BS10, DRW^{+}18, DLP16, FQ19, KM11, NKL18, RS11, RPC^{+}19]. Release [BGHL10, BT16, CvPW15, CI10, CFWC11, CCC16, DKWG14, DLSW12, ES17, EBS18, FHC16, FSW18, GBL^{+}13, GDP14, GKT11, HR11b, ITY13, IMM13, LM16a, LFD^{+}17, MMS^{+}17, McC16, MH11a, PB13, RZ10, RSQ^{+}19, Sam19, SVFG10, SMP^{+}14, SBL^{+}18b, TCG^{+}17, TLC^{+}19, TRB^{+}13, TEB^{+}19]. Released [CFW^{+}10, GMB10, KSMN18, LM16a, MGWC13, NCJ^{+}11, OOWW12, RHC^{+}14, RSQH18]. Releases [BSG10, JCB^{+}18]. Releasing [CL12, SMK13]. Relevant [CI10]. Reliability [RJC11]. Reliably [SHL16]. Relief [ES17, HR11a, SCKZ17, SMW16]. Relying [MB12]. Remediated [PN19]. Remnant [BZ13, BHG^{+}13]. Remote [BWB18, HAS^{+}19]. Remotely [SPW^{+}12]. Removal [AHVJ11, BPT^{+}19, CCMA^{+}10, CA11, DVWS10, DKSW15, HB14a, HHB12, MPL^{+}15, MH11b, PDD^{+}18, RHMH15, SFW11, SRSS12, SH18]. Removals [DHB^{+}10, RH14, TCB18]. Repeat [CHDW18]. Repeat-Spawning [CHDW18]. Repeated [RJP^{+}15]. Repellent [BWHM16]. Report [KFW12]. Reported [BGCH17, KCCW19]. Reporting [MQS13b, MEL^{+}12, MS14, VBT^{+}12]. Representative [WAW^{+}19]. Reproducibility [RWCP11]. Reproduction [AJR^{+}19, BSS^{+}17, RTMJ10, SRFS15, SL14]. Reproductive [CBHB16, DZS10, DWCK16, HPA^{+}11, TS13, VJK13, WMMB19]. Require [Rey15, SL15b]. Requirements [MSW^{+}14]. Resampling [MSW^{+}14]. Rescue [TPF^{+}13]. Research [HO13, HK17, MI10, TSNH10, WLK^{+}16].

Reserves [CVD^{+}14]. Reservoir [BPB13, BS10, CSH^{+}17, DSB19, DBK11, DDAS14, FSHR17, GBFR18, HTH^{+}19, MMP12, MCP15, NTJ15, NKC16, PJLB14, PBA^{+}17, RCGF18, SSK19, SDDC16, STP19, SPL18, SDF12, SK18b, THZ^{+}17, WBT17, WCWA10, BBO16, BCMB14]. Reservoir-River [DSB19]. Reservoirs [BOM16, CFG16, CF18, DM12, DBK11, FBB19, GWB^{+}10, HPTH18, HHS^{+}13, KML14, KNC14, KM11, LN12b, MKBB13, MSK18, NP12, NKC16, PLCN17, PCP^{+}19, SL15a, SL12]. Residence [SLB10]. Resident [BKP^{+}14, HHP^{+}12b, WHC12]. Residential [LGC11]. Residualism [TLC^{+}19]. Residualization [HM12]. Resistances [MK13]. Resistivity [SB15]. Resolution [KIST16]. Resolving [ZEM11]. Resource [CSF12, CGL18, DCS^{+}16, EWS^{+}17, JM15]. Resource-Conserving [CGL18]. Resources [VEIW^{+}16]. Respond [OOWW12]. Response [ADMC10, ABP^{+}18, AJ13, BCRM^{+}14b, BJZ11, BRSA15, BHG^{+}13, Dag17, DFMP18, DHB^{+}10, FSV12, FMA^{+}12, FBB18, GGCF14, HGW10, HMW^{+}10, HBS14b, KML14, Kra18, ORC^{+}16, PPR11, RFW^{+}11, RTJJ16, RV12, RJP^{+}15, SBL^{+}12, SL15b, VPV18, WLK^{+}16, WBK^{+}13]. Responses [BPT^{+}19, CEH^{+}15, CMC^{+}19, HIS16, JTB16, KMB^{+}12, KPP^{+}18, MSC13, MGWC13, MKWM18, PDD^{+}18, PCP^{+}19, SWG^{+}16, WWS^{+}14, WHB11]. Responsible [DCP^{+}19]. Restoration [BZ13, BBG19, BRA18a, CBWB14, CGG^{+}18, GAH^{+}18, LSL^{+}13, MODC10, ORC^{+}16, RTJJ16, RPBM10, RÅW18, RABK18, WWP18]. Restore [CCMA^{+}10]. Restored [RGH^{+}13, WHC12]. Restoring [HFRH17]. Restrictions [LS10, Sam16, SL17a]. Restrictive [DAKP13, VPV18]. Restructure [MCS^{+}17]. Result [HT11]. Resulting [CPM^{+}10, DHA12]. Results [CC15, GWCA10, HSC^{+}16, RHMH15]. Resurrection [SJM15]. Retardation [SL14]. Retention [BSL18, BF14, BA11, BRSB^{+}16, BR12, CSV13, CF13, DBKG14, FBL^{+}19, GM12, HHP12a, Han19, HW10, HQ10, HBLL12, HUST17, KWMB11, Kro19, LXM^{+}12, MM16, MLHS13, MHG^{+}11, NLK17, NBP15, PFWG16, RWPH11, RH16, SNT^{+}16, SH11, TPC^{+}15, WRNJ^{+}19, WF17]. Retrospective [Der14]. Return [McC16, SL15a, SS16]. Returns [HLM^{+}18, KSMN18, QA16]. Reveal [LR12, PCP^{+}19]. Reveals [CFG16, JDG10, VWB^{+}11, WAW^{+}19, WWP18]. Reversion [ZGW11]. Review [ACRBB10, JBRWH18, RÅW18, RABK18]. Reviewer [Ano18h, DDD^{+}19]. Revisit [DRG10]. Reward [MEL^{+}12]. Reward-Recovery [MEL^{+}12]. Rewards [SC17]. RFID [FAW14]. Rho [Der14]. Rich [DCL19]. Rico [ASS17, NBFL17, PLCN17]. Right [HBS14b]. Ring [GCLC18, RH16, SBS14, SCW^{+}17]. Rio [AR12, GHL12, ZRR^{+}19]. Riparian [CBM13, DFMP18, RIMR19]. Risk [CJ14, CAS16, HW12, HH12, HL15, PCT^{+}14, TRB11]. Risk-Based [PCT^{+}14]. Rita [PKKR10]. River [ACRBB10, ASP^{+}17, AT18, ALCD17, AD13, AHAC15, AM17, BFP12, BRS^{+}12, BWLB12, BCRM^{+}14a, BCRM^{+}14b, BBD19, BSS^{+}17, BSS^{+}14, BSH10, BGD^{+}19, BSVB19, BWT^{+}13, BGW^{+}17, BQR^{+}19, BSG10, BSBF13, BBS18, CSED10, CMMR19, CCC16, Cle15, CDBW17, CJH16, DHLD11, DSB19, DB13, DRW^{+}18, DFPS10, DCS^{+}16, DMC^{+}19, DR19, EJL10, FBD^{+}19, FRL16, FWR^{+}15b, GD10, GRSF^{+}16, GAH^{+}18, GKS13, GHK19, Gur12, HDKB10, HP14, HHR^{+}18, HPD^{+}19, HH11a, HH11b, HLM^{+}18, HNPB15, Hax15, HGW16, HPM^{+}18, HH15, HHB12, ITY13, JM12a, JPWR10, JKSK13, JNC^{+}16, KFM12, KCF15, KRJT13, KTHM13, KDMH10, KEC11, KACS18, KD10, Kra18, LAA^{+}18, LSQ12, MAM^{+}12, MNYV12, MCK^{+}15, MPL^{+}15, MRA14, MSM^{+}14, MCSF^{+}17, NCJ^{+}11, ND11, NEH10, NEC^{+}13, OHSP17, POWF11, PKKR10, PSB^{+}10b, PFD17, PTH^{+}10, PRTS14, PDL^{+}18]. River [PP11, QK11, RCGF18, RSHS15, RS18, RÅW18, RHPH14, Sam16, SBW^{+}17a, SPT^{+}12, SB15, SZHC15, SMK13, SZKK10, SEP11, SD11, SDL^{+}15, SCB^{+}19, SPWC15, SCWP17, SBB19, TBH^{+}14, TGS^{+}15, VWB^{+}11, WKS^{+}10, WPS^{+}19, WBZ19, WBH13, WP11, WHKB17, WTB^{+}15, WBK^{+}10, WBK^{+}13, WBB10, ZLB10, ZOC^{+}14, APAC12, ABBG13, AKP15, BRN16, BQHR17, BFL^{+}10, BFS^{+}19, BPT^{+}19, Bur11, BHR^{+}15, CFG16, CJ14, CFW^{+}10, CBDW15, CAWB17, CHPP11, DPK^{+}17, DZS10, DG12, EMK^{+}17, EHP^{+}16, FHC16, FBC^{+}10, FCE^{+}12, GDP14, GMG^{+}15, GWB^{+}10, GHK17, GKT11, HD14, HFB^{+}13, HWFN14, HB10, HHRB15, HFRH17, HGK15, JKN^{+}10, JBZ^{+}10, JF13, JF14, KJN^{+}17, KSM^{+}18, KSV^{+}10, KB11, LKV^{+}17, LKM^{+}17, LN12c, LO18, MMP12, MH11a, MB12, MSK19a, NTJ^{+}14, OLOI13, PTH^{+}10, PVH^{+}13, PHFK10, PDH^{+}19, RH14, RWAJ16, SVPD13, Sin18]. River [SEP^{+}14, SWH^{+}16, SM19, SHMJ19a, SHMJ19b, SPK10, SPP12, SEM14, TPF^{+}13, TFH^{+}16, TEB^{+}19, WA16, WBH13, WCWA10]. Riverine [Ano12b, EPM16, JY11, MLEK15, Sam19, SBL^{+}18a]. Rivers [CKBP19, ESL^{+}15, FAW14, KSMN18, MI10, MASN16, NQ11, ORA17, PSB^{+}10a, SP13, SS13b, TMC10, WS11a, WSH^{+}14, KYD18, LDP^{+}10]. Riverwide [KSM16]. RNAlater [FRH10]. Road [PQB10]. Roads [MRA14]. Roanoke [HH11b, HH15]. Robust [VEIW^{+}16]. Rock [ABVS14, MC10, RD10, Sta12, WALL17]. Rockfish [HR11a, HR11b, TS13]. Rocks [LSL^{+}13]. Rocky [HR11a, RFHP17, MH11b, PN19, SNTZ14]. Roe [CBS13]. Roof [MHBR19]. Roof-Hooking [MHBR19]. Rotary [BKD^{+}10, MHC10, TRB^{+}13]. Rotenone [BKSH^{+}12, BJZ11, Erm12, FSS^{+}10, FSV10, FSV12, FBB18, HMW^{+}10, HSF18, RB15]. Round [NCMT12, PCP^{+}13]. Roundtail [UMC14]. Route [BSBF13, KJN^{+}17, MAP14, MSK19a, PSB^{+}10b]. Roving [MQS13b, SPW^{+}12]. Royale [SQGK10]. Rudd [KFW12]. Rules [WWM13, ZC12]. Run [BZ13, BWLB12, BCRM^{+}14a, BCRM^{+}14b, BBS18, CA14, HD14, HWFN14, MS18, PFD17, PRTS14, SM19, TPF^{+}13]. Runs [Bou10].

S [STP19, WMD^{+}19, CGM13, FGS13]. Sablefish [WHGC16]. Sacramento [BGD^{+}19, HPM^{+}18, NB10, PSB^{+}10b, TPF^{+}13]. Safe [HHCH15]. Saginaw [BSB^{+}15, FB14, FJB16]. Saint [ALCD17, LAA^{+}18, SDL^{+}15]. Sales [HBD^{+}17]. Salish [AHVJ11]. Salmon [APC^{+}14, AT18, AZE^{+}18, BMM^{+}12, BCRM^{+}14a, BCRM^{+}14b, BCD^{+}10, BS11b, BG12, BG19, BRBP10, BSM17, BGP10, BHC^{+}10, BC17, BSG10, BSBF13, BBS18, BKP^{+}14, CBWB14, CBS^{+}12, CA14, CJ14, CWH^{+}14, CSH^{+}17, CBC^{+}16, CBC^{+}17, CCC16, CC10, CHBY13, CHJ^{+}16, CMD12, CHPP11, DMRJ10, DBE11, FHC16, FBL^{+}19, FDM^{+}13, FBC^{+}10, FCE^{+}12, FOH^{+}15, GLCD13, GLH^{+}18, GBHO11, GBL15, GDLM17, GL16, GKT11, GCA^{+}13, HTF^{+}13, HD14, HLM^{+}18, HJB14, HPM^{+}18, HWFN14, HQ10, HLS^{+}17, HRH10, HB11, JM12b, JKN^{+}10, JDG10, JF13, JF14, JNC^{+}16, JBHM19, JCB^{+}18, KHBH19, KSMN18, KJN^{+}17, KAZ15, KRB^{+}16, KTHM13, KLR^{+}12, KPP^{+}18, KSL^{+}19, MMP12, MS18, MNYV12, MFLS18, MQS12, MQS13a, MQS13b, McC17, MSD11, MSS^{+}12, MGB^{+}14, MSHC15, MCTM14, MEJ10, MPM10, MHC10, NKC^{+}18, NB10, NMG^{+}17, OMT^{+}15, POWF11, PRM15]. Salmon [PSB^{+}10b, PFD17, PG14, PHFK10, PRTS14, PPFC18, RGH^{+}13, RFQ14, RM14, RPBM10, RABK18, RCP^{+}13, SHMS16, SMW16, SZHC15, STSH10, SEP^{+}14, SOS^{+}15, SM19, SLG^{+}15, SCA11, SLP17, TFP^{+}11, TBH^{+}14, TPC^{+}15, VWB^{+}11, VEW^{+}13, VPV18, VJK13, WS11a, WWS^{+}14, WS16, WHMS14, YC11, ZWS^{+}19, ZOC^{+}14, JM15, JNC^{+}16, VWB^{+}11]. Salmonid [BSM^{+}15, GWCA10, HTH^{+}19, HW10, HEL^{+}15, JBRWH18, ORC^{+}16, PRR^{+}13, SFW11, SHMJ19b, VEIW^{+}16]. Salmonid-Dominated [HTH^{+}19]. Salmonids [ACMB^{+}14, BGHL10, BSM17, BBC^{+}13, CSWW18, ERC^{+}11, FOH^{+}12, HTPS10, HJK11, JKSK13, JNC^{+}16, MAM^{+}12, MC10, MSK19a, PK18, RSQH18, SLB10, SBL11, SD11, WBT17, YS12]. Saltwater [BW17, CL12, CHDW18, LS10, LL12, SL17a]. Sample [AG15, BvPCH17, BKME17, KKSH12, KQDC19, KNC14, MM17, VKBL11]. Sampled [SBW^{+}17b, YS12]. Samples [ASP^{+}17, PP11, WMB^{+}15]. Sampling [ABP^{+}18, BFN16, BDSB15, BSBC14, BGW^{+}17, BFL^{+}10, BE11, CDBW17, CKBP19, DCS11, DWW12, DP13, DRW^{+}18, EJL10, ETD16, FK16, FSW16, GRSF^{+}16, HKMH10, HTPS10, HBS14a, JJ10, KQDC19, KNC14, KSM16, LCC^{+}15, MQS13a, MJCM15, MSK18, NQ11, NKC16, PBA^{+}17, RTMJ10, RPC^{+}19, RJC11, SSK19, SBDB16, SEP^{+}19, UBR^{+}17, UWL13, WKP14, WCY^{+}18, WBGM13, XM14]. Sampling- [UWL13]. San [BGD^{+}19, NB10, PSB^{+}10b, BWT^{+}13, BSBF13, BBS18, CMMR19, ERC^{+}11, RGOR10]. Sand [AD13, AHAC15, DHB^{+}10, SCB^{+}19, WBK^{+}10, WBK^{+}13]. Sand-Bed [AD13, AHAC15, SCB^{+}19, WBK^{+}10, WBK^{+}13]. Sand-Laden [DHB^{+}10]. sapidissima [WBZ19]. sapidus [AOS^{+}17, ASD^{+}18]. Saprolegnia [SMP^{+}14]. Sardine [DZ14, DZ17]. Satisfaction [BW17, HMD16, MP14, TBIJ19]. Saturation [LJR11]. Sauger [EBS18, LO18, PDH^{+}19, WD10]. Saugers [GWB^{+}10, Hax15, KB11, MH11a, SVFG10]. Saugeye [KSG^{+}18, KNS18, SDF12]. SCA [FB14]. Scale [AW10, Aka10, BSVB19, CZH^{+}19, CVD^{+}14, HCL^{+}12, HMP10, KTN14, MAP14, MSC13, PLCG13, PT11, Pto13, RHG12, SCS17b]. Scale- [HCL^{+}12]. Scales [BKM19, CHDW18, ETD16, IWB10, KMPP12, SQP12, TO10, TP17, WS16].

Scallop [SMW12, SBG^{+}19, WHC14b]. Scan [GHK17, RSI16, VKH^{+}19, WA16]. Scaphirhynchus [CGG^{+}18, PTH^{+}10]. Scenarios [BGD^{+}19, BRA18a, LSC^{+}16, SGC13]. Schedules [TO10]. Schizothoracin [WLL^{+}15, WALL17]. Schizothorax [WLL^{+}15, WALL17]. School [MG14]. School-Size [MG14]. Science [CH17]. Scientific [BM14]. Scotia [GBHO11]. Scotty [PPFC18]. Screen [BKD^{+}10, GRR^{+}17, MRC12, MJH17, RM12, ZS16]. Screening [HL15]. Screens [BP14, MRC12]. Screw [MHC10, TRB^{+}13]. Sculpin [CC10, RHMH15]. Sculpins [TWL^{+}17]. Scylla [RJMG13]. Sea [BDH^{+}14, BT16, BWHM16, CS17, HDB^{+}12, ILBJ12, JSDJ16, Lee12, MJH17, RWAJ16, SHM^{+}10, SMW12, SVD^{+}18, SBL^{+}12, AHVJ11, SGF^{+}13, Sta12, SC14]. Season [Bal15, BPB13, CJ14, CMC^{+}19, MSK18, PA16, Sam19]. Season-Specific [MSK18]. Seasonal [BA18, BATK16, HH17, MCK^{+}15, RBB10a, SSS15]. Seasonality [BRN16]. Seasons [RMTW16, TSFS19]. Seatrout [DLP16, WBP13]. Seawalls [MCTM14]. Seawater [GLCD13]. Secondary [BRN16, WHC12]. Sectioned [WI17]. Sectioning [WKTQ15]. Sedate [FVH^{+}17]. Sedation [FMA^{+}12, JTB16, KAZ15]. Sedative [FGS13]. Sediment [HWZ19, SM14, ZW12]. Sedimentation [PPFC18, WHC12]. Segregation [Bri17, Hax15]. Seine [DRW^{+}18, TSNH10]. Seines [TSNH10]. Seining [Mea12, NQ11]. Seismic [CDQG12]. Select [MSK^{+}16]. Selection [CJMD10, CHPP11, DTA18, GDP14, HH10, HH11c, HCL^{+}12, HLS^{+}17, MHAO11, ND15, Ols19, SL15a, SZHC15, SSP18, WWW15, WP11, WS11b]. Selective [CBHS16, CHBY13, DHB^{+}10, YC11]. Selectivity [BCBW18, EF10, GCLC18, HB12, KQDC19, SBDB16, SR16, SBW^{+}17b, VKBL11, ZM17]. Self [AJ13, BGCH17, JDG10, MQS13b, RJC11]. Self-Regulation [AJ13]. Self-Reported [BGCH17]. Self-Reporting [MQS13b]. Self-Sampling [RJC11]. Semi [WA14]. Semi-Automated [WA14]. Seminole [PTD10]. Sensitivity [DBK11, MM18a, SC14]. Sensors [IH11]. Separation [DZS10]. Septicemia [PCT^{+}14, TPBF15]. Sequences [AES11]. serrata [RJMG13]. Serve [APD16]. Service [ESL^{+}15]. Sets [SS11a]. Setting [HHCH15, WP11]. Settlement [LCC^{+}15, McG14]. Seven [RBB^{+}10b]. Several [Erm12, FSV10, FSV12]. Severity [MHBR19, PD13, SCG11a]. Sex [CBHB16, DBS19, DHK^{+}19, DRC17, FG13, HPA^{+}11, ITT10, ISC10, Ise10, JF13, KHH13, LS11, MBPG12, MMHM17, SLMM10]. Sex-Balanced [MMHM17]. Sex-Specific [DBS19, ITT10]. Sexual [MAH^{+}13, SLMM10]. Shad [ABBG13, BFP12, BZ13, CA11, DSB19, FSHR17, HH11a, HH11b, HH11c, MHR^{+}18, NCSR17, NP12, NKL18, TO10, WHB^{+}14, WSH^{+}14, WBZ19]. Shadow [JHB10]. Shallow [BSBC14, BMMMLML15, CWB16, CGG^{+}18, DCL19, RJP^{+}15, SKBSA11, TBH^{+}14, TKH^{+}16]. Shallow-Water [BSBC14, CGG^{+}18]. Shape [CF13]. Share [BWHM16]. Shark [CAS16]. Sharks [RGC^{+}12]. Shasta [CSH^{+}17, HHB12]. Shedding [LMB^{+}18, VBT^{+}12]. Sheepshead [RCMB18]. Shelf [BGCH17]. Shifting [LMD^{+}19]. Shifts [HLS^{+}17, SGCB19, WMB19]. Shiner [BPQ13, SL14]. Shoal [BBS19, GKS13, ITY13, Sam16, TTS^{+}18]. Shocking [CC15]. Shore [SPW^{+}12]. Shore-Based [SPW^{+}12]. Shoreline [BHR^{+}15, DV17, SCKZ17]. Shoreline-Spawning [BHR^{+}15]. Short [BEK^{+}18, Bur11, CvPW15, FWWTJD16, GM12, HUST17, LBL11, MQS13a, PPR11, RLC12, SS11a, SBL^{+}18b, VMC^{+}11, YHP10]. Short- [YHP10]. Short-Duration [MQS13a]. Short-Term [BEK^{+}18, Bur11, FWWTJD16, GM12, HUST17, LBL11, PPR11, RLC12, SS11a, SBL^{+}18b, VMC^{+}11]. Shortnose [BFP12, CHL13, CHW12, EHP^{+}16, SL17b, SBL^{+}18b, UWL13]. Should [HHP^{+}12b, Hen15, HBS14b]. Shovelnose [HHP12a, HHR^{+}18, KCF15, PVH^{+}13, RHPH14, SPT^{+}12, TGKI10]. Show [GMB10, OHSP17]. Shows [LMC^{+}16]. Shrimp [GGC17, HB12, Mat10, MCH12, RGC^{+}12]. Siberian [BDC^{+}13]. Side [GHK17, RSI16, SD11, VKH^{+}19, WA16]. Side-Channel [SD11]. Side-Scan [GHK17, RSI16, VKH^{+}19, WA16]. Signal [HSM^{+}18]. Signatures [GPP^{+}10, HTPS10]. Significantly [See11]. Siltation [CF13]. Silver [BGBG17, CSWP19, CDBW17, CVD^{+}14, HPD^{+}19, LDP^{+}10, LKV^{+}17, RSLG10, RJP^{+}15, SP13, SPWC15, SCWP17, WSH^{+}14]. Silvery [AR12, GHL12, NBP15]. Similarities [HEL^{+}15]. Simple [DSJS16, FDHWP16, RBB10a]. Simulated [CBS^{+}12, GLH^{+}18, GRCR10, KB10, KDA^{+}18, LMS^{+}18, LBL11, RSQ^{+}19, SMH17, SFH12, WHB11]. Simulating [PNWC11]. Simulation [CJ14, DMC^{+}19, FHH16, MR16, RHMH15, SLS16, WHC14a, AHS15]. Simulation-Based [CJ14]. Simultaneous [YC11]. Simultaneously [Sin18]. Single [Ger10, SHL16, SB15, SOS^{+}15]. Single-Channel [SB15]. Single-Unit [SHL16]. Site [BvPCH17, CSF12, ETD16, HR11a, HWNR14, JHK11, KWMB11, KB10, MM17, ND15, SM10, WF17, WS11b, McC15]. Site-Specific [HWNR14]. Sites [DHH17, HH10, HH11a, HW19, RMTW16]. Situ [RWT17, TGKZ19]. Six [HW10, MG10, RCGF18]. Six-Decade [MG10]. Size [ASD^{+}18, Ask16, BCBW18, BvPCH17, BCZ17, BSS18, Bou10, BJZ11, BMMMLML15, CSWW18, CF13, DHK^{+}19, DN11, EF10, FH16, FHC16, GSCR11, GZ13, GKT11, HD14, HDB^{+}12, Han19, HB12, HS12, JWC10, JHS15a, KQDC19, LXM^{+}12, MHAO11, MG14, MM17, MWS15, MM18b, OEZ17, ORL^{+}16, PW13, PPO18, RSLG10, RLC12, RH16, Ryp15, SOB18, Sam19, SBDB16, SKH^{+}16, SR16, SZHC15, SCW^{+}17, SPP12, SDF^{+}13, TWL^{+}17, TLC^{+}19, THZ^{+}17, WF17, ZM17]. Size-Bias [SR16]. Size-Dependent [GZ13, Han19, LXM^{+}12]. Size-Graded [Ask16]. Size-Sorting [HB12]. Size-Specific [BSS18]. Sized [KDMH10, LMS^{+}18, MSK^{+}16]. Sizes [BSTC13, KWMB11, KNC14, PFWG16]. Skate [DC13, HCH13]. Skeena [BWLB12, BCRM^{+}14a, BCRM^{+}14b, PRTS14]. Skews [DHK^{+}19]. Skipper [SF16]. Sleeper [CAS16]. Slimy [CC10]. Slot [WBZ19]. Slow [ZWH^{+}15]. Slow-Growing [ZWH^{+}15]. Sluiceway [JKSK13, KRJT13]. Small [Aka10, AD13, BS12, BACRB14, BGW^{+}11, FBL^{+}19, HWDA12, HHRB15, HKP^{+}14, KCS^{+}17, LN12a, LDW10, LCR19, MC10, MOJ12, NKL18, OHL10, OL17, RGC^{+}12, RS17, RPC^{+}19, SCS17a, SFW11, SNTZ14, SBL11, SE12, SEP^{+}19, TO10, WS11b, WBK^{+}10, WBK^{+}13, WWD13]. Small-Bodied [AD13, KCS^{+}17, RPC^{+}19, WBK^{+}10, WBK^{+}13, BGW^{+}11]. Small-Scale [Aka10]. Smallmouth [BKME17, BKM19, BE11, Bre13, BEH^{+}19, DWB^{+}15, HGW10, HB14b, KB10, MMB19, MB17, SBW^{+}17a, SBL^{+}18a, SHK^{+}17, SSP^{+}19, TLSB18, WMB19, WCWA10]. Smartphone [McC17]. Smelt [CMF^{+}14, CSH^{+}18, FRL16, GZ13, LC14, LSLC15, OEZ17, WHG^{+}16]. Smoky [GKMC15]. Smolt [AZE^{+}18, BCD^{+}10, GKT11, HLM^{+}18, MHC10, RPBM10]. Smolt-to-Adult [HLM^{+}18]. Smolting [SLG^{+}15, ZOC^{+}14]. Smolts [FHC16, GKT11, JPWR10, NMG^{+}17, STSH10]. Snake [KEC11, KACS18, MSM^{+}14, RS18, BSG10, CFW^{+}10, CAWB17, GKT11, HGK15, KD10, MCK^{+}15, MSK19a, NTJ^{+}14, PHFK10]. Snakehead [hGgXfC^{+}13, LN12c, OLOI13]. Snapper [FWWTJD16, GGC17, GGCF14, GPP^{+}10, GSCR11, JHS15a, PA16, RKB15, SC17, SCS17a, SM14, TSFS19, Vin14, WGS16]. Snook [BA11]. Snorkel [SCPC14, WKP14]. Snorkeling [HDKB10, KDMH10, PLUH19, UBR^{+}17]. Snorkeling-Based [PLUH19]. Snow [SGF^{+}13]. SNPs [SOS^{+}15]. Soak [CS17, LJR11]. Social [ASS17, CH17, HSG16, LTG^{+}11]. Society [PBA^{+}17, UBR^{+}17]. Sociodemographics [SDKS16]. Sockeye [BCRM^{+}14a, BCRM^{+}14b, BRBP10, CHPP11, GKT11, HTF^{+}13, KHBH19, MMP12, PHFK10, PRTS14, SCA11]. Soft [MLHS13, SPP^{+}18]. Sole [DDD^{+}18, Sta12]. Solea [DDD^{+}18]. Solution [FRH10]. Some [Bea17, Dag17, SC14, ZEM11]. Something [Can17]. Sonar [BOM16, GHK17, HHMQ15, MLEK15, PNHA12, RSI16, VKH^{+}19, WA16]. Sonars [HH15]. Sonic [CHL13]. Sorting [DHK^{+}19, HB12]. Sound [ABP^{+}18, LPY16, LSLC15, MJH17, WCY^{+}18, PQB10, RFQ14, TWL^{+}17]. South [BKME17, JWC10, SDDC16, ASP^{+}17, CWB16, DWB^{+}15, DBJD12, DCS^{+}16, GRCR10, HH11a, HB14a, LPY16, MNYV12, SGCB19, WHC12]. Southeast [CL12, DWCK16, MS17a, MBM13, RIMR19]. Southeastern [FBB19, GMB10, ADD14, FMP19, ORD19, RKB15]. Southern [AK12, FTSF13, PFWG16, SCS17b, SS11b, BMMMLML15, DK16, FWG15, GBHO11, HADG10, KKM16, PSR16, RFHP17, TDM19]. Southwest [CJA^{+}13]. Southwestern [BKSH^{+}12]. Space [CWH^{+}14, MQJK19]. Spanish [SPM15]. Spatial [AOS^{+}17, BFL^{+}10, Bri17, Hax15, HB14a, KTHM13, KSL^{+}19, MSW^{+}14, MSBP10, PDH^{+}19, RHG12, SC17, SSK19, SCS17b, SQP12, TO10, WDD^{+}14, WSH^{+}14, WHC14a, WHC14b, WBF13]. Spatially [CBC^{+}16, RPJ18, WQP13]. Spatiotemporal [DDAS14, DGF11, FTSF13, HK17, NP12]. Spawn [CHDW18, LPY16, MHAP13]. Spawned [GZ13]. Spawner [GAWC10, MNYV12]. Spawners [SLR^{+}16, TFH^{+}16]. Spawning [AOS^{+}17, AKP15, BZ13, Bal15, BPB13, BACRB14, BC17, BWR12, BHR^{+}15, CJMD10, CFBB15, CHDW18, CMD12, CF13, DB13, DHA12, FDM^{+}13, GHL12, GHK17, GCA^{+}13, HH10, HH11a, HH11c, JM12a, KHBH19, KBH^{+}19, LKM^{+}17, LBF17, MRB^{+}11, MKRH14, MPM10, ND15, PK18, PRRD15, RSI16, STP19, SRSS12, SM19, SSK18, TPF^{+}13, VJK13, VKH^{+}19, WA16, WBH13, WGJ10, ZDC14]. Spearers [SF14]. Spearing [MSE^{+}18]. Specialization [HM18]. Specialized [BHHA11, HBS^{+}19]. Species [BPQ13, BFN16, CAS16, DCP^{+}19, FMP19, GPLT12, GTB^{+}19, HB15, HR11a, HMP10, HL15, HEL^{+}15, LMC^{+}16, LSC^{+}16, LM16b, McB14, MJCM15, MOJ12, PLCG13, RWAJ16, RPC^{+}19, SPQ12, SQP12, SBW^{+}17b, SLP17, UMC14, ZWS^{+}19]. Specific [BSS18, CCDF18, DBS19, FWR^{+}15a, HWNR14, HWFN14, ITT10, MS18, MSK18, SB11, SBL11, TBH^{+}14]. Specified [OI17]. Speckled [NBFL17]. Spectra [SWG^{+}16]. Spill [SM14]. Spines [BKM19, CHW12, IWB10, MH13, WKTQ15, WI17, WD10]. Spiny [BR14]. Split [HH15]. Split-Beam [HH15]. Splitter [SE12]. Sport [AJ10, CL12, FMA^{+}12, GBL^{+}13, HADG10, HPTH18, LDW10, LL12, MQS13b, MM18a, MS14, MOJ12, NKC16, PKKR10, RD10]. Sportfishing [CB10, LS10, LM16b, OM10, SL17a]. Spot [BRSB^{+}16]. Spotted [BKSH^{+}12, DLP16, GKS13, WBP13]. spp [DKSW15, DBE11, FOH^{+}15]. Spread [ACMB^{+}14]. Spreading [HW12]. Spring [CMD12, ISC10, Ise10, JHK11, JF13, JF14, KPP^{+}18, MPM10, PSC15, RM14, SLG^{+}15, WBF^{+}15]. Springtime [HAS^{+}19]. Squalius [GPLT12, GFP^{+}12]. St [BSB^{+}15]. St. [KYD18, RWAJ16]. Stability [HK17, VWB^{+}11]. Stable [BSH10, FDM^{+}13, HTPS10, THQ15, WGM^{+}11]. Stage [CSED10, CBHB16, PTH^{+}10]. Stages [CMD12, DDD^{+}18, DSD^{+}13]. Standard [BBC10, BBC11, BBS19, CK10, DP13, Ger10, Ger11, GPLT12, GFP^{+}12, MSK18, PBA^{+}17, PCP^{+}19, RFW^{+}10, RFW^{+}11, RV12, SR16, SBL11, SBW^{+}17b, UBR^{+}17]. Standardization [RFW^{+}10]. Standardized [ETD16, GRSF^{+}16, KNC14, MH13, THZ^{+}17]. State [Aka10, CWH^{+}14, SJM15, AHVJ11, KHDA14, PQB10]. State-Space [CWH^{+}14]. Stated [LL12]. States [ADD14, BW17, McB14, MAG^{+}15, ORD19]. Statewide [HRKB19, MS14, SSP^{+}19]. Static [Ask16, SFH12]. Station [LCC^{+}15, MHAP13, MAP14, PMP^{+}14]. Stationary [BRB^{+}17, BC18]. Statistical [FC17, FB14, HP14]. Status [BG12, DFPS10, FDM^{+}13, GWCA10, GSCCP18, Gre11, HER^{+}12, HB11, JHU^{+}13, MSM^{+}14, MAG^{+}15, SPQ12, SWH^{+}16, TLC^{+}19]. Stay [HHP^{+}12b]. Steady [GDP14]. Steelhead [APPR14, APC^{+}14, BS11a, BWLB12, BKHS14, Bou10, Bri17, CMWM18, CSH^{+}17, CFW^{+}10, CFWC11, Cle15, CAWB17, CHDW18, DK16, FBD^{+}19, FCH^{+}13, HDKB10, HFB^{+}13, HM12, HQ10, HLS^{+}17, JPWR10, KRJT13, KDMH10, MR16, MPL^{+}15, NWZ14, PNWC11, PNHA12, PDL^{+}18, Pto13, RS18, RPBM10, RABK18, SPS16, SMK13, TRB11, TCG^{+}17, TLC^{+}19, THQ15, TFH^{+}16, TEB^{+}19, WSS13, WWB^{+}17, WPJ^{+}18, WMMB19, WBB10]. Stem [DB13, KJN^{+}17, SEP^{+}14, SM19, SS16]. Sterile [BT16]. Sterile-Male [BT16]. Sterility [ZGW11]. Stewardship [LKvR^{+}18]. Still [CI10, DMRJ10]. Stimulation [SVD^{+}18]. Stock [ADD14, BZ13, BFLT17, BSB^{+}15, CDD^{+}11, DBS19, DZ14, Der14, DLP16, DC13, FB14, FHH16, FBC^{+}10, GSCCP18, HMBT16, HJB14, HWFN14, LBB16, McB14, PLCG13, PLUH19, SRCC^{+}18, SLG^{+}15, Sta12, TBH^{+}14, WHB^{+}14, WAW^{+}19, ZKM13]. Stock-Specific [HWFN14, TBH^{+}14]. Stocked [BSS^{+}17, BTD^{+}12, CGDB14, CMK^{+}19, CCDF18, DHK^{+}19, FHE^{+}12, HB15, HMD16, HM18, KRB^{+}16, KML14, KKMS11, KM11, LAM16, LN12b, LSK^{+}14, ORA17, PNWC11, PSR16, QA16, SDDC16, SKBSA11, SDF^{+}13, ZOC^{+}14]. Stocked-Trout [HMD16, HM18]. Stocking [APP13, CVD^{+}14, DBJD12, DWW17, FSHR17, HS19, HBA^{+}12, HADG10, LOS^{+}11, LC14, LSL^{+}13, LP17, Ols19, PT11, RS17, SMH17, SKBSA11, SSPP19, SDF12, TPB^{+}16, WSS13]. Stocks [APD^{+}18, DZ17, DWW17, VEW^{+}13, WFWSV14]. Stomach [LMS^{+}18, NSW16]. Stone [DKSW15, GCLC18]. Stonecats [PPO18]. Storage [MHR^{+}18]. Strain [PNWC11, SBVB12]. Strains [PDL^{+}18]. Stranded [BS11a]. Stranding [TPF^{+}13]. Strategic [Kra18, LTG^{+}11]. Strategies [APP13, BB19, CA14, CA11, CBHS16, CPSG12, DS12, HS19, KFM12, NFQ16, SH11, WHMS14]. Strategy [BDSB15, RJMG13, TRB^{+}13]. Stratified [SSK19]. Straying [CFWC11, CCC16, HWFN14]. streaked [SNT^{+}16]. Stream [ACMB^{+}14, ACS19, BACRB14, BP15, CBM13, DHB^{+}10, ETD16, Erm12, FDHWP16, FSV10, FSV12, GDLM17, HKMH10, HMP10, HWZ19, HK17, HHRB15, JSDJ16, KSH^{+}16, KCS^{+}17, MSFS12, MS17a, Mea12, MRC12, MHG^{+}11, MMB19, MB17, MKK^{+}12, Nak13, NWZ14, ORC^{+}16, POWF11, PFRM12, RGH^{+}13, RTJJ16, RD10, RBB^{+}10b, RIMR19, SRSS12, SBC^{+}11, SLB10, SBL11, SMS13, TRB11, WHMS14, WKP14, ZSW11]. Stream-Dwelling [ETD16, JSDJ16, KSH^{+}16, MHG^{+}11, MB17, PFRM12, SMS13]. Streamflow [HWNR14, HB16]. Streamflows [KLR^{+}12]. Streams [AJ10, BB19, BS11a, BBB^{+}14, BHN11, BE11, BG13, CEH^{+}15, CBWB14, DP13, DKWG14, DWCK16, DTA18, DCHC13, FSS^{+}10, HM17, HMW^{+}10, HAS^{+}19, HW10, HWZ19, IH11, JHB10, KKM16, KW13, MRA14, MH11b, MKK^{+}12, NCMT12, OHL10, PN19, PP10, QA16, RFKB17, RHMH15, SFW11, SLMM10, SHL16, SNTZ14, SPQ12, SQP12, SSP^{+}19, TWL^{+}17, WMB19, WWP18, ZW12]. Streamside [MHAO11]. Streamside-Reared [MHAO11]. Strength [FC17, Gur12, HFH12, HFH18, HSM^{+}18, LMS^{+}18, LOV17, NTJ15]. Stress [NBP15, SCG11b, WWS^{+}14]. Striped [ADD14, ALCD17, AZE^{+}18, BRA^{+}18b, BRA18a, CGDB14, JHU^{+}13, KHDA14, LAA^{+}18, LHS^{+}14, SHMS16, SG13]. Stripes [WBT17]. Strobe [FWT14, SWW14]. Strong [DS12]. Structural [DDAS14, KDP^{+}10, SS16, WWW15]. Structure [APPR14, ADD14, BWLB12, BCRM^{+}14a, BCRM^{+}14b, BS10, BSVB19, BRBP10, CSWW18, DPK^{+}17, EWS^{+}17, FH16, FQ19, HMBT16, HW16, HMCT16, JWC10, JF14, LSMD15, LR12, LTG^{+}11, MNYV12, McB14, NEC^{+}13, ORD19, OEZ17, ORL^{+}16, OHSP17, PW13, PRTS14, PRRD15, Ryp15, SBW^{+}17a, SQGK10, TTS^{+}18, TLSB18, THZ^{+}17, WSA18]. Structured [FC17, FJB16, GWB^{+}10, RZ10, WQP13]. Structures [ABVS14, BCZ^{+}18, BKM16, BMW^{+}19, CPM^{+}10, KNS18, OLM15, SP13, SBTW17, SS13b, TP17, WMD^{+}19, ZWH^{+}15]. Studies [HMP10, Rey15, SL15b]. Study [ACMB^{+}14, BACRB14, BT16, BWR12, CWH^{+}14, DASB16, ES17, FJB16, GWCA10, GHK19, MKBB13, MEL^{+}12, MWS15, MCH12, MSK19a, MEJ10, MHC10, RWAJ16, RGOR10, SBL^{+}18a, SDF12, ZEM11]. Sturgeon [BFP12, BRS^{+}12, Bal15, BBG19, BGD^{+}19, BDC^{+}13, BDT^{+}16, BFL^{+}10, CSED10, CBHB16, CGG^{+}18, CHL13, CJH16, CHW12, DFPS10, DJS^{+}16, FGS13, GMG^{+}15, HHP12a, HHR^{+}18, HFCH15, HUB^{+}18, JTB16, KCF15, KSM^{+}18, KD10, KYD18, LMS^{+}18, LMB^{+}18, MHAO11, MCB^{+}19, MHAP13, MAP14, MLEK15, PPR11, PTH^{+}10, PVH^{+}13, RHPH14, SH11, SPT^{+}12, SKBB15, SPK10, SPP12, SWHA15, SH18, SDL^{+}15, SBL^{+}18b, TPF^{+}13, TGKI10, UWL13, VKH^{+}19, WA16, WBF^{+}15, WEC^{+}19, WMD^{+}19]. Sturgeon-Sized [LMS^{+}18]. Style [RLC12]. Subadult [PPR11]. Subarctic [ZWH^{+}15]. Subestuaries [SEB^{+}19, UML^{+}11]. Subject [SRFS15]. Subjected [KDA^{+}18, PD13]. Subjective [CGL18]. Sublegal [MFRS11, SS11b]. Submerged [DMRJ10, MKWM18]. Submergence [CMC^{+}19]. Submersible [BSH10]. Subsistence [GD10]. Substrate [BSH10, CF13, GHK17, PP11, SCKZ17, TFP^{+}11]. Substrates [HWZ19]. Subyearling [CCC16, SEP^{+}14]. Subyearlings [RCP^{+}13]. Success [AJ10, BA11, CJMD10, DB13, DWW17, MP14, MSK19a, PRR^{+}13, RSLG10, RW11, See11, SKH^{+}16, SMS13, VJK13, WP11, WGJ10, WMMB19, WWP18, WGM^{+}11]. Successful [FWG15, SNTZ14]. Sucker [BC14, BCZ17, HPA^{+}11, PC12]. Suckers [Bur11, BHR^{+}15, EHP^{+}16, HUST17]. Suffocation [TGKZ19]. Suggest [CBHS16]. Suggests [TPBF15]. Suitability [KYD18, LS13, RBP16, ZSW11]. Suitable [Kae12]. Summer [BBB^{+}14, CFW^{+}10, DCS^{+}16, HGW10, HLM^{+}18, HF13, JKN^{+}10, JM15, MS18, MBPG12, MMHM17, RSQH18, SOB18, SG13, SM19, SD11, UML^{+}11]. Summer-Run [MS18, SM19]. Superimposition [WGJ10]. Superior [SBHJ17, AHS15, JHS15b, LSL^{+}13, LOV17, PNWC11, RHG12, RHG13, SQGK10, WSS13]. Supersaturated [WLL^{+}15, WALL17]. Supplemental [WWD13]. Supplementation [BZ13, MNYV12, SH11]. Supply [VBBK16]. Support [HMP10, Ols19]. Suppression [CSWW18, DGF11, FGS13, KQRS16, KACS18, SMH17, SGC13, TGKZ19]. Surf [FWG15, LSLC15]. Surface [APPR14, BP15, BHR^{+}15, UML^{+}11]. Surfaces [SM13]. Surgery [DBW13, HKP^{+}14]. Surgery-Based [HKP^{+}14]. Surgical [CSV13, CHL13, WWS^{+}14]. Surgically [GM12, HBLL12]. Surplus [NW12]. Surrogacy [BSG10]. Surrogates [BKM19]. Survey [BvPCH17, Cle15, DPS^{+}17, GBL^{+}13, GAWC10, GCA^{+}13, HWDA12, HWB11, KD10, LSLC15, MSW^{+}14, MQS12, MQS13a, OM10, PZC^{+}12, Qui10, SCS17a, SCPC14, SKS^{+}15, WLK^{+}16, ZSW11, ZW12]. Survey-Derived [PZC^{+}12]. Surveys [AMQSK15, APF18, BvPCH17, CFBB15, FWG15, MQS13b, McC15, MM17, SPW^{+}12, SLP13, SM10]. Survival [APPR14, APAC12, ABČ^{+}15, APD^{+}18, BS11a, BK16, BQHR17, BKD^{+}10, BP14, BB12, BRSB^{+}16, BSBF13, BSM^{+}15, BBS18, BWR12, BKP^{+}14, CBS^{+}12, CSV13, CCDF18, CFWC11, CCC16, CDQG12, DVWS10, DDEW18, ES17, FBL^{+}19, FHE^{+}12, GM12, GLCD13, GLH^{+}18, GKT11, GMB10, HGW10, HNH11, HFB^{+}13, HR11b, HEL^{+}15, ITY13, JPWR10, KWMB11, KDA^{+}18, KB11, KBH^{+}19, LBF17, MCB^{+}19, MAP14, MSD11, MFRS11, NB10, NASG12, OL17, PFWG16, PSB^{+}10b, QK11, RLC12, RSQH18, RSQ^{+}19, SFH^{+}13, SBVB12, SSE19, SPS16, SQS^{+}17, STSH10, SEP^{+}14, SWH^{+}16, SS11b, SPK10, SJM15, TRB11, TPC^{+}15, TGKI10, TPL18, TLL^{+}18, WH12, WF17, WMMB19, WWD13]. Suspected [AZE^{+}18]. Susquehanna [SBW^{+}17a, SEM14]. Sustainability [FBM^{+}19, PNWC11, SCA11]. Sustainable [HCH13]. Suwannee [BBC10, BBC11, MASN16]. Swamp [SGCB19]. Swan [SGC13]. Swim [PRR^{+}13]. Swim-Up [PRR^{+}13]. Swimming [BLSS19, MCB^{+}19, UMC14, WALL17]. Sylvia [GD10]. Sympatric [Hax15, JBZ^{+}10]. System [BSVB19, BQR^{+}19, CJA^{+}13, FWR^{+}15a, HB12, HB14a, HJK11, KBK13a, KBK13b, KPL^{+}16, MI10, MS14, PMP^{+}14, RCGF18, SWH^{+}16]. System-Specific [FWR^{+}15a]. Systematic [XM14]. Systems [BRB^{+}17, CGDB14, FAW14, KADH15, LMB^{+}18, PSC15].

T [BF14, BR12, HHP12a, RWPH11]. T-Bar [BF14, BR12, HHP12a, RWPH11]. Table [ABVS14]. Tabulation [WH11, WH12]. Tackle [LMB^{+}18]. Tactics [CJ14]. Tag [AAYC14, BSL18, BA11, BRSB^{+}16, BHC^{+}10, Bur11, CBS^{+}12, CSV13, DBW13, FBL^{+}19, FOH^{+}12, HHP12a, HUST17, KWMB11, KBK13b, KPL^{+}16, LKB19, LXM^{+}12, McC16, MM18a, MS14, NKC^{+}18, PFWG16, RWPH11, SC17, SAYJ17, SKBK12, TPC^{+}15, VBT^{+}12, WBP13, WRNJ^{+}19, WPY^{+}15, WH12, WF17, YHP10]. Tag-Return [McC16]. Tagged [CMD12, DHLD11, KPL^{+}16, KBH^{+}19, LKB19, MEL^{+}12, SJM15, VBT^{+}12, WH11, WS13, YS12, FAW14, OHL10, RFKB17, SBL11]. Tagged-Fish [MEL^{+}12]. Tagging [Bou10, Bur11, DR10, HKP^{+}14, JSDJ16, MCB^{+}19, SC17, WPY^{+}15, WH11, WH12, WHG^{+}16]. Tagging-Related [WPY^{+}15]. Tags [BRB^{+}17, BF14, BR12, DBKG14, ERC^{+}11, GL16, HHP12a, LN12a, LXM^{+}12, MM16, MLHS13, MHG^{+}11, MWS15, NKC^{+}18, NBP15, OL17, PFWG16, SNT^{+}16, SLSS16, SKBB15, SCB^{+}19, TPC^{+}15, WS13]. Tailwater [QK11]. Tailwaters [MODC10]. Taken [HWB11]. Tamarac [GHK19]. Target [DZ14, DZ17, DCL19, Gur12]. Target-Rich [DCL19]. Targeting [HAW^{+}18]. Taxonomic [KIST16]. Technique [BFLT17, BF13, BP15, BGBG17, BT16, CKBP19, ERC^{+}11, HD14, MODC10, MH13, SBB19, TL15]. Techniques [BE11, CBWB14, DVWS10, DCHC13, GAWC10, HWDA12, HO13, HKP^{+}14, PBA^{+}17, PC12, RTMJ10, RMTW16, SQS^{+}17, SBL11, SPL18, TMC10, UBR^{+}17, WD10]. Technological [ASS17]. Technologies [Aka10]. Technology [BDT^{+}16, CHJ^{+}16, SAYJ17]. TEDs [RGC^{+}12]. Teeth [LKV^{+}17]. Telemetric [BGP10]. Telemetry [HSM^{+}18, HBLL12, JPWR10, STP19, TL15, WGS16, WEC^{+}19]. Teleost [SWG^{+}16]. temensis [NBFL17]. Temperate [SCKZ17, WS11b]. Temperature [AJR^{+}19, BS10, CSH^{+}17, CHW12, FDHWP16, HP14, HLM^{+}18, HB16, IH11, KDA^{+}18, KMPV13, MMP12, PCP^{+}13, PTH^{+}10, WCWA10]. Temperature-Driven [KMPV13]. Temperatures [BP15, BGHL10, CBM13, DCS^{+}16, HB14b, Kae12, SK18a, SHS15]. Temporal [AOS^{+}17, Bri17, DZS10, IMM13, MS17a, SGCB19, SBL^{+}18a, SSK19, SCS17b, VEW^{+}13, WS11a, WDD^{+}14, WSH^{+}14, WBF13]. Temporally [KSL^{+}19]. Ten [FBB18]. Tennessee [KB11, WA16, HRKB19, MB12]. Tenualosa [NCSR17]. Term [APC^{+}14, BCD^{+}10, BEK^{+}18, BDC^{+}13, Bur11, FWWTJD16, GM12, HWZ19, HHRB15, HUST17, KML14, KCRR19, LBL11, PPR11, Pie10, PRRD15, RLC12, RWPH11, SS11a, SBC^{+}11, SBL^{+}18b, VMC^{+}11, YHP10]. Termination [JCB^{+}18]. Tern [ERC^{+}11]. Terrapin [ASD^{+}18, ASD^{+}18]. Territories [ASP^{+}17, CEH^{+}15]. Territory [HS19, MSE^{+}18]. Test [GFM^{+}18, HVW^{+}17, MJH17]. Testing [DCS11, KDP^{+}10, MRC12, MEJ10, SWH^{+}16, SCG11a]. Tetanized [SVD^{+}18]. Texas [BSO^{+}16, DDAS14, FTSF13, Mat10, PSB^{+}10a, SDKS16]. Their [FDM^{+}13, DMC^{+}19, PPO18, TTS^{+}18, WQF15, WBT17, ZC15]. Theoretical [ASD^{+}18]. Thermal [BBB^{+}14, BP15, SD11]. Thermally [SPL18]. Thiamine [FDM^{+}13, RRH^{+}11]. Those [LMC^{+}16]. Threadfin [NP12]. Threat [BB19, HTH^{+}19, KACS18]. Threaten [RFHP17]. Threatened [BBC^{+}13, CWH^{+}14, DGS^{+}11, GMG^{+}16a, GTB^{+}19, VWB^{+}11, VEW^{+}13]. Three [BBG19, BMW^{+}19, BSTC13, DS12, DBW13, DP13, FSW16, GWB^{+}10, GKT11, HKMH10, HB15, Kae13, KPL12, KNS18, RGC^{+}12, SLB10, SS17, SBTW17, SM13, SS13b, SQGK10, WSH^{+}14, ZWH^{+}15]. Three-Dimensional [BBG19]. Three-Pass [DP13, HKMH10]. Threshold [LCR19, Rey15, SL15b]. Thresholds [BBB^{+}14]. Throat [SBW^{+}16]. Thunder [FBM^{+}19]. Tidal [GRCR10, TBH^{+}14]. Tide [WBZ19]. Tiger [KMK^{+}15, SSP18, WBT17]. Timber [RIMR19]. Time [AWG^{+}18, CS17, GKT11, HRH10, KAZ15, KDA^{+}18, KCCW19, LM16a, LJR11, LLW^{+}18, MQJK19, NW12, RMTW16, SLB10, SK18b, WH12, WMMB19]. Time-Lapse [AWG^{+}18, SK18b]. Time-Varying [NW12]. Timers [SEP11]. Times [CMWM18, LMB^{+}18, McC16, TGA15]. Timing [BWLB12, BCRM^{+}14a, BCRM^{+}14b, BC14, BK12, BGP10, HWFN14, JKN^{+}10, LPY16, PRTS14]. Tissue [ASP^{+}17, LLW^{+}18]. Tissues [HTPS10]. Toe [PK18]. Toe-Width [PK18]. Tolerance [GLCD13, WLL^{+}15]. Tomography [MHBR19]. Tool [CvPW15, CCMA^{+}10, CDQG12, FDHWP16, HBS14b, MBBS10, Ols19, PLCG13]. Tools [RABK18, ZSW11]. Topanga [DK16]. Topeka [BPQ13]. Topminnow [RW11]. Total [DZ17, GLCD13, LKV^{+}17, MQS12, MS14, PSCS11, SKBB15, SEP^{+}19, WLL^{+}15, WALL17]. Tournament [BRSA15, DLSW12, HAW^{+}18, KB10, KDA^{+}18]. Tournaments [BEH^{+}19, EBS18, MMS^{+}17, Qui10, SVFG10, SHS15]. Toxicity [Erm12, FSV10, FSV12]. Tracking [BSTC13, HHRB15, NTJ^{+}14, SCB^{+}19, WPS^{+}19]. Trade [SLR^{+}16]. Trade-Offs [SLR^{+}16]. Tradeoff [RWAJ16]. Traditional [RPC^{+}19]. Train [OOWW12]. Training [DIK19]. Traits [DWCK16, HFB^{+}13]. Trammel [WKS^{+}10]. Transitions [BBB^{+}14]. Translocated [ASL^{+}16]. Translocation [HW19]. Translocations [GMG^{+}16a]. Transmitter [GLH^{+}18, HO13, WRNJ^{+}19, WPS^{+}19]. Transmitters [BF13, BDC^{+}13, BHC^{+}10, CHL13, FVH^{+}17, HBLL12, NKC^{+}18, TGA15]. Transplanted [HVW^{+}17]. Transponder [AAYC14, BRB^{+}17, BR12, Bur11, CHJ^{+}16, FOH^{+}12, HHP12a, HUST17, KWMB11, KZ12, MM16, MHG^{+}11, NKC^{+}18, OL17, SNT^{+}16, SKBK12, WF17, WHG^{+}16, YHP10]. Transponders [BSTC13, GM12, SLSS16]. Transport [BEH^{+}19, MHAP13, MSD11, SZHC15, SEM14]. Transported [DB13, HH11b]. Trap [ASS17, AHVJ11, ASD^{+}18, BDS^{+}11, BC18, CS17, GCLC18, HWZ19, JM12b, KPP^{+}18, MHAP13, MCH12, SCW^{+}17, TRB^{+}13, ZM17]. Trap-and-Haul [KPP^{+}18]. Trap-and-Transport [MHAP13]. Trapping [BA18]. Traps [AHVJ11, DBE11, MHC10, PDM12, PSC15, RJMG13, RH16, SBL^{+}12, ZW12]. Trauma [PD13]. Travel [GKT11, KAZ15]. Traveling [BP14]. Traverse [BGCF11]. Trawl [BSS^{+}14, GGC17, HPD^{+}19, RGOR10, TSNH10]. Trawling [DDD^{+}18, NQ11]. Trawls [HB12, SWHA15]. Treating [HSF18]. Treatment [CHW12, FBB18, RFLP14]. Treatments [CST^{+}18, LP12, TFH^{+}16]. Trend [HS12]. Trends [BG12, GWCA10, GHL12, IMM13, MS17a, MGS14, SBL^{+}18a, WS11a]. Trial [SC18]. Triangulation [TL15]. Tribal [MSE^{+}18]. Tributaries [BBO16, BWT^{+}13, BSM^{+}15, CMMR19, CFW^{+}10, GAH^{+}18, HPD^{+}19, HBA^{+}12, JBZ^{+}10, LSL^{+}13, SQGK10]. Tributary [ACS19, BQHR17, BT16, BJ19, CJMD10, FDM^{+}13, LBF17, LPY16, MFLS18, WSS13, WBH13]. Tribute [CMMR19]. Triggerfish [AES11]. Trip [BHHA11, HHSB13, LM16b]. Triploid [BTD^{+}12, CMK^{+}19, KML14, KSG^{+}18, KKMS11, KM11, MKBB13, ZGW11]. Trips [SM10]. Troll [MG14]. Trolling [MDM12]. Trophic [BRBP10, FBB19, WQF15]. Trophodynamic [MG10]. Trophy [DAKP13, FILH15, FH16, HBS^{+}19, HPTH18, PJLB14, SPHK16]. Trophy-Oriented [FH16]. Tropical [LFD^{+}17, NP12]. Trotline [SEP11]. Trout [AW10, AHS15, ACRBB10, ACS19, APAC12, ASP^{+}17, ASL^{+}13, ASL^{+}16, AKP15, BB19, BPB13, BBD19, BACRB14, BHN11, BATK16, BBO16, Blo13, BSVB19, BB12, BS11b, BBL^{+}11, BPT^{+}19, BJZ11, BGM17, BTD^{+}12, BWR12, BKP^{+}14, BHG^{+}13, BJ19, CEP15, CCMA^{+}10, Car13, CMK^{+}19, CDQG12, CBM13, DRG10, DBW13, DBKG14, DB13, DZS10, DKWG14, DWCK16, DTA18, DCS^{+}16, DSD^{+}13, DHB^{+}10, DGS^{+}11, DGF11, EKM10, ES14, Erm12, EMK^{+}17, FVH^{+}17, FSV10, FSV12, GMG^{+}16a, GBFR18, Gre11, GFC^{+}15, HW12, HKMH10, HB15, HNH11, HH17, HM17, HHT^{+}13, HAS^{+}19, HW16, HHMQ15, HER^{+}12, HWZ19, HK17, HHRB15, HFRH17, HGK15, HS12, HDM15, HBA^{+}12, HMD16, HM18, JWC10, JBZ^{+}10, JM15, JHS15b, Kae12, Kae13, KKM16, KHH13, KSH^{+}16, KKMS11, KM11, KMK^{+}15, KPL^{+}16, KBH^{+}19, KEC11, KACS18, KW13, Kra18]. Trout [LM16a, LKB19, LOS^{+}11, LLP10, LN12a, LWJS16, LSL^{+}13, LLW^{+}18, LN12b, LPY16, LP17, LMD^{+}19, MD10, MM16, MAM^{+}12, MKRH14, MP14, MHDD16, MJ11, MSC13, MPL^{+}15, MB12, MGWC13, MH11b, MHG^{+}11, MGS14, MSM^{+}14, MSK^{+}16, MKHC17, MAG^{+}15, ND11, NFQ16, NWZ14, NEC^{+}13, OL17, PT15, PNWC11, PJLB14, PSR16, PFRM12, PSCL13, PN19, PVST11, Pto13, QK11, RSHS15, RTMJ10, RRH^{+}11, RZ10, RFHP17, RSQ^{+}19, SHSCA16, SDDC16, SLMM10, SMH17, SMP^{+}14, SRSS12, SBVB12, SBC^{+}11, SKH^{+}16, SNTZ14, SBHJ17, SSPP19, SQGK10, SMS13, SGC13, TCM19, TBIJ19, TGS^{+}15, TGKZ19, THQ15, UACZ^{+}16, VBBK16, VMA^{+}16, WDD^{+}14, WQF15, WMSV18, WGJ10, WHV15, WSA18, WBT17, WWP18, ZRR^{+}19]. Truckee [APAC12]. Truman [STP19]. Truncated [CSWW18]. Tuna [GSG18, MG14]. Turbid [BOM16, PP11, TKH^{+}16]. Turbidity [SL15a, SSP18]. Turbine [ABČ^{+}15]. Turbines [GBTL14]. Turkey [GPLT12]. Turtle [RGC^{+}12, SHM^{+}10, SMW12]. Two [APD^{+}18, APF18, BCBW18, BA18, BFL^{+}10, BE11, BGM17, CA14, CPM^{+}10, DSB19, DT11, DWCK16, FK16, FSS^{+}10, HKMH10, HMW^{+}10, HEL^{+}15, ISC10, JHB10, JBZ^{+}10, Kae13, KWMB11, KFM12, KCS^{+}17, LCR19, ORA17, PFWG16, PSB^{+}10a, PC12, PDL^{+}18, RB15, RS17, RTMJ10, STSH10, SEP^{+}19, UACZ^{+}16, VMC^{+}11, VKBL11, WHB^{+}14, WI17]. Two-Month [APF18]. Two-Way [FK16]. Type [CKCF12, GRSF^{+}16, JHS15a, LLW^{+}18, SM14, TEB^{+}19]. Types [BA11, DR19, LMB^{+}18, NCMT12, SBTW17].

U.S [RKB15]. U.S. [AES11, BGCH17, BM14, CDD^{+}11, CH17, DTS19, GSCR11, GSG18, GMG16b, LSMD15, LS17, MG14, SP13, SMW12, TCB18, WS11a]. U.S.-Built [DTS19]. ui [SRFW14, SRFS15]. Ultrasonic [BSH10, BWB18, DG12, GBTL14, STP19, TL15, WGS16]. Ultrasonography [CBHB16]. Ultrasound [DR10, LS11]. Umatilla [CCC16, JM12a]. Unanticipated [PM14]. Unbaited [Sin18]. Unbiased [CHBY13, SM10]. Uncertain [GSCCP18, ZRR^{+}19]. Uncertainty [CA11, HJB14, HSM^{+}18, SJM15]. Under-Ice [FHL18]. Undersized [GMB10]. Understand [SEP11]. Understanding [CGL18, DHH17, FQ19, HSC^{+}16, SF16]. Underwater [CFBB15, IH11, MI10, NTJ^{+}14]. Unexplained [EBS18]. Unexploited [HVW^{+}17, HHH14, LWJS16, MS16]. Unimpounded [MCK^{+}15]. Unique [MRC12, MGWC13]. Unit [Bre13, LJR11, MQJK19, SHL16, WHC14a, WHC14b]. United [ADD14, MAG^{+}15, ORD19, BW17, McB14]. Units [CKBP19, DC13]. Unrecovered [BRA^{+}18b, BRA18a]. Unreported [APD16]. Unsectioned [MH13]. Unstocked [PSR16]. Unsuccessful [BBG19]. Unsuitability [MRB^{+}11]. Untagged [NKC^{+}18]. Update [PK18]. Updated [GGC17]. Upland [GBHO11]. Upper [DM12, BSS^{+}17, BWT^{+}13, CJH16, EMK^{+}17, EHP^{+}16, FSHR17, GKS13, HLM^{+}18, HGW16, JF13, JF14, KEC11, KACS18, MSM^{+}14, OHSP17, OH15, WA16]. Upstream [HH11b, LKM^{+}17, MHAP13, MJH17, MSK19a, PFD17, PN19, SM19, SEM14, WA14]. Upwelling [HSF18]. Urban [BLM^{+}11, MBBS10]. USA [BACRB14, CL12, CMMR19, GRCR10, HHB12, HM18, JBRWH18, JKSK13, KRJT13, Sam16, SHL16, SCWP17, WHB^{+}14]. Usage [BWT^{+}13]. Use [APAC12, ABVS14, AWG^{+}18, BSL18, BDH^{+}14, BSBC14, BG19, Bre13, BSBF13, BJ19, BWHM16, CKC^{+}12, CHJ^{+}16, CDQG12, DDAS14, DCS^{+}16, DR19, EWS^{+}17, FG13, FVH^{+}17, FSW18, FJB16, FDM^{+}13, GFC^{+}15, GTB^{+}19, HGW10, HO13, HH11a, HW16, HPM^{+}18, LS11, McC15, MS14, MKHC17, MMB19, MKK^{+}12, OHL10, ORA17, Pto13, SG13, SCKZ17, SPM15, SDF12, SK18b, SD11, SWW14, SLP17, SF16, TFP^{+}11, TMC10, VEIW^{+}16, WHMS14, WHKB17]. Used [BCZ^{+}18, BFLT17, BBC^{+}13, HTF^{+}13, HSF18, KQDC19, PVST11, RBB^{+}10b, SVD^{+}18, SEP^{+}19, TP17, VKBL11]. User [ZSW11]. User-Friendly [ZSW11]. Users [SDKS16]. Using [ASP^{+}17, AR12, BRN16, BS12, BB19, Bal15, BBG19, BOM16, BCD^{+}10, BDT^{+}16, BRWQ18, BGP10, BSTC13, CSWP19, CSH^{+}18, CWH^{+}14, CPM^{+}10, CBS13, CDD^{+}11, DPS^{+}17, DBJD12, FOH^{+}15, GRSF^{+}16, GHL12, GL16, GCA^{+}13, HBS^{+}19, HS19, HSM^{+}18, HF13, HWFN14, HBS14b, KIST16, KNC14, LKV^{+}17, LKB19, LSC^{+}16, MS16, MKBB13, MC10, MFLS18, MSW^{+}14, MJCM15, McC16, McC17, McG14, MPL^{+}15, MWS15, MLEK15, MM18b, MPM10, NKC^{+}18, NEH10, ORC^{+}16, PNHA12, PA16, PRRD15, RGH^{+}13, RZ10, SBL^{+}18a, SLK15, STP19, SNTZ14, SBL11, SHM^{+}10, SK18b, SEP11, SL12, SMCS14, THQ15, UWL13, VKH^{+}19, WS16, WBZ19, WHV15, WD10, ZEM11, CA14, CBHB16, Der14, DLP16, FOH^{+}12, GHK17, HM17, JPWR10, SC17, SFW11, WA16, WTB^{+}15, ZM17]. Utah [BWT^{+}13]. Utility [CvPW15, DCS11, DAKP13, NKC16, RCGF18, SB15]. Utilization [RFKB17]. Utilizing [DRC17].

Vacant [ASL^{+}13]. Validating [SM10]. Validation [ASD^{+}18, AM17, CHDW18, DSD^{+}13, FBC^{+}10, HM17, HB16, HBS14b, HUB^{+}18, Ise10, LOV17, MJ11, PVST11, RSI16, RHPH14, SBS14]. Valley [SHMS16, MGB^{+}14]. Valuable [PLUH19]. Value [BRN16, BM14, CL12, FJB16, JBHM19]. Values [BLM^{+}11, GSG18, HHS^{+}13, LL12, LN12b]. Valuing [ML13]. Varden [SS13b, SMCS14]. Variability [FWR^{+}15a, HSC^{+}16, HB16, KTN14, LOV17, SSK19, SCWP17]. Variable [BSM^{+}15, SC17]. Variables [GRSF^{+}16, MOJ12]. Variation [AES11, BACRB14, DBK11, DASB16, EBS18, KSL^{+}19, PP11, SCS17b, SVPD13, SEP^{+}19, WSH^{+}14, WBF13]. Variations [VBBK16]. Varies [RH16]. Various [DT11]. Vary [TO10]. Varying [AJ13, NW12, WS11b]. Vegetation [CBM13, DFMP18, MKWM18, SB18, TFP^{+}11]. Venting [ES17, SFH^{+}13]. Verify [PN19]. Verifying [BS11b]. Vermont [KW13, Kra18]. versus [BW18, Blo13, FH16, MQJK19, TKH^{+}16]. Vertical [GH17, WBZ19]. Vessel [BRS^{+}12]. Veterinary [LS11]. via [MG10]. Viability [APC^{+}14, BBC^{+}13, JCB^{+}18, PHS^{+}13, SZKK10]. Video [AM17, CKC^{+}12, MI10, MJCM15, NTJ^{+}14]. Video-Based [AM17]. Videography [CFBB15]. VIE [LN12a, ZC10]. Viewing [WI17]. Viral [PCT^{+}14, TPBF15]. Virgin [CH17]. Virginia [ABBG13, BRS^{+}12, DMC^{+}19, HH11b, HM18, SEB^{+}19]. Virile [RB15]. virilis [RB15]. Virus [PCT^{+}14, TPBF15]. Visibility [WH12]. Visible [DBKG14, LN12a, MM16, MLHS13, NBP15, YHP10, ZC10]. Visual [BGW^{+}11, KHH13]. Visualizing [SLR^{+}16]. Vital [BRA18a, UACZ^{+}16]. Vitellogenin [HPA^{+}11]. Volitional [CFWC11]. Voltage [SPS16]. volume [Ano12b, BBC11, MFRS11]. Volunteer [Cle15]. Vortex [MC10]. Vulnerabilities [CDD^{+}11].

Wabash [KCF15, MCSF^{+}17]. Wadeable [AJ10, DCHC13, NCMT12, PP10, SHL16, SPQ12, SQP12]. Wading [SM13]. Waikato [DHLD11]. Walker [SBVB12]. Walleye [BEK^{+}18, BSB^{+}15, CSF12, CF13, DIK19, EBS18, FB14, FWT14, FHE^{+}12, GHK17, HIS16, HS19, HFH12, HHCH15, HFH18, HNPB15, HSHH17, KKSH12, KIST16, KBK13a, KBK13b, LSM^{+}16, LC14, LBF17, LJR11, LO18, MLHS13, MH13, MSE^{+}18, Ran18, RS17, RSI16, SS11a, SFH12, SSK18, THI^{+}19, VMC^{+}11, WFS13, WF17, WI17, ZEM11, ZKM13]. Walleyes [BBW18, CFG16, CJMD10, FJB16, GWB^{+}10, GHK19, HFN11, Hax15, HB10, LLP10, LP12, LAM16, MSK18, RS11, SVFG10, TS11, TLL^{+}18, VKBL11, VBT^{+}12, WCWA10]. Warmer [HW12]. Warmwater [BE11, BG13, PP10, SWG^{+}16, WMB19]. Washington [APPR14, AT18, AHVJ11, BCD^{+}10, BBD19, BPT^{+}19, HFB^{+}13, KTHM13, LPY16, LP17, LSLC15, MS18, PQB10, SMK13, SM19, TFP^{+}11, WBB10]. Water [BP14, BSBC14, BGHL10, BEH^{+}19, CFBB15, CGG^{+}18, CSH^{+}17, CHW12, DBK11, DT11, DMRJ10, DCS^{+}16, GRR^{+}17, HHR^{+}18, HB16, HB14b, KDA^{+}18, LSM^{+}16, LBL11, MC10, NB10, PTH^{+}10, RSLG10, Rey15, RTMJ10, RJP^{+}15, SL15b, SHS15, VBBK16, WLL^{+}15, WALL17]. Waterbird [EHP^{+}16]. Waters [AHVJ11, AES11, BR14, BMMMLML15, Car13, FWWTJD16, GSCR11, KHDA14, LSMD15, MSK^{+}16, PNWC11, RKB15, RMTW16, TS11]. Watershed [AJ10, ASP^{+}17, GKMC15, HMP10, Kra18, MMP12, RPBM10, TCB18, UML^{+}11]. Watershed-Scale [HMP10]. Watersheds [BC17].

Waveform [HRH10, SPS16]. Waveforms [BDB^{+}19]. Way [FK16]. Weak [CF18]. Web [HTH^{+}19, SOSL10]. Weigh [KB10]. Weigh-In [KB10]. Weight [BBC10, BBC11, BBS19, CK10, Ger10, Ger11, GPLT12, GFP^{+}12, HDB^{+}12, ITT10, LRHR17, RFW^{+}10, RFW^{+}11, Ran18, RV12, SKBB15]. Weir [APD^{+}19]. Weirs [FK16, MC10]. Well [BEH^{+}19, SHS15]. Wells [MS18, SM19]. West [CDD^{+}11, BPQ13, FWWTJD16, NEH10, WFWSV14]. West-Central [BPQ13, FWWTJD16, WFWSV14]. Western [MKWM18, NBP15, AHS15, AOS^{+}17, AES11, BACRB14, GPLT12, GSCR11, HHT^{+}13, JBRWH18, JHS15b, LP17, MAG^{+}15, ORA17, RMTW16, WBT17]. Westslope [CMK^{+}19, DCS^{+}16, WQF15]. Wetland [CKCF12]. Wetlands [BA18, CKCF12, SBTW17]. Where [WPS^{+}19]. While [RCP^{+}13]. Whirling [CEP15]. White [BC14, BCZ17, BDS^{+}11, BGD^{+}19, BDT^{+}16, CJH16, EJL10, FBB19, GBFR18, GHK19, GMG16b, KD10, LMS^{+}18, LMB^{+}18, MSK^{+}16, MSK18, ORD19, PPR11, SNT^{+}16, SRFW14, SPL18, TGS^{+}15, WKS^{+}10, WBH13]. White-streaked [SNT^{+}16]. Whitefish [BATK16, BGW^{+}17, EWS^{+}17, GZ13, HMS11, HHH14, PCP^{+}13, PM13, ZM17, ZWH^{+}15]. Whitespotted [Nak13]. Whole [WSA18]. Width [DC13, PK18]. Wild [BKHS14, BEK^{+}18, Bri17, BSG10, FDM^{+}13, GPP^{+}10, HHT^{+}13, HUST17, MM16, MHAO11, MS18, OMT^{+}15, PTD10, PHFK10, SHSCA16, SKBSA11, SPP12, TPB^{+}16, TEB^{+}19, WPJ^{+}18]. Wild- [WPJ^{+}18]. Wild-Caught [BEK^{+}18]. Will [LN12c]. Willamette [JF13, JF14, Cle15, KJN^{+}17]. William [SC18]. Williamson [HHB12]. Wind [SCKZ17]. Winnebago [KBK13a, KBK13b]. Winter [Bri17, DTA18, HLM^{+}18, KLR^{+}12, MH11a, WMB19]. Winter-Rearing [HLM^{+}18]. Wintertime [SSPP19]. Wire [AAYC14, BSL18, BR12, ERC^{+}11, GL16, JSDJ16, KPL^{+}16, KBH^{+}19, LKB19, LXM^{+}12]. Wisconsin [LO18, BHN11, CSED10, CBM13, DHB^{+}10, EDN10, FILH15, FH16, GBL^{+}13, HIS16, HM17, HS19, HFN11, HFH12, HHCH15, HFH18, KIST16, MIH15, MSE^{+}18, ND15, ORL^{+}16, RSI16, Ryp15, SSE19, SEP^{+}19, THI^{+}19]. Within [BCBW18, CMMR19, CAWB17, HTF^{+}13, KTHM13, LAA^{+}18, LTG^{+}11, LSL^{+}13, TTS^{+}18, VWB^{+}11]. without [DTA18]. Wood [CBWB14, DHB^{+}10, GDLM17, Kra18, MRA14]. Woody [LGC11, WWW15, WS11b]. Working [SRCC^{+}18]. Worthy [ASD^{+}18]. Wound [FSW18]. Wyoming [BQR^{+}19, KEC11].

X [BDT^{+}16]. X-Ray [BDT^{+}16].

Yakima [HFB^{+}13, TFH^{+}16]. Year [DS12, FC17, HFH12, HFH18, KFM12, KCCW19, LSM^{+}16, LOV17, LO18, MCS^{+}17, NTJ15, SB11, SC18, WH11, WH12]. Year-Class [FC17, HFH12, HFH18, LOV17, LO18, NTJ15]. Year-Classes [DS12]. Year-Specific [SB11]. Yearling [BSG10, HLM^{+}18]. Years [ES14, KTHM13, RIMR19, SWH^{+}16]. Yellow [BK12, DWB^{+}15, DRC17, FRH10, HTH^{+}19, HT11, KWMB11, LSM^{+}16, LJR11, LSK^{+}14, MIH15, MSK19b, RBB10a, SB11, TCB18, WA14]. Yellow-Phase [TCB18, WA14]. Yelloweye [HR11b]. Yellowstone [EMK^{+}17, SZKK10, ACS19, DZS10, EMK^{+}17, Gre11, Kae12, Kae13, KEC11, MKHC17, RSQ^{+}19, TGS^{+}15, TGKZ19, UACZ^{+}16]. Yield [CBS13, LRHR17, SLR^{+}16]. Yield-per-Recruit [CBS13]. Yields [CBS13]. York [CJMD10, DJS^{+}16, Gur12, JNC^{+}16, MSC13]. Young [KFM12, LSM^{+}16, RSHS15, WMD^{+}19]. Young-of-the-Year [KFM12, LSM^{+}16]. Yukon [CJ14, FBC^{+}10, FCE^{+}12, HD14]. YY [SMH17]. YY-Male [SMH17].

Zealand [DHLD11, HWB11, SM13, SAF16]. Zone [ABBG13, EJL10, KHDA14, RSI16, SS17]. Zones [GH17]. Zooplankton [NSW16].

References

Ashton:2014:CWT

[AAYC14] Neil K. Ashton, Paul J. Anders, Shawn P. Young, and Kenneth D. Cain. Coded wire tag and passive integrated transponder tag implantations in juvenile burbot. North American Journal of Fisheries Management, 34(2):391–400, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Aunins:2013:MMA

[ABBG13] Aaron W. Aunins, Bonnie L. Brown, Matt Balazik, and Greg C. Garman. Migratory movements of American shad in the James River fall zone, Virginia. North American Journal of Fisheries Management, 33(3):569–575, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Amaral:2015:EBS

[ABČ^{+}15] Stephen V. Amaral, Mark S. Bevelhimer, Glenn F. Čada, Daniel J. Giza, Paul T. Jacobson, Brian J. McMahon, and Brenda M. Pracheil. Evaluation of behavior and survival of fish exposed to an axial-flow hydrokinetic turbine. North American Journal of Fisheries Management, 35(1):97–113, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Amberg:2018:ISM

[ABP^{+}18] Jon J. Amberg, Scott A. Bonar, Christina Perez, Christopher B. Rees, Craig Jackson, and Roy M. Ulibarri. The importance of sound methodology in environmental DNA sampling: Response to comment. North American Journal of Fisheries Management, 38(3):597–600, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Allen:2014:BBC

[ABVS14] Michael J. Allen, Shane C. Bush, Ivan Vining, and Michael J. Siepker. Black bass and crappie use of installed habitat structures in Table Rock Lake, Missouri. North American Journal of Fisheries Management, 34(2):223–231, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Al-Chokhachy:2014:QEC

[ACMB^{+}14] Robert Al-Chokhachy, Clint C. Muhlfeld, Matthew C. Boyer, Leslie A. Jones, Amber Steed, and Jeffrey L. Kershner. Quantifying the effectiveness of conservation measures to control the spread of anthropogenic hybridization in stream salmonids: a climate adaptation case study. North American Journal of Fisheries Management, 34(3):642–652, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Al-Chokhachy:2010:RBT

[ACRBB10] Robert Al-Chokhachy, Brett B. Roper, Tracy Bowerman, and Phaedra Budy. A review of bull trout habitat associations and exploratory analyses of patterns across the interior Columbia River Basin. North American Journal of Fisheries Management, 30(2):464–480, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Al-Chokhachy:2019:INB

[ACS19] Robert Al-Chokhachy and Adam J. Sepulveda. Impacts of nonnative brown trout on Yellowstone cutthroat trout in a tributary stream. North American Journal of Fisheries Management, 39(1):17–28, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Archdeacon:2013:CDP

[AD13] Thomas P. Archdeacon and Stephen R. Davenport. Comment: Detection and population estimation for small-bodied fishes in a sand-bed river. North American Journal of Fisheries Management, 33(2):446–452, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2014:GSS

[ADD14] Andrew P. Anderson, Michael R. Denson, and Tanya L. Darden. Genetic structure of striped bass in the Southeastern United States and effects from stock enhancement. North American Journal of Fisheries Management, 34(3):653–667, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Able:2010:EBB

[ADMC10] Kenneth W. Able, Joseph Dobarro, and Angela M. Muzeni-Corino. An evaluation of boat basin dredging effects: Response of fishes and crabs in a New Jersey Estuary. North American Journal of Fisheries Management, 30(4):1001–1015, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Antoni:2011:GVG

[AES11] Luca Antoni, Nicholas Emerick, and Eric Saillant. Genetic variation of gray triggerfish in U.S. waters of the Gulf of Mexico and Western Atlantic Ocean as inferred from mitochondrial DNA sequences. North American Journal of Fisheries Management, 31(4):714–721, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Acheson:2011:MDF

[AG11] James M. Acheson and Roy Gardner. Modeling disaster: The failure of the management of the New England groundfish industry. North American Journal of Fisheries Management, 31(6):1005–1018, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Acre:2015:DPP

[AG15] Matthew R. Acre and Timothy B. Grabowski. Deployment of paired push nets from jet-propelled kayaks to sample ichthyoplankton. North American Journal of Fisheries Management, 35(5):925–929, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Archdeacon:2015:CFC

[AHAC15] Thomas P. Archdeacon, Kjetil R. Henderson, Tristan J. Austring, and Rebecca L. Cook. Comparison of fish communities at random and nonrandom locations in a sand-bed river. North American Journal of Fisheries Management, 35(3):578– 585, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Akins:2015:ERL

[AHS15] Andrea L. Akins, Michael J. Hansen, and Michael J. Seider. Effectiveness of a refuge for lake trout in Western Lake Superior II: Simulation of future performance. North American Journal of Fisheries Management, 35(5):1003–1018, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Antonelis:2011:DCM

[AHVJ11] Kyle Antonelis, Daniel Huppert, Don Velasquez, and Jeffrey June. Dungeness crab mortality due to lost traps and a cost– benefit analysis of trap removal in Washington State waters of the Salish Sea. North American Journal of Fisheries Management, 31(5):880–893, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Alford:2010:ABW

[AJ10] J. Brian Alford and Donald C. Jackson. Associations between watershed characteristics and angling success for sport fishes in Mississippi wadeable streams. North American Journal of Fisheries Management, 30(1):112–120, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Askey:2013:SRO

[AJ13] P. J. Askey and N. T. Johnston. Self-regulation of the Okanagan Lake Kokanee recreational fishery: Dynamic angler effort response to varying fish abundance and productivity. North American Journal of Fisheries Management, 33(5):926–939, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ashton:2019:TMA

[AJR^{+}19] Neil K. Ashton, Nathan R. Jensen, Tyler J. Ross, Shawn P. Young, Ryan S. Hardy, and Kenneth D. Cain. Temperature and maternal age effects on burbot reproduction. North American Journal of Fisheries Management, 39(6):1192–1206, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2012:PGS

[AK12] J. D. Anderson and W. J. Karel. Population genetics of southern flounder with implications for management. North American Journal of Fisheries Management, 32(4):656–662, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Akanni:2010:FTC

[Aka10] K. A. Akanni. Fishing technologies and catch levels among small-scale fishers in Lagos State, Nigeria. North American Journal of Fisheries Management, 30(1):309–315, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Avery:2015:EID

[AKP15] Luke A. Avery, Josh Korman, and William R. Persons. Effects of increased discharge on spawning and age-0 recruitment of rainbow trout in the Colorado River at Lees Ferry, Arizona. North American Journal of Fisheries Management, 35(4):671–680, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Andrews:2017:MSB

[ALCD17] S. N. Andrews, T. Linnansaari, R. A. Curry, and M. J. Dadswell. The misunderstood striped bass of the Saint John River, New Brunswick: Past, present, and future. North American Journal of Fisheries Management, 37(1):235–254, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Arntzen:2017:VBE

[AM17] Evan V. Arntzen and Robert P. Mueller. Video-based electroshocking platform to identify lamprey ammocoete habitats: Field validation and new discoveries in the Columbia River Basin. North American Journal of Fisheries Management, 37 (3):676–681, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Alexiades:2015:MEA

[AMQSK15] Alexander V. Alexiades, Benjamin Marcy-Quay, Patrick J. Sullivan, and Clifford E. Kraft. Measurement error in angler creel surveys. North American Journal of Fisheries Management, 35(2):253–261, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2010:Ea

[Ano10a] Anonymous. Errata. North American Journal of Fisheries Management, 30(3):851–852, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2010:Eb

[Ano10b] Anonymous. Erratum. North American Journal of Fisheries Management, 30(6):1599, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2011:GA

[Ano11] Anonymous. Guide for authors. North American Journal of Fisheries Management, 31(1):182–186, February 2011. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2012:E

[Ano12a] Anonymous. ERRATUM. North American Journal of Fisheries Management, 32(5):1032, October 2012. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2012:EDR

[Ano12b] Anonymous. Erratum: Development of a riverine index netting protocol: Comparisons of net orientation, height, panel order, and line diameter (volume 31: 23–31). North American Journal of Fisheries Management, 32(2):413, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2012:GA

[Ano12c] Anonymous. Guide for authors. North American Journal of Fisheries Management, 32(1):186–190, February 2012. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2013:GA

[Ano13] Anonymous. Guide for authors. North American Journal of Fisheries Management, 33(1):230–235, February 2013. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2014:Ea

[Ano14a] Anonymous. Erratum. North American Journal of Fisheries Management, 34(1):215, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2014:Eb

[Ano14b] Anonymous. Erratum. North American Journal of Fisheries Management, 34(6):1063, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2014:Ec

[Ano14c] Anonymous. Erratum. North American Journal of Fisheries Management, 34(6):1282, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2015:GA

[Ano15] Anonymous. Guide for authors. North American Journal of Fisheries Management, 35(1):178–183, February 2015. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2016:E

[Ano16a] Anonymous. Erratum. North American Journal of Fisheries Management, 36(6):1489, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2016:NAJ

[Ano16b] Anonymous. North American journal of fisheries management. North American Journal of Fisheries Management, 36(1):207– 212, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2017:Ea

[Ano17a] Anonymous. ERRATUM. North American Journal of Fisheries Management, 37(1):255, February 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2017:Eb

[Ano17b] Anonymous. Erratum. North American Journal of Fisheries Management, 37(5):1181, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2018:E

[Ano18a] Anonymous. ERRATUM. North American Journal of Fisheries Management, 38(3):777, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2018:IIa

[Ano18b] Anonymous. Issue information. North American Journal of Fisheries Management, 38(1):1–2, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2018:IIb

[Ano18c] Anonymous. Issue information. North American Journal of Fisheries Management, 38(2):269–270, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2018:IIc

[Ano18d] Anonymous. Issue information. North American Journal of Fisheries Management, 38(3):505–506, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2018:IId

[Ano18e] Anonymous. Issue information. North American Journal of Fisheries Management, 38(4):781–782, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2018:IIe

[Ano18f] Anonymous. Issue information. North American Journal of Fisheries Management, 38(5):977–978, October 2018. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2018:IIf

[Ano18g] Anonymous. Issue information. North American Journal of Fisheries Management, 38(6):1201–1202, December 2018. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2018:RA

[Ano18h] Anonymous. Reviewer acknowledgments. North American Journal of Fisheries Management, 38(6):1423–1425, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2019:E

[Ano19a] Anonymous. ERRATUM. North American Journal of Fisheries Management, 39(6):1426, December 2019. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2019:IIa

[Ano19b] Anonymous. Issue information. North American Journal of Fisheries Management, 39(1):1–2, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2019:IIb

[Ano19c] Anonymous. Issue information. North American Journal of Fisheries Management, 39(2):229–230, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2019:IIc

[Ano19d] Anonymous. Issue information. North American Journal of Fisheries Management, 39(3):413–414, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2019:IId

[Ano19e] Anonymous. Issue information. North American Journal of Fisheries Management, 39(4):605–606, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2019:IIe

[Ano19f] Anonymous. Issue information. North American Journal of Fisheries Management, 39(5):817–818, October 2019. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2019:IIf

[Ano19g] Anonymous. Issue information. North American Journal of Fisheries Management, 39(6):1117–1118, December 2019. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anonymous:2019:NAJ

[Ano19h] Anonymous. North American journal of fisheries management: Guide for authors. North American Journal of Fisheries Management, 39(1):219–225, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2017:EDS

[AOS^{+}17] Joel Anderson, Zachary Olsen, Tom Wagner Glen Sutton, Carey Gelpi, and Darin Topping. Environmental drivers of the spatial and temporal distribution of spawning blue crabs Callinectes sapidus in the Western Gulf of Mexico. North American Journal of Fisheries Management, 37(4):920–934, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Alexiades:2012:MPH

[APAC12] Alexander V. Alexiades, Mary M. Peacock, and Robert Al-Chokhachy. Movement patterns, habitat use, and survival of Lahontan cutthroat trout in the Truckee River. North American Journal of Fisheries Management, 32(5):974–983, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2014:PPS

[APC^{+}14] Joseph H. Anderson, George R. Pess, Richard W. Carmichael, Michael J. Ford, Thomas D. Cooney, Casey M. Baldwin, and Michelle M. McClure. Planning Pacific salmon and steelhead reintroductions aimed at long-term viability and recovery. North American Journal of Fisheries Management, 34(1): 72–93, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Addis:2016:PUA

[APD16] Dustin T. Addis, William F. Patterson III, and Michael A. Dance. The potential for unreported artificial reefs to serve as refuges from fishing mortality for reef fishes. North American Journal of Fisheries Management, 36(1):131–139, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Andree:2018:ECG

[APD^{+}18] Sara R. Andree, Joseph J. Parkos III, Matthew J. Diana, Curtis P. Wagner, and David H. Wahl. Evidence for cogradient growth and survival between two geographically distinct stocks of muskellunge. North American Journal of Fisheries Management, 38(4):922–929, August 2018. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ackerman:2019:PEP

[APD^{+}19] Nicklaus K. Ackerman, Brian J. Pyper, Margaret M. David, Garth J. Wyatt, Dan P. Cramer, and Timothy M. Shibahara. Passage effectiveness at a pool-and-weir fishway designed to accommodate Pacific lampreys. North American Journal of Fisheries Management, 39(3):426–440, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Andrews:2018:CRE

[APF18] William R. Andrews, Katherine J. Papacostas, and John Foster. A comparison of recall error in recreational fisheries surveys with one- and two-month reference periods. North American Journal of Fisheries Management, 38(6):1284–1298, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Askey:2013:LFA

[APP13] Paul J. Askey, Eric A. Parkinson, and John R. Post. Linking fish and angler dynamics to assess stocking strategies for hatchery-dependent, open-access recreational fisheries. North American Journal of Fisheries Management, 33(3):557–568, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Adams:2014:PSB

[APPR14] Noah S. Adams, John M. Plumb, Russell W. Perry, and Dennis W. Rondorf. Performance of a surface bypass structure to enhance juvenile steelhead passage and survival at Lower Granite Dam, Washington. North American Journal of Fisheries Management, 34(3):576–594, June 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Archdeacon:2012:OHR

[AR12] Thomas P. Archdeacon and W. Jason Remshardt. Observations of hatchery-reared Rio Grande silvery minnow using a fishway. North American Journal of Fisheries Management, 32(4):648–655, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ahrenholz:2010:EHP

[AS10] Dean W. Ahrenholz and Joseph W. Smith. Effect of hang-in percentage on catch rates of flounder in the North Carolina inshore gill-net fishery. North American Journal of Fisheries Management, 30(6):1401–1407, December 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Arendt:2018:BHW

[ASD^{+}18] Michael D. Arendt, Jeffrey A. Schwenter, Julie Dingle, Christopher A. Evans, Ellen Waldrop, Brooke Czwartacki, Amy E. Fowler, and J. David Whitaker. A “BRD” in the hand worthy of four in the trap: Validation of optimal bycatch reduction device (BRD) size to maximize blue crab Callinectes sapidus entry and diamondback terrapin Malaclemys terrapin exclusion through theoretical modeling and application. North American Journal of Fisheries Management, 38 (2):411–423, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Askey:2016:MDF

[Ask16] P. J. Askey. Managing dynamic fisheries with static regulations: an assessment of size-graded bag limits for recreational Kokanee fisheries. North American Journal of Fisheries Management, 36(2):241–253, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Andrews:2013:JMA

[ASL^{+}13] Tessa M. Andrews, Bradley B. Shepard, Andrea R. Litt, Carter G. Kruse, Alexander V. Zale, and Steven T. Kalinowski. Juvenile movement among different populations of cutthroat trout introduced as embryos to vacant habitat. North American Journal of Fisheries Management, 33(4):

795–805, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Andrews:2016:PJC

[ASL^{+}16] Tessa C. Andrews, Bradley B. Shepard, Andrea R. Litt, Carter G. Kruse, M. Lee Nelson, Patrick Clancey, Alexander V. Zale, Mark L. Taper, and Steven T. Kalinowski. Performance of juvenile cutthroat trout translocated as embryos from five populations into a common habitat. North American Journal of Fisheries Management, 36(4):926–941, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2017:QEA

[ASP^{+}17] Julie Anderson, Garry Scrimgeour, Vince Palace, Michael Suitor, and John Wilcockson. Quantifying elements in Arctic grayling and bull trout in the South Nahanni River Watershed, Northwest Territories, using nonlethal tissue samples. North American Journal of Fisheries Management, 37 (1):50–63, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Agar:2017:CTF

[ASS17] J. Agar, M. Shivlani, and D. Solís. The commercial trap fishery in the Commonwealth of Puerto Rico: an economic, social, and technological profile. North American Journal of Fisheries Management, 37(4):778–788, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Anderson:2018:JLH

[AT18] Joseph H. Anderson and Peter C. Topping. Juvenile life history diversity and freshwater productivity of Chinook salmon in the Green River, Washington. North American Journal of Fisheries Management, 38(1):180–193, February 2018. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Addison:2010:MMB

[AW10] Peter A. Addison and Chris C. Wilson. Matching management to biological scale: Connectivity among Lacustrine brook trout populations. North American Journal of Fisheries Management, 30(5):1132–1142, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Askey:2018:AEE

[AWG^{+}18] Paul J. Askey, Hillary Ward, Theresa Godin, Marcus Boucher, and Sara Northrup. Angler effort estimates from instantaneous aerial counts: Use of high-frequency time-lapse camera data to inform model-based estimators. North American Journal of Fisheries Management, 38(1):194–209, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Andrews:2018:DSB

[AZE^{+}18] S. N. Andrews, K. Zelman, T. Ellis, T. Linnansaari, and R. A. Curry. Diet of striped bass and muskellunge downstream of a large hydroelectric dam: a preliminary investigation into suspected Atlantic salmon smolt predation. North American Journal of Fisheries Management, 38(3):734–746, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boucek:2011:CRS

[BA11] Ross E. Boucek and Aaron J. Adams. Comparison of retention success for multiple tag types in common snook. North American Journal of Fisheries Management, 31(4):693–699, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barnett:2018:CTC

[BA18] Zanethia C. Barnett and Susan B. Adams. Comparison of two crayfish trapping methods in coastal plain seasonal wetlands. North American Journal of Fisheries Management, 38 (4):911–921, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bennett:2014:AVS

[BACRB14] Stephen Bennett, Robert Al-Chokhachy, Brett B. Roper, and Phaedra Budy. Annual variation of spawning cutthroat trout in a small Western USA stream: a case study with implications for the conservation of potamodromous trout life history diversity. North American Journal of Fisheries Management, 34(5):1033–1046, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Balazik:2015:CBI

[Bal15] Matthew T. Balazik. Capture and brief invasive procedures using electronarcosis does not appear to affect postrelease habits in male Atlantic sturgeon during the spawning season. North American Journal of Fisheries Management, 35(2):398–402, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bergstedt:2016:SDB

[BATK16] Roger A. Bergstedt, Ray L. Argyle, William W. Taylor, and Charles C. Krueger. Seasonal and diel bathythermal distributions of lake whitefish in Lake Huron: Potential implications for lake trout bycatch in commercial fisheries. North American Journal of Fisheries Management, 36(4):705–719, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bowerman:2012:IMP

[BB12] Tracy Bowerman and Phaedra Budy. Incorporating movement patterns to improve survival estimates for juvenile bull trout. North American Journal of Fisheries Management, 32 (6):1123–1136, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baker:2019:UMM

[BB19] Joy Price Baker and Scott A. Bonar. Using a mechanistic model to develop management strategies to cool Apache trout streams under the threat of climate change. North American Journal of Fisheries Management, 39(5):849–867, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beauchene:2014:STT

[BBB^{+}14] Mike Beauchene, Mary Becker, Christopher J. Bellucci, Neal Hagstrom, and Yoichiro Kanno. Summer thermal thresholds of fish community transitions in Connecticut streams. North American Journal of Fisheries Management, 34(1):119–131, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2010:PSW

[BBC10] Timothy F. Bonvechio, Kimberly I. Bonvechio, and Richard L. Cailteux. Proposed standard weight (W_{s}) equation and standard length categories for suwannee bass. North American Journal of Fisheries Management, 30(4):983–988, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2011:EPS

[BBC11] Timothy F. Bonvechio, Kimberly I. Bonvechio, and Richard L. Cailteux. Erratum: Proposed standard weight (W_{s}) equation and standard length categories for Suwannee bass (volume 30:983–988). North American Journal of Fisheries Management, 31(2):417, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Busch:2013:PCV

[BBC^{+}13] D. Shallin Busch, David A. Boughton, Thomas Cooney, Peter Lawson, Steven T. Lindley, Michelle McClure, Mary H. Ruckelshaus, Norma Jean Sands, Brian C. Spence, Thomas C. Wainwright, Thomas H. Williams, and Paul McElhany. A practical comparison of viability models used for management of endangered and threatened anadromous Pacific salmonids. North American Journal of Fisheries Management, 33(6): 1125–1141, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Benjamin:2019:DAR

[BBD19] Joseph R. Benjamin, William R. Brignon, and Jason B. Dunham. Decision analysis for the reintroduction of bull trout into the Lower Pend Oreille River, Washington. North American Journal of Fisheries Management, 39(5):1026–1045, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baril:2019:AUR

[BBG19] André-Marcel Baril, Pascale M. Biron, and James W. A. Grant. An assessment of an unsuccessful restoration project for lake sturgeon using three-dimensional numerical modelling. North American Journal of Fisheries Management, 39(1):69– 81, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brenden:2011:PDL

[BBL^{+}11] Travis O. Brenden, James R. Bence, Brian F. Lantry, Jana R. Lantry, and Ted Schaner. Population dynamics of Lake Ontario lake trout during 1985–2007. North American Journal of Fisheries Management, 31(5):962–979, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bingham:2016:IML

[BBO16] Daniel M. Bingham, Preston Buckskin, and Hunter Osborne. Isolation, migration, and local recruitment drive persistence of cutthroat trout in tributaries near American Falls Reservoir. North American Journal of Fisheries Management, 36(3):590– 605, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buchanan:2018:SJF

[BBS18] Rebecca A. Buchanan, Patricia L. Brandes, and John R. Skalski. Survival of juvenile fall-run Chinook salmon through the San Joaquin River Delta, California, 2010–2015. North American Journal of Fisheries Management, 38(3):663–679, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2019:PSW

[BBS19] Timothy F. Bonvechio, Kimberly I. Bonvechio, and Steven M. Sammons. Proposed standard weight (W_{s}) equation and standard length categories for shoal bass. North American Journal of Fisheries Management, 39(5):933–938, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blankman:2018:EEP

[BBW18] Trevor Blankman, Emily E. Ball, and Michael J. Weber. Evaluating the efficiency of pulsed gastric lavage for age-0 walleyes. North American Journal of Fisheries Management, 38(4):953–958, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beckman:2014:TFA

[BC14] Daniel W. Beckman and John W. Calfee. Timing of first annulus formation in white sucker otoliths. North American Journal of Fisheries Management, 34(6):1187–1189, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brumbaugh:2017:EBN

[BC17] Steven M. Brumbaugh and Ronald Coleman. Evaluation of Bayesian networks for predicting spawning habitat quality of Chinook salmon in data-poor watersheds. North American Journal of Fisheries Management, 37(1):176–188, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boenish:2018:QSA

[BC18] Robert Boenish and Yong Chen. Quasi-stationary Atlantic cod bycatch estimation in the Maine American lobster Homarus americanus trap fishery. North American Journal of Fisheries Management, 38(1):3–17, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bahr:2018:RRS

[BCBW18] Derek L. Bahr, Andrew J. Carlson, Owen E. Baird, and Matthew C. Ward. Recapture rates and size selectivity of muskellunge by anglers within two Minnesota fisheries. North American Journal of Fisheries Management, 38(1):152–158, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beecher:2010:EAP

[BCD^{+}10] Hal A. Beecher, Brad A. Caldwell, S. Brett DeMond, Dave Seiler, and Steven N. Boessow. An empirical assessment of PHABSIM using long-term monitoring of Coho salmon smolt production in Bingham Creek, Washington. North American Journal of Fisheries Management, 30(6):1529–1543, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Budnik:2014:APR

[BCMB14] Richard R. Budnik, Mike Clancy, Jeffrey G. Miner, and William D. Brown. Assessment of paddlefish reintroduction into Allegheny Reservoir. North American Journal of Fisheries Management, 34(6):1055–1062, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beacham:2014:PSRa

[BCRM^{+}14a] Terry D. Beacham, Steven Cox-Rogers, Cathy MacConnachie, Brenda McIntosh, and Colin G. Wallace. Population structure and run timing of sockeye salmon in the Skeena River, British Columbia. North American Journal of Fisheries Management, 34(2):335–348, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beacham:2014:PSRb

[BCRM^{+}14b] Terry D. Beacham, Steven Cox-Rogers, Cathy MacConnachie, Brenda McIntosh, and Colin G. Wallace. Population structure and run timing of sockeye salmon in the Skeena River, British Columbia: Response to comment. North American Journal of Fisheries Management, 34(6):1171–1176, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Begley:2017:CAS

[BCZ17] Meg Begley, Stephen M. Coghlan, Jr., and Joseph Zydlewski. A comparison of age, size, and fecundity of harvested and reference white sucker populations. North American Journal of Fisheries Management, 37(3):510–523, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bauerlien:2018:PCS

[BCZ^{+}18] Cory J. Bauerlien, Marinda R. Cornett, Emily A. Zielonka, Derek P. Crane, and James S. Bulak. Precision of calcified structures used for estimating age of chain pickerel. North American Journal of Fisheries Management, 38(4):930–939, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Briggs:2019:OEW

[BDB^{+}19] Andrew S. Briggs, Jan C. Dean, James C. Boase, Patrick M. Kocovsky, and James A. Luoma. Optimum electrofishing waveforms and parameters to induce immobilization of juvenile grass carp. North American Journal of Fisheries Management, 39(4):705–713, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boone:2013:PPE

[BDC^{+}13] S. Shaun Boone, Stephen J. Divers, Alvin C. Camus, Douglas L. Peterson, Cecil A. Jennings, James L. Shelton, and Sonia M. Hernandez. Pathologic and physiologic effects associated with long-term intracoelomic transmitters in captive Siberian sturgeon. North American Journal of Fisheries Management, 33(5):869–877, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bergmann:2014:EAC

[BDH^{+}14] Charles Bergmann, William B. Driggers III, Eric R. Hoffmayer, Matthew D. Campbell, and Gilmore Pellegrin. Effects of appendaged circle hook use on catch rates and deep hooking of black sea bass in a recreational fishery. North American Journal of Fisheries Management, 34(6):1199–1203, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Binion:2011:ELA

[BDS^{+}11] Greg R. Binion, Daniel J. Daugherty, J. Warren Schlechte, Richard A. Ott, Jr., and Timothy J. Bister. Efficacy of a light attractant for increasing trap net catches of white crappies. North American Journal of Fisheries Management, 31(3):455– 460, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bodine:2015:SIG

[BDSB15] Kristopher A. Bodine, Daniel J. Daugherty, J. Warren Schlechte, and Greg R. Binion. A strategy for increasing gillnet catch rates and minimizing sampling mortality of alligator gars. North American Journal of Fisheries Management, 35 (3):611–615, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bowersox:2016:DPH

[BDT^{+}16] Brett J. Bowersox, Joseph M. DuPont, Russell Tucker, Larry Barrett, and James A. Lamansky, Jr. Determining the presence of hooks inside white sturgeon using metal detector and portable X-ray technology. North American Journal of Fisheries Management, 36(5):1045–1052, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brewer:2011:ETO

[BE11] Shannon K. Brewer and Mark R. Ellersieck. Evaluating two observational sampling techniques for determining the distribution and detection probability of age-0 smallmouth bass in clear, warmwater streams. North American Journal of Fisheries Management, 31(5):894–904, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beaumont:2017:CRE

[Bea17] W. R. C. Beaumont. Comment: Rates and effects of branding due to electroshock: Some additional perspectives. North American Journal of Fisheries Management, 37(2):429–430, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brooke:2019:ILW

[BEH^{+}19] Thomas C. Brooke, Connor W. Elliott, Jeremy Holden, Yuxiang Wang, Rachael L. Hornsby, and Bruce L. Tufts. The importance of live-well transport in the physiological disturbance experienced by smallmouth bass in tournaments on large water bodies. North American Journal of Fisheries Management, 39(6):1260–1268, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:2018:AST

[BEK^{+}18] Brian G. Blackwell, Mark J. Ermer, Todd M. Kaufman, Tyrel S. Moos, Steven J. Kennedy, and Ryan J. Braun. Assessment of short-term mortality of wild-caught female walleye broodstock. North American Journal of Fisheries Management, 38(4):803–810, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bodine:2013:EAT

[BF13] Kristopher A. Bodine and B. Paul Fleming. Evaluation of an alternative technique for attaching external transmitters to blue catfish. North American Journal of Fisheries Management, 33(5):950–955, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bodine:2014:RPB

[BF14] Kristopher A. Bodine and Paul Fleming. Retention of PIT and T-bar anchor tags in blue catfish. North American Journal of Fisheries Management, 34(1):68–71, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Braaten:2010:SDD

[BFL^{+}10] Patrick J. Braaten, David B. Fuller, Ryan D. Lott, Michael P. Ruggles, and Robert J. Holm. Spatial distribution of drifting pallid sturgeon larvae in the Missouri River inferred from two net designs and multiple sampling locations. North American Journal of Fisheries Management, 30(4):1062–1074, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Becker:2017:DCM

[BFLT17] Alistair Becker, D. Stewart Fielder, Michael B. Lowry, and Matthew D. Taylor. Development of a calcein marking technique for juvenile mulloway Argyrosomus japonicus to be used in stock enhancement programs. North American Journal of Fisheries Management, 37(1):207–210, February 2017. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bies:2016:CEG

[BFN16] Jason M. Bies, Cynthia N. Fox, and J. Wesley Neal. Comparison of electrofishing and gill nets for sampling cichlid species. North American Journal of Fisheries Management, 36(5):975– 981, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bahn:2012:BSS

[BFP12] Robert A. Bahn, Joel E. Fleming, and Douglas L. Peterson. Bycatch of shortnose sturgeon in the commercial American shad fishery of the Altamaha River, Georgia. North American Journal of Fisheries Management, 32(3):557–562, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brandenburg:2019:DGC

[BFS^{+}19] W. Howard Brandenburg, Travis A. Francis, Darrel E. Snyder, Kevin R. Bestgen, Brian A. Hines, Wade D. Wilson, Sandra Bohn, Alexis S. Harrison, and Stephani L. Clark Barkalow. Discovery of grass carp larvae in the Colorado River arm of Lake Powell. North American Journal of Fisheries Management, 39(1):166–171, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bowlby:2012:IAS

[BG12] Heather D. Bowlby and A. Jamie F. Gibson. Inferring adult status and trends from juvenile density data for Atlantic salmon. North American Journal of Fisheries Management, 32(6):1225–1236, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Briggs:2013:FPT

[BG13] Andrew S. Briggs and Tracy L. Galarowicz. Fish passage through culverts in Central Michigan warmwater streams. North American Journal of Fisheries Management, 33(3):652– 664, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bowlby:2019:LLU

[BG19] Heather D. Bowlby and A. Jamie F. Gibson. Linking land use to Atlantic salmon production to guide recovery planning. North American Journal of Fisheries Management, 39 (4):664–675, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bouska:2017:RET

[BGBG17] Wesley W. Bouska, David C. Glover, Kristen L. Bouska, and James E. Garvey. A refined electrofishing technique for collecting silver carp: Implications for management. North American Journal of Fisheries Management, 37(1):101–107, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barton:2011:CEF

[BGCF11] Nathan T. Barton, Tracy L. Galarowicz, Randall M. Claramunt, and John D. Fitzsimons. A comparison of egg funnel and egg bag estimates of egg deposition in Grand Traverse Bay, Lake Michigan. North American Journal of Fisheries Management, 31(3):580–587, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bell:2017:DES

[BGCH17] Richard J. Bell, Brian Gervelis, Glenn Chamberlain, and John Hoey. Discard estimates from self-reported catch data: an example from the U.S. northeast shelf. North American Journal of Fisheries Management, 37(5):1130–1144, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barber:2011:EHF

[BGCS11] Jarrett J. Barber, Kenneth G. Gerow, Patrick J. Connolly, and Sarabdeep Singh. Estimating the hatchery fraction of a natural population: a Bayesian approach. North American Journal of Fisheries Management, 31(5):934–942, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackburn:2019:PDE

[BGD^{+}19] Shannon E. Blackburn, Marty L. Gingras, Jason DuBois, Zachary J. Jackson, and Michael C. Quist. Population dynamics and evaluation of management scenarios for white sturgeon in the Sacramento–San Joaquin River Basin. North American Journal of Fisheries Management, 39(5):896–912, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boyd:2010:ECR

[BGHL10] James W. Boyd, Christopher S. Guy, Travis B. Horton, and Stephen A. Leathe. Effects of catch-and-release angling on salmonids at elevated water temperatures. North American Journal of Fisheries Management, 30(4):898–907, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:2017:CTM

[BGM17] Peter J. Brown, Christopher S. Guy, and Michael H. Meeuwig. A comparison of two mobile electrode arrays for increasing mortality of lake trout embryos. North American Journal of Fisheries Management, 37(2):363–369, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bromaghin:2010:LFJ

[BGP10] Jeffrey F. Bromaghin, Kenneth S. Gates, and Douglas E. Palmer. A likelihood framework for joint estimation of salmon abundance and migratory timing using telemetric mark — recapture. North American Journal of Fisheries Management, 30(6):1385–1394, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bower:2011:APV

[BGW^{+}11] Michael R. Bower, Daniel B. Gaines, Kevin P. Wilson, John G. Wullschleger, Maria C. Dzul, Michael C. Quist, and Stephen J. Dinsmore. Accuracy and precision of visual estimates and photogrammetric measurements of the length of a small-bodied fish. North American Journal of Fisheries Management, 31 (1):138–143, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boyer:2017:GCS

[BGW^{+}17] Jan K. Boyer, Christopher S. Guy, Molly A. H. Webb, Travis B. Horton, and Thomas E. McMahon. Gear comparison for sampling age-0 mountain whitefish in the Madison River, Montana. North American Journal of Fisheries Management, 37(1):189–195, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:2010:EMT

[BHC^{+}10] Richard S. Brown, Ryan A. Harnish, Kathleen M. Carter, James W. Boyd, Katherine A. Deters, and M. Brad Eppard. An evaluation of the maximum tag burden for implantation of acoustic transmitters in juvenile Chinook salmon. North American Journal of Fisheries Management, 30(2):499–505, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buktenica:2013:ENB

[BHG^{+}13] Mark W. Buktenica, David K. Hering, Scott F. Girdner, Brian D. Mahoney, and Bruce D. Rosenlund. Eradication of nonnative brook trout with electrofishing and antimycin-A and the response of a remnant bull trout population. North American Journal of Fisheries Management, 33(1):117–129, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beardmore:2011:ITC

[BHHA11] Ben Beardmore, Wolfgang Haider, Len M. Hunt, and Robert Arlinghaus. The importance of trip context for determining primary angler motivations: Are more specialized anglers more catch-oriented than previously believed? North American Journal of Fisheries Management, 31(5):861–879, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bergman:2011:RBE

[BHN11] Paul S. Bergman, Michael J. Hansen, and Nancy A. Nate. Relationship between electrofishing catch rate and adult trout abundance in Wisconsin streams. North American Journal of Fisheries Management, 31(5):952–961, October 2011. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burdick:2015:ELS

[BHR^{+}15] Summer M. Burdick, David A. Hewitt, Josh E. Rasmussen, Brian S. Hayes, Eric C. Janney, and Alta C. Harris. Effects of lake surface elevation on shoreline-spawning Lost River suckers. North American Journal of Fisheries Management, 35 (3):478–490, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Breton:2013:CLF

[BHW13] André R. Breton, John A. Hawkins, and Dana L. Winkelman. Correcting length–frequency distributions for imperfect detection. North American Journal of Fisheries Management, 33(6):1156–1165, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bunt:2019:RTM

[BJ19] Christopher Bunt and Bailey Jacobson. Rainbow trout migration and use of a nature-like fishway at a Great Lakes tributary. North American Journal of Fisheries Management, 39 (3):460–467, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:2011:ERT

[BJZ11] Peter J. Brown, Heather Johnson, and Alexander V. Zale. Effect of rainbow trout size on response to rotenone and antimycin. North American Journal of Fisheries Management, 31(6):1146–1152, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:2012:TYP

[BK12] Brian G. Blackwell and Todd M. Kaufman. Timing of yellow perch otolith annulus formation and relationship between fish and otolith lengths. North American Journal of Fisheries Management, 32(2):239–248, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baumhoer:2016:SAA

[BK16] Brandon M. Baumhoer and Anita M. Kelly. Survival, age-0 abundance, and growth of black crappie and hybrid crappie in 0.1-ha earthen ponds. North American Journal of Fisheries Management, 36(3):447–451, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bigbee:2010:SFI

[BKD^{+}10] Daniel L. Bigbee, Ronald G. King, Kent M. Dixon, Douglas A. Dixon, and Elgin S. Perry. Survival of fish impinged on a rotary disk screen. North American Journal of Fisheries Management, 30(6):1420–1433, December 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bohn:2017:NHL

[BKDB17] Sandra Bohn, Brian R. Kreiser, Daniel J. Daugherty, and Kristopher A. Bodine. Natural hybridization of lepisosteids: Implications for managing the alligator gar. North American Journal of Fisheries Management, 37(2):405–413, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bingham:2014:LGI

[BKHS14] Daniel M. Bingham, Benjamen M. Kennedy, Kyle C. Hanson, and Christian T. Smith. Loss of genetic integrity in hatchery steelhead produced by juvenile-based broodstock and wild integration: Conflicts in production and conservation goals. North American Journal of Fisheries Management, 34(3):609– 620, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:2016:ACS

[BKM16] Brian G. Blackwell, Todd M. Kaufman, and Tyrel S. Moos. An assessment of calcified structures for estimating northern pike ages. North American Journal of Fisheries Management, 36(4):964–974, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:2019:EAS

[BKM19] Brian G. Blackwell, Todd M. Kaufman, and Tyrel S. Moos. Evaluation of anal spines, dorsal spines, and scales as potential nonlethal surrogates to otoliths for estimating ages of largemouth bass and smallmouth bass. North American Journal of Fisheries Management, 39(3):596–603, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Blackwell:2017:CDN

[BKME17] Brian G. Blackwell, Todd M. Kaufman, Tyrel S. Moos, and Mark J. Ermer. Comparison of day and night electrofishing to sample smallmouth bass in natural lakes of Eastern South Dakota. North American Journal of Fisheries Management, 37(6):1191–1198, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buehrens:2014:IJC

[BKP^{+}14] T. W. Buehrens, P. Kiffney, G. R. Pess, T. R. Bennett, S. M. Naman, G. Brooks, and T. P. Quinn. Increasing juvenile Coho salmon densities during early recolonization have not affected resident coastal cutthroat trout growth, movement, or survival. North American Journal of Fisheries Management, 34 (5):892–907, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Billman:2012:ERC

[BKSH^{+}12] Hilary G. Billman, Carter G. Kruse, Sophie St-Hilaire, Todd M. Koel, Jeffrey L. Arnold, and Charles R. Peterson. Effects of rotenone on Columbia spotted frogs Rana luteiventris during field applications in lentic habitats of Southwestern Montana. North American Journal of Fisheries Management, 32(4):781–789, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bunnell:2010:BID

[BLM10] David B. Bunnell, Douglas W. Lipton, and Thomas J. Miller. The bioeconomic impact of different management regulations on the Chesapeake Bay blue crab fishery. North American Journal of Fisheries Management, 30(6):1505–1521, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bingham:2011:ABM

[BLM^{+}11] Matthew F. Bingham, Zhimin Li, Kristy E. Mathews, Colleen M. Spagnardi, Jennifer S. Whaley, Sara G. Veale, and Jason C. Kinnell. An application of behavioral modeling to characterize urban angling decisions and values. North American Journal of Fisheries Management, 31(2):257–268, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bloom:2013:CEB

[Blo13] Roger K. Bloom. Capture efficiency of barbed versus barbless artificial flies for trout. North American Journal of Fisheries Management, 33(3):493–498, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bieber:2019:IIA

[BLSS19] John F. Bieber, Michael J. Louison, Jeffrey A. Stein, and Cory D. Suski. Impact of ice-angling and handling on swimming performance in bluegill and largemouth bass. North American Journal of Fisheries Management, 39(6):1301–1310, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bisack:2014:MEV

[BM14] Kathryn D. Bisack and Gisele Magnusson. Measuring the economic value of increased precision in scientific estimates of marine mammal abundance and bycatch: Harbor porpoise Phocoena phocoena in the Northeast U.S. gill-net fishery. North American Journal of Fisheries Management, 34(2):311–321, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bacon:2012:FLM

[BMM^{+}12] P. J. Bacon, I. A. Malcolm, J. C. MacLean, I. McLaren, and A. Thorne. Female length and mass quantify the difference between Atlantic salmon ova fecundity and ova deposition. North American Journal of Fisheries Management, 32(3):491– 498, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brule:2015:ICH

[BMMMLML15] Thierry Brulé, Jorge Montero-Muñoz, Natalia Morales-López, and Andy Mena-Loria. Influence of circle hook size on catch rate and size of red grouper in shallow waters of the Southern Gulf of Mexico. North American Journal of Fisheries Management, 35(6):1196–1208, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Branigan:2019:APA

[BMW^{+}19] Philip R. Branigan, Kevin A. Meyer, Nicholas C. Wahl, Matthew P. Corsi, and Andrew M. Dux. Accuracy and precision of age estimates obtained from three calcified structures for known-age Kokanee. North American Journal of Fisheries Management, 39(3):498–508, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baumann:2016:EEA

[BOM16] Jessica R. Baumann, N. Corey Oakley, and Brian J. McRae. Evaluating the effectiveness of artificial fish habitat designs in turbid reservoirs using sonar imagery. North American Journal of Fisheries Management, 36(6):1437–1444, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Boughton:2010:ESS

[Bou10] David A. Boughton. Estimating the size of steelhead runs by tagging juveniles and monitoring migrants. North American Journal of Fisheries Management, 30(1):89–101, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Black:2014:LES

[BP14] Jonathan L. Black and Elgin S. Perry. Laboratory evaluation of the survival of fish impinged on modified traveling water screens. North American Journal of Fisheries Management, 34(2):359–372, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonar:2015:GBT

[BP15] Scott A. Bonar and Sally J. Petre. Ground-based thermal imaging of stream surface temperatures: Technique and evaluation. North American Journal of Fisheries Management, 35(6):1209–1218, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barnett:2013:IRE

[BPB13] Heidy K. Barnett, Dwayne K. Paige, and William C. Belknap. Impact of reservoir elevation during the spawning season on the distribution of bull trout redds. North American Journal of Fisheries Management, 33(5):917–925, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bakevich:2013:HFS

[BPQ13] Bryan D. Bakevich, Clay L. Pierce, and Michael C. Quist. Habitat, fish species, and fish assemblage associations of the Topeka shiner in West-Central Iowa. North American Journal of Fisheries Management, 33(6):1258–1268, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brenkman:2019:RRL

[BPT^{+}19] Samuel J. Brenkman, Roger J. Peters, Roger A. Tabor, Joshua J. Geffre, and Kathryn T. Sutton. Rapid recolonization and life history responses of bull trout following dam removal in Washington’s Elwha River. North American Journal of Fisheries Management, 39(3):560–573, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beard:2017:SMD

[BQHR17] Zachary S. Beard, Michael C. Quist, Ryan S. Hardy, and Tyler J. Ross. Survival, movement, and distribution of juvenile burbot in a tributary of the Kootenai River. North American Journal of Fisheries Management, 37(6):1274–1288, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brauer:2019:PCM

[BQR^{+}19] Tucker A. Brauer, Michael C. Quist, Darren T. Rhea, Troy W. Laughlin, and John D. Walrath. Population characteristics and management of lentic populations of nonnative burbot in the Green River System, Wyoming. North American Journal of Fisheries Management, 39(1):45–57, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buckmeier:2012:RPI

[BR12] David L. Buckmeier and Kerry S. Reeves. Retention of passive integrated transponder, T-bar anchor, and coded wire tags in lepisosteids. North American Journal of Fisheries Management, 32(3):573–576, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bangley:2014:FHD

[BR14] Charles W. Bangley and Roger A. Rulifson. Feeding habits, daily ration, and potential predatory impact of mature female spiny dogfish in North Carolina coastal waters. North American Journal of Fisheries Management, 34(3):668–677, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bradley:2018:MEV

[BRA18a] Caitlin E. Bradley, James A. Rice, and D. Derek Aday. Modeling the effects of vital rate manipulation and management scenarios to predict the population impact of restoration programs on an unrecovered coastal population of striped bass. North American Journal of Fisheries Management, 38(3):639– 649, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bradley:2018:JAS

[BRA^{+}18b] Caitlin E. Bradley, James A. Rice, D. Derek Aday, Joseph E. Hightower, Jason Rock, and Kelsey L. Lincoln. Juvenile and adult striped bass mortality and distribution in an unrecovered coastal population. North American Journal of Fisheries Management, 38(1):104–119, February 2018. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Baker:2017:EMH

[BRB^{+}17] Cindy F. Baker, Kathryn Reeve, Dylan Baars, Don Jellyman, and Paul Franklin. Efficacy of 12-mm half-duplex passive integrated transponder tags in monitoring fish movements through stationary antenna systems. North American Journal of Fisheries Management, 37(6):1289–1298, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brahney:2010:CPT

[BRBP10] Janice Brahney, Richard Routledge, Darren G. Bos, and Marlow G. Pellatt. Changes to the productivity and trophic structure of a sockeye salmon rearing lake in British Columbia. North American Journal of Fisheries Management, 30(2):433– 444, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brewer:2013:CUU

[Bre13] Shannon K. Brewer. Channel unit use by smallmouth bass: Do land-use constraints or quantity of habitat matter? North American Journal of Fisheries Management, 33(2):351–358, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brignon:2017:STS

[Bri17] William R. Brignon. Spatial and temporal segregation of wild and hatchery winter steelhead populations in Eagle Creek, Oregon. North American Journal of Fisheries Management, 37(3):667–675, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bair:2016:EVA

[BRN16] Lucas S. Bair, David L. Rogowski, and Chris Neher. Economic value of angling on the Colorado River at Lees Ferry: Using secondary data to estimate the influence of seasonality. North American Journal of Fisheries Management, 36(6):1229–1239, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Balazik:2012:PVI

[BRS^{+}12] Matthew T. Balazik, Kevin J. Reine, Albert J. Spells, Charles A. Fredrickson, Michael L. Fine, Greg C. Garman, and Stephen P. McIninch. The potential for vessel interactions with adult Atlantic sturgeon in the James River, Virginia. North American Journal of Fisheries Management, 32 (6):1062–1069, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brown:2015:DPC

[BRSA15] Daniel T. Brown, James A. Rice, Cory D. Suski, and D. Derek Aday. Dispersal patterns of coastal largemouth bass in response to tournament displacement. North American Journal of Fisheries Management, 35(3):431–439, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brewer:2016:STR

[BRSB^{+}16] Melissa A. Brewer, Paul J. Rudershausen, Blair D. Sterba-Boatwright, Jeffery H. Merrell, and Jeffrey A. Buckel. Survival, tag retention, and growth of spot and mummichog following PIT tag implantation. North American Journal of Fisheries Management, 36(3):639–651, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brauer:2018:EUD

[BRWQ18] Tucker A. Brauer, Darren T. Rhea, John D. Walrath, and Michael C. Quist. Efficacy of using data from angler-caught burbot to estimate population rate functions. North American Journal of Fisheries Management, 38(2):346–354, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bodine:2010:CEB

[BS10] Kristopher A. Bodine and Daniel E. Shoup. Capture efficiency of blue catfish electrofishing and the effects of temperature, habitat, and reservoir location on electrofishing-derived length structure indices and relative abundance. North American Journal of Fisheries Management, 30(2):613–621, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barnett:2011:FSS

[BS11a] Lewis A. K. Barnett and Brian C. Spence. Freshwater survival of stranded steelhead kelts in coastal Central California streams. North American Journal of Fisheries Management, 31(4):757–764, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bowlby:2011:VIS

[BS11b] James N. Bowlby and Paul J. Savoie. Verifying identification of salmon and trout by boat anglers in Lake Ontario. North American Journal of Fisheries Management, 31(3):468–473, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bajer:2012:UBE

[BS12] Przemyslaw G. Bajer and Peter W. Sorensen. Using boat electrofishing to estimate the abundance of invasive common carp in small Midwestern lakes. North American Journal of Fisheries Management, 32(5):817–822, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brenden:2015:CLE

[BSB^{+}15] Travis O. Brenden, Kim T. Scribner, James R. Bence, Iyob Tsehaye, Jeannette Kanefsky, Christopher S. Vandergoot, and David G. Fielder. Contributions of Lake Erie and Lake St. Clair walleye populations to the Saginaw Bay, Lake Huron, recreational fishery: Evidence from genetic stock identification. North American Journal of Fisheries Management, 35 (3):567–577, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bonvechio:2014:UMF

[BSBC14] K. I. Bonvechio, R. E. Sawyers, R. Bitz, and S. Crawford. Use of mini-fyke nets for sampling shallow-water fish communities in Florida lakes. North American Journal of Fisheries Management, 34(4):693–701, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buchanan:2013:RUS

[BSBF13] Rebecca A. Buchanan, John R. Skalski, Patricia L. Brandes, and Andrea Fuller. Route use and survival of juvenile Chinook salmon through the San Joaquin River Delta. North American Journal of Fisheries Management, 33(1):216–229, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buchanan:2010:ESH

[BSG10] Rebecca A. Buchanan, John R. Skalski, and Albert E. Giorgi. Evaluating surrogacy of hatchery releases for the performance of wild yearling Chinook salmon from the Snake River Basin. North American Journal of Fisheries Management, 30(5): 1258–1269, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bettoli:2010:ASU

[BSH10] Phillip W. Bettoli, George D. Scholten, and Don W. Hubbs. Anchoring submersible ultrasonic receivers in river channels with stable substrate. North American Journal of Fisheries Management, 30(4):989–992, August 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Becher:2018:CWT

[BSL18] Cory Becher, Madelyn G. Strahan, and Stuart A. Ludsin. Coded wire tag use with juvenile channel catfish: Evaluation of mortality, retention, and growth. North American Journal of Fisheries Management, 38(6):1367–1374, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Buchanan:2015:ECS

[BSM^{+}15] Rebecca A. Buchanan, John R. Skalski, Gregory Mackey, Charles Snow, and Andrew R. Murdoch. Estimating cohort survival through tributaries for salmonid populations with variable ages at migration. North American Journal of Fisheries Management, 35(5):958–973, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brenkman:2017:LHO

[BSM17] Samuel J. Brenkman, Kathryn T. Sutton, and Anne R. Marshall. Life history observations of adfluvial Chinook salmon prior to reintroduction of anadromous salmonids. North American Journal of Fisheries Management, 37(6):1220–1230, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bodine:2016:EEM

[BSO^{+}16] Kristopher A. Bodine, J. Warren Schlechte, Richard A. Ott, Daniel L. Bennett, and Jacob D. Norman. Estimating exploitation and modeling the effects of hand fishing on a flathead catfish population in East Texas. North American Journal of Fisheries Management, 36(6):1416–1424, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bethoney:2014:CRH

[BSS^{+}14] N. David Bethoney, Kevin D. E. Stokesbury, Bradley P. Schondelmeier, William S. Hoffman, and Michael P. Armstrong. Characterization of river herring bycatch in the Northwest Atlantic midwater trawl fisheries. North American Journal of Fisheries Management, 34(4):828–838, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bestgen:2017:FRS

[BSS^{+}17] Kevin R. Bestgen, Robert C. Schelly, Richard R. Staffeldt, Matthew J. Breen, Darrel E. Snyder, and M. Tildon Jones. First reproduction by stocked bonytail in the Upper Colorado River Basin. North American Journal of Fisheries Management, 37(2):445–455, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bodine:2018:IME

[BSS18] Kristopher A. Bodine, J. Warren Schlechte, and Daniel E. Shoup. An indirect method for estimating size-specific exploitation. North American Journal of Fisheries Management, 38(5):1085–1090, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burnett:2013:CDE

[BSTC13] Nicholas J. Burnett, Keith M. Stamplecoskie, Jason D. Thiem, and Steven J. Cooke. Comparison of detection efficiency among three sizes of half-duplex passive integrated transponders using manual tracking and fixed antenna arrays. North American Journal of Fisheries Management, 33 (1):7–13, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bohling:2019:DFS

[BSVB19] Justin H. Bohling, Steven J. Starcevich, Jennifer Von Bargen, and Elizabeth J. Bailey. Describing fine-scale patterns of genetic structure and introgression of redband trout in a complex river system. North American Journal of Fisheries Management, 39(3):509–523, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bravener:2016:ESM

[BT16] Gale Bravener and Michael Twohey. Evaluation of a sterilemale release technique: a case study of invasive sea lamprey control in a tributary of the Laurentian Great Lakes. North American Journal of Fisheries Management, 36(5):1125–1138, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Budy:2012:CEE

[BTD^{+}12] Phaedra Budy, Gary P. Thiede, Andrew Dean, Devin Olsen, and Gilbert Rowley. A comparative and experimental evaluation of performance of stocked diploid and triploid brook trout. North American Journal of Fisheries Management, 32 (6):1211–1224, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Burdick:2011:TLS

[Bur11] Summer M. Burdick. Tag loss and short-term mortality associated with passive integrated transponder tagging of juvenile Lost River suckers. North American Journal of Fisheries Management, 31(6):1088–1092, December 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Barrett:2017:CAS

[BvPCH17] Boyd N. Barrett, Brett van Poorten, Andrew B. Cooper, and Wolfgang Haider. Concurrently assessing survey mode and sample size in off-site angler surveys. North American Journal of Fisheries Management, 37(4):756–767, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brinson:2017:DSA

[BW17] Ayeisha A. Brinson and Kristy Wallmo. Determinants of saltwater anglers’ satisfaction with fisheries management: Regional perspectives in the United States. North American Journal of Fisheries Management, 37(1):225–234, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bevil:2018:RED

[BW18] Matthew C. Bevil and Michael J. Weber. Relative effects of density dependence versus environmental factors on bluegill populations. North American Journal of Fisheries Management, 38(1):140–151, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Brizendine:2018:EUI

[BWB18] Morgan E. Brizendine, David L. Ward, and Scott A. Bonar. Effectiveness of ultrasonic imaging for evaluating presence and maturity of eggs in fishes in remote field locations. North American Journal of Fisheries Management, 38(5):1017–1026, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Byford:2016:DNP

[BWHM16] Greg J. Byford, C. Michael Wagner, John B. Hume, and Mary L. Moser. Do native Pacific lamprey and invasive sea lamprey share an alarm cue? Implications for use of a natural repellent to guide imperiled Pacific lamprey into fishways. North American Journal of Fisheries Management, 36 (5):1090–1096, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Beacham:2012:PSR

[BWLB12] Terry D. Beacham, Colin G. Wallace, Khai D. Le, and Mark Beere. Population structure and run timing of steelhead in the Skeena River, British Columbia. North American Journal of Fisheries Management, 32(2):262–275, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Budy:2012:SSE

[BWR12] Phaedra Budy, Sara Wood, and Brett Roper. A study of the spawning ecology and early life history survival of Bonneville cutthroat trout. North American Journal of Fisheries Management, 32(3):436–449, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bottcher:2013:FUT

[BWT^{+}13] Jared L. Bottcher, Timothy E. Walsworth, Gary P. Thiede, Phaedra Budy, and David W. Speas. Frequent usage of tributaries by the endangered fishes of the Upper Colorado River Basin: Observations from the San Rafael River, Utah. North American Journal of Fisheries Management, 33(3):585–594, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Bailey:2013:SSA

[BZ13] Michael M. Bailey and Joseph D. Zydlewski. To stock or not to stock? Assessing the restoration potential of a remnant American shad spawning run with hatchery supplementation. North American Journal of Fisheries Management, 33(3):459– 467, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Catalano:2011:ESG

[CA11] Matthew J. Catalano and Micheal S. Allen. Exploring strategies for gizzard shad removal that account for compensatory density dependence and uncertainty. North American Journal of Fisheries Management, 31(6):1153–1162, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carney:2014:EPT

[CA14] Justin M. Carney and Milo D. Adkison. Evaluating the performance of two salmon management strategies using run reconstruction. North American Journal of Fisheries Management, 34(1):159–174, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Canfield:2017:CDS

[Can17] Daniel E. Canfield, Jr. Comment: Do something or do nothing — the fisheries management conundrum. North American Journal of Fisheries Management, 37(4):809–815, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Caroffino:2013:ACL

[Car13] David C. Caroffino. Angler compliance with lake trout length limit regulations in Great Lakes waters. North American Journal of Fisheries Management, 33(6):1203–1209, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Caroffino:2014:GA

[Car14] David C. Caroffino. Guide for authors. North American Journal of Fisheries Management, 34(1):216–221, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Courtney:2016:RAP

[CAS16] Dean L. Courtney, Milo D. Adkison, and Michael F. Sigler. Risk analysis of plausible incidental exploitation rates for the Pacific sleeper shark, a data-poor species in the Gulf of Alaska. North American Journal of Fisheries Management, 36(3):523– 548, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Copeland:2017:LHD

[CAWB17] Timothy Copeland, Michael W. Ackerman, Kristin K. Wright, and Alan Byrne. Life history diversity of Snake River steelhead populations between and within management categories. North American Journal of Fisheries Management, 37(2):395– 404, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Connelly:2010:SPL

[CB10] Nancy A. Connelly and Tommy L. Brown. Sportfishing participation on Lake Ontario: Modeling the past, predicting the future. North American Journal of Fisheries Management, 30 (3):821–830, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2016:SEA

[CBC^{+}16] Richard D. Clark, Jr., James R. Bence, Randall M. Claramunt, James E. Johnson, David Gonder, Nicholas D. Legler, Steve R. Robillard, and Ben D. Dickinson. A spatially explicit assessment of changes in Chinook salmon fisheries in Lakes Michigan and Huron from 1986 to 2011. North American Journal of Fisheries Management, 36(5):1068–1083, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2017:CMC

[CBC^{+}17] Richard D. Clark, Jr., James R. Bence, Randall M. Claramunt, John A. Clevenger, Matthew S. Kornis, Charles R. Bronte, Charles P. Madenjian, and Edward F. Roseman. Changes in movements of Chinook salmon between Lakes Huron and Michigan after alewife population collapse. North American Journal of Fisheries Management, 37(6):1311–1331, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Collins:2015:CRC

[CBDW15] Scott F. Collins, Steven E. Butler, Matthew J. Diana, and David H. Wahl. Catch rates and cost effectiveness of entrapment gears for Asian carp: a comparison of pound nets, hoop nets, and fyke nets in backwater lakes of the Illinois River. North American Journal of Fisheries Management, 35 (6):1219–1225, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chiotti:2016:ASR

[CBHB16] Justin A. Chiotti, James C. Boase, Darryl W. Hondorp, and Andrew S. Briggs. Assigning sex and reproductive stage to adult lake sturgeon using ultrasonography and common morphological measurements. North American Journal of Fisheries Management, 36(1):21–29, February 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chizinski:2016:DMS

[CBHS16] Christopher J. Chizinski, Przemyslaw G. Bajer, Mary E. Headrick, and Peter W. Sorensen. Different migratory strategies of invasive common carp and native northern pike in the American Midwest suggest an opportunity for selective management strategies. North American Journal of Fisheries Management, 36(4):769–779, August 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cross:2013:IRV

[CBM13] Benjamin K. Cross, Michael A. Bozek, and Matthew G. Mitro. Influences of riparian vegetation on trout stream temperatures in Central Wisconsin. North American Journal of Fisheries Management, 33(4):682–692, August 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carlson:2012:ITP

[CBS^{+}12] Thomas J. Carlson, Richard S. Brown, John R. Stephenson, Brett D. Pflugrath, Alison H. Colotelo, Andrew J. Gingerich, Piper L. Benjamin, Mike J. Langeslay, Martin L. Ahmann, Robert L. Johnson, John R. Skalski, Adam G. Seaburg, and Richard L. Townsend. The influence of tag presence on the mortality of juvenile Chinook salmon exposed to simulated hydroturbine passage: Implications for survival estimates and management of hydroelectric facilities. North American Journal of Fisheries Management, 32(2):249–261, April 2012. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colvin:2013:PPR

[CBS13] Michael E. Colvin, Phillip W. Bettoli, and George D. Scholten. Predicting paddlefish Roe yields using an extension of the Beverton–Holt equilibrium yield-per-recruit model. North American Journal of Fisheries Management, 33(5):940–949, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carah:2014:LCR

[CBWB14] Jennifer K. Carah, Christopher C. Blencowe, David W. Wright, and Lisa A. Bolton. Low-cost restoration techniques for rapidly increasing wood cover in coastal Coho salmon streams. North American Journal of Fisheries Management, 34(5):1003–1013, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clement:2010:PIJ

[CC10] Marie Clément and Richard A. Cunjak. Physical injuries in juvenile Atlantic salmon, slimy sculpin, and blacknose dace attributable to electrofishing. North American Journal of Fisheries Management, 30(3):840–850, June 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Culver:2015:SRA

[CC15] Edward F. Culver and John H. Chick. Shocking results: Assessing the rates of fish injury from Pulsed-DC electrofishing. North American Journal of Fisheries Management, 35 (5):1055–1063, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clarke:2016:NEI

[CCC16] Lance R. Clarke, William A. Cameron, and Richard W. Carmichael. No evidence of increased survival or decreased straying from acclimating subyearling fall Chinook salmon to release locations in the Umatilla River of Oregon. North American Journal of Fisheries Management, 36(1):161–166, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clark:2018:ASE

[CCDF18] Scott R. Clark, Mary M. Conner, Scott L. Durst, and Nathan R. Franssen. Age-specific estimates indicate potential deleterious capture effects and low survival of stocked juvenile Colorado pikeminnow. North American Journal of Fisheries Management, 38(5):1059–1074, October 2018. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carmona-Catot:2010:BTR

[CCMA^{+}10] Gerard Carmona-Catot, Peter B. Moyle, Enric Aparicio, Patrick K. Crain, Lisa C. Thompson, and Emili García-Berthou. Brook trout removal as a conservation tool to restore Eagle Lake rainbow trout. North American Journal of Fisheries Management, 30(5):1315–1323, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Collins:2017:CSG

[CDBW17] Scott F. Collins, Matthew J. Diana, Steven E. Butler, and David H. Wahl. A comparison of sampling gears for capturing juvenile silver carp in river–floodplain ecosystems. North American Journal of Fisheries Management, 37(1):94–100, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cope:2011:ADS

[CDD^{+}11] Jason M. Cope, John DeVore, E. J. Dick, Kelly Ames, John Budrick, Daniel L. Erickson, Joanna Grebel, Gretchen Hanshew, Robert Jones, Lynn Mattes, Corey Niles, and Sarah Williams. An approach to defining stock complexes for U.S. West Coast groundfishes using vulnerabilities and ecological distributions. North American Journal of Fisheries Management, 31(4):589–604, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cox:2012:USA

[CDQG12] Benjamin S. Cox, Andrew M. Dux, Michael C. Quist, and Christopher S. Guy. Use of a seismic air gun to reduce survival of nonnative lake trout embryos: a tool for conservation? North American Journal of Fisheries Management, 32(2):292– 298, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cahill:2015:ARF

[CEH^{+}15] Christopher L. Cahill, Andrea C. Erwin, Kimberly L. Howland, Mark F. Hulsman, Brianne D. Lunn, Fred Noddin, William M. Tonn, Abul Basar Baki, Gregory Courtice, and David Z. Zhu. Assessing responses of fish to habitat enhancement in barrenlands streams of the Northwest Territories. North American Journal of Fisheries Management, 35 (4):755–764, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carim:2015:DWD

[CEP15] Kellie Carim, Lisa Eby, and Ron Pierce. Does whirling disease mediate hybridization between a native and nonnative trout? North American Journal of Fisheries Management, 35(2):337– 351, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crane:2013:SSS

[CF13] Derek P. Crane and John M. Farrell. Spawning substrate size, shape, and siltation influence walleye egg retention. North American Journal of Fisheries Management, 33(2):329–337, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Catalano:2018:WED

[CF18] Matthew J. Catalano and Troy M. Farmer. Weak effects of density on mean length at age and condition of largemouth bass at large Alabama reservoirs. North American Journal of Fisheries Management, 38(2):281–294, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chaudoin:2015:UVO

[CFBB15] Ambre L. Chaudoin, Olin G. Feuerbacher, Scott A. Bonar, and Paul J. Barrett. Underwater videography outperforms above-water videography and in-person surveys for monitoring the spawning of devils hole pupfish. North American Journal of Fisheries Management, 35(6):1252–1262, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carlson:2016:OMR

[CFG16] Andrew K. Carlson, Mark J. Fincel, and Brian D. S. Graeb. Otolith microchemistry reveals natal origins of walleyes in Missouri River reservoirs. North American Journal of Fisheries Management, 36(2):341–350, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clarke:2010:PPA

[CFW^{+}10] Lance R. Clarke, Michael W. Flesher, Timothy A. Whitesel, Gary R. Vonderohe, and Richard W. Carmichael. Postrelease performance of acclimated and directly released hatchery summer steelhead into Oregon tributaries of the Snake River. North American Journal of Fisheries Management, 30 (5):1098–1109, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clarke:2011:SSH

[CFWC11] Lance R. Clarke, Michael W. Flesher, Shelby M. Warren, and Richard W. Carmichael. Survival and straying of hatchery steelhead following forced or volitional release. North American Journal of Fisheries Management, 31(1):116–123, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Callihan:2014:GMH

[CGDB14] Jody L. Callihan, Charlton H. Godwin, Kevin J. Dockendorf, and Jeffrey A. Buckel. Growth and mortality of hatcheryreared striped bass stocked into nonnatal systems. North American Journal of Fisheries Management, 34(6):1131–1139, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Civiello:2018:DCA

[CGG^{+}18] Anthony P. Civiello, Nathan J. C. Gosch, Todd R. Gemeinhardt, Marcus L. Miller, Joseph L. Bonneau, Kimberly A. Chojnacki, Aaron J. DeLonay, and James M. Long. Diet and condition of age-0 Scaphirhynchus sturgeon: Implications for shallow-water habitat restoration. North American Journal of Fisheries Management, 38(6):1324–1338, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crandall:2018:URC

[CGL18] Chelsey A. Crandall, Taryn M. Garlock, and Kai Lorenzen. Understanding resource-conserving behaviors among fishers: Barotrauma mitigation and the power of subjective norms in Florida’s reef fisheries. North American Journal of Fisheries Management, 38(2):271–280, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Christiansen:2013:AJP

[CGM13] Helena E. Christiansen, Lisa P. Gee, and Matthew G. Mesa. Anesthesia of juvenile Pacific lampreys with MS-222, BEN-ZOAK, AQUI-S 20E, and aquacalm. North American Journal of Fisheries Management, 33(2):269–276, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crosson:2017:ICF

[CH17] Scott Crosson and Lia Hibbert. Integrating commercial fisheries registration, education, and social science in the U.S. Virgin Islands. North American Journal of Fisheries Management, 37(2):349–352, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Conrad:2013:UMC

[CHBY13] Robert Conrad, Angelika Hagen-Breaux, and Henry Yuen. Unbiased methods for calculating mortality in mark-selective fisheries models for ocean salmon. North American Journal of Fisheries Management, 33(2):253–264, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Copeland:2018:VSC

[CHDW18] Timothy Copeland, Kala Hernandez, Micah T. Davison, and Kristin Wright. Validation of spawn checks and saltwater age assignments based on scales from known repeat-spawning steelhead. North American Journal of Fisheries Management, 38(5):1050–1058, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Conrad:2016:API

[CHJ^{+}16] J. Louise Conrad, Eric Holmes, Carson Jeffres, Lynn Takata, Naoaki Ikemiyagi, Jacob Katz, and Ted Sommer. Application of passive integrated transponder technology to juvenile salmon habitat use on an experimental agricultural floodplain.

North American Journal of Fisheries Management, 36(1):30– 39, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crossman:2013:EES

[CHL13] James A. Crossman, K. Larry Hammell, and Matthew K. Litvak. Experimental examination of surgical procedures for implanting sonic transmitters in juvenile shortnose sturgeon and Atlantic sturgeon. North American Journal of Fisheries Management, 33(3):549–556, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cummings:2011:IDP

[CHPP11] Jonathan W. Cummings, Merran J. Hague, David A. Patterson, and Randall M. Peterman. The impact of different performance measures on model selection for Fraser River sockeye salmon. North American Journal of Fisheries Management, 31(2):323–334, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crumpton:2012:MOF

[CHW12] Roman L. Crumpton, James P. Henne, and Kent M. Ware. Marking otoliths and fin spines of juvenile shortnose sturgeon with oxytetracycline and the effects of water temperature during treatment. North American Journal of Fisheries Management, 32(3):523–527, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carlson:2010:MCR

[CI10] Andrew J. Carlson and Daniel A. Isermann. Mandatory catch and release and maximum length limits for largemouth bass in Minnesota: Is exploitation still a relevant concern? North American Journal of Fisheries Management, 30(1):209–220, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Catalano:2014:SBE

[CJ14] Matthew J. Catalano and Michael L. Jones. A simulationbased evaluation of in-season management tactics for anadromous fisheries: Accounting for risk in the Yukon River fall chum salmon fishery. North American Journal of Fisheries Management, 34(6):1227–1241, December 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Caldwell:2013:PFE

[CJA^{+}13] Colleen A. Caldwell, Gerald Z. Jacobi, Michael C. Anderson, Robert R. Parmenter, Jeanine McGann, William R. Gould, Robert DuBey, and M. Donna Jacobi. Prescribed-fire effects on an aquatic community of a southwest montane grassland system. North American Journal of Fisheries Management, 33(5):1049–1062, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Crossman:2016:DDJ

[CJH16] James A. Crossman, Kathleen J. Jay, and Larry R. Hildebrand. Describing the diet of juvenile white sturgeon in the Upper Columbia River Canada with lethal and nonlethal methods. North American Journal of Fisheries Management, 36(2):421–432, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chalupnicki:2010:HSS

[CJMD10] Marc A. Chalupnicki, James H. Johnson, James E. McKenna, Jr., and Dawn E. Dittman. Habitat selection and spawning success of walleyes in a tributary to Owasco Lake, New York. North American Journal of Fisheries Management, 30 (1):170–178, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cooney:2010:DSW

[CK10] Patrick B. Cooney and Thomas J. Kwak. Development of standard weight equations for Caribbean and Gulf of Mexico amphidromous fishes. North American Journal of Fisheries Management, 30(5):1203–1209, October 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cope:2019:EAC

[CKBP19] W. Robert Cope, Thomas J. Kwak, Tyler R. Black, and Krishna Pacifici. Evaluation of artificial cover units as a sampling technique and habitat enhancement for madtoms in rivers. North American Journal of Fisheries Management, 39 (4):778–787, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clabough:2012:UNV

[CKC^{+}12] Tami S. Clabough, Matthew L. Keefer, Christopher C. Caudill, Eric L. Johnson, and Christopher A. Peery. Use of night video to enumerate adult Pacific lamprey passage at hydroelectric dams: Challenges and opportunities to improve escapement estimates. North American Journal of Fisheries Management, 32(4):687–695, August 2012. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cvetkovic:2012:GTI

[CKCF12] Maja Cvetkovic, Kristina Kostuk, and Patricia Chow-Fraser. Gear-type influences on fish catch and a wetland fish index in Georgian Bay wetlands. North American Journal of Fisheries Management, 32(2):313–324, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carter:2012:EVC

[CL12] David W. Carter and Christopher Liese. The economic value of catching and keeping or releasing saltwater sport fish in the Southeast USA. North American Journal of Fisheries Management, 32(4):613–625, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clemens:2015:SSA

[Cle15] Benjamin J. Clemens. A survey of steelhead age and iteroparity rates from a volunteer angler program in the Willamette River Basin, Oregon. North American Journal of Fisheries Management, 35(5):1046–1054, October 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Coppola:2019:SRC

[CMC^{+}19] Giancarlo Coppola, Leandro E. Miranda, Michael E. Colvin, Hunter R. Hatcher, and Marcus A. Lashley. Submergence responses of cool-season annual plants and potential for fish habitat. North American Journal of Fisheries Management, 39(6):1269–1276, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Corbett:2012:EEA

[CMD12] Stephen C. Corbett, Mary L. Moser, and Andrew H. Dittman. Experimental evaluation of adult spring Chinook salmon radio-tagged during the late stages of spawning migration.

North American Journal of Fisheries Management, 32(5):853– 858, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Castillo:2014:ECP

[CMF^{+}14] Gonzalo Castillo, Jerry Morinaka, Robert Fujimura, Jason DuBois, Bradd Baskerville-Bridges, Joan Lindberg, Galen Tigan, Luke Ellison, and James Hobbs. Evaluation of calcein and photonic marking for cultured delta smelt. North American Journal of Fisheries Management, 34(1):30–38, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cassinelli:2019:PDT

[CMK^{+}19] John D. Cassinelli, Kevin A. Meyer, Martin K. Koenig, Ninh V. Vu, and Matthew R. Campbell. Performance of diploid and triploid westslope cutthroat trout fry stocked into Idaho alpine lakes. North American Journal of Fisheries Management, 39(1):112–123, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cathcart:2019:TTE

[CMMR19] Charles N. Cathcart, Mark C. McKinstry, Peter D. MacKinnon, and Claire M. Ruffing. A tribute to tributaries: Endangered fish distributions within critical habitat of the San Juan River, USA. North American Journal of Fisheries Management, 39(5):1015–1025, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chiaramonte:2018:AEF

[CMWM18] Luciano V. Chiaramonte, Kevin A. Meyer, Don W. Whitney, and Joshua L. McCormick. Air exposure, fight times, and deep-hooking rates of steelhead caught in Idaho fisheries. North American Journal of Fisheries Management, 38 (5):1114–1121, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colombo:2010:CCC

[CPM^{+}10] Robert E. Colombo, Quinton E. Phelps, Candice M. Miller, James E. Garvey, Roy C. Heidinger, and Nathaniel S. Richards. Comparison of channel catfish age estimates and resulting population demographics using two common structures. North American Journal of Fisheries Management, 30 (1):305–308, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Colvin:2012:SCC

[CPSG12] Michael E. Colvin, Clay L. Pierce, Timothy W. Stewart, and Scott E. Grummer. Strategies to control a common carp population by pulsed commercial harvest. North American Journal of Fisheries Management, 32(6):1251–1264, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cullen:2017:EBS

[CS17] Daniel W. Cullen and Bradley G. Stevens. Examination of black sea bass trap catches in relation to soak time in the Middle Atlantic Bight. North American Journal of Fisheries Management, 37(1):9–15, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Caroffino:2010:ELS

[CSED10] David C. Caroffino, Trent M. Sutton, Robert F. Elliott, and Michael C. Donofrio. Early life stage mortality rates of lake sturgeon in the Peshtigo River, Wisconsin. North American Journal of Fisheries Management, 30(1):295–304, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carlin:2012:SCA

[CSF12] Caroline Carlin, Susan A. Schroeder, and David C. Fulton. Site choice among Minnesota walleye anglers: The influence of resource conditions, regulations and catch orientation on lake preference. North American Journal of Fisheries Management, 32(2):299–312, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Clancey:2017:EHR

[CSH^{+}17] Katherine Clancey, Laurel Saito, Kurt Hellmann, Connie Svoboda, John Hannon, and Randy Beckwith. Evaluating headof-reservoir water temperature for juvenile Chinook salmon and steelhead at Shasta lake with modeled temperature curtains. North American Journal of Fisheries Management, 37 (5):1161–1175, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Castillo:2018:UNM

[CSH^{+}18] Gonzalo C. Castillo, Marade E. Sandford, Tien-Chieh Hung, Galen Tigan, Joan C. Lindberg, Wan-Ru Yang, and Erwin E. Van Nieuwenhuyse. Using natural marks to identify individual cultured adult delta smelt. North American Journal of Fisheries Management, 38(3):698–705, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Cupp:2018:ACD

[CST^{+}18] Aaron R. Cupp, Justin R. Smerud, John A. Tix, Jose M. Rivera, Stacie A. Kageyama, Christopher M. Merkes, Richard A. Erickson, Jon J. Amberg, and Mark P. Gaikowski. Assessment of carbon dioxide piscicide treatments. North American Journal of Fisheries Management, 38(6):1241–1250, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Castro-Santos:2013:PSP

[CSV13] Theodore Castro-Santos and Volney Vono. Posthandling survival and PIT tag retention by alewives — a comparison of gastric and surgical implants. North American Journal of Fisheries Management, 33(4):790–794, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Camacho:2019:MIB

[CSWP19] Carlos A. Camacho, Christopher J. Sullivan, Michael J. Weber, and Clay L. Pierce. Morphological identification of bighead carp, silver carp, and grass carp eggs using random forests machine learning classification. North American Journal of Fisheries Management, 39(6):1373–1384, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Courtney:2018:QNB

[CSWW18] Michael B. Courtney, Erik R. Schoen, Andy Wizik, and Peter A. H. Westley. Quantifying the net benefits of suppression: Truncated size structure and consumption of native salmonids by invasive northern pike in an Alaska lake. North American Journal of Fisheries Management, 38(6):1306–1315, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Couillard:2014:AFR

[CVD^{+}14] Catherine M. Couillard, Guy Verreault, Pierre Dumont, David Stanley, and Ron W. Threader. Assessment of fat reserves adequacy in the first migrant silver American eels of a largescale stocking experiment. North American Journal of Fisheries Management, 34(4):802–813, August 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Camp:2015:ESO

[CvPW15] Edward V. Camp, Brett T. van Poorten, and Carl J. Walters. Evaluating short openings as a management tool to maximize catch-related utility in catch-and-release fisheries. North American Journal of Fisheries Management, 35(6):1106–1120, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Carl:2016:EAI

[CWB16] Dray D. Carl, Michael J. Weber, and Michael L. Brown. An evaluation of attractants to increase catch rates and deplete age-0 common carp in shallow South Dakota lakes. North American Journal of Fisheries Management, 36(3):506–513, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Chasco:2014:EAP

[CWH^{+}14] Brandon E. Chasco, Eric J. Ward, Jay A. Hesse, Craig Rabe, Ryan Kinzer, Jason L. Vogel, and Rick Orme. Evaluating the accuracy and precision of multiple abundance estimators using state-space models: a case study for a threatened population of Chinook salmon in Johnson Creek, Idaho. North American Journal of Fisheries Management, 34(5):945–954, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Capizzano:2019:FSD

[CZH^{+}19] Connor W. Capizzano, Douglas R. Zemeckis, William S. Hoffman, Hugues P. Benoît, Emily Jones, Micah J. Dean, Nate Ribblett, James A. Sulikowski, and John W. Mandelman. Fishery-scale discard mortality rate estimate for haddock in the Gulf of Maine recreational fishery. North American Journal of Fisheries Management, 39(5):964–979, October 2019.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dorow:2012:RBP

[DA12] Malte Dorow and Robert Arlinghaus. The relationship between personal commitment to angling and the opinions and attitudes of German anglers towards the conservation and management of the European eel Anguilla anguilla. North American Journal of Fisheries Management, 32(3):466–479, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dagit:2017:REB

[Dag17] Rosi Dagit. Rates and effects of branding due to electroshock: Some additional perspectives: Response to comment. North American Journal of Fisheries Management, 37(2):431, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dotson:2013:URH

[DAKP13] Jason R. Dotson, Micheal S. Allen, Janice A. Kerns, and William F. Pouder. Utility of restrictive harvest regulations for trophy largemouth bass management. North American Journal of Fisheries Management, 33(3):499–507, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DeHaan:2016:IIH

[DASB16] Patrick W. DeHaan, Brice A. Adams, Paul D. Scheerer, and Brian L. Bangs. Influence of introduction history on genetic variation in introduced populations: a case study of Oregon chub. North American Journal of Fisheries Management, 36 (6):1278–1291, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DeHaan:2013:SSB

[DB13] Patrick W. DeHaan and Shana R. Bernall. Spawning success of bull trout transported above main-stem Clark Fork River Dams in Idaho and Montana. North American Journal of Fisheries Management, 33(6):1269–1282, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Duffy:2011:PJP

[DBE11] Walter G. Duffy, Eric P. Bjorkstedt, and Christopher S. Ellings. Predation on juvenile Pacific salmon Oncorhynchus spp. in downstream migrant traps in Prairie Creek, California. North American Journal of Fisheries Management, 31 (1):151–164, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Denson:2012:ARD

[DBJD12] Michael R. Denson, Karl Brenkert IV, Wallace E. Jenkins, and Tanya L. Darden. Assessing red drum juvenile stocking in a South Carolina Estuary using genetic identification. North American Journal of Fisheries Management, 32(1):32– 43, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2011:SRA

[DBK11] Daniel J. Daugherty, David L. Buckmeier, and Praveen K. Kokkanti. Sensitivity of recreational access to reservoir water level variation: An approach to identify future access needs in reservoirs. North American Journal of Fisheries Management, 31(1):63–69, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Davis:2014:EFL

[DBKG14] Jacob L. Davis, Michael E. Barnes, Jeremy L. Kientz, and Austin G. Galinat. Effects of fish length and anatomical placement on retention of visible implant alpha tags in hatcheryreared rainbow trout. North American Journal of Fisheries Management, 34(5):932–937, October 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2019:SSD

[DBS19] Daniel J. Daugherty, David L. Buckmeier, and Nathan G. Smith. Sex-specific dynamic rates in the alligator gar: Implications for stock assessment and management. North American Journal of Fisheries Management, 39(3):535–542, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Davis:2013:ETC

[DBW13] Jacob L. Davis, Michael E. Barnes, and Jerry W. Wilhite. Effectiveness of three compounds to anesthetize rainbow trout during PIT tag implantation surgery. North American Journal of Fisheries Management, 33(3):482–487, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Downs:2013:LLW

[DC13] Donna E. Downs and Yuk W. Cheng. Length–length and width–length conversion of longnose skate and big skate off the Pacific Coast: Implications for the choice of alternative measurement units in fisheries stock assessment. North American Journal of Fisheries Management, 33(5):887–893, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunham:2013:CET

[DCHC13] Jason B. Dunham, Nathan D. Chelgren, Michael P. Heck, and Steven M. Clark. Comparison of electrofishing techniques to detect larval lampreys in wadeable streams in the Pacific Northwest. North American Journal of Fisheries Management, 33(6):1149–1155, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dillon:2019:HDA

[DCL19] Rebecca A. Dillon, Joseph D. Conroy, and Stuart A. Ludsin. Hydroacoustic data-analysis recommendations to quantify prey-fish abundance in shallow, target-rich ecosystems. North American Journal of Fisheries Management, 39(2):270–288, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Drymon:2019:GIS

[DCP^{+}19] J. Marcus Drymon, Pearce T. Cooper, Sean P. Powers, Molly M. Miller, Sharon Magnuson, Evan Krell, and Chris Bird. Genetic identification of species responsible for depredation in commercial and recreational fisheries. North American Journal of Fisheries Management, 39(3):524–534, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Davis:2011:TUA

[DCS11] Johnathan G. Davis, S. Bradford Cook, and David D. Smith. Testing the utility of an adaptive cluster sampling method for monitoring a rare and imperiled darter. North American Journal of Fisheries Management, 31(6):1123–1132, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dobos:2016:ISW

[DCS^{+}16] Marika E. Dobos, Matthew P. Corsi, Daniel J. Schill, Joseph M. DuPont, and Michael C. Quist. Influences of summer water temperatures on the movement, distribution, and resource use of fluvial westslope cutthroat trout in the South Fork Clearwater River Basin. North American Journal of Fisheries Management, 36(3):549–567, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2014:ESS

[DDAS14] Daniel J. Daugherty, M. Todd Driscoll, Daniel E. Ashe, and J. Warren Schlechte. Effects of structural and spatiotemporal factors on fish use of artificial habitat in a Texas reservoir. North American Journal of Fisheries Management, 34(2):453– 462, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Desender:2018:PTI

[DDD^{+}18] Marieke Desender, Lisa Dumolein, Luc Duchateau, Dominique Adriaens, Daan Delbare, Hans Polet, Koen Chiers, and Annemie Decostere. Pulse trawling: The impact of pulsed direct current on early life stages of sole Solea solea. North American Journal of Fisheries Management, 38(2):432–438, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Desender:2019:RA

[DDD^{+}19] Marieke Desender, Lisa Dumolein, Luc Duchateau, Dominique Adriaens, Daan Delbare, Hans Polet, Koen Chiers, and Annemie Decostere. Reviewer acknowledgments. North American Journal of Fisheries Management, 39(6):1427–1429, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Diana:2018:ERE

[DDEW18] Matthew J. Diana, Brett J. Diffin, Lisa M. Einfalt, and David H. Wahl. Effect of rearing experience on the survival, growth, and behavior of hatchery-reared largemouth bass. North American Journal of Fisheries Management, 38 (4):794–802, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Deroba:2014:ECA

[Der14] Jonathan J. Deroba. Evaluating the consequences of adjusting fish stock assessment estimates of biomass for retrospective patterns using Mohn’s rho. North American Journal of Fisheries Management, 34(2):380–390, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dauwalter:2018:RRV

[DFMP18] Daniel C. Dauwalter, Kurt A. Fesenmyer, Scott W. Miller, and Tucker Porter. Response of riparian vegetation, instream habitat, and aquatic biota to riparian grazing exclosures. North American Journal of Fisheries Management, 38(5):1187–1200, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dieterman:2010:LSP

[DFPS10] Douglas J. Dieterman, John Frank, Nate Painovich, and David F. Staples. Lake sturgeon population status and demography in the Kettle River, Minnesota, 1992–2007. North American Journal of Fisheries Management, 30(2):337–351, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunning:2012:DPJ

[DG12] Dennis J. Dunning and Christopher W. D. Gurshin. Downriver passage of juvenile blueback herring near an ultrasonic field in the Mohawk River. North American Journal of Fisheries Management, 32(2):365–380, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dux:2011:SDP

[DGF11] Andrew M. Dux, Christopher S. Guy, and Wade A. Fredenberg. Spatiotemporal distribution and population characteristics of a nonnative lake trout population, with implications for suppression. North American Journal of Fisheries Management, 31(2):187–196, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunham:2011:AFN

[DGS^{+}11] Jason Dunham, Kirsten Gallo, Dan Shively, Chris Allen, and Brad Goehring. Assessing the feasibility of native fish reintroductions: a framework applied to threatened bull trout. North American Journal of Fisheries Management, 31(1):106–115, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dean:2012:DAC

[DHA12] Micah J. Dean, William S. Hoffman, and Michael P. Armstrong. Disruption of an Atlantic cod spawning aggregation resulting from the opening of a directed gill-net fishery. North American Journal of Fisheries Management, 32(1):124–134, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dumke:2010:CMR

[DHB^{+}10] Joshua D. Dumke, Thomas R. Hrabik, Valerie J. Brady, Karen B. Gran, Ronald R. Regal, and Michael J. Seider. Channel morphology response to selective wood removals in a sandladen Wisconsin trout stream. North American Journal of Fisheries Management, 30(3):776–790, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dabrowksa:2017:UHA

[DHH17] Kora Dabrowksa, Len M. Hunt, and Wolfgang Haider. Understanding how angler characteristics and context influence angler preferences for fishing sites. North American Journal of Fisheries Management, 37(6):1350–1361, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Diver:2019:SSB

[DHK^{+}19] Tracy Diver, Alexis Harrison, William Knight, Manuel Ulibarri, and Wade Wilson. Size sorting in bonytail skews the sex ratio of stocked fish. North American Journal of Fisheries Management, 39(3):452–459, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daniel:2011:MRA

[DHLD11] Adam J. Daniel, Brendan J. Hicks, Nicholas Ling, and Bruno O. David. Movements of radio- and acoustic-tagged adult koi carp in the Waikato River, New Zealand. North American Journal of Fisheries Management, 31(2):352–362, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dembkowski:2019:PIA

[DIK19] Daniel J. Dembkowski, Daniel A. Isermann, and Ryan P. Koenigs. Potential for improving among-reader precision and accuracy of walleye age estimates with minimal training. North American Journal of Fisheries Management, 39 (4):625–636, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunton:2016:AGA

[DJS^{+}16] Keith J. Dunton, Adrian Jordaan, David H. Secor, Christopher M. Martinez, Thomas Kehler, Kathy A. Hattala, Joel P. Van Eenennaam, Matthew T. Fisher, Kim A. McKown, David O. Conover, and Michael G. Frisk. Age and growth of Atlantic sturgeon in the New York bight. North American Journal of Fisheries Management, 36(1):62–73, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dagit:2016:REB

[DK16] Rosi Dagit and Jenna Krug. Rates and effects of branding due to electroshock observed in Southern California steelhead in Topanga Creek, California. North American Journal of Fisheries Management, 36(4):888–899, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Duermit:2015:CCR

[DKSW15] Elizabeth Duermit, Peter R. Kingsley-Smith, and Dara H. Wilber. The consequences of claw removal on stone crabs Menippe spp. and the ecological and fishery implications. North American Journal of Fisheries Management, 35(5):895– 905, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Detar:2014:ECR

[DKWG14] Jason Detar, David Kristine, Tyler Wagner, and Tom Greene. Evaluation of catch-and-release regulations on brook trout in Pennsylvania streams. North American Journal of Fisheries Management, 34(1):49–56, February 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dumont:2011:EGC

[DLC11] Spencer C. Dumont and Dijar J. Lutz-Carrillo. Estimating genetic composition in introgressed largemouth bass populations: Does age matter? North American Journal of Fisheries Management, 31(1):176–181, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dippold:2016:EMR

[DLP16] David A. Dippold, Robert T. Leaf, and Mark S. Peterson. Evaluating management regimes using per-recruit models and relative stock density for Mississippi’s spotted seatrout. North American Journal of Fisheries Management, 36(5):1178–1189, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Diana:2012:ETC

[DLSW12] Matthew J. Diana, Aaron L. Larsen, Michael J. Siepker, and David H. Wahl. Effects of tournament compared with catch and release angling on nest abandonment of largemouth bass. North American Journal of Fisheries Management, 32(5):832– 837, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dagel:2012:BUR

[DM12] Jonah D. Dagel and L. E. Miranda. Backwaters in the upper reaches of reservoirs produce high densities of age-0 crappies. North American Journal of Fisheries Management, 32 (4):626–634, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Doss:2019:FES

[DMC^{+}19] Sasha S. Doss, Brian R. Murphy, Leandro Castello, Joseph A. Williams, John Copeland, and Victor J. DiCenzo. Field evaluation and simulation modeling of length limits and their effects on fishery quality for muskellunge in the New River, Virginia. North American Journal of Fisheries Management, 39 (1):3–16, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Deng:2010:IMJ

[DMRJ10] Zhiqun Deng, Robert P. Mueller, Marshall C. Richmond, and Gary E. Johnson. Injury and mortality of juvenile salmon entrained in a submerged jet entering still water. North American Journal of Fisheries Management, 30(3):623–628, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dumont:2011:PCP

[DN11] Spencer C. Dumont and Ben C. Neely. A proposed change to palmetto bass proportional size distribution length categories. North American Journal of Fisheries Management, 31 (4):722–725, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Densmore:2013:EDS

[DP13] Christine L. Densmore and Frank M. Panek. Effects of depletion sampling by standard three-pass pulsed DC electrofishing on blood chemistry parameters of fishes from Appalachian streams. North American Journal of Fisheries Management, 33(2):298–306, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2017:PSA

[DPK^{+}17] Daniel J. Daugherty, Kevin L. Pangle, William Karel, Francesca Baker, Clinton R. Robertson, David L. Buckmeier, Nathan G. Smith, and Norman Boyd. Population structure of alligator gar in a Gulf Coast River: Insights from otolith microchemistry and genetic analyses. North American Journal of Fisheries Management, 37(2):337–348, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DeWeber:2017:HPM

[DPS^{+}17] J. Tyrell DeWeber, James T. Peterson, Cameron Sharpe, Michael L. Kent, Michael E. Colvin, and Carl B. Schreck. A hidden-process model for estimating prespawn mortality using carcass survey data. North American Journal of Fisheries Management, 37(1):162–175, February 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dunning:2010:ERT

[DR10] Dennis J. Dunning and Quentin E. Ross. Effect of radiotagging on escape reactions of adult blueback herring to ultrasound. North American Journal of Fisheries Management, 30(1):26–32, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dugan:2019:UNA

[DR19] Jessica A. Dugan and Frank J. Rahel. Use of natural and added cover types by game and nongame fishes in a Great Plains river. North American Journal of Fisheries Management, 39(5):980–988, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dub:2017:UEF

[DRC17] Joshua D. Dub, Charles R. Roswell, and Sergiusz Czesny. Utilizing external features and a consensus field-based approach to determine sex of Lake Michigan yellow perch. North American Journal of Fisheries Management, 37(5):1107–1111, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dauwalter:2010:PRM

[DRG10] Daniel C. Dauwalter, Frank J. Rahel, and Kenneth G. Gerow. Power of revisit monitoring designs to detect forestwide declines in trout populations. North American Journal of Fisheries Management, 30(6):1462–1468, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Devine:2018:PRE

[DRW^{+}18] Matthew T. Devine, Allison H. Roy, Andrew R. Whiteley, Benjamin I. Gahagan, Michael P. Armstrong, and Adrian Jordaan. Precision and relative effectiveness of a purse seine for sampling age-0 river herring in lakes. North American Journal of Fisheries Management, 38(3):650–662, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Daugherty:2012:FSY

[DS12] Daniel J. Daugherty and Nathan G. Smith. Frequency of strong year-classes: Implications on fishery dynamics for three life history strategies of fishes. North American Journal of Fisheries Management, 32(6):1191–1200, December 2012. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dattilo:2019:AGF

[DSB19] John Dattilo, Daniel E. Shoup, and Shannon K. Brewer. Age and growth of freshwater drum and gizzard shad occupying two reservoir-river complexes with different groundwater contributions. North American Journal of Fisheries Management, 39(6):1132–1142, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dodson:2013:OMD

[DSD^{+}13] Julian J. Dodson, Pascal Sirois, Gaétan Daigle, Philippe Gaudin, and Agnès Bardonnet. Otolith microstructure during the early life-history stages of brown trout: Validation and interpretation. North American Journal of Fisheries Management, 33(1):108–116, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Docker:2016:SGA

[DSJS16] Margaret F. Docker, Gregory S. Silver, Jeffrey C. Jolley, and Erin K. Spice. Simple genetic assay distinguishes lamprey genera Entosphenus and Lampetra: Comparison with existing genetic and morphological identification methods. North American Journal of Fisheries Management, 36(4):780–787, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dean:2011:MOP

[DT11] Jan C. Dean and Alan J. Temple. Maximum output of peak power for two backpack electrofishers operated at various pulsed direct current duty cycles and water conductivity levels. North American Journal of Fisheries Management, 31 (3):520–529, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dieterman:2018:WHS

[DTA18] Douglas J. Dieterman, William C. Thorn, and Charles S. Anderson. Winter habitat selection by large brown trout in streams with and without habitat rehabilitation. North American Journal of Fisheries Management, 38(1):253–266, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dean:2019:OPU

[DTS19] Jan C. Dean, Alan J. Temple, and Jeremiah W. Smith. Output performance of U.S.-built backpack electrofishers. North American Journal of Fisheries Management, 39(6):1321–1328, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dustin:2017:NHF

[DV17] Donna L. Dustin and Bruce Vondracek. Nearshore habitat and fish assemblages along a gradient of shoreline development. North American Journal of Fisheries Management, 37(2):432– 444, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DeBoom:2010:EFH

[DVWS10] Corey S. DeBoom, Matthew M. VanLandeghem, David H. Wahl, and Michael J. Siepker. Effects of four hook removal techniques on feeding, growth, and survival of deeply hooked largemouth bass. North American Journal of Fisheries Management, 30(4):956–963, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dembkowski:2015:ISB

[DWB^{+}15] Daniel J. Dembkowski, D. W. Willis, B. G. Blackwell, S. R. Chipps, T. D. Bacula, and M. R. Wuellner. Influence of smallmouth bass predation on recruitment of age-0 yellow perch in South Dakota glacial lakes. North American Journal of Fisheries Management, 35(4):736–747, August 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dieterman:2016:RTB

[DWCK16] Douglas J. Dieterman, Thomas S. Walker, Philip A. Cochran, and Melissa Konsti. Reproductive traits of brown trout in two contrasting streams of Southeast Minnesota. North American Journal of Fisheries Management, 36(3):465–476, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Dembkowski:2012:SGL

[DWW12] Daniel J. Dembkowski, Melissa R. Wuellner, and David W. Willis. Sampling glacial lake littoral fish assemblages with four gears. North American Journal of Fisheries Management, 32(6):1160–1166, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Diana:2017:DSS

[DWW17] Matthew J. Diana, Curtis P. Wagner, and David H. Wahl. Differences in stocking success among geographically distinct stocks of juvenile muskellunge in Illinois lakes. North American Journal of Fisheries Management, 37(3):633–643, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Demer:2014:OFQ

[DZ14] David A. Demer and Juan P. Zwolinski. Optimizing fishing quotas to meet target fishing fractions of an internationally exploited stock of Pacific sardine. North American Journal of Fisheries Management, 34(6):1119–1130, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Demer:2017:MCA

[DZ17] David A. Demer and Juan P. Zwolinski. A method to consistently approach the target total fishing fraction of Pacific sardine and other internationally exploited fish stocks. North American Journal of Fisheries Management, 37(2):284–293, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

DeRito:2010:TRS

[DZS10] James N. DeRito, Alexander V. Zale, and Bradley B. Shepard. Temporal reproductive separation of fluvial Yellowstone cutthroat trout from rainbow trout and hybrids in the Yellowstone River. North American Journal of Fisheries Management, 30(4):866–886, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eberts:2018:UVM

[EBS18] Rebecca L. Eberts, Jessica C. Butt, and Christopher M. Somers. Unexplained variation in movement by walleye and sauger after catch-and-release angling tournaments. North American Journal of Fisheries Management, 38(6):1350–1366, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eslinger:2010:FAR

[EDN10] Lawrence D. Eslinger, David M. Dolan, and Steven P. Newman. Factors affecting recruitment of age-0 muskellunge in Escanaba Lake, Wisconsin, 1987–2006. North American Journal of Fisheries Management, 30(4):908–920, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Etnier:2010:SSM

[EF10] Michael A. Etnier and Charles W. Fowler. Size selectivity in marine mammal diets as a guide to evolutionarily enlightened fisheries management. North American Journal of Fisheries Management, 30(2):588–603, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Evans:2016:CWP

[EHP^{+}16] Allen F. Evans, David A. Hewitt, Quinn Payton, Bradley M. Cramer, Ken Collis, and Daniel D. Roby. Colonial waterbird predation on Lost River and shortnose suckers in the Upper Klamath Basin. North American Journal of Fisheries Management, 36(6):1254–1268, December 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eggleton:2010:CGS

[EJL10] Michael A. Eggleton, John R. Jackson, and Benjamin J. Lubinski. Comparison of gears for sampling littoral-zone fishes in floodplain lakes of the Lower White River, Arkansas. North American Journal of Fisheries Management, 30(4):928–939, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Elle:2010:ECM

[EKM10] F. Steven Elle, Martin K. Koenig, and Kevin A. Meyer. Evaluation of calcein as a mass mark for rainbow trout raised in outdoor hatchery raceways. North American Journal of Fisheries Management, 30(6):1408–1412, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ertel:2017:LHM

[EMK^{+}17] Brian D. Ertel, Thomas E. McMahon, Todd M. Koel, Robert E. Gresswell, and Jason C. Burckhardt. Life history migrations of adult Yellowstone cutthroat trout in the Upper Yellowstone River. North American Journal of Fisheries Management, 37(4):743–755, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eder:2016:MEL

[EPM16] Brandon L. Eder, Mark A. Pegg, and Gerald E. Mestl. Modeling effects of length limit regulations on riverine populations of channel catfish. North American Journal of Fisheries Management, 36(1):140–146, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Evans:2011:RCW

[ERC^{+}11] Allen F. Evans, Daniel D. Roby, Ken Collis, Bradley M. Cramer, John A. Sheggeby, Lindsay J. Adrean, Daniel S. Battaglia, and Donald E. Lyons. Recovery of coded wire tags at a Caspian tern colony in San Francisco Bay: a technique to evaluate impacts of avian predation on juvenile salmonids. North American Journal of Fisheries Management, 31(1):79– 87, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Erman:2012:CRT

[Erm12] Don C. Erman. Comment: Rotenone toxicity to rainbow trout and several mountain stream insects. North American Journal of Fisheries Management, 32(1):53–59, February 2012. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Erhardt:2014:PCA

[ES14] John M. Erhardt and Dennis L. Scarnecchia. Population changes after 14 years of harvest closure on a migratory population of bull trout in Idaho. North American Journal of Fisheries Management, 34(3):482–492, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eberts:2017:VDP

[ES17] Rebecca L. Eberts and Christopher M. Somers. Venting and descending provide equivocal benefits for catch-and-release survival: Study design influences effectiveness more than barotrauma relief method. North American Journal of Fisheries Management, 37(3):612–623, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Esselman:2015:LPM

[ESL^{+}15] Peter C. Esselman, R. Jan Stevenson, Frank Lupi, Catherine M. Riseng, and Michael J. Wiley. Landscape prediction and mapping of game fish biomass, an ecosystem service of Michigan rivers. North American Journal of Fisheries Management, 35(2):302–320, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Engelbert:2016:DSS

[ETD16] Bryan S. Engelbert, Christopher A. Taylor, and Robert J. DiStefano. Development of standardized stream-dwelling crayfish sampling methods at site and drainage scales. North American Journal of Fisheries Management, 36(1):104–115, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Eberts:2017:ISL

[EWS^{+}17] Rebecca L. Eberts, Björn Wissel, Gavin L. Simpson, Stephen S. Crawford, Wendylee Stott, Robert H. Hanner, Richard G. Manzon, Joanna Y. Wilson, Douglas R. Boreham, and Christopher M. Somers. Isotopic structure of lake whitefish in Lake Huron: Evidence for regional and local populations based on resource use. North American Journal of Fisheries Management, 37(1):133–148, February 2017. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fetherman:2014:RFR

[FAW14] Eric R. Fetherman, Brian W. Avila, and Dana L. Winkelman. Raft and floating radio frequency identification (RFID) antenna systems for detecting and estimating abundance of PIT-tagged fish in rivers. North American Journal of Fisheries Management, 34(6):1065–1077, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fielder:2014:IAI

[FB14] David G. Fielder and James R. Bence. Integration of auxiliary information in statistical catch-at-age (SCA) analysis of the Saginaw Bay stock of walleye in Lake Huron. North American Journal of Fisheries Management, 34(5):970–987, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fried:2018:ARR

[FBB18] Lynda M. Fried, Matthew C. Boyer, and Meriel J. Brooks. Amphibian response to rotenone treatment of ten alpine lakes in Northwest Montana. North American Journal of Fisheries Management, 38(1):237–246, February 2018. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Feiner:2019:CHT

[FBB19] Zachary S. Feiner, Marybeth K. Brey, and Caroline Burgett. Consistently high trophic overlap between invasive white perch and native black crappies in southeastern reservoirs. North American Journal of Fisheries Management, 39(1):135–149, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flannery:2010:MSA

[FBC^{+}10] Blair G. Flannery, Terry D. Beacham, John R. Candy, Russell R. Holder, Gerald F. Maschmann, Eric J. Kretschmer, and John K. Wenburg. Mixed-stock analysis of Yukon River chum salmon: Application and validation in a complex fishery. North American Journal of Fisheries Management, 30 (5):1324–1338, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Feeken:2019:DMS

[FBD^{+}19] Stacey F. Feeken, Brett J. Bowersox, Marika E. Dobos, Matthew P. Corsi, Michael C. Quist, and Timothy Copeland. Distribution and movement of steelhead and anglers in the Clearwater River, Idaho. North American Journal of Fisheries Management, 39(5):1056–1072, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fischer:2019:HSC

[FBL^{+}19] Eric S. Fischer, Shannon E. Blackburn, Stephanie A. Liss, James S. Hughes, Huidong Li, and Zhiqun Daniel Deng. How small can we go? Evaluating survival, tag retention, and growth of juvenile Chinook salmon implanted with a new acoustic microtag. North American Journal of Fisheries Management, 39(6):1329–1336, December 2019. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fisch:2019:ESC

[FBM^{+}19] Nicholas C. Fisch, James R. Bence, Jared T. Myers, Eric K. Berglund, and Daniel L. Yule. Evaluating the sustainability of a cisco fishery in thunder Bay, Ontario, under alternative harvest policies. North American Journal of Fisheries Management, 39(3):543–559, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Feltz:2017:SCF

[FC17] Nicholas G. Feltz and Matthew J. Catalano. Statistical catchfree age-structured assessment to estimate year-class strength for inland recreational fisheries: a largemouth bass example. North American Journal of Fisheries Management, 37 (1):64–77, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flannery:2012:CRM

[FCE^{+}12] Blair G. Flannery, Penny A. Crane, John H. Eiler, Terry D. Beacham, Nick A. Decovich, William D. Templin, Ora L. Schlei, and John K. Wenburg. Comparison of radiotelemetry and microsatellites for determining the origin of Yukon River Chinook salmon. North American Journal of Fisheries Management, 32(4):720–730, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frechette:2013:BAA

[FCH^{+}13] Danielle Frechette, Alison L. Collins, James T. Harvey, Sean A. Hayes, David D. Huff, Andrew W. Jones, Nicolas A. Retford, Alina E. Langford, Jonathan W. Moore, Ann-Marie K. Osterback, William H. Satterthwaite, and Scott A. Shaffer. A bioenergetics approach to assessing potential impacts of avian predation on juvenile steelhead during freshwater rearing. North American Journal of Fisheries Management, 33(5):1024–1038, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Falke:2016:SPT

[FDHWP16] Jeffrey A. Falke, Jason B. Dunham, David Hockman-Wert, and Randy Pahl. A simple prioritization tool to diagnose impairment of stream temperature for coldwater fishes in the Great Basin. North American Journal of Fisheries Management, 36(1):147–160, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fitzsimons:2013:USI

[FDM^{+}13] John D. Fitzsimons, Alex Dalton, Brydon MacVeigh, Mark Heaton, Chris Wilson, and Dale C. Honeyfield. Use of stable isotopes to identify redds of putative hatchery and wild Atlantic salmon and evaluate their spawning habitat and egg thiamine status in a Lake Ontario tributary. North American Journal of Fisheries Management, 33(4):741–753, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Farmer:2015:FRF

[FF15] Nicholas A. Farmer and John T. Froeschke. Forecasting for recreational fisheries management: What’s the catch? North American Journal of Fisheries Management, 35(4):720–735, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Falahatkar:2013:ULS

[FG13] Bahram Falahatkar and Mohammad H. Tolouei Gilani. Use of laparoscopy for sex identification of Asp Aspius aspius. North American Journal of Fisheries Management, 33(5):894–899, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fenn:2013:EAS

[FGS13] Carlin M. Fenn, David C. Glover, and Brian C. Small. Efficacy of AQUI-S 20E as a sedative for handling and cortisol suppression in pallid sturgeon. North American Journal of Fisheries Management, 33(6):1172–1178, December 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Faust:2016:ECO

[FH16] Matthew D. Faust and Michael J. Hansen. Effects of consumption-oriented versus trophy-oriented fisheries on muskellunge population size structure in Northern Wisconsin. North American Journal of Fisheries Management, 36 (6):1336–1346, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Feldhaus:2016:ISR

[FHC16] Joseph W. Feldhaus, Timothy L. Hoffnagle, and Richard W. Carmichael. The influence of size at release on performance of Imnaha River Chinook salmon hatchery smolts. North American Journal of Fisheries Management, 36(2):363–374, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Freedman:2012:LBP

[FHE^{+}12] Jonathan A. Freedman, R. John H. Hoxmeier, Lisa M. Einfalt, Ronald C. Brooks, and David H. Wahl. Largemouth bass predation effect on stocked walleye survival in Illinois impoundments. North American Journal of Fisheries Management, 32(6):1039–1045, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fisher:2016:EEF

[FHH16] Emily A. Fisher, S. Alex Hesp, and Norman G. Hall. Exploring the effectiveness of a fisheries simulation model for communicating stock assessment information. North American Journal of Fisheries Management, 36(4):813–827, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fraley:2018:UIP

[FHL18] Kevin M. Fraley, Trevor B. Haynes, and J. Andrés López. Under-ice population density estimation of Alaska blackfish. North American Journal of Fisheries Management, 38(2):454– 461, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Faust:2015:MGP

[FILH15] Matthew D. Faust, Daniel A. Isermann, Mark A. Luehring, and Michael J. Hansen. Muskellunge growth potential in Northern Wisconsin: Implications for trophy management. North American Journal of Fisheries Management, 35(4): 765–774, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fielder:2016:USA

[FJB16] David G. Fielder, Michael L. Jones, and James R. Bence. Use of a structured approach to the analysis of management options and value of information for a recreationally exploited fish population: a case study of walleyes in Saginaw Bay. North American Journal of Fisheries Management, 36(2):407– 420, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Favrot:2016:ETW

[FK16] Scott D. Favrot and Thomas J. Kwak. Efficiency of two-way weirs and prepositioned electrofishing for sampling potamodromous fish migrations. North American Journal of Fisheries Management, 36(1):167–182, February 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fredricks:2012:FRS

[FMA^{+}12] Kim T. Fredricks, Jeffery R. Meinertz, Ryan D. Ambrose, Leanna M. Jackan, Jeremy K. Wise, and Mark P. Gaikowski. Feeding response of sport fish after electrical immobilization, chemical sedation, or both. North American Journal of Fisheries Management, 32(4):679–686, August 2012. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fisk:2019:ERH

[FMP19] J. Michael Fisk II, Clinton W. Morgeson, and Madison E. Polera. Evaluation of recreational hand-Crank electrofishing on introduced catfish species in Southeastern North Carolina. North American Journal of Fisheries Management, 39(1):150– 165, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frechette:2012:AAP

[FOH^{+}12] Danielle Frechette, Ann-Marie K. Osterback, Sean A. Hayes, Morgan H. Bond, Jonathan W. Moore, Scott A. Shaffer, and James T. Harvey. Assessing avian predation on juvenile salmonids using passive integrated transponder tag recoveries and mark–recapture methods. North American Journal of Fisheries Management, 32(6):1237–1250, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Frechette:2015:ARB

[FOH^{+}15] Danielle Frechette, Ann-Marie K. Osterback, Sean A. Hayes, Jonathan W. Moore, Scott A. Shaffer, Mark Pavelka, Clark Winchell, and James T. Harvey. Assessing the relationship between gulls Larus spp. and Pacific salmon in Central California using radiotelemetry. North American Journal of Fisheries Management, 35(4):775–788, August 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fischer:2019:UFA

[FQ19] Jesse R. Fischer and Michael C. Quist. Understanding fish assemblage structure in lentic ecosystems: Relative effects of abiotic factors and management legacies. North American Journal of Fisheries Management, 39(4):607–624, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Foley:2010:LRL

[FRH10] Carolyn J. Foley, Daniel J. Ryan, and Tomas O. Höök. Length reduction of larval yellow perch and freshwater amphipods in RNAlater solution. North American Journal of Fisheries Management, 30(5):1143–1148, October 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fincel:2016:ERS

[FRL16] Mark J. Fincel, William J. Radigan, and Christopher M. Longhenry. Entrainment of rainbow smelt through Oahe Dam during the 2011 Missouri River Flood. North American Journal of Fisheries Management, 36(4):844–851, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fincel:2017:RSG

[FSHR17] Mark J. Fincel, Mike J. Smith, Robert P. Hanten, and William J. Radigan. Recommendations for stocking gizzard shad in a large Upper Midwest reservoir. North American Journal of Fisheries Management, 37(3):599–604, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Finlayson:2010:CCE

[FSS^{+}10] Brian J. Finlayson, Rosalie Schnick, Don Skaar, William D. Horton, Daniel Duffield, Jon D. Anderson, Leo DeMong, and Jarle Steinkjer. Comment: Comparative effects of rotenone and antimycin on macroinvertebrate diversity in two streams in Great Basin National Park, Nevada. North American Journal of Fisheries Management, 30(5):1126–1128, October 2010.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Finlayson:2010:RTR

[FSV10] Brian Finlayson, William L. Somer, and Mark R. Vinson. Rotenone toxicity to rainbow trout and several mountain stream insects. North American Journal of Fisheries Management, 30(1):102–111, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Finlayson:2012:RTR

[FSV12] Brian Finlayson, William L. Somer, and Mark R. Vinson. Rotenone toxicity to rainbow trout and several mountain stream insects: Response to comment. North American Journal of Fisheries Management, 32(1):60–64, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flammang:2016:CTM

[FSW16] Mark Flammang, Randall D. Schultz, and Michael J. Weber. Comparison of three methods for sampling panfish in Iowa impoundments. North American Journal of Fisheries Management, 36(6):1347–1357, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Fernholz:2018:HWL

[FSW18] Savannah Fernholz, Andrea Sylvia, and Michael J. Weber. Hook wound longevity and use as an indicator of largemouth bass catch-and-release angling pressure. North American Journal of Fisheries Management, 38(3):759–768, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Froeschke:2013:SPM

[FTSF13] Bridgette F. Froeschke, Philippe Tissot, Gregory W. Stunz, and John T. Froeschke. Spatiotemporal predictive models for juvenile southern flounder in Texas estuaries. North American Journal of Fisheries Management, 33(4):817–828, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Faust:2017:UES

[FVH^{+}17] Matthew D. Faust, Christopher S. Vandergoot, Eric T. Hostnik, Thomas R. Binder, Julia L. Mida Hinderer, Jessica T. Ives, and Charles C. Krueger. Use of electricity to sedate lake trout for intracoelomic implantation of electronic transmitters. North American Journal of Fisheries Management, 37(4):768–777, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ferguson:2015:RSC

[FWG15] Greg J. Ferguson, Tim M. Ward, and Daniel Gorman. Recovery of a surf clam Donax deltoides population in Southern Australia: Successful outcomes of fishery-independent surveys. North American Journal of Fisheries Management, 35 (6):1185–1195, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Forbes:2015:SSV

[FWR^{+}15a] Jamin P. Forbes, Robyn J. Watts, Wayne A. Robinson, Lee J. Baumgartner, Micheal S. Allen, Prue McGuffie, Leo M. Cameron, and David A. Crook. System-specific variability in Murray cod and golden perch maturation and growth influences fisheries management options. North American Journal of Fisheries Management, 35(6):1226–1238, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Forbes:2015:RFE

[FWR^{+}15b] Jamin P. Forbes, Robyn J. Watts, Wayne A. Robinson, Lee J. Baumgartner, Aldo S. Steffe, and Jeff J. Murphy. Recreational fishing effort, catch, and harvest for Murray cod and golden perch in the Murrumbidgee River, Australia. North American Journal of Fisheries Management, 35(4):649–658, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flammang:2014:LEB

[FWT14] Mark K. Flammang, Michael J. Weber, and Megan D. Thul. Laboratory evaluation of a bioacoustic bubble strobe light barrier for reducing walleye escapement. North American Journal of Fisheries Management, 34(6):1047–1054, October 2014.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Flaherty-Walia:2016:STD

[FWWTJD16] Kerry E. Flaherty-Walia, Brent L. Winner, Amanda J. Tyler-Jedlund, and John P. Davis. Short-term discard mortality estimates for gray snapper in a West-Central Florida Estuary and adjacent nearshore Gulf of Mexico waters. North American Journal of Fisheries Management, 36(2):329–340, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goble:2018:FDH

[GAH^{+}18] Cameron W. Goble, Nancy A. Auer, Casey J. Huckins, Brian M. Danhoff, J. Marty Holtgren, and Stephanie A. Ogren. Fish distributions and habitat associations in Manistee River, Michigan, tributaries: Implications for Arctic grayling restoration. North American Journal of Fisheries Management, 38(2):469–486, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallagher:2010:PSS

[GAWC10] Sean P. Gallagher, Peter B. Adams, David W. Wright, and Barry W. Collins. Performance of spawner survey techniques at low abundance levels. North American Journal of Fisheries Management, 30(5):1086–1097, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

George:2018:CGR

[GBFR18] Scott D. George, Barry P. Baldigo, Michael J. Flaherty, and Eileen A. Randall. Changes in growth of rainbow trout in a Catskill Mountain reservoir following alewife and white perch introductions. North American Journal of Fisheries Management, 38(5):1027–1038, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gibson:2011:PBL

[GBHO11] A. Jamie F. Gibson, Heather D. Bowlby, David C. Hardie, and Patrick T. O’Reilly. Populations on the brink: Low abundance of Southern Upland Atlantic salmon in Nova Scotia, Canada. North American Journal of Fisheries Management, 31(4):733–741, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gaeta:2013:CRR

[GBL^{+}13] Jereme W. Gaeta, Ben Beardmore, Alexander W. Latzka, Bill Provencher, and Stephen R. Carpenter. Catch-and-release rates of sport fishes in Northern Wisconsin from an angler diary survey. North American Journal of Fisheries Management, 33(3):606–614, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gibson:2015:DEE

[GBL15] A. Jamie F. Gibson, Heather D. Bowlby, and Alex L. Levy. Dynamics of endangered Eastern Cape Breton Atlantic salmon populations. North American Journal of Fisheries Management, 35(2):372–387, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gurshin:2014:IUF

[GBTL14] Christopher W. D. Gurshin, Matthew P. Balge, Michael M. Taylor, and Benjamin E. Lenz. Importance of ultrasonic field direction for guiding juvenile blueback herring past hydroelectric turbines. North American Journal of Fisheries Management, 34(6):1242–1258, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Groves:2013:ESS

[GCA^{+}13] Phillip A. Groves, James A. Chandler, Brad Alcorn, Tracy J. Richter, William P. Connor, Aaron P. Garcia, and Steven M. Bradbury. Evaluating salmon spawning habitat capacity using redd survey data. North American Journal of Fisheries Management, 33(4):707–716, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gandy:2018:ISS

[GCLC18] Ryan L. Gandy, Claire E. Crowley, Erin H. Leone, and Charles R. Crawford. Increasing the selectivity of the stone crab Menippe mercenaria trap by the addition of a cull ring. North American Journal of Fisheries Management, 38(6): 1275–1283, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallagher:2010:HCP

[GD10] Colin P. Gallagher and Terry A. Dick. Historical and current population characteristics and subsistence harvest of Arctic char from the Sylvia Grinnell River, Nunavut, Canada. North American Journal of Fisheries Management, 30(1):126–141, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gonzalez:2017:LWI

[GDLM17] Rosalinda Gonzalez, Jason Dunham, Scott Lightcap, and Jeff McEnroe. Large wood and instream habitat for juvenile Coho salmon and larval lampreys in a Pacific Northwest stream. North American Journal of Fisheries Management, 37(4): 683–699, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerig:2014:HSM

[GDP14] Brandon Gerig, Michael J. Dodrill, and William E. Pine III. Habitat selection and movement of adult humpback chub in the Colorado River in Grand Canyon, Arizona, during an experimental steady flow release. North American Journal of Fisheries Management, 34(1):39–48, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerow:2010:BRL

[Ger10] Kenneth G. Gerow. Biases with the regression line percentile method and the fallacy of a single standard weight. North American Journal of Fisheries Management, 30(3):679–690, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gerow:2011:CAL

[Ger11] Kenneth G. Gerow. Comment: Assessing length-related biases in standard weight equations. North American Journal of Fisheries Management, 31(4):656–660, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Greszkiewicz:2018:EPF

[GF18] Martyna Greszkiewicz and Dariusz P. Fey. Effect of preservation in formalin and alcohol on the growth rate estimates of larval northern pike. North American Journal of Fisheries Management, 38(3):601–605, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gross:2015:PUD

[GFC^{+}15] Jackson Gross, Bahram Farokhkish, Matthew Cornacione, Steven Shaw, Phong L. Nguyen, and Theodore B. Henry. Potential use of direct current electric fields to eradicate rainbow trout embryos from freshwater ecosystems. North American Journal of Fisheries Management, 35(5):871–879, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grubbs:2018:BBT

[GFM^{+}18] S. P. Grubbs, H. Funkhouser, P. Myer, M. Arendt, J. Schwenter, and R. M. Chambers. To BRD or not to BRD? A test of bycatch reduction devices (BRDs) for the blue crab fishery. North American Journal of Fisheries Management, 38 (1):18–23, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Giannetto:2012:PES

[GFP^{+}12] Daniela Giannetto, Elisabetta Franchi, Laura Pompei, Massimo Lorenzoni, Stefano Porcellotti, and Lorenzo Tancioni. Proposed empirical standard weight equation for brook chub Squalius lucumonis. North American Journal of Fisheries Management, 32(3):428–435, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallaway:2017:UDB

[GGC17] Benny J. Gallaway, W. J. Gazey, and J. G. Cole. An updated description of the benefits and consequences of red snapper shrimp trawl bycatch management actions in the Gulf of Mexico. North American Journal of Fisheries Management, 37 (2):414–419, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gazey:2014:AFM

[GGCF14] William J. Gazey, Benny J. Gallaway, John G. Cole, and David A. Fournier. Accounting for fishing mortality when comparing density-dependent with density-independent mortality in Gulf of Mexico red snapper: Response to comment. North American Journal of Fisheries Management, 34(2):352– 358, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goulette:2017:AVP

[GH17] Graham S. Goulette and James P. Hawkes. Altering vertical placement of hydroacoustic receivers for improved efficiency in coldwater estuary zones. North American Journal of Fisheries Management, 37(5):981–988, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Graham:2017:QWS

[GHK17] Jake D. Graham, Andrew W. Hafs, and Anthony J. Kennedy. Quantification of walleye spawning substrate in a Northern Minnesota River using side-scan sonar. North American Journal of Fisheries Management, 37(2):420–428, April 2017. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Graham:2019:DBM

[GHK19] Jake D. Graham, Andrew W. Hafs, and Anthony J. Kennedy. Does a bioenergetics model accurately predict fish consumption by American white pelicans? A case study on walleyes in the Tamarac River, Minnesota. North American Journal of Fisheries Management, 39(3):485–497, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gonzales:2012:UFN

[GHL12] Eric J. Gonzales, Grace M. Haggerty, and Anders Lundahl. Using fyke-net capture data to assess daily trends in abundance of spawning Rio Grande silvery minnow. North American Journal of Fisheries Management, 32(3):544–547, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gibbs:2015:EAN

[GKMC15] W. Keith Gibbs, Matt A. Kulp, Steve E. Moore, and S. Bradford Cook. Effects of antimycin and its neutralizing agent, potassium permanganate, on aquatic macroinvertebrates in a Great Smoky Mountains National Park watershed. North American Journal of Fisheries Management, 35(6):1239–1251, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goclowski:2013:MHD

[GKS13] Matthew R. Goclowski, Adam J. Kaeser, and Steven M. Sammons. Movement and habitat differentiation among adult shoal bass, largemouth bass, and spotted bass in the Upper Flint River, Georgia. North American Journal of Fisheries Management, 33(1):56–70, February 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Griswold:2011:SES

[GKT11] Robert G. Griswold, André E. Kohler, and Doug Taki. Survival of endangered Snake River sockeye salmon smolts from three Idaho lakes: Relationships with parr size at release, parr growth rate, smolt size, discharge, and travel time. North American Journal of Fisheries Management, 31(5):813–825, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goulette:2016:NBI

[GL16] Graham S. Goulette and Christine A. Lipsky. Nonlethal batch identification of Atlantic salmon using coded wire tags. North American Journal of Fisheries Management, 36(5):1084–1089, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Geist:2013:ETD

[GLCD13] David R. Geist, Timothy J. Linley, Valerie Cullinan, and Zhiqun Deng. The effects of total dissolved gas on chum salmon fry survival, growth, gas bubble disease, and seawater tolerance. North American Journal of Fisheries Management, 33(1):200–215, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Geist:2018:JCS

[GLH^{+}18] David R. Geist, Stephanie A. Liss, Ryan A. Harnish, Katherine A. Deters, Richard S. Brown, Zhiqun Daniel Deng, Jayson J. Martinez, Robert P. Mueller, and John R. Stephenson. Juvenile Chinook salmon survival when exposed to simulated dam passage after being implanted with a new microacoustic transmitter. North American Journal of Fisheries Management, 38(4):940–952, August 2018. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gardunio:2012:STR

[GM12] E. I. Gardunio and C. A. Myrick. Short-term retention rates of passive integrated transponders surgically implanted in burbot and the effects on survival. North American Journal of Fisheries Management, 32(5):1000–1004, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Grixti:2010:UAB

[GMB10] Daniel Grixti, Alexander Morison, and Justin D. Bell. Undersized Acanthopagrus butcheri caught and released from commercial gill nets show high survival rates in the Gippsland Lakes, Southeastern Australia. North American Journal of Fisheries Management, 30(3):723–731, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gosch:2015:ASA

[GMG^{+}15] Nathan J. C. Gosch, Marcus L. Miller, Todd R. Gemeinhardt, Schuyler J. Sampson, and Joseph L. Bonneau. Age-0 sturgeon accessibility to constructed and modified chutes in the Lower Missouri River. North American Journal of Fisheries Management, 35(1):75–85, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Galloway:2016:FAF

[GMG^{+}16a] Benjamin T. Galloway, Clint C. Muhlfeld, Christopher S. Guy, Christopher C. Downs, and Wade A. Fredenberg. A framework for assessing the feasibility of native fish conservation translocations: Applications to threatened bull trout. North American Journal of Fisheries Management, 36(4):754–768, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Graves:2016:EAE

[GMG16b] John E. Graves, Benjamin J. Marcek, and William M. Goldsmith. Effects of air exposure on postrelease mortality rates of white marlin caught in the U.S. offshore recreational fishery. North American Journal of Fisheries Management, 36(6): 1221–1228, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Giannetto:2012:ESW

[GPLT12] Daniela Giannetto, Laura Pompei, Massimo Lorenzoni, and Ali Serhan Tarkan. Empirical standard weight equation for the Aegean chub Squalius fellowesii, an endemic freshwater fish species of Western Anatolia, Turkey. North American Journal of Fisheries Management, 32(6):1102–1107, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gibson:2010:DWH

[GPP^{+}10] Suzanne M. Gibson, William F. Patterson III, Ronald P. Phelps, William P. Patterson, and Zhongxing Chen. Distinguishing wild from hatchery-produced juvenile red snapper with otolith chemical signatures. North American Journal of Fisheries Management, 30(5):1176–1186, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goldman:2010:ESC

[GRCR10] Sarah F. Goldman, Angela R. Rourk, Mark R. Collins, and Marcel J. M. Reichert. Effects of simulated commercial harvest on mummichogs in tidal creeks in South Carolina, USA. North American Journal of Fisheries Management, 30(4):921–927, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gresswell:2011:BSM

[Gre11] Robert E. Gresswell. Biology, status, and management of the Yellowstone cutthroat trout. North American Journal of Fisheries Management, 31(5):782–812, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goodman:2017:SEI

[GRR^{+}17] Damon H. Goodman, Stewart B. Reid, Rene C. Reyes, Brandon J. Wu, and Brent B. Bridges. Screen efficiency and implications for losses of lamprey macrophthalmia at California’s largest water diversions. North American Journal of Fisheries Management, 37(1):30–40, February 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gibson-Reinemer:2016:EEE

[GRSF^{+}16] Daniel K. Gibson-Reinemer, David R. Stewart, Mark W. Fritts, Jason A. DeBoer, and Andrew F. Casper. Estimating the effects of environmental variables and gear type on the detection and occupancy of large-river fishes in a standardized sampling program using multiseason Bayesian mixture models. North American Journal of Fisheries Management, 36(6): 1445–1456, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goethel:2018:ESS

[GSCCP18] Daniel R. Goethel, Matthew W. Smith, Shannon L. Cass-Calay, and Clay E. Porch. Establishing stock status determination criteria for fisheries with high discards and uncertain recruitment. North American Journal of Fisheries Management, 38(1):120–139, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gold:2011:GDE

[GSCR11] John R. Gold, Eric Saillant, Nancie J. Cummings, and Mark A. Renshaw. Genetic divergence and effective size among lane snapper in U.S. waters of the Western Atlantic Ocean. North American Journal of Fisheries Management, 31(2):209–223, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Goldsmith:2018:CPV

[GSG18] William M. Goldsmith, Andrew M. Scheld, and John E. Graves. Characterizing the preferences and values of U.S. recreational Atlantic bluefin tuna anglers. North American Journal of Fisheries Management, 38(3):680–697, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gwinn:2019:ATF

[GTB^{+}19] Daniel C. Gwinn, Charles R. Todd, Paul Brown, Taylor L. Hunt, Gavin Butler, Adrian Kitchingman, John D. Koehn, and Brett Ingram. Assessing a threatened fish species under budgetary constraints: Evaluating the use of existing monitoring data. North American Journal of Fisheries Management, 39(2):315–327, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gurshin:2012:TSM

[Gur12] Christopher W. D. Gurshin. Target strength measurements of juvenile blueback herring from the Mohawk River, New York. North American Journal of Fisheries Management, 32(2):381– 386, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Graeb:2010:ASA

[GWB^{+}10] Brian D. S. Graeb, David W. Willis, Neil Billington, Rachael N. Koigi, and Justin A. VanDeHey. Age-structured assessment of walleyes, saugers, and naturally produced hybrids in three Missouri River reservoirs. North American Journal of Fisheries Management, 30(4):887–897, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gallagher:2010:RAM

[GWCA10] Sean P. Gallagher, David W. Wright, Barry W. Collins, and Peter B. Adams. A regional approach for monitoring salmonid status and trends: Results from a pilot study in Coastal Mendocino County, California. North American Journal of Fisheries Management, 30(5):1075–1085, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gorsky:2013:EES

[GZ13] Dimitry Gorsky and Joseph Zydlewski. Experimental evaluation of size-dependent predation by adult post-spawned rainbow smelt on larval lake whitefish. North American Journal of Fisheries Management, 33(1):163–169, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunt:2010:ESF

[HADG10] Taylor L. Hunt, Micheal S. Allen, John Douglas, and Anne Gason. Evaluation of a sport fish stocking program in lakes of the Southern Murray–Darling Basin, Australia. North American Journal of Fisheries Management, 30(3):805–811, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2019:SDR

[Han19] Adam G. Hansen. Size-dependent retention of pelagic-oriented Kokanee in multimesh gill nets. North American Journal of Fisheries Management, 39(5):921–932, October 2019. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hartman:2019:SEB

[HAS^{+}19] Kyle J. Hartman, Ross G. Andrew, Christopher W. Schwinghamer, David W. Thorne, and Jeremy S. Webster. Springtime exploitation of brook trout by anglers in remote headwater streams of central appalachia. North American Journal of Fisheries Management, 39(2):403–411, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hargrove:2018:GPM

[HAW^{+}18] John S. Hargrove, Micheal S. Allen, Olaf L. F. Weyl, Chelsey A. C. Crandall, and James D. Austin. Global patterns in the motivations and behaviors of tournament anglers targeting bedding bass. North American Journal of Fisheries Management, 38(2):334–345, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haxton:2015:CSS

[Hax15] Tim Haxton. Characteristics and spatial segregation of sympatric saugers and walleyes in the Ottawa River, Canada. North American Journal of Fisheries Management, 35(3):418– 430, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoagstrom:2010:NRW

[HB10] C. W. Hoagstrom and C. R. Berry, Jr. The native range of walleyes in the Missouri River drainage. North American Journal of Fisheries Management, 30(3):642–654, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holt:2011:EBP

[HB11] Carrie A. Holt and Michael J. Bradford. Evaluating benchmarks of population status for Pacific salmon. North American Journal of Fisheries Management, 31(2):363–378, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

He:2012:ISS

[HB12] Pingguo He and Vincent Balzano. Improving size selectivity of shrimp trawls in the Gulf of Maine with a modified dualgrid size-sorting system. North American Journal of Fisheries Management, 32(6):1113–1122, December 2012. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hennen:2014:MSD

[HB14a] Matthew J. Hennen and Michael L. Brown. Movement and spatial distribution of common carp in a South Dakota glacial lake system: Implications for management and removal. North American Journal of Fisheries Management, 34(6):1270–1281, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hill:2014:ODI

[HB14b] Angela A. Hill and Kevin R. Bestgen. Otolith daily increment deposition in age-0 smallmouth bass reared in constant and fluctuating water temperatures. North American Journal of Fisheries Management, 34(4):774–779, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hafen:2015:ABA

[HB15] Konrad Hafen and Phaedra Budy. Agonistic behavior among three stocked trout species. North American Journal of Fisheries Management, 35(3):551–556, June 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haworth:2016:DIV

[HB16] Matthew R. Haworth and Kevin R. Bestgen. Daily increment validation and effects of streamflow variability and water temperature on growth of age-0 flathead chub. North American Journal of Fisheries Management, 36(4):744–753, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Humston:2012:CSH

[HBA^{+}12] Robert Humston, Kelly A. Bezold, Nathaniel D. Adkins, Ryan J. Elsey, Jessica Huss, Brent A. Meekins, Paul R. Cabe, and Timothy L. King. Consequences of stocking headwater impoundments on native populations of brook trout in tributaries. North American Journal of Fisheries Management, 32 (1):100–108, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hunt:2017:DFD

[HBD^{+}17] Len M. Hunt, Allison E. Bannister, D. Andrew R. Drake, Shannon A. Fera, and Timothy B. Johnson. Do fish drive recreational fishing license sales? North American Journal of Fisheries Management, 37(1):122–132, February 2017. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hockersmith:2013:LBB

[HBDA13] Eric E. Hockersmith, Gabriel Brooks, Nathan D. Dumdei, and Stephen Achord. A lightweight battery for backpack electrofishing. North American Journal of Fisheries Management, 33(2):265–268, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holbrook:2012:RPE

[HBLL12] S. Chad Holbrook, William D. Byars, Scott D. Lamprecht, and Jean K. Leitner. Retention and physiological effects of surgically implanted telemetry transmitters in blue catfish. North American Journal of Fisheries Management, 32(2):276–281, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hessenauer:2014:GPR

[HBS14a] Jan-Michael Hessenauer, Mary Tate Bremigan, and Kim T. Scribner. Genetic pedigree reconstruction detects bias in largemouth bass nest sampling procedures. North American Journal of Fisheries Management, 34(1):175–183, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hessenauer:2014:URT

[HBS14b] Jan-Michael Hessenauer, Mary Tate Bremigan, and Kim T. Scribner. Using the right tool for the right job should include validation when possible: Response to comment. North American Journal of Fisheries Management, 34(6):1207–1210, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hall:2019:USA

[HBS^{+}19] Edward P. Hall, Timothy F. Bonvechio, Stephanie L. Shaw, Micheal S. Allen, Jerry Brown, and Larry Pugh. Using specialized angling to assess a trophy Florida bass fishery at calling Panther Lake, Mississippi. North American Journal of Fisheries Management, 39(3):589–595, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hogan:2013:FMC

[HCH13] Fiona Hogan, Steve Cadrin, and Alyssa Haygood. Fishery management complexes: An impediment or aid to sustainable harvest? A discussion based on the northeast skate complex. North American Journal of Fisheries Management, 33(2):406– 421, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hartill:2012:SCD

[HCL^{+}12] Bruce W. Hartill, Martin Cryer, Jeremy M. Lyle, Eugene B. Rees, Karina L. Ryan, Aldo S. Steffe, Stephen M. Taylor, Laurie West, and Brent S. Wise. Scale- and context-dependent selection of recreational harvest estimation methods: The Australasian experience. North American Journal of Fisheries Management, 32(1):109–123, February 2012. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hamazaki:2014:AGM

[HD14] Toshihide Hamazaki and Nick DeCovich. Application of the genetic mark–recapture technique for run size estimation of Yukon River Chinook salmon. North American Journal of Fisheries Management, 34(2):276–286, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hannah:2012:GRC

[HDB^{+}12] Lucie Hannah, Nicholas Duprey, John Blackburn, Claudia M. Hand, and Christopher M. Pearce. Growth rate of the California sea cucumber Parastichopus californicus: Measurement accuracy and relationships between size and weight metrics. North American Journal of Fisheries Management, 32(1):167– 176, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hagen:2010:ENS

[HDKB10] John Hagen, Scott Decker, Josh Korman, and Robert G. Bison. Effectiveness of night snorkeling for estimating steelhead parr abundance in a large river basin. North American Journal of Fisheries Management, 30(5):1303–1314, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hoxmeier:2015:BTD

[HDM15] R. John H. Hoxmeier, Douglas J. Dieterman, and Loren M. Miller. Brook trout distribution, genetics, and population characteristics in the Driftless Area of Minnesota. North American Journal of Fisheries Management, 35(4):632–648, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hostetter:2015:IIF

[HEL^{+}15] Nathan J. Hostetter, Allen F. Evans, Frank J. Loge, Rolland R. O’Connor, Bradley M. Cramer, Derek Fryer, and Ken Collis. The influence of individual fish characteristics on survival and detection: Similarities across two salmonid species. North American Journal of Fisheries Management, 35(5):1034–1045, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hennen:2015:HSW

[Hen15] Daniel R. Hennen. How should we harvest an animal that can live for centuries? North American Journal of Fisheries Management, 35(3):512–527, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

He:2012:LTS

[HER^{+}12] Ji X. He, Mark P. Ebener, Stephen C. Riley, Adam Cottrill, Adam Kowalski, Scott Koproski, Lloyd Mohr, and James E. Johnson. Lake trout status in the main basin of Lake Huron, 1973–2010. North American Journal of Fisheries Management, 32(2):402–412, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Henderson:2013:DNS

[HF13] Mark J. Henderson and Mary C. Fabrizio. Detecting noncompliance in the summer flounder recreational fishery using a mark–recapture growth model. North American Journal of Fisheries Management, 33(5):1039–1048, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hatch:2013:STR

[HFB^{+}13] Douglas R. Hatch, David E. Fast, William J. Bosch, Joseph W. Blodgett, John M. Whiteaker, Ryan Branstetter, and Andrew L. Pierce. Survival and traits of reconditioned kelt steelhead Oncorhynchus mykiss in the Yakima River, Washington. North American Journal of Fisheries Management, 33(3):615– 625, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haxton:2015:AME

[HFCH15] Tim Haxton, Mike Friday, Tim Cano, and Charles Hendry. Assessing the magnitude of effect of hydroelectric production on lake sturgeon abundance in Ontario. North American Journal of Fisheries Management, 35(5):930–941, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2012:RBA

[HFH12] Jonathan F. Hansen, Andrew H. Fayram, and Joseph M. Hennessy. The relationship between age-0 walleye density and adult year-class strength across Northern Wisconsin. North American Journal of Fisheries Management, 32(4):663–670, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2018:ERB

[HFH18] Jonathan F. Hansen, Andrew H. Fayram, and Joseph M. Hennessy. Erratum: The relationship between age-0 walleye density and adult year-class strength across Northern Wisconsin. North American Journal of Fisheries Management, 38(4): 971–974, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2011:NMR

[HFN11] Michael J. Hansen, Andrew H. Fayram, and Steven P. Newman. Natural mortality in relation to age and fishing mortality on walleyes in Escanaba Lake, Wisconsin, during 1956–2009.

North American Journal of Fisheries Management, 31(3):506– 514, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hodge:2017:EFR

[HFRH17] Brian W. Hodge, Eric R. Fetherman, Kevin B. Rogers, and Richard Henderson. Effectiveness of a fishway for restoring passage of Colorado River cutthroat trout. North American Journal of Fisheries Management, 37(6):1332–1340, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Gu:2013:CDA

[hGgXfC^{+}13] Pang hua Gu, Jian guo Xiang, Yi fang Chen, Yan liang Li, Jun Tang, Song guang Xie, and Yushun Chen. A comparison of different age estimation methods for the northern snakehead. North American Journal of Fisheries Management, 33 (5):994–999, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Homel:2015:LHD

[HGK15] K. M. Homel, R. E. Gresswell, and J. L. Kershner. Life history diversity of Snake River finespotted cutthroat trout: Managing for persistence in a rapidly changing environment. North American Journal of Fisheries Management, 35(4):789–801, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hafs:2010:SBS

[HGW10] Andrew W. Hafs, Charles J. Gagen, and J. Keith Whalen. Smallmouth bass summer habitat use, movement, and survival in response to low flow in the Illinois Bayou, Arkansas. North American Journal of Fisheries Management, 30(2):604–612, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hedden:2016:IFC

[HGW16] Skyler C. Hedden, Keith B. Gido, and James E. Whitney. Introduced flathead catfish consumptive demand on native fishes of the Upper Gila River, New Mexico. North American Journal of Fisheries Management, 36(1):55–61, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:2010:EMI

[HH10] Julianne E. Harris and Joseph E. Hightower. Evaluation of methods for identifying spawning sites and habitat selection for alosines. North American Journal of Fisheries Management, 30(2):386–399, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:2011:IAS

[HH11a] Julianne E. Harris and Joseph E. Hightower. Identification of American shad spawning sites and habitat use in the pee Dee River, North Carolina and South Carolina. North American Journal of Fisheries Management, 31(6):1019–1033, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:2011:MPA

[HH11b] Julianne E. Harris and Joseph E. Hightower. Movement patterns of American shad transported upstream of dams on the Roanoke River, North Carolina and Virginia. North American Journal of Fisheries Management, 31(2):240–256, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:2011:SHS

[HH11c] Julianne E. Harris and Joseph E. Hightower. Spawning habitat selection of hickory shad. North American Journal of Fisheries Management, 31(3):495–505, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hardin:2012:RAB

[HH12] Scott Hardin and Jeffrey E. Hill. Risk analysis of barramundi perch Lates calcarifer aquaculture in Florida. North American Journal of Fisheries Management, 32(3):577–585, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hughes:2015:CSB

[HH15] Jacob B. Hughes and Joseph E. Hightower. Combining splitbeam and dual-frequency identification sonars to estimate abundance of anadromous fishes in the Roanoke River, North Carolina. North American Journal of Fisheries Management, 35(2):229–240, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hafs:2017:SCC

[HH17] Andrew W. Hafs and Kyle J. Hartman. Seasonal changes in condition of Appalachian brook trout. North American Journal of Fisheries Management, 37(1):196–206, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hurst:2012:DRI

[HHB12] Charlene N. Hurst, Richard A. Holt, and Jerri L. Bartholomew. Dam removal and implications for fish health: Ceratomyxa shasta in the Williamson River, Oregon, USA. North American Journal of Fisheries Management, 32(1):14–23, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2015:IMP

[HHCH15] Gretchen J. A. Hansen, Joseph M. Hennessy, Thomas A. Cichosz, and Steven W. Hewett. Improved models for predicting walleye abundance and setting safe harvest quotas in Northern Wisconsin lakes. North American Journal of Fisheries Management, 35(6):1263–1277, December 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hosack:2014:DIU

[HHH14] Michael A. Hosack, Michael J. Hansen, and Ned J. Horner. Dynamics of an introduced and unexploited lake whitefish population in Lake Pend Oreille, Idaho. North American Journal of Fisheries Management, 34(5):1014–1027, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hayes:2015:ETA

[HHMQ15] John W. Hayes, Joe Hay, Iain Maxwell, and Aaron Quarterman. Estimating trout abundance with cataraft-mounted dual-frequency identification sonar: a comparison with drift diving. North American Journal of Fisheries Management, 35 (3):528–536, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hamel:2012:TRB

[HHP12a] Martin J. Hamel, Jeremy J. Hammen, and Mark A. Pegg. Tag retention of T-bar anchor tags and passive integrated transponder tags in shovelnose sturgeon. North American Journal of Fisheries Management, 32(3):533–538, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hayes:2012:SSS

[HHP^{+}12b] Sean A. Hayes, Chad V. Hanson, Devon E. Pearse, Morgan H. Bond, John Carlos Garza, and R. Bruce MacFarlane. Should I stay or should I go? The influence of genetic origin on emigration behavior and physiology of resident and anadromous juvenile Oncorhynchus mykiss. North American Journal of Fisheries Management, 32(4):772–780, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hammen:2018:PCS

[HHR^{+}18] Jeremy J. Hammen, Martin J. Hamel, Matthew L. Rugg, Edward J. Peters, and Mark A. Pegg. Population characteristics of shovelnose sturgeon during low- and high-water conditions in the Lower Platte River, Nebraska. North American Journal of Fisheries Management, 38(2):308–315, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hodge:2015:EPP

[HHRB15] Brian W. Hodge, Richard Henderson, Kevin B. Rogers, and Kyle D. Battige. Efficacy of portable PIT detectors for tracking long-term movement of Colorado River cutthroat trout in a small montane stream. North American Journal of Fisheries Management, 35(3):605–610, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hutt:2013:EVR

[HHS^{+}13] Clifford P. Hutt, Kevin M. Hunt, Susan F. Steffen, Stephen C. Grado, and L. E. Miranda. Economic values and regional economic impacts of recreational fisheries in Mississippi reservoirs. North American Journal of Fisheries Management, 33 (1):44–55, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hutt:2013:ECA

[HHSB13] Clifford P. Hutt, Kevin M. Hunt, J. Warren Schlechte, and David L. Buckmeier. Effects of catfish angler catch-related attitudes on fishing trip preferences. North American Journal of Fisheries Management, 33(5):965–976, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanson:2013:EWJ

[HHT^{+}13] S. Dale Hanson, Mark E. Holey, Ted J. Treska, Charles R. Bronte, and Ted H. Eggebraaten. Evidence of wild juvenile lake trout recruitment in Western Lake Michigan. North American Journal of Fisheries Management, 33(1):186–191, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haglund:2016:WPF

[HIS16] Justin M. Haglund, Daniel A. Isermann, and Greg G. Sass. Walleye population and fishery responses after elimination of legal harvest on Escanaba Lake, Wisconsin. North American Journal of Fisheries Management, 36(6):1315–1324, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harvey:2014:QUJ

[HJB14] Brett N. Harvey, David P. Jacobson, and Michael A. Banks. Quantifying the uncertainty of a juvenile Chinook salmon race identification method for a mixed-race stock. North American Journal of Fisheries Management, 34(6):1177–1186, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hudson:2011:PES

[HJK11] J. Michael Hudson, Jeffrey R. Johnson, and Boyd Kynard. A portable electronarcosis system for anesthetizing salmonids and other fish. North American Journal of Fisheries Management, 31(2):335–339, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hilderbrand:2017:SSP

[HK17] Robert H. Hilderbrand and David C. Kazyak. Spatiotemporal stability patterns of brook trout abundance and implications for stream research and monitoring. North American Journal of Fisheries Management, 37(2):353–362, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Habera:2010:TPD

[HKMH10] James W. Habera, Matt A. Kulp, Stephen E. Moore, and Theodore B. Henry. Three-pass depletion sampling accuracy of two electric fields for estimating trout abundance in a lowconductivity stream with limited habitat complexity. North American Journal of Fisheries Management, 30(3):757–766, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Huhn:2014:IES

[HKP^{+}14] Daniel Hühn, Thomas Klefoth, Thilo Pagel, Petr Zajicek, and Robert Arlinghaus. Impacts of external and surgery-based tagging techniques on small northern pike under field conditions. North American Journal of Fisheries Management, 34 (2):322–334, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hill:2015:RSA

[HL15] Jeffrey E. Hill and Katelyn M. Lawson. Risk screening of arapaima, a new species proposed for aquaculture in Florida. North American Journal of Fisheries Management, 35(5):885– 894, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harstad:2018:WRT

[HLM^{+}18] Deborah L. Harstad, Donald A. Larsen, Joseph Miller, Ian Adams, Dina K. Spangenberg, Shelly Nance, Larissa Rohrbach, Joshua G. Murauskas, and Brian R. Beckman. Winter-rearing temperature affects growth profiles, age of maturation, and smolt-to-adult returns for yearling summer Chinook salmon in the Upper Columbia River Basin. North American Journal of Fisheries Management, 38(4):867–885, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hines:2017:DSM

[HLS^{+}17] David Hines, Martin Liermann, Tiffany Seder, Brian Cluer, George Pess, and Casey Schoenebeck. Diel shifts in microhabitat selection of steelhead and Coho salmon fry. North American Journal of Fisheries Management, 37(5):989–998, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hausch:2012:RHS

[HM12] Stephen J. Hausch and Michael C. Melnychuk. Residualization of hatchery steelhead: a meta-analysis of hatchery practices. North American Journal of Fisheries Management, 32 (5):905–921, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haglund:2017:AVB

[HM17] Justin M. Haglund and Matthew G. Mitro. Age validation of brown trout in Driftless Area streams in Wisconsin using otoliths. North American Journal of Fisheries Management, 37(4):829–835, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hyman:2018:SCS

[HM18] Amanda A. Hyman and Steve L. McMullin. Specialization and characterization of stocked-trout anglers in Virginia, USA. North American Journal of Fisheries Management, 38(6): 1394–1403, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harris:2016:GSS

[HMBT16] Les N. Harris, Jean-Sébastien Moore, Robert Bajno, and Ross F. Tallman. Genetic stock structure of anadromous Arctic char in Canada’s Central Arctic: Potential implications for the management of Canada’s largest Arctic char commercial fishery. North American Journal of Fisheries Management, 36(6):1473–1488, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hastings:2016:CID

[HMCT16] Ryan P. Hastings, Scott J. Meiners, Robert E. Colombo, and Trent E. Thomas. Contrasting impacts of dams on the metacommunity structure of fish and macroinvertebrate assemblages. North American Journal of Fisheries Management, 36(6):1358–1367, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hyman:2016:DAA

[HMD16] Amanda A. Hyman, Steve L. McMullin, and Vic DiCenzo. Dispelling assumptions about stocked-trout fisheries and angler satisfaction. North American Journal of Fisheries Management, 36(6):1395–1404, December 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hense:2010:ECE

[HMP10] Zina Hense, Roy W. Martin, and J. Todd Petty. Electrofishing capture efficiencies for common stream fish species to support watershed-scale studies in the Central Appalachians. North American Journal of Fisheries Management, 30(4):1041–1050, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herbst:2011:LWL

[HMS11] Seth J. Herbst, J. Ellen Marsden, and Stephen J. Smith. Lake whitefish in Lake Champlain after commercial fishery closure and ecosystem changes. North American Journal of Fisheries Management, 31(6):1106–1115, December 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hamilton:2010:CER

[HMW^{+}10] Bryan T. Hamilton, Steve E. Moore, Tod B. Williams, Neal Darby, and Mark R. Vinson. Comparative effects of rotenone and antimycin on macroinvertebrate diversity in two streams in Great Basin National Park, Nevada: Response to comment. North American Journal of Fisheries Management, 30 (5):1129–1131, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hafs:2011:EGL

[HNH11] Andrew W. Hafs, Jonathan M. Niles, and Kyle J. Hartman. Efficiency of gastric lavage on age-0 brook trout and the influence on growth and survival. North American Journal of Fisheries Management, 31(3):530–534, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haxton:2015:AWM

[HNPB15] Tim Haxton, Sarah Nienhuis, Kirby Punt, and Tania Baker. Assessing walleye movement among reaches of a large, fragmented river. North American Journal of Fisheries Management, 35(3):537–550, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanson:2013:ETA

[HO13] Kyle C. Hanson and Kenneth G. Ostrand. Evaluation of transmitter application techniques for use in research of adult eulachon. North American Journal of Fisheries Management, 33(6):1119–1124, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hague:2014:ESR

[HP14] Merran J. Hague and David A. Patterson. Evaluation of statistical river temperature forecast models for fisheries management. North American Journal of Fisheries Management, 34 (1):132–146, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hinck:2011:CPV

[HPA^{+}11] Jo Ellen Hinck, Diana M. Papoulias, Mandy L. Annis, Donald E. Tillitt, Carrie Marr, Nancy D. Denslow, Kevin J. Kroll, and Jason Nachtmann. Characterization of plasma vitellogenin and sex hormone concentrations during the annual reproductive cycle of the endangered razorback sucker. North American Journal of Fisheries Management, 31(5):765–781, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hammen:2019:CBC

[HPD^{+}19] Jeremy J. Hammen, Emily Pherigo, Wyatt Doyle, Jeff Finley, Kevin Drews, and Jason M. Goeckler. A comparison between conventional boat electrofishing and the electrified dozer trawl for capturing silver carp in tributaries of the Missouri River, Missouri. North American Journal of Fisheries Management, 39(3):582–588, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hellmair:2018:PCI

[HPM^{+}18] Michael Hellmair, Matthew Peterson, Brian Mulvey, Kip Young, John Montgomery, and Andrea Fuller. Physical characteristics influencing nearshore habitat use by juvenile Chinook salmon in the Sacramento River, California. North American Journal of Fisheries Management, 38(4):959–970, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hupfeld:2018:QEP

[HPTH18] Ryan N. Hupfeld, Quinton E. Phelps, Sara J. Tripp, and David P. Herzog. Quantitative evaluation of paddlefish sport fisheries in Missouri’s large reservoirs: Implications for the management of trophy sport fisheries. North American Journal of Fisheries Management, 38(2):295–307, April 2018. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hill:2010:CMR

[HQ10] Megan S. Hill and Cory J. Quesada. Calcein mark retention in Chinook salmon and steelhead fry in artificial and natural rearing environments. North American Journal of Fisheries Management, 30(6):1370–1375, December 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hannah:2011:SFM

[HR11a] Robert W. Hannah and Polly S. Rankin. Site fidelity and movement of eight species of Pacific rockfish at a high-relief rocky reef on the Oregon Coast. North American Journal of Fisheries Management, 31(3):483–494, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hochhalter:2011:EDR

[HR11b] Samuel J. Hochhalter and Daniel J. Reed. The effectiveness of deepwater release at improving the survival of discarded yelloweye rockfish. North American Journal of Fisheries Management, 31(5):852–860, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Holliman:2010:EHT

[HRH10] F. M. Holliman, J. B. Reynolds, and J. A. S. Holmes. Effects of healing time and Pulsed-DC waveform on injury detection and incidence in electroshocked juvenile Chinook salmon. North American Journal of Fisheries Management, 30(6):1413–1419, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hargrove:2019:SEF

[HRKB19] John S. Hargrove, Mark W. Rogers, Phillip T. Kacmar, and Pat Black. A statewide evaluation of Florida bass genetic introgression in Tennessee. North American Journal of Fisheries Management, 39(4):637–651, August 2019. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Howell:2012:ERC

[HS12] Philip J. Howell and Paul M. Sankovich. An evaluation of redd counts as a measure of bull trout population size and trend. North American Journal of Fisheries Management, 32 (1):1–13, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hammen:2019:WGC

[HS19] J. J. Hammen and B. L. Sloss. Walleye genetic characterization in the northern ceded territory of Wisconsin: Implications for stocking using conservation strategies. North American Journal of Fisheries Management, 39(4):693–704, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hamel:2016:VAE

[HSC^{+}16] M. J. Hamel, J. J. Spurgeon, C. J. Chizinski, K. D. Steffensen, and M. A. Pegg. Variability in age estimation results in ambiguity and false understanding of population persistence. North American Journal of Fisheries Management, 36(3):514–522, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haukebo:2018:III

[HSF18] Trond Haukebø, Jarle Steinkjer, and Brian Finlayson. Infrared imagery and inert media used in treating upwelling groundwater with rotenone. North American Journal of Fisheries Management, 38(6):1299–1305, December 2018. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heck:2016:IES

[HSG16] Nadine Heck, Richard C. Stedman, and Marc Gaden. Indicators to evaluate the social dimensions of the recreational fishery in the Great Lakes. North American Journal of Fisheries Management, 36(3):477–484, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Herbst:2017:IMD

[HSHH17] Seth J. Herbst, Bryan S. Stevens, Daniel B. Hayes, and Patrick A. Hanchin. Influence of movement dynamics on walleye harvest management in intermixed fisheries in a chain of lakes. North American Journal of Fisheries Management, 37 (3):467–479, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Heim:2018:QUA

[HSM^{+}18] Kurt C. Heim, Michael E. Steeves, Thomas E. McMahon, Brian D. Ertel, and Todd M. Koel. Quantifying uncertainty in aquatic telemetry: Using received signal strength to estimate telemetry error. North American Journal of Fisheries Management, 38(5):979–990, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hutchinson:2011:ADY

[HT11] Charles E. Hutchinson and Todd T. TenBrink. Age determination of the yellow Irish lord: Management implications as a result of new estimates of maximum age. North American Journal of Fisheries Management, 31(6):1116–1122, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Habicht:2013:HSS

[HTF^{+}13] Christopher Habicht, Terri M. Tobias, Gary Fandrei, Nathan Webber, Bert Lewis, and W. Stewart Grant. Homing of sockeye salmon within Hidden Lake, Alaska, can be used to achieve hatchery management goals. North American Journal of Fisheries Management, 33(4):777–782, August 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2019:PTI

[HTH^{+}19] Adam G. Hansen, Jedidiah S. Thompson, Lauren N. Hargis, Daniel Brauch, and Brett M. Johnson. Predatory threat of introduced yellow perch in a salmonid-dominated reservoir food Web. North American Journal of Fisheries Management, 39 (1):172–190, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hanisch:2010:DCD

[HTPS10] Justin R. Hanisch, William M. Tonn, Cynthia A. Paszkowski, and Garry J. Scrimgeour. δ ^{13}C and δ ^{15}N signatures in muscle and fin tissues: Nonlethal sampling methods for stable isotope analysis of salmonids. North American Journal of Fisheries Management, 30(1):1–11, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hessenauer:2018:MRV

[HUB^{+}18] Jan-Michael Hessenauer, Brad E. Utrup, Andrew S. Briggs, Todd C. Wills, and Michael V. Thomas. Mark–recapture validation of pectoral fin ray age estimation for lake sturgeon. North American Journal of Fisheries Management, 38 (6):1251–1257, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hooley-Underwood:2017:STP

[HUST17] Zachary E. Hooley-Underwood, Summer B. Stevens, and Kevin G. Thompson. Short-term passive integrated transponder tag retention in wild populations of bluehead and flannelmouth suckers. North American Journal of Fisheries Management, 37(3):582–586, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hansen:2018:MRB

[HV18] Jon M. Hansen and Robert W. Van Kirk. A mark–recapturebased approach for estimating angler harvest. North American Journal of Fisheries Management, 38(2):400–410, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hessenauer:2017:CLB

[HVW^{+}17] Jan-Michael Hessenauer, Jason Vokoun, Amy Welsh, Justin Davis, Robert Jacobs, and Eileen O’ Donnell. Can largemouth bass transplanted from an unexploited population genetically contribute to an active fishery? A test case for genetic management of exploited fish populations. North American Journal of Fisheries Management, 37(2):271–283, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harvey:2010:CAD

[HW10] Bret C. Harvey and Margaret A. Wilzbach. Carcass addition does not enhance juvenile salmonid biomass, growth, or retention in six Northwestern California streams. North American Journal of Fisheries Management, 30(6):1445–1451, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haak:2012:SRN

[HW12] Amy L. Haak and Jack E. Williams. Spreading the risk: Native trout management in a warmer and less-certain future. North American Journal of Fisheries Management, 32(2):387– 401, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harvey:2016:UCC

[HW16] Bret C. Harvey and Jason L. White. Use of cover for concealment behavior by rainbow trout: Influences of cover structure and area. North American Journal of Fisheries Management, 36(6):1308–1314, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hickerson:2019:EPT

[HW19] Brian T. Hickerson and Annika W. Walters. Evaluation of potential translocation sites for an imperiled cyprinid, the hornyhead chub. North American Journal of Fisheries Management, 39(1):205–218, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hartill:2011:RAM

[HWB11] Bruce W. Hartill, Tim G. Watson, and Richard Bian. Refining and applying a maximum-count aerial-access survey design to estimate the harvest taken from New Zealand’s largest recreational fishery. North American Journal of Fisheries Management, 31(6):1197–1210, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Haley:2012:POS

[HWDA12] Norman V. Haley III, Russell A. Wright, Dennis R. DeVries, and Micheal S. Allen. Privately owned small impoundments in Central Alabama: a survey and evaluation of management techniques for largemouth bass and bluegill. North American Journal of Fisheries Management, 32(6):1180–1190, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hess:2014:MSS

[HWFN14] Jon E. Hess, John M. Whiteaker, Jeffrey K. Fryer, and Shawn R. Narum. Monitoring stock-specific abundance, run timing, and straying of Chinook salmon in the Columbia River using genetic stock identification (GSI). North American Journal of Fisheries Management, 34(1):184–201, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Harvey:2014:ESD

[HWNR14] Bret C. Harvey, Jason L. White, Rodney J. Nakamoto, and Steven F. Railsback. Effects of streamflow diversion on a fish population: Combining empirical data and individual-based models in a site-specific evaluation. North American Journal of Fisheries Management, 34(2):247–257, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Huang:2015:MEC

[HWWT15] Pei Huang, Richard T. Woodward, Michael J. Wilberg, and David Tomberlin. Management evaluation for the Chesapeake Bay blue crab fishery: An integrated bioeconomic approach. North American Journal of Fisheries Management, 35(2):216– 228, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Hessenauer:2019:LTA

[HWZ19] Jan-Michael Hessenauer, Todd C. Wills, and Troy G. Zorn. A long-term analysis of channel morphology and stream substrates before and after sediment trap construction in Michigan trout streams. North American Journal of Fisheries Management, 39(2):379–389, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isaak:2011:EUE

[IH11] Daniel J. Isaak and Dona L. Horan. An evaluation of underwater epoxies to permanently install temperature sensors in mountain streams. North American Journal of Fisheries Management, 31(1):134–137, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Irwin:2012:DEI

[ILBJ12] Brian J. Irwin, Weihai Liu, James R. Bence, and Michael L. Jones. Defining economic injury levels for sea lamprey control in the Great Lakes Basin. North American Journal of Fisheries Management, 32(4):760–771, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2013:TRT

[IMM13] Daniel A. Isermann, Joshua B. Maxwell, and Michael C. McInerny. Temporal and regional trends in black bass release rates in Minnesota. North American Journal of Fisheries Management, 33(2):344–350, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2010:SRB

[ISC10] Daniel A. Isermann, Douglas W. Schultz, and Andrew J. Carlson. Sex ratios of black crappies harvested during spring fisheries on two Minnesota lakes: Are males in the majority? North American Journal of Fisheries Management, 30 (3):812–820, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2010:VNS

[Ise10] Daniel A. Isermann. Validation of nonlethal sex determination for black crappies during spring. North American Journal of Fisheries Management, 30(2):352–353, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2010:CSS

[ITT10] Daniel A. Isermann, Andrew L. Thompson, and Philip J. Talmage. Comparisons of sex-specific growth and weight–length relationships in Minnesota black crappie populations. North American Journal of Fisheries Management, 30(2):354–360, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ingram:2013:PRS

[ITY13] Travis R. Ingram, Josh E. Tannehill, and Shawn P. Young. Post-release survival and behavior of adult shoal bass in the Flint River, Georgia. North American Journal of Fisheries Management, 33(4):717–722, August 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Isermann:2010:EBC

[IWB10] Daniel A. Isermann, Max H. Wolter, and Jason J. Breeggemann. Estimating black crappie age: An assessment of dorsal spines and scales as nonlethal alternatives to otoliths. North American Journal of Fisheries Management, 30(6):1591–1598, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2019:QMV

[JBHM19] Adelaide C. Johnson, J. Ryan Bellmore, Stormy Haught, and Ronald Medel. Quantifying the monetary value of Alaska National Forests to commercial Pacific salmon fisheries. North American Journal of Fisheries Management, 39(6):1119–1131, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson-Bice:2018:RBS

[JBRWH18] Sean M. Johnson-Bice, Kathryn M. Renik, Steve K. Windels, and Andrew W. Hafs. A review of Beaver–salmonid relationships and history of management actions in the Western Great Lakes (USA) region. North American Journal of Fisheries Management, 38(6):1203–1225, December 2018. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2010:EPB

[JBZ^{+}10] Jeffrey R. Johnson, Jason Baumsteiger, Joseph Zydlewski, J. Michael Hudson, and William Ardren. Evidence of panmixia between sympatric life history forms of coastal cutthroat trout in two Lower Columbia River tributaries. North American Journal of Fisheries Management, 30(3):691–701, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:2018:PVI

[JCB^{+}18] Kim K. Jones, Trevan J. Cornwell, Daniel L. Bottom, Staci Stein, and Kara J. Anlauf-Dunn. Population viability improves following termination of Coho salmon hatchery releases. North American Journal of Fisheries Management, 38(1):39–55, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2010:MMR

[JDG10] James E. Johnson, Steven P. DeWitt, and David J. A. Gonder. Mass-marking reveals emerging self regulation of the Chinook salmon population in Lake Huron. North American Journal of Fisheries Management, 30(2):518–529, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jin:2016:APM

[JDH16] Di Jin, Geret DePiper, and Porter Hoagland. Applying portfolio management to implement ecosystem-based fishery management (EBFM). North American Journal of Fisheries Management, 36(3):652–669, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2013:AMF

[JF13] Marc A. Johnson and Thomas A. Friesen. Age at maturity, fork length, and sex ratio of Upper Willamette River hatchery spring Chinook salmon. North American Journal of Fisheries Management, 33(2):318–328, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2014:GDP

[JF14] Marc A. Johnson and Thomas A. Friesen. Genetic diversity and population structure of spring Chinook salmon from the Upper Willamette River, Oregon. North American Journal of Fisheries Management, 34(4):853–862, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2010:DPC

[JHB10] Ronald L. Johnson, Sam D. Henry, and Sam W. Barkley. Distribution and population characteristics of shadow bass in two Arkansas Ozark streams. North American Journal of Fisheries Management, 30(6):1522–1528, December 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jennings:2011:SCS

[JHK11] Martin J. Jennings, Gene R. Hatzenbeler, and Jeffrey M. Kampa. Spring capture site fidelity of adult muskellunge in inland lakes. North American Journal of Fisheries Management, 31(3):461–467, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jaxion-Harm:2015:DAR

[JHS15a] J. Jaxion-Harm and S. T. Szedlmayer. Depth and artificial reef type effects on size and distribution of red snapper in the Northern Gulf of Mexico. North American Journal of Fisheries Management, 35(1):86–96, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2015:ERL

[JHS15b] Melissa J. Johnson, Michael J. Hansen, and Michael J. Seider. Effectiveness of a refuge for lake trout in Western Lake Superior I: Empirical analysis of past performance. North American Journal of Fisheries Management, 35(5):988–1002, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jacobs:2013:BRP

[JHU^{+}13] John M. Jacobs, Reginal M. Harrell, Jim Uphoff, Howard Townsend, and Kyle Hartman. Biological reference points for the nutritional status of Chesapeake Bay striped bass. North American Journal of Fisheries Management, 33(3):468–481, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Janac:2010:MSD

[JJ10] Michal Janáč and Pavel Jurajda. Modified sampling design for age-0 fish electrofishing at beach habitats. North American Journal of Fisheries Management, 30(5):1210–1220, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jepson:2010:PCM

[JKN^{+}10] M. A. Jepson, M. L. Keefer, G. P. Naughton, C. A. Peery, and B. J. Burke. Population composition, migration timing, and harvest of Columbia River Chinook salmon in late summer and fall. North American Journal of Fisheries Management, 30(1):72–88, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2013:SOP

[JKSK13] Gary E. Johnson, Fenton Khan, John R. Skalski, and Bernard A. Klatte. Sluiceway operations to pass juvenile salmonids at The Dalles Dam, Columbia River, USA. North American Journal of Fisheries Management, 33(5):1000–1012, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jackson:2012:LED

[JM12a] Aaron Jackson and Mary Moser. Low-elevation dams are impediments to adult Pacific lamprey spawning migration in the Umatilla River, Oregon. North American Journal of Fisheries Management, 32(3):548–556, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jeffres:2012:WGF

[JM12b] Carson Jeffres and Peter Moyle. When good fish make bad decisions: Coho salmon in an ecological trap. North American Journal of Fisheries Management, 32(1):87–92, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2015:DRP

[JM15] James H. Johnson and James E. McKenna, Jr. Diel resource partitioning among juvenile Atlantic Salmon, Brown trout, and rainbow trout during summer. North American Journal of Fisheries Management, 35(3):586–597, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2016:PCS

[JNC^{+}16] James H. Johnson, Christopher C. Nack, Marc A. Chalupnicki, Ross Abbett, and J. E. McKenna, Jr. Predation on Chinook salmon parr by hatchery salmonids and fallfish in the Salmon River, New York. North American Journal of Fisheries Management, 36(1):74–84, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2010:CSM

[JPWR10] Steven L. Johnson, James Harry Power, Derek R. Wilson, and James Ray. A comparison of the survival and migratory behavior of hatchery-reared and naturally reared steelhead smolts in the Alsea River and Estuary, Oregon, using acoustic telemetry. North American Journal of Fisheries Management, 30(1): 55–71, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2016:ECW

[JSDJ16] Nicholas S. Johnson, William D. Swink, Heather A. Dawson, and Michael L. Jones. Effects of coded-wire-tagging on streamdwelling sea lamprey larvae. North American Journal of Fisheries Management, 36(5):1059–1067, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Johnson:2016:IRH

[JTB16] Jennifer L. Johnson, Jesse T. Trushenski, and James D. Bowker. Induction, recovery, and hematological responses of pallid sturgeon to chemical and electrical sedation. North American Journal of Fisheries Management, 36(3):568–575, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

James:2010:IDC

[JWC10] Daniel A. James, Jerry W. Wilhite, and Steven R. Chipps. Influence of drought conditions on brown trout biomass and size structure in the Black Hills, South Dakota. North American Journal of Fisheries Management, 30(3):791–798, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Jones:2011:DRI

[JY11] Nicholas E. Jones and Geoffrey B. Yunker. Development of a riverine index netting protocol: Comparisons of net orientation, height, panel order, and line diameter. North American Journal of Fisheries Management, 31(1):23–31, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kovach:2018:PRT

[KACS18] Ryan P. Kovach, Robert Al-Chokhachy, and Tracy Stephens. Proactive rainbow trout suppression reduces threat of hybridization in the Upper Snake River Basin. North American Journal of Fisheries Management, 38(4):811–819, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kerns:2015:ERF

[KADH15] Janice A. Kerns, Micheal S. Allen, Jason R. Dotson, and Joseph E. Hightower. Estimating regional fishing mortality for freshwater systems: a Florida largemouth bass example. North American Journal of Fisheries Management, 35(4):681–689, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kaeding:2012:YLT

[Kae12] Lynn R. Kaeding. Are Yellowstone Lake temperatures more suitable to nonnative lake trout than to native cutthroat trout? North American Journal of Fisheries Management, 32(5):848–852, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kaeding:2013:ICA

[Kae13] Lynn R. Kaeding. Informative correlations among metrics of Yellowstone Lake cutthroat trout caught by two quantitative methods across three recent decades. North American Journal of Fisheries Management, 33(5):882–886, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keep:2015:CEC

[KAZ15] Shane G. Keep, M. Brady Allen, and Joseph S. Zendt. Comparison of electronarcosis and carbon dioxide sedation effects on travel time in adult Chinook and Coho salmon. North American Journal of Fisheries Management, 35(5):906–912, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kaintz:2010:DSB

[KB10] Melissa A. Kaintz and Phillip W. Bettoli. Dispersal of smallmouth bass from a simulated tournament weigh-in site. North American Journal of Fisheries Management, 30(4):976–982, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kitterman:2011:SAS

[KB11] Christy L. Kitterman and Phillip W. Bettoli. Survival of angled saugers in the Lower Tennessee River. North American Journal of Fisheries Management, 31(3):567–573, June 2011.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kornis:2019:FAP

[KBH^{+}19] Matthew S. Kornis, Charles R. Bronte, Mark E. Holey, S. Dale Hanson, Theodore J. Treska, Jory L. Jonas, Charles P. Madenjian, Randall M. Claramunt, Steven R. Robillard, Brian Breidert, Kevin C. Donner, Stephen J. Lenart, Archie W. Martell, Jr., Patrick C. McKee, and Erik J. Olsen. Factors affecting postrelease survival of coded-wire-tagged lake trout in Lake Michigan at four historical spawning locations. North American Journal of Fisheries Management, 39(5):868–895, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koenigs:2013:IAE

[KBK13a] Ryan P. Koenigs, Ronald M. Bruch, and Kendall K. Kamke. Impacts of aging error on walleye management in the winnebago system. North American Journal of Fisheries Management, 33(5):900–908, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koenigs:2013:IAT

[KBK13b] Ryan P. Koenigs, Ronald M. Bruch, and Kendall K. Kamke. Impacts of anchor tag loss on walleye management in the winnebago system. North American Journal of Fisheries Management, 33(5):909–916, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kroboth:2019:BCN

[KCCW19] P. T. Kroboth, C. L. Cox, D. C. Chapman, and G. W. Whitledge. Black carp in North America: a description of range, habitats, time of year, and methods of reported captures. North American Journal of Fisheries Management, 39 (5):1046–1055, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kc:2015:DSS

[KCF15] Vaskar Nepal Kc, Robert E. Colombo, and Leslie D. Frankland. Demographics of shovelnose sturgeon in the Lower Wabash River, Illinois. North American Journal of Fisheries Management, 35(4):835–844, August 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krall:2019:LLL

[KCRR19] Michelle Krall, Christopher Clark, Phil Roni, and Kai Ross. Lessons learned from long-term effectiveness monitoring of instream habitat projects. North American Journal of Fisheries Management, 39(6):1395–1411, December 2019. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kelly:2017:DEP

[KCS^{+}17] Brett B. Kelly, Joshua B. Cary, Alisha D. Smith, Kasey C. Pregler, Seoghyun Kim, and Yoichiro Kanno. Detection efficiency of a portable PIT antenna for two small-bodied fishes in a Piedmont stream. North American Journal of Fisheries Management, 37(6):1362–1369, December 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kozfkay:2010:CSM

[KD10] Joseph R. Kozfkay and Jeff C. Dillon. Creel survey methods to assess catch, loss, and capture frequency of white sturgeon in the Snake River, Idaho. North American Journal of Fisheries Management, 30(1):221–229, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keretz:2018:EWT

[KDA^{+}18] Kevin R. Keretz, Colin P. Dinken, Peter J. Allen, Michael E. Colvin, and Harold L. Schramm, Jr. The effect of water temperature, angling time, and dissolved oxygen on the survival of largemouth bass subjected to simulated angling and tournament handling procedures. North American Journal of Fisheries Management, 38(3):606–622, June 2018. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Korman:2010:CES

[KDMH10] Josh Korman, A. Scott Decker, Brent Mossop, and John Hagen. Comparison of electrofishing and snorkeling mark– recapture estimation of detection probability and abundance of juvenile steelhead in a medium-sized river. North American Journal of Fisheries Management, 30(5):1280–1302, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keefer:2010:TAP

[KDP^{+}10] Matthew L. Keefer, William R. Daigle, Christopher A. Peery, Howard T. Pennington, Steven R. Lee, and Mary L. Moser. Testing adult Pacific lamprey performance at structural challenges in fishways. North American Journal of Fisheries Management, 30(2):376–385, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kovach:2011:HBY

[KEC11] Ryan P. Kovach, Lisa A. Eby, and Matthew P. Corsi. Hybridization between Yellowstone cutthroat trout and rainbow trout in the Upper Snake River Basin, Wyoming. North American Journal of Fisheries Management, 31(6):1077–1087, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kapuscinski:2012:FSD

[KFM12] Kevin L. Kapuscinski, John M. Farrell, and Brent A. Murry. Feeding strategies and diets of young-of-the-year muskellunge from two large river ecosystems. North American Journal of Fisheries Management, 32(4):635–647, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kapuscinski:2012:FRA

[KFW12] Kevin L. Kapuscinski, John M. Farrell, and Michael A. Wilkinson. First report of abundant rudd populations in North America. North American Journal of Fisheries Management, 32(1):82–86, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kanigan:2019:GNF

[KHBH19] Adam M. Kanigan, Scott G. Hinch, Arthur L. Bass, and William L. Harrower. Gill-net fishing effort predicts physical injuries on sockeye salmon captured near spawning grounds. North American Journal of Fisheries Management, 39(3):441– 451, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kneebone:2014:MSB

[KHDA14] Jeff Kneebone, William S. Hoffman, Micah J. Dean, and Michael P. Armstrong. Movements of striped bass between the exclusive economic zone and Massachusetts State waters.

North American Journal of Fisheries Management, 34(3):524– 534, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kazyak:2013:RVA

[KHH13] David C. Kazyak, Robert H. Hilderbrand, and Amanda E. Holloway. Rapid visual assessment to determine sex in brook trout. North American Journal of Fisheries Management, 33 (3):665–668, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kelling:2016:DOP

[KIST16] Craig J. Kelling, Daniel A. Isermann, Brian L. Sloss, and Keith N. Turnquist. Diet overlap and predation between largemouth bass and walleye in Wisconsin lakes using DNA barcoding to improve taxonomic resolution. North American Journal of Fisheries Management, 36(3):621–629, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Keefer:2017:CDR

[KJN^{+}17] Matthew L. Keefer, Michael A. Jepson, George P. Naughton, Timothy J. Blubaugh, Tami S. Clabough, and Christopher C. Caudill. Condition-dependent en route migration mortality of adult Chinook salmon in the Willamette River main stem. North American Journal of Fisheries Management, 37(2):370– 379, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kanno:2016:RNB

[KKM16] Yoichiro Kanno, Matt A. Kulp, and Stephen E. Moore. Recovery of native brook trout populations following the eradication of nonnative rainbow trout in Southern Appalachian mountains streams. North American Journal of Fisheries Management, 36(6):1325–1335, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koenig:2011:PDT

[KKMS11] Martin K. Koenig, Joseph R. Kozfkay, Kevin A. Meyer, and Daniel J. Schill. Performance of diploid and triploid rainbow trout stocked in Idaho alpine lakes. North American Journal of Fisheries Management, 31(1):124–133, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Katt:2012:ANG

[KKSH12] Jordan D. Katt, Keith D. Koupal, Casey W. Schoenebeck, and W. Wyatt Hoback. Assessment of a New Gear to sample walleye eggs. North American Journal of Fisheries Management, 32(1):44–48, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kock:2012:ESI

[KLR^{+}12] Tobias J. Kock, Theresa L. Liedtke, Dennis W. Rondorf, John D. Serl, Mike Kohn, and Karin A. Bumbaco. Elevated streamflows increase dam passage by juvenile Coho salmon during winter: Implications of climate change in the Pacific Northwest. North American Journal of Fisheries Management, 32(6):1070–1079, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koenig:2011:RPD

[KM11] Martin K. Koenig and Kevin A. Meyer. Relative performance of diploid and triploid catchable rainbow trout stocked in Idaho lakes and reservoirs. North American Journal of Fisheries Management, 31(4):605–613, August 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Knapp:2012:FCR

[KMB^{+}12] Michael L. Knapp, Steven W. Mero, David J. Bohlander, David F. Staples, and Jerry A. Younk. Fish community responses to the introduction of muskellunge into Minnesota lakes. North American Journal of Fisheries Management, 32 (2):191–201, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koenig:2015:EAT

[KMK^{+}15] Martin K. Koenig, Kevin A. Meyer, Joseph R. Kozfkay, Joseph M. DuPont, and Edward B. Schriever. Evaluating the ability of tiger muskellunge to eradicate brook trout in Idaho alpine lakes. North American Journal of Fisheries Management, 35(4):659–670, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kirk:2014:LTP

[KML14] James P. Kirk, Kenneth L. Manuel, and Scott D. Lamprecht. Long-term population response of triploid grass carp stocked in Piedmont and coastal plain reservoirs to control hydrilla. North American Journal of Fisheries Management, 34(4): 795–801, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kowalewski:2012:LIA

[KMPP12] Lucas K. Kowalewski, Alexis P. Maple, Mark A. Pegg, and Kevin L. Pope. Latitudinal influence on age estimates derived from scales and otoliths for bluegills. North American Journal of Fisheries Management, 32(6):1175–1179, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kumar:2013:TDD

[KMPV13] Rajeev Kumar, Steven J. Martell, Tony J. Pitcher, and Divya A. Varkey. Temperature-driven decline of a cisco population in Mille Lacs Lake, Minnesota. North American Journal of Fisheries Management, 33(4):669–681, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koch:2014:EPS

[KNC14] Jeff D. Koch, Ben C. Neely, and Michael E. Colvin. Evaluation of precision and sample sizes using standardized sampling in Kansas reservoirs. North American Journal of Fisheries Management, 34(6):1211–1220, December 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koch:2018:PTS

[KNS18] Jeff Koch, Ben Neely, and Bryan Sowards. Precision of three structures for saugeye age estimation. North American Journal of Fisheries Management, 38(1):31–38, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Karam:2012:ECB

[KPL12] Abraham P. Karam, Michael S. Parker, and Lindsey T. Lyons. Ecological comparison between three artificial refuges and the natural habitat for devils hole pupfish. North American Journal of Fisheries Management, 32(2):224–238, April 2012. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kornis:2016:FAP

[KPL^{+}16] Matthew S. Kornis, Kevin W. Pankow, Allen A. Lane, James L. Webster, and Charles R. Bronte. Factors affecting prestocking coded wire tag loss in lake trout tagged by an automated system. North American Journal of Fisheries Management, 36(3):670–680, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kock:2018:RHN

[KPP^{+}18] Tobias J. Kock, Russell W. Perry, Adam C. Pope, John D. Serl, Mike Kohn, and Theresa L. Liedtke. Responses of hatchery- and natural-origin adult spring Chinook salmon to a trap-and-haul reintroduction program. North American Journal of Fisheries Management, 38(5):1004–1016, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klein:2019:SSS

[KQDC19] Zachary B. Klein, Michael C. Quist, Andrew M. Dux, and Matthew P. Corsi. Size selectivity of sampling gears used to sample Kokanee. North American Journal of Fisheries Management, 39(2):343–352, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klein:2016:PCS

[KQRS16] Zachary B. Klein, Michael C. Quist, Darren T. Rhea, and Anna C. Senecal. Population characteristics and the suppression of nonnative burbot. North American Journal of Fisheries Management, 36(5):1006–1017, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kratzer:2018:RBT

[Kra18] Jud F. Kratzer. Response of brook trout biomass to strategic wood additions in the East Branch Nulhegan River Watershed, Vermont. North American Journal of Fisheries Management, 38(6):1415–1422, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kerns:2016:CLH

[KRB^{+}16] Janice A. Kerns, Mark W. Rogers, David B. Bunnell, Randall M. Claramunt, and Paris D. Collingsworth. Comparing life history characteristics of Lake Michigan’s naturalized and stocked Chinook salmon. North American Journal of Fisheries Management, 36(5):1106–1118, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Khan:2013:SOA

[KRJT13] Fenton Khan, Ida M. Royer, Gary E. Johnson, and Sean C. Tackley. Sluiceway operations for adult steelhead downstream passage at The Dalles Dam, Columbia River, USA. North American Journal of Fisheries Management, 33(5):1013–1023, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krogman:2019:CPC

[Kro19] Rebecca M. Krogman. Comparison of panfish catch and retention by four modified fyke-net designs. North American Journal of Fisheries Management, 39(6):1277–1287, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Koch:2018:CDT

[KSG^{+}18] Jeff Koch, Chris Steffen, Jason Goeckler, Ron Marteney, Josh Jagels, and Bonnie Brown. Comparison of diploid and triploid saugeye recruitment, growth, and condition in Kansas impoundments. North American Journal of Fisheries Management, 38(2):446–453, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kazyak:2016:CCM

[KSH^{+}16] David C. Kazyak, Matthew T. Sell, Robert H. Hilderbrand, Alan A. Heft, and Ryan M. Cooper. A comparison of catchability and mortality with circle and J hooks for streamdwelling brook trout. North American Journal of Fisheries Management, 36(2):259–266, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kornis:2019:GHR

[KSL^{+}19] Matthew S. Kornis, Darin G. Simpkins, Allen A. Lane, David M. Warner, and Charles R. Bronte. Growth of hatcheryreared Chinook salmon in Lakes Michigan and Huron exhibits limited spatial variation but is temporally linked to alewife abundance. North American Journal of Fisheries Management, 39(6):1155–1174, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Korman:2016:ERA

[KSM16] Josh Korman, Jody Schick, and Brent Mossop. Estimating riverwide abundance of juvenile fish populations: How much sampling is enough? North American Journal of Fisheries Management, 36(2):213–229, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kehler:2018:AGJ

[KSM^{+}18] Thomas Kehler, John A. Sweka, Jerre Mohler, Amanda Higgs, and Gregg Kenney. Age and growth of juvenile Atlantic sturgeon in the Lower Hudson River. North American Journal of Fisheries Management, 38(1):84–95, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kasugai:2018:RCS

[KSMN18] Kiyoshi Kasugai, Hiroyuki Sakamoto, Yasuyuki Miyakoshi, and Mitsuhiro Nagata. Returns of chum salmon released from net-pens and rivers in Eastern Hokkaido, Japan. North American Journal of Fisheries Management, 38(1):24–30, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

King:2010:FII

[KSV^{+}10] Ronald G. King, Greg Seegert, Joe Vondruska, Elgin S. Perry, and Douglas A. Dixon. Factors influencing impingement at 15 Ohio River power plants. North American Journal of Fisheries Management, 30(5):1149–1175, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Klett:2013:SCC

[KTHM13] Katherine J. C. Klett, Christian E. Torgersen, Julie A. Henning, and Christopher J. Murray. Spatial consistency of Chinook salmon redd distribution within and among years in the Cowlitz River, Washington. North American Journal of Fisheries Management, 33(3):508–518, June 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kimmel:2014:RBI

[KTN14] David G. Kimmel, Mitchell Tarnowski, and Roger I. E. Newell. The relationship between interannual climate variability and juvenile Eastern oyster abundance at a regional scale in Chesapeake Bay. North American Journal of Fisheries Management, 34(1):1–15, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kratzer:2013:FLB

[KW13] Jud F. Kratzer and Dana R. Warren. Factors limiting brook trout biomass in Northeastern Vermont streams. North American Journal of Fisheries Management, 33(1):130–139, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kaemingk:2011:EPI

[KWMB11] Mark A. Kaemingk, Michael J. Weber, Paul R. McKenna, and Michael L. Brown. Effect of passive integrated transponder tag implantation site on tag retention, growth, and survival of two sizes of juvenile bluegills and yellow perch. North American Journal of Fisheries Management, 31(4):726–732, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Krieger:2018:ECH

[KYD18] Joseph R. Krieger, Ryan T. Young, and James S. Diana. Evaluation and comparison of a habitat suitability model for postdrift larval lake sturgeon in the St. Clair and Detroit Rivers. North American Journal of Fisheries Management, 38(5):1091–1104, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Kazyak:2012:HDP

[KZ12] David Kazyak and Joseph Zydlewski. High-density polyethylene pipe: a new material for pass-by passive integrated transponder antennas. North American Journal of Fisheries Management, 32(1):49–52, February 2012. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leblanc:2018:EGD

[LAA^{+}18] Nathalie M. Leblanc, Samuel N. Andrews, Trevor S. Avery, Gregory N. Puncher, Benjamin I. Gahagan, Andrew R. Whiteley, R. Allen Curry, and Scott A. Pavey. Evidence of a genetically distinct population of striped bass within the Saint John River, New Brunswick, Canada. North American Journal of Fisheries Management, 38(6):1339–1349, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Logsdon:2016:CPS

[LAM16] Dale E. Logsdon, Charles S. Anderson, and Loren M. Miller. Contribution and performance of stocked walleyes in the recovery of the Red Lakes, Minnesota, fishery. North American Journal of Fisheries Management, 36(4):828–843, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Li:2016:ISA

[LBB16] Yang Li, James R. Bence, and Travis O. Brenden. The influence of stock assessment frequency on the achievement of fishery management objectives. North American Journal of Fisheries Management, 36(4):793–812, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leblanc:2017:IWE

[LBF17] John Paul Leblanc, Brandeis L. Brown, and John M. Farrell. Increased walleye egg-to-larvae survival following spawning habitat enhancement in a tributary of Eastern Lake Ontario. North American Journal of Fisheries Management, 37 (5):999–1009, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lynch:2011:SST

[LBL11] Patrick D. Lynch, Mark J. Brush, and Robert J. Latour. Simulated short-term impacts of the Atlantic menhaden reduction fishery on Chesapeake Bay water quality. North American Journal of Fisheries Management, 31(1):70–78, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lawson:2014:MAS

[LC14] Zach J. Lawson and Stephen R. Carpenter. A morphometric approach for stocking walleye fingerlings in lakes invaded by rainbow smelt. North American Journal of Fisheries Management, 34(5):998–1002, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Li:2015:EEF

[LCC^{+}15] Bai Li, Jie Cao, Jui-Han Chang, Carl Wilson, and Yong Chen. Evaluation of effectiveness of fixed-station sampling for monitoring American lobster settlement. North American Journal of Fisheries Management, 35(5):942–957, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lutnesky:2019:ITF

[LCR19] Marvin M. F. Lutnesky, Kenwyn R. Cradock, and James B. Reynolds. Immobilization threshold and fish conductivity of two small fishes. North American Journal of Fisheries Management, 39(4):788–792, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lamer:2010:IHB

[LDP^{+}10] James Thomas Lamer, Chad Ryan Dolan, Jessica Lynn Petersen, John Howard Chick, and John Michael Epifanio. Introgressive hybridization between bighead carp and silver carp in the Mississippi and Illinois Rivers. North American Journal of Fisheries Management, 30(6):1452–1461, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leonard:2010:IIB

[LDW10] D. Michael Leonard, Dennis R. DeVries, and Russell A. Wright. Investigating interactions between channel catfish and other sport fishes in small impoundments. North American Journal of Fisheries Management, 30(3):732–741, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lee:2012:EBP

[Lee12] Min-Yang Lee. Examining bargaining power in the northeast multispecies days-at-sea market. North American Journal of Fisheries Management, 32(5):1017–1031, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lennox:2017:FIP

[LFD^{+}17] Robert J. Lennox, Alexander Filous, Sascha Clark Danylchuk, Steven J. Cooke, Jacob W. Brownscombe, Alan M. Friedlander, and Andy J. Danylchuk. Factors influencing postrelease predation for a catch-and-release tropical flats fishery with a high predator burden. North American Journal of Fisheries Management, 37(5):1045–1053, October 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lawson:2011:CWH

[LGC11] Zach J. Lawson, Jereme W. Gaeta, and Stephen R. Carpenter. Coarse woody habitat, lakeshore residential development, and largemouth bass nesting behavior. North American Journal of Fisheries Management, 31(4):666–670, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lothrop:2014:EIR

[LHS^{+}14] Ryan L. Lothrop, Terry R. Hanson, Steven M. Sammons, Diane Hite, and Michael J. Maceina. Economic impact of a recreational striped bass fishery. North American Journal of Fisheries Management, 34(2):301–310, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Li:2011:GNS

[LJR11] Yan Li, Yan Jiao, and Kevin Reid. Gill-net saturation in Lake Erie: Effects of soak time and fish accumulation on catch per unit effort of walleye and yellow perch. North American Journal of Fisheries Management, 31(2):280–290, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lane:2019:ECW

[LKB19] Allen A. Lane, Matthew S. Kornis, and Charles R. Bronte. Evaluation of coded wire tag loss in brook trout tagged using automated methods. North American Journal of Fisheries Management, 39(1):29–35, February 2019. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Larson:2017:EAC

[LKM^{+}17] James H. Larson, Brent C. Knights, S. Grace McCalla, Emy Monroe, Maren Tuttle-Lau, Duane C. Chapman, Amy E. George, Jon M. Vallazza, and Jon Amberg. Evidence of Asian carp spawning upstream of a key choke point in the Mississippi River. North American Journal of Fisheries Management, 37(4):903–919, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lampo:2017:UPT

[LKV^{+}17] Eli G. Lampo, Brent C. Knights, Jon M. Vallazza, Cory A. Anderson, Will T. Rechkemmer, Levi E. Solomon, Andrew F. Casper, Richard M. Pendleton, and James T. Lamer. Using pharyngeal teeth and chewing pads to estimate juvenile silver carp total length in the la grange reach, Illinois River. North American Journal of Fisheries Management, 37(5):1145–1150, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Landon:2018:EPD

[LKvR^{+}18] Adam C. Landon, Gerard T. Kyle, Carena J. van Riper, Michael A. Schuett, and Jihee Park. Exploring the psychological dimensions of stewardship in recreational fisheries. North American Journal of Fisheries Management, 38(3):579–591, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lew:2012:EVS

[LL12] Daniel K. Lew and Douglas M. Larson. Economic values for saltwater sport fishing in Alaska: a stated preference analysis. North American Journal of Fisheries Management, 32(4): 745–759, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lavigne:2010:RBM

[LLP10] Mélyssa Lavigne, Marc Lucotte, and Serge Paquet. Relationship between mercury concentration and growth rates for walleyes, northern pike, and lake trout from Quebec lakes. North American Journal of Fisheries Management, 30(5): 1221–1237, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lizee:2018:ILN

[LLW^{+}18] Teah W. Lizée, Robert J. Lennox, Taylor D. Ward, Jacob W. Brownscombe, Jacqueline M. Chapman, Andy J. Danylchuk, Liane B. Nowell, and Steven J. Cooke. Influence of landing net mesh type on handling time and tissue damage of angled brook trout. North American Journal of Fisheries Management, 38 (1):76–83, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lamansky:2016:AET

[LM16a] James A. Lamansky, Jr. and Kevin A. Meyer. Air exposure time of trout released by anglers during catch and release. North American Journal of Fisheries Management, 36 (5):1018–1023, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Long:2016:MRB

[LM16b] James M. Long and Richard T. Melstrom. Measuring the relationship between sportfishing trip expenditures and anglers’ species preferences. North American Journal of Fisheries Management, 36(4):731–737, August 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lamansky:2018:ITS

[LMB^{+}18] James A. Lamansky, Jr., Kevin A. Meyer, Brett J. Bowersox, Joe M. DuPont, Brandon Bentz, and Ken B. Lepla. Incidence, types, and shedding and ingestion times of metallic fishing tackle in the digestive systems of white sturgeon. North American Journal of Fisheries Management, 38(5):1152–1159, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Larson:2016:GAS

[LMC^{+}16] James H. Larson, S. Grace McCalla, Duane C. Chapman, Christopher Rees, Brent C. Knights, Jon M. Vallazza, Amy E. George, William B. Richardson, and Jon Amberg. Genetic analysis shows that morphology alone cannot distinguish Asian carp eggs from those of other cyprinid species. North American Journal of Fisheries Management, 36(5):1053–1058, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Luo:2019:SDL

[LMD^{+}19] Miles K. Luo, Charles P. Madenjian, James S. Diana, Matthew S. Kornis, and Charles R. Bronte. Shifting diets of lake trout in Northeastern Lake Michigan. North American Journal of Fisheries Management, 39(4):793–806, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lamansky:2018:CRC

[LMS^{+}18] James A. Lamansky, Jr., Kevin A. Meyer, Brett Spaulding, Brian J. Jaques, and Darryl P. Butt. Corrosion rates and compression strength of white sturgeon-sized fishing hooks exposed to simulated stomach conditions. North American Journal of Fisheries Management, 38(4):896–902, August 2018.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leblanc:2012:VIE

[LN12a] C. A. Leblanc and D. L. Noakes. Visible implant elastomer (VIE) tags for marking small rainbow trout. North American Journal of Fisheries Management, 32(4):716–719, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Loomis:2012:CEV

[LN12b] John Loomis and Kawa Ng. Comparing economic values of trout anglers and nontrout anglers in Colorado’s stocked public reservoirs. North American Journal of Fisheries Management, 32(2):202–210, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Love:2012:WEN

[LN12c] Joseph W. Love and Joshua J. Newhard. Will the expansion of northern snakehead negatively affect the fishery for largemouth bass in the Potomac River (Chesapeake Bay)? North American Journal of Fisheries Management, 32(5):859–868, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lyons:2018:SWY

[LO18] John Lyons and Daniel L. Oele. Sauger and walleye year-class formation in the Lower Wisconsin River, 1987–2016. North American Journal of Fisheries Management, 38(4):842–855, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lantry:2011:EOS

[LOS^{+}11] B. F. Lantry, R. O’Gorman, T. G. Strang, J. R. Lantry, M. J. Connerton, and T. Schaner. Evaluation of offshore stocking of lake trout in Lake Ontario. North American Journal of Fisheries Management, 31(4):671–682, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lepak:2017:AYC

[LOV17] T. A. Lepak, D. H. Ogle, and M. R. Vinson. Age, year-class strength variability, and partial age validation of kiyis from Lake Superior. North American Journal of Fisheries Management, 37(5):1151–1160, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Logsdon:2012:EOI

[LP12] Dale E. Logsdon and Bruce J. Pittman. Evaluation of osmotic induction of calcein treatments for marking juvenile walleyes. North American Journal of Fisheries Management, 32(4):796–805, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Losee:2017:BBO

[LP17] James P. Losee and Larry Phillips. Bigger is better: Optimizing trout stocking in Western Washington lakes. North American Journal of Fisheries Management, 37(3):489–496, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Losee:2016:STR

[LPY16] James P. Losee, Larry Phillips, and William C. Young. Spawn timing and redd morphology of anadromous coastal cutthroat trout Oncorhynchus clarkii clarkii in a tributary of South Puget Sound, Washington. North American Journal of Fisheries Management, 36(2):375–384, April 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lemay:2012:NLR

[LR12] Matthew A. Lemay and Michael A. Russello. Neutral loci reveal population structure by geography, not ecotype, in Kootenay Lake Kokanee. North American Journal of Fisheries Management, 32(2):282–291, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lyons:2017:FWF

[LRHR17] John Lyons, Andrew L. Rypel, Jonathan F. Hansen, and David C. Rowe. Fillet weight and fillet yield: New metrics for the management of panfish and other consumption-oriented recreational fisheries. North American Journal of Fisheries Management, 37(3):550–557, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lew:2010:EIS

[LS10] Daniel K. Lew and Chang K. Seung. The economic impact of saltwater sportfishing harvest restrictions in Alaska: An empirical analysis of nonresident anglers. North American Journal of Fisheries Management, 30(2):538–551, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Loher:2011:UVU

[LS11] Timothy Loher and Sarah M. Stephens. Use of veterinary ultrasound to identify sex and assess female maturity of Pacific halibut in nonspawning condition. North American Journal of Fisheries Management, 31(6):1034–1042, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lee:2013:CHS

[LS13] Pey-Yi Lee and Jian-Ping Suen. Comparing habitat suitability indices (HSIs) based on abundance and occurrence data. North American Journal of Fisheries Management, 33(1):89– 96, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leonard:2017:IEI

[LS17] Jerry Leonard and Erin Steiner. Initial economic impacts of the U.S. Pacific Coast groundfish fishery individual fishing quota program. North American Journal of Fisheries Management, 37(4):862–881, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lauber:2016:USA

[LSC^{+}16] T. Bruce Lauber, Richard C. Stedman, Nancy A. Connelly, Lars G. Rudstam, Richard C. Ready, Gregory L. Poe, David B. Bunnell, Tomas O. Höök, Marten A. Koops, Stuart A. Ludsin, and Edward S. Rutherford. Using scenarios to assess possible future impacts of invasive species in the Laurentian Great Lakes. North American Journal of Fisheries Management, 36(6):1292–1307, December 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lundgren:2014:QEF

[LSK^{+}14] Seth A. Lundgren, Casey W. Schoenebeck, Keith D. Koupal, Jared A. Lorensen, and Caleb G. Huber. Quantification and evaluation of factors influencing largemouth bass predation of stocked advanced fingerling yellow perch. North American Journal of Fisheries Management, 34(3):595–601, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leonard:2013:BCC

[LSL^{+}13] Jill B. K. Leonard, Wendylee Stott, Delora M. Loope, Paul C. Kusnierz, and Ashwin Sreenivasan. Biological consequences of the coaster brook trout restoration stocking program in Lake Superior tributaries within pictured rocks national lakeshore. North American Journal of Fisheries Management, 33(2):359– 372, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lowry:2015:ECS

[LSLC15] Dayv Lowry, Kurt Stick, Adam Lindquist, and Yuk W. Cheng. Evaluation of creel survey methods to estimate recreational harvest of surf smelt in Puget Sound, Washington. North American Journal of Fisheries Management, 35(3):403–417, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Larson:2016:DWL

[LSM^{+}16] James H. Larson, David F. Staples, Ryan P. Maki, Jon M. Vallazza, Brent C. Knights, and Kevin E. Peterson. Do water level fluctuations influence production of walleye and yellow perch young-of-the-year in large northern lakes? North American Journal of Fisheries Management, 36(6):1425–1436, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leidig:2015:GPS

[LSMD15] Jacqueline M. Leidig, Virginia R. Shervette, Christopher J. McDonough, and Tanya L. Darden. Genetic population structure of black drum in U.S. waters. North American Journal of Fisheries Management, 35(3):464–477, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leone:2012:DPP

[LSQ12] Frank J. Leone, Joseph N. Stoeckel, and Jeffrey W. Quinn. Differences in paddlefish populations among impoundments of the Arkansas River, Arkansas. North American Journal of Fisheries Management, 32(4):731–744, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Leonard:2011:IFW

[LTG^{+}11] Nancy J. Leonard, William W. Taylor, Chris I. Goddard, Kenneth A. Frank, Ann E. Krause, and Michael G. Schechter. Information flow within the social network structure of a joint strategic plan for management of Great Lakes fisheries. North American Journal of Fisheries Management, 31(4):629–655, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lenker:2016:DRH

[LWJS16] Melissa A. Lenker, Brian C. Weidel, Olaf P. Jensen, and Christopher T. Solomon. Developing recreational harvest regulations for an unexploited lake trout population. North American Journal of Fisheries Management, 36(2):385–397, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Lin:2012:SDE

[LXM^{+}12] Mingli Lin, Yuguo Xia, Brian R. Murphy, Zhongjie Li, Jiashou Liu, Tanglin Zhang, and Shaowen Ye. Size-dependent effects of coded wire tags on mortality and tag retention in redtail culter Culter mongolicus. North American Journal of Fisheries Management, 32(5):968–973, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Muhlfeld:2015:SCI

[MAG^{+}15] Clint C. Muhlfeld, Shannon E. Albeke, Stephanie L. Gunckel, Benjamin J. Writer, Bradley B. Shepard, and Bruce E. May. Status and conservation of interior redband trout in the Western United States. North American Journal of Fisheries Management, 35(1):31–53, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McDonald:2013:SDA

[MAH^{+}13] D. L. McDonald, J. D. Anderson, C. Hurley, B. W. Bumguardner, and C. R. Robertson. Sexual dimorphism in alligator gar. North American Journal of Fisheries Management, 33(4):811–816, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marcot:2012:EPA

[MAM^{+}12] Bruce G. Marcot, Chris S. Allen, Steve Morey, Dan Shively, and Rollie White. An expert panel approach to assessing potential effects of bull trout reintroduction on federally listed salmonids in the Clackamas River, Oregon. North American Journal of Fisheries Management, 32(3):450–465, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McDougall:2014:DPL

[MAP14] C. A. McDougall, W. G. Anderson, and S. J. Peake. Downstream passage of lake sturgeon through a hydroelectric generating station: Route determination, survival, and fine-scale movements. North American Journal of Fisheries Management, 34(3):546–558, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Middaugh:2016:RES

[MASN16] Christopher R. Middaugh, Ted Alfermann, P. Andrew Strickland, and Phong Nguyen. A regional evaluation of Suwannee bass and largemouth bass exploitation in North Florida rivers. North American Journal of Fisheries Management, 36 (4):958–963, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matlock:2010:DTC

[Mat10] Gary C. Matlock. Did the Texas closure increase brown shrimp catches off Texas? North American Journal of Fisheries Management, 30(3):799–804, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meerbeek:2012:RFE

[MB12] Jonathan R. Meerbeek and Phillip W. Bettoli. Relying on fin erosion to identify hatchery-reared brown trout in a Tennessee River. North American Journal of Fisheries Management, 32 (5):922–928, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mollenhauer:2017:MMM

[MB17] Robert Mollenhauer and Shannon K. Brewer. Multinomial N-mixture models improve the applicability of electrofishing for developing population estimates of stream-dwelling smallmouth bass. North American Journal of Fisheries Management, 37(1):211–224, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mahasuweerachai:2010:EDU

[MBBS10] Phumsith Mahasuweerachai, Tracy A. Boyer, Dane M. Balsman, and Daniel E. Shoup. Estimating demand for urban fisheries management: an illustration of conjoint analysis as a tool for fisheries managers. North American Journal of Fisheries Management, 30(5):1339–1351, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maselko:2013:GFS

[MBM13] Jacek Maselko, Gretchen Bishop, and Peter Murphy. Ghost fishing in the Southeast Alaska commercial Dungeness crab fishery. North American Journal of Fisheries Management, 33(2):422–431, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morson:2012:SLS

[MBPG12] Jason M. Morson, Eleanor A. Bochenek, Eric N. Powell, and Jennifer E. Gius. Sex at length of summer flounder landed in the New Jersey recreational party boat fishery. North American Journal of Fisheries Management, 32(6):1201–1210, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martens:2010:ERW

[MC10] Kyle D. Martens and Patrick J. Connolly. Effectiveness of a redesigned water diversion using rock vortex weirs to enhance longitudinal connectivity for small salmonids. North American Journal of Fisheries Management, 30(6):1544–1552, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McBride:2014:MMS

[McB14] Richard S. McBride. Managing a marine stock portfolio: Stock identification, structure, and management of 25 fishery species along the Atlantic Coast of the United States. North American Journal of Fisheries Management, 34(4):710–734, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCabe:2019:AAT

[MCB^{+}19] M. M. McCabe, J. A. Chiotti, J. C. Boase, A. T. Fisk, and T. E. Pitcher. Assessing acoustic tagging effects on survival, growth, and swimming ability of juvenile lake sturgeon. North American Journal of Fisheries Management, 39(3):574–581, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2015:FEP

[McC15] Joshua L. McCormick. Feasibility and evaluation of probabilistic on-site surveys for estimating angler use and catch in high mountain lakes. North American Journal of Fisheries Management, 35(6):1152–1163, December 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2016:UTR

[McC16] Joshua L. McCormick. Using tag-return models to estimate the number of times fish are captured in fisheries with high catch-and-release rates. North American Journal of Fisheries Management, 36(3):584–589, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2017:EME

[McC17] Joshua L. McCormick. Evaluation of methods to estimate salmon harvest using angler harvest permits available on a smartphone application. North American Journal of Fisheries Management, 37(1):1–8, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McGregor:2014:UPE

[McG14] Andrea M. McGregor. Using Ecopath modeling to describe historical conditions for a large, boreal lake ecosystem prior to European settlement. North American Journal of Fisheries Management, 34(1):16–29, February 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moffett:2012:PST

[MCH12] Cinamon Moffett, Yong Chen, and Margaret Hunter. Preliminary study of trap bycatch in the Gulf of Maine’s northern shrimp fishery. North American Journal of Fisheries Management, 32(4):704–715, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McIlraith:2015:SMB

[MCK^{+}15] Brian J. McIlraith, Christopher C. Caudill, Brian P. Kennedy, Christopher A. Peery, and Matthew L. Keefer. Seasonal migration behaviors and distribution of adult Pacific lampreys in unimpounded reaches of the Snake River Basin. North American Journal of Fisheries Management, 35(1):123–134, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martin:2015:RAI

[MCP15] Dustin R. Martin, Christopher J. Chizinski, and Kevin L. Pope. Reservoir area of influence and implications for fisheries management. North American Journal of Fisheries Management, 35(2):185–190, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miranda:2017:LLF

[MCS^{+}17] L. E. Miranda, M. E. Colvin, A. C. Shamaskin, L. A. Bull, T. Holman, and R. Jones. Length limits fail to restructure a largemouth bass population: a 28-year case history. North American Journal of Fisheries Management, 37(3):624–632, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moody-Carpenter:2017:DIM

[MCSF^{+}17] Cassi J. Moody-Carpenter, Greg G. Sass, Leslie D. Frankland, Eric K. Bollinger, and Robert E. Colombo. Demographics and interstate management implications for flathead catfish in the Lower Wabash River, Illinois and Indiana. North American Journal of Fisheries Management, 37(6):1265–1273, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Munsch:2014:ESP

[MCTM14] Stuart H. Munsch, Jeffery R. Cordell, Jason D. Toft, and Erin E. Morgan. Effects of seawalls and piers on fish assemblages and juvenile salmon feeding behavior. North American Journal of Fisheries Management, 34(4):814–827, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madenjian:2010:LTP

[MD10] Charles P. Madenjian and Timothy J. Desorcie. Lake trout population dynamics in the northern refuge of Lake Michigan: Implications for future rehabilitation. North American Journal of Fisheries Management, 30(3):629–641, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meals:2012:TMI

[MDM12] K. O. Meals, A. W. Dunn, and L. E. Miranda. Trolling may intensify exploitation in crappie fisheries. North American Journal of Fisheries Management, 32(2):325–332, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meador:2012:ESA

[Mea12] Michael R. Meador. Effectiveness of seining after electrofishing to characterize stream fish communities. North American Journal of Fisheries Management, 32(1):177–185, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Munson:2010:MBK

[MEJ10] A. Douglas Munson, Diane G. Elliott, and Keith Johnson. Management of bacterial kidney disease in Chinook salmon hatcheries based on broodstock testing by enzyme-linked immunosorbent assay: a multiyear study. North American Journal of Fisheries Management, 30(4):940–955, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2012:RRS

[MEL^{+}12] Kevin A. Meyer, F. Steven Elle, James A. Lamansky, Jr., Elizabeth R. J. M. Mamer, and Arthur E. Butts. A rewardrecovery study to estimate tagged-fish reporting rates by Idaho anglers. North American Journal of Fisheries Management, 32(4):696–703, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matter:2018:RAM

[MFLS18] Allison N. Matter, Jeffrey A. Falke, J. Andrés López, and James W. Savereide. A rapid-assessment method to estimate the distribution of juvenile Chinook salmon in tributary habitats using eDNA and occupancy estimation. North American Journal of Fisheries Management, 38(1):223–236, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Milliken:2011:ESD

[MFRS11] Henry O. Milliken, Marianne Farrington, Tom Rudolph, and Melissa Sanderson. Erratum: Survival of discarded sublegal Atlantic cod in the Northwest Atlantic demersal longline fishery (volume 29:985–995). North American Journal of Fisheries Management, 31(2):416, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Munyandorero:2010:SDP

[MG10] Joseph Munyandorero and Cameron B. Guenther. A sixdecade portrait of Florida marine fisheries via landings-based trophodynamic indicators. North American Journal of Fisheries Management, 30(1):259–280, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marcek:2014:EPM

[MG14] Benjamin J. Marcek and John E. Graves. An estimate of postrelease mortality of school-size bluefin tuna in the U.S. recreational troll fishery. North American Journal of Fisheries Management, 34(3):602–608, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Merz:2014:MDG

[MGB^{+}14] Joseph E. Merz, Thomas M. Garrison, Paul S. Bergman, Scott Blankenship, and John Carlos Garza. Morphological discrimination of genetically distinct Chinook salmon populations: an example from California’s Central Valley. North American Journal of Fisheries Management, 34(6):1259–1269, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2014:BTT

[MGS14] Kevin A. Meyer, Edward O. Garton, and Daniel J. Schill. Bull trout trends in abundance and probabilities of persistence in Idaho. North American Journal of Fisheries Management, 34 (1):202–214, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mesa:2013:PRA

[MGWC13] Matthew G. Mesa, Lisa P. Gee, Lisa K. Weiland, and Helena E. Christiansen. Physiological responses of adult rainbow trout experimentally released through a unique fish conveyance device. North American Journal of Fisheries Management, 33(6):1179–1183, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meerbeek:2011:WCR

[MH11a] Jonathan R. Meerbeek and R. John H. Hoxmeier. Winter catch-and-release hooking mortality of saugers below lock and Dam 3 of the Mississippi River. North American Journal of Fisheries Management, 31(2):197–202, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2011:ARE

[MH11b] Kevin A. Meyer and Brett High. Accuracy of removal electrofishing estimates of trout abundance in Rocky Mountain streams. North American Journal of Fisheries Management, 31(5):923–933, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meerbeek:2013:STB

[MH13] Jonathan R. Meerbeek and Kimberly A. Hawkins. A standardized technique to back-calculate length at age from unsectioned walleye spines. North American Journal of Fisheries Management, 33(4):861–868, August 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mann:2011:CSM

[MHAO11] Kevin A. Mann, J. Marty Holtgren, Nancy A. Auer, and Stephanie A. Ogren. Comparing size, movement, and habitat selection of wild and streamside-reared lake sturgeon. North American Journal of Fisheries Management, 31(2):305–314, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McDougall:2013:RUM

[MHAP13] C. A. McDougall, C. L. Hrenchuk, W. G. Anderson, and S. J. Peake. The rapid upstream migration of pre-spawn lake sturgeon following trap-and-transport over a hydroelectric generating station. North American Journal of Fisheries Management, 33(6):1236–1242, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mikles:2019:ESR

[MHBR19] Chloe S. Mikles, Craig A. Harms, Jeffrey A. Buckel, and Paul J. Rudershausen. Examining the severity of roof-hooking injuries in dolphinfish: a comparison between computed tomography and Gross necropsy. North American Journal of Fisheries Management, 39(1):36–44, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Music:2010:MCS

[MHC10] Paul A. Music, James P. Hawkes, and Michael S. Cooperman. Magnitude and causes of smolt mortality in rotary screw traps: An Atlantic salmon case study. North American Journal of Fisheries Management, 30(3):713–722, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCubbins:2016:DCA

[MHDD16] Jonathan L. McCubbins, Michael J. Hansen, Joseph M. Dos-Santos, and Andrew M. Dux. Demographic characteristics of an adfluvial bull trout population in Lake Pend Oreille, Idaho. North American Journal of Fisheries Management, 36(6):1269–1277, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2011:RPI

[MHG^{+}11] Kevin A. Meyer, Brett High, Nick Gastelecutto, Elizabeth R. J. Mamer, and F. Steven Elle. Retention of passive integrated transponder tags in stream-dwelling rainbow trout. North American Journal of Fisheries Management, 31(2):236– 239, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mathur:2018:EJA

[MHR^{+}18] Dilip Mathur, Paul G. Heisey, Doug D. Royer, Eric J. White, Adam C. Slowik, Ray A. Bleistine, Brenda M. Pracheil, Kimberly A. Long, and Thomas J. Sullivan. Entrainment of juvenile and adult American shad at a pumped storage facility. North American Journal of Fisheries Management, 38(1): 56–75, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martin:2010:DUV

[MI10] Benjamin M. Martin and Elise R. Irwin. A digital underwater video camera system for aquatic research in regulated rivers. North American Journal of Fisheries Management, 30 (6):1365–1369, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mosel:2015:EDC

[MIH15] Kyle J. Mosel, Daniel A. Isermann, and Jonathan F. Hansen. Evaluation of daily creel and minimum length limits for black crappie and yellow perch in Wisconsin. North American Journal of Fisheries Management, 35(1):1–13, February 2015. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKenna:2011:LMB

[MJ11] James E. McKenna, Jr. and James H. Johnson. Landscape models of brook trout abundance and distribution in lotic habitat with field validation. North American Journal of Fisheries Management, 31(4):742–756, August 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2015:EPS

[MJCM15] Joshua L. McCormick, Laura S. Jackson, Fabian M. Carr, and Daniel H. Meyer III. Evaluation of probabilistic sampling designs for estimating abundance of multiple species of migratory fish using video recordings at fishways. North American Journal of Fisheries Management, 35(4):818–826, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miehls:2017:TNB

[MJH17] Scott M. Miehls, Nicholas S. Johnson, and Pete J. Hrodey. Test of a nonphysical barrier consisting of light, sound, and bubble screen to block upstream movement of sea lampreys in an experimental raceway. North American Journal of Fisheries Management, 37(3):660–666, June 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martinez:2013:PFB

[MK13] Patrick J. Martinez and A. Lawrence Kolz. Performance of four boat electrofishers with measured electrode resistances for electrofishing boats and rafts. North American Journal of Fisheries Management, 33(1):32–43, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Manuel:2013:HMP

[MKBB13] Kenneth L. Manuel, James P. Kirk, D. Hugh Barwick, and Tommy W. Bowen. Hydrilla management in Piedmont reservoirs using herbicides and triploid grass carp: a case study. North American Journal of Fisheries Management, 33(3):488– 492, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2017:DYC

[MKHC17] Kevin A. Meyer, Patrick Kennedy, Brett High, and Matthew R. Campbell. Distinguishing Yellowstone cutthroat trout, rainbow trout, and hybrids by use of field-based phenotypic characteristics. North American Journal of Fisheries Management, 37(2):456–466, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morgan:2012:SCR

[MKK^{+}12] Raymond P. Morgan II, Kathleen M. Kline, Matthew J. Kline, Susan F. Cushman, Matthew T. Sell, Roy E. Weitzell, Jr., and John B. Churchill. Stream conductivity: Relationships to land use, chloride, and fishes in Maryland streams. North American Journal of Fisheries Management, 32(5):941–952, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marsden:2014:ECE

[MKRH14] J. Ellen Marsden, Kevin P. Kelsey, Jacob W. Riley, and Joanna Hatt. Evaluation of calcein for estimating abundance of lake trout alevins on a spawning reef. North American Journal of Fisheries Management, 34(2):270–275, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:2018:FCR

[MKWM18] Jacob W. Miller, Patrick M. Kocovsky, Daniel Wiegmann, and Jeffrey G. Miner. Fish community responses to submerged aquatic vegetation in Maumee Bay, Western Lake Erie. North American Journal of Fisheries Management, 38(3):623–629, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Melstrom:2013:VRF

[ML13] Richard T. Melstrom and Frank Lupi. Valuing recreational fishing in the Great Lakes. North American Journal of Fisheries Management, 33(6):1184–1193, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mora:2015:ERA

[MLEK15] E. A. Mora, S. T. Lindley, D. L. Erickson, and A. P. Klimley. Estimating the riverine abundance of green sturgeon using a dual-frequency identification sonar. North American Journal of Fisheries Management, 35(3):557–566, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meerbeek:2013:RLF

[MLHS13] Jonathan R. Meerbeek, Joseph G. Larscheid, Michael J. Hawkins, and George D. Scholten. Retention of large-format, soft visible implant alphanumeric tags in walleye. North American Journal of Fisheries Management, 33(1):26–31, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mamer:2016:RRP

[MM16] Elizabeth R. J. M. Mamer and Kevin A. Meyer. Retention rates of passive integrated transponder tags, visible implant elastomer tags, and maxillary marks in wild trout. North American Journal of Fisheries Management, 36(5):1119–1124, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2017:SSE

[MM17] Joshua L. McCormick and Kevin A. Meyer. Sample size estimation for on-site creel surveys. North American Journal of Fisheries Management, 37(5):970–980, October 2017. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2018:SEE

[MM18a] Joshua L. McCormick and Kevin A. Meyer. Sensitivity of exploitation estimates to tag loss estimation methods in Idaho sport fisheries. North American Journal of Fisheries Management, 38(1):170–179, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Moreau:2018:ULD

[MM18b] C. M. Moreau and B. G. Matthias. Using limited data to identify optimal bag and size limits to prevent overfishing. North American Journal of Fisheries Management, 38(3):747– 758, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:2019:MDH

[MMB19] Andrew D. Miller, Robert Mollenhauer, and Shannon K. Brewer. Movement and diel habitat use of juvenile neosho smallmouth bass in an Ozark stream. North American Journal of Fisheries Management, 39(2):240–253, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morson:2017:EPS

[MMHM17] Jason M. Morson, Daphne Munroe, Ryan Harner, and Rachel Marshall. Evaluating the potential for a sex-balanced harvest approach in the recreational summer flounder fishery. North American Journal of Fisheries Management, 37(6):1231–1242, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Macdonald:2012:ERF

[MMP12] J. S. Macdonald, J. Morrison, and D. A. Patterson. The efficacy of reservoir flow regulation for cooling migration temperature for sockeye salmon in the Nechako River watershed of British Columbia. North American Journal of Fisheries Management, 32(3):415–427, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maynard:2017:BBD

[MMS^{+}17] George A. Maynard, Timothy B. Mihuc, V. Alex Sotola, Danielle E. Garneau, and Mark H. Malchoff. Black bass dispersal patterns following catch-and-release tournaments on Lake Champlain. North American Journal of Fisheries Management, 37(3):524–535, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Matala:2012:IHS

[MNYV12] Andrew P. Matala, Shawn R. Narum, William Young, and Jason L. Vogel. Influences of hatchery supplementation, spawner distribution, and habitat on genetic structure of Chinook salmon in the South Fork Salmon River, Idaho. North American Journal of Fisheries Management, 32(2):346–359, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McManamay:2010:GAH

[MODC10] Ryan A. McManamay, D. J. Orth, Charles A. Dolloff, and Mark A. Cantrell. Gravel addition as a habitat restoration technique for tailwaters. North American Journal of Fisheries Management, 30(5):1238–1257, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Michaletz:2012:IEV

[MOJ12] Paul H. Michaletz, Daniel V. Obrecht, and John R. Jones. Influence of environmental variables and species interactions on sport fish communities in small Missouri impoundments. North American Journal of Fisheries Management, 32(6): 1146–1159, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2014:EFS

[MP14] Joshua L. McCormick and Timothy K. Porter. Effect of fishing success on angler satisfaction on a Central Oregon rainbow trout fishery: Implications for establishing management objectives. North American Journal of Fisheries Management, 34 (5):938–944, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMillan:2015:URA

[MPL^{+}15] John R. McMillan, George R. Pess, Martin Liermann, Sarah A. Morley, Michael L. McHenry, Lance A. Campbell, and Thomas P. Quinn. Using redd attributes, fry density, and otolith microchemistry to distinguish the presence of steelhead and rainbow trout in the Elwha River Dam Removal Project. North American Journal of Fisheries Management, 35(5):1019–1033, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Murdoch:2010:ESE

[MPM10] Andrew R. Murdoch, Todd N. Pearsons, and Travis W. Maitland. Estimating the spawning escapement of hatchery- and natural-origin spring Chinook salmon using redd and carcass data. North American Journal of Fisheries Management, 30 (2):361–375, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Marcy-Quay:2019:TVS

[MQJK19] Benjamin Marcy-Quay, Kurt J. Jirka, and Clifford E. Kraft. Time versus space: Choice of effort metric can avoid bias in boat electrofishing catch per unit effort. North American Journal of Fisheries Management, 39(6):1207–1216, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2012:ESD

[MQS12] Joshua L. McCormick, Michael C. Quist, and Daniel J. Schill. Effect of survey design and catch rate estimation on total catch estimates in Chinook salmon fisheries. North American Journal of Fisheries Management, 32(6):1090–1101, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2013:CSS

[MQS13a] Joshua L. McCormick, Michael C. Quist, and Daniel J. Schill. Creel survey sampling designs for estimating effort in shortduration Chinook salmon fisheries. North American Journal of Fisheries Management, 33(5):977–993, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2013:SRB

[MQS13b] Joshua L. McCormick, Michael C. Quist, and Daniel J. Schill. Self-reporting bias in Chinook salmon sport fisheries in Idaho: Implications for roving creel surveys. North American Journal of Fisheries Management, 33(4):723–731, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McCormick:2016:SAM

[MR16] Joshua L. McCormick and James R. Ruzycki. A simulation assessment of model-assisted estimation of steelhead redd abundance. North American Journal of Fisheries Management, 36 (4):913–925, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meredith:2014:RIW

[MRA14] Christy Meredith, Brett Roper, and Eric Archer. Reductions in instream wood in streams near roads in the interior Columbia River Basin. North American Journal of Fisheries Management, 34(3):493–506, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Madenjian:2011:SHU

[MRB^{+}11] Charles P. Madenjian, Edward S. Rutherford, Marc A. Blouin, Bryan J. Sederberg, and Jeff R. Elliott. Spawning habitat unsuitability: An impediment to cisco rehabilitation in Lake Michigan? North American Journal of Fisheries Management, 31(5):905–913, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mesa:2012:FBE

[MRC12] Matthew G. Mesa, Brien P. Rose, and Elizabeth S. Copeland. Field-based evaluations of horizontal flat-plate fish screens, II: Testing of a unique off-stream channel device — the farmers screen. North American Journal of Fisheries Management, 32 (3):604–612, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2014:USA

[MS14] Kevin A. Meyer and Daniel J. Schill. Use of a statewide angler tag reporting system to estimate rates of exploitation and total mortality for Idaho sport fisheries. North American Journal of Fisheries Management, 34(6):1145–1158, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Maceina:2016:AAP

[MS16] Michael J. Maceina and Steven M. Sammons. Assessing the accuracy of published natural mortality estimators using rates determined from five unexploited freshwater fish populations. North American Journal of Fisheries Management, 36(2):433– 446, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Martin:2017:TTS

[MS17a] Douglas J. Martin and Alice Shelly. Temporal trends in stream habitat on managed forestlands in coastal Southeast Alaska. North American Journal of Fisheries Management, 37(4):882–902, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2017:NAJ

[MS17b] Kevin A. Meyer and Daniel J. Schill. North American journal of fisheries management guide for authors. North American Journal of Fisheries Management, 37(1):256–261, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mann:2018:PSM

[MS18] Ryan D. Mann and Charles G. Snow. Population-specific migration patterns of wild adult summer-run Chinook salmon passing Wells Dam, Washington. North American Journal of Fisheries Management, 38(2):377–392, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Monroy:2010:DFG

[MSBP10] Carmen Monroy, Silvia Salas, and Javier Bello-Pineda. Dynamics of fishing gear and spatial allocation of fishing effort in a multispecies fleet. North American Journal of Fisheries Management, 30(5):1187–1202, October 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McKenna:2013:BSP

[MSC13] James E. McKenna, Jr., Michael T. Slattery, and Kean M. Clifford. Broad-scale patterns of brook trout responses to introduced brown trout in New York. North American Journal of Fisheries Management, 33(6):1221–1235, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

McMichael:2011:SJC

[MSD11] Geoffrey A. McMichael, John R. Skalski, and Katherine A. Deters. Survival of juvenile Chinook salmon during barge transport. North American Journal of Fisheries Management, 31 (6):1187–1196, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mrnak:2018:CAT

[MSE^{+}18] Joseph T. Mrnak, Stephanie L. Shaw, Lawrence D. Eslinger, Thomas A. Cichosz, and Greg G. Sass. Characterizing the angling and tribal spearing walleye fisheries in the ceded territory of Wisconsin, 1990–2015. North American Journal of Fisheries Management, 38(6):1381–1393, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

MacPherson:2012:ECS

[MSFS12] Laura M. MacPherson, Michael G. Sullivan, A. Lee Foote, and Cameron E. Stevens. Effects of culverts on stream fish assemblages in the Alberta foothills. North American Journal of Fisheries Management, 32(3):480–490, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mocq:2015:IEP

[MSHC15] J. Mocq, A. St-Hilaire, and R. A. Cunjak. Influences of experts’ personal experiences in fuzzy logic modeling of Atlantic salmon habitat. North American Journal of Fisheries Management, 35(2):271–280, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2016:PAW

[MSK^{+}16] Kevin A. Meyer, Christopher L. Sullivan, Patrick Kennedy, Daniel J. Schill, David M. Teuscher, Arnie F. Brimmer, and D. Tommy King. Predation by American white pelicans and double-crested cormorants on catchable-sized hatchery rainbow trout in select Idaho lentic waters. North American Journal of Fisheries Management, 36(2):294–308, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:2018:GSS

[MSK18] Brett T. Miller, Casey W. Schoenebeck, and Keith D. Koupal. Gear- and season-specific catch rates of age-0 walleyes and white bass: Standard sampling recommendations for Great Plains reservoirs. North American Journal of Fisheries Management, 38(4):903–910, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Morrisett:2019:PRU

[MSK19a] Christina N. Morrisett, John R. Skalski, and Russell B. Kiefer. Passage route and upstream migration success: a case study of Snake River salmonids ascending Lower Granite Dam. North American Journal of Fisheries Management, 39(1):58– 68, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Munter:2019:ERC

[MSK19b] R. A. Munter, C. W. Schoenebeck, and K. D. Koupal. Experimentally reducing competition and other biotic limitations can impact growth, food habits, and activity and lead to age-1 maturation in yellow perch. North American Journal of Fisheries Management, 39(2):362–371, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meyer:2014:SRT

[MSM^{+}14] Kevin A. Meyer, Daniel J. Schill, Elizabeth R. J. M. Mamer, Christine C. Kozfkay, and Matthew R. Campbell. Status of redband trout in the Upper Snake River Basin of Idaho. North American Journal of Fisheries Management, 34(3):507–523, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Merz:2012:OMP

[MSS^{+}12] Joseph E. Merz, Paul Skvorc, Susan M. Sogard, Clark Watry, Scott M. Blankenship, and Erwin E. Van Nieuwenhuyse. Onset of melanophore patterns in the head region of Chinook salmon: a natural marker for the reidentification of individual fish. North American Journal of Fisheries Management, 32(4):806–816, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Mayfield:2014:USR

[MSW^{+}14] M. P. Mayfield, L. D. Schultz, L. A. Wyss, M. E. Colvin, and C. B. Schreck. Using spatial resampling to assess redd count survey length requirements for Pacific lamprey. North American Journal of Fisheries Management, 34(5):923–931, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Meerbeek:2019:EHB

[MW19] Jonathan R. Meerbeek and Michael J. Weber. Effects of hatchery broodstock collection on adult muskellunge populations. North American Journal of Fisheries Management, 39(4):807–816, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Miller:2015:UGM

[MWS15] Loren M. Miller, Matthew C. Ward, and Douglas W. Schultz. Using genetic markers as individual tags: a case study of a mark–recapture estimate of adult muskellunge population size. North American Journal of Fisheries Management, 35(2):210– 215, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nakamura:2013:EFR

[Nak13] Tomoyuki Nakamura. Effects of flow reduction on a whitespotted char population in a Japanese mountain stream. North American Journal of Fisheries Management, 33(6):1142–1148, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nutile:2012:EME

[NASG12] Sam Nutile, Jon J. Amberg, María S. Sepúlveda, and Reuben R. Goforth. Effects of multiple electrical field exposures on cyprinid embryo survival. North American Journal of Fisheries Management, 32(5):875–879, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Newman:2010:HMJ

[NB10] Ken B. Newman and Patricia L. Brandes. Hierarchical modeling of juvenile Chinook salmon survival as a function of Sacramento–San Joaquin Delta water exports. North American Journal of Fisheries Management, 30(1):157–169, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neal:2017:EPS

[NBFL17] J. Wesley Neal, Jason M. Bies, Cynthia N. Fox, and Craig G. Lilyestrom. Evaluation of proposed speckled peacock bass Cichla temensis introduction to Puerto Rico. North American Journal of Fisheries Management, 37(5):1093–1106, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neufeld:2015:RSE

[NBP15] Kenton Neufeld, Salvatore Blair, and Mark Poesch. Retention and stress effects of visible implant tags when marking western silvery minnow and its application to other cyprinids (family Cyprinidae). North American Journal of Fisheries Management, 35(5):1070–1076, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neufeld:2011:MLO

[NCJ^{+}11] Matthew D. Neufeld, Kenneth D. Cain, Nathan R. Jensen, Susan C. Ireland, and Vaughn L. Paragamian. Movement of lake-origin burbot reared in a hatchery environment and released into a large river drainage. North American Journal of Fisheries Management, 31(1):56–62, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nett:2012:DIR

[NCMT12] Joseph H. G. Nett, Timothy B. Campbell, Nicholas E. Mandrak, and Scott D. Tiegs. Detecting invasive round goby in wadeable streams: a comparison of gear types. North American Journal of Fisheries Management, 32(2):360–364, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Naskar:2017:MFQ

[NCSR17] Malay Naskar, Ganesh Chandra, S. K. Sahu, and R. K. Raman. A modeling framework to quantify the influence of hydrology on the abundance of a migratory Indian shad, the Hilsa Tenualosa ilisha. North American Journal of Fisheries Management, 37(6):1208–1219, December 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neville:2011:PHN

[ND11] Helen M. Neville and Jason B. Dunham. Patterns of hybridization of nonnative cutthroat trout and hatchery rainbow trout with native redband trout in the Boise River, Idaho. North American Journal of Fisheries Management, 31(6):1163–1176, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nohner:2015:MSS

[ND15] Joel K. Nohner and James S. Diana. Muskellunge spawning site selection in Northern Wisconsin lakes and a GIS-Based predictive habitat model. North American Journal of Fisheries Management, 35(1):141–157, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nyce:2013:GPS

[NEC^{+}13] Leslie G. Nyce, Lisa Eby, Christopher G. Clancy, Sally Painter, and Robb F. Leary. Genetic population structure of bull trout in the East Fork Bitterroot River Drainage, Montana. North American Journal of Fisheries Management, 33 (2):432–445, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nwosu:2010:AAF

[NEH10] Francis M. Nwosu, Udeme I. Enin, and Sieghard Holzlöhner. Assessment of artisanal fishery impacts on West African croaker Pseudotolithus elongatus in the Cross River Estuary, Nigeria, using length-based models. North American Journal of Fisheries Management, 30(4):860–865, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ng:2016:PDE

[NFQ16] Elizabeth L. Ng, Jim P. Fredericks, and Michael C. Quist. Population dynamics and evaluation of alternative management strategies for nonnative lake trout in Priest Lake, Idaho. North American Journal of Fisheries Management, 36(1):40– 54, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neely:2016:URC

[NKC16] Ben C. Neely, Jeff D. Koch, and Michael E. Colvin. Utility of reservoir characteristics to determine minimum sampling effort needed to assess sport fish populations in Kansas reservoirs. North American Journal of Fisheries Management, 36 (2):285–293, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Naughton:2018:TEP

[NKC^{+}18] George P. Naughton, Matthew L. Keefer, Tami S. Clabough, Matthew J. Knoff, Timothy J. Blubaugh, and Christopher C. Caudill. Tag effects on prespawn mortality of Chinook salmon: a field experiment using passive integrated transponder tags, radio transmitters, and untagged controls. North American Journal of Fisheries Management, 38(1):96–103, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neely:2018:EGS

[NKL18] Ben C. Neely, Jeff D. Koch, and Sean T. Lynott. Effects of gizzard shad reduction on relative abundance, growth, and body condition of bluegills and largemouth bass in a small Kansas impoundment. North American Journal of Fisheries Management, 38(4):856–866, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neely:2017:FBR

[NLK17] Ben C. Neely, Sean T. Lynott, and Jeff D. Koch. Freeze brand retention in channel catfish and channel catfish × blue catfish hybrids. North American Journal of Fisheries Management, 37(6):1299–1303, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nyqvist:2017:DMM

[NMG^{+}17] D. Nyqvist, S. D. McCormick, L. Greenberg, W. R. Ardren, E. Bergman, O. Calles, and T. Castro-Santos. Downstream migration and multiple dam passage by Atlantic salmon smolts. North American Journal of Fisheries Management, 37(4):816–828, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neal:2012:SDT

[NP12] J. Wesley Neal and Marie Prchalová. Spatiotemporal distributions of threadfin shad in tropical reservoirs. North American Journal of Fisheries Management, 32(5):929–940, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Neebling:2011:CBE

[NQ11] Travis E. Neebling and Michael C. Quist. Comparison of boat electrofishing, trawling, and seining for sampling fish assemblages in Iowa’s nonwadeable rivers. North American Journal of Fisheries Management, 31(2):390–402, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nannini:2016:LSC

[NSW16] Michael Nannini, Daniel Schermerhorn, and David Wahl. Lethal stomach content collection methods reduce detection of zooplankton prey of juvenile largemouth bass compared with nonlethal methods. North American Journal of Fisheries Management, 36(4):738–743, August 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Negrea:2014:ADT

[NTJ^{+}14] Cristi Negrea, Donald E. Thompson, Steven D. Juhnke, Derek S. Fryer, and Frank J. Loge. Automated detection and tracking of adult Pacific lampreys in underwater video collected at Snake and Columbia River fishways. North American Journal of Fisheries Management, 34(1):111–118, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nagid:2015:ERD

[NTJ15] Eric J. Nagid, Travis Tuten, and Kevin G. Johnson. Effects of reservoir drawdowns and the expansion of hydrilla coverage on year-class strength of largemouth bass. North American Journal of Fisheries Management, 35(1):54–61, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nesslage:2012:PSP

[NW12] Geneviève M. Nesslage and Michael J. Wilberg. Performance of surplus production models with time-varying parameters for assessing multispecies assemblages. North American Journal of Fisheries Management, 32(6):1137–1145, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Nuhfer:2014:CBT

[NWZ14] Andrew J. Nuhfer, Todd C. Wills, and Troy G. Zorn. Changes to a brown trout population after introducing steelhead in a Michigan stream. North American Journal of Fisheries Management, 34(2):411–423, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

OMalley:2017:SAS

[OEZ17] Andrew J. O’Malley, Claire Enterline, and Joseph Zydlewski. Size and age structure of anadromous and landlocked populations of rainbow smelt. North American Journal of Fisheries Management, 37(2):326–336, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ogren:2015:AIF

[OH15] Stephanie A. Ogren and Casey J Huckins. Applicability and interpretation of fish indices of biotic integrity (IBI) for bioassessment in the Upper Midwest. North American Journal of Fisheries Management, 35(2):281–295, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

ODonnell:2010:UPA

[OHL10] Matthew J. O’Donnell, Gregg E. Horton, and Benjamin H. Letcher. Use of portable antennas to estimate abundance of PIT-Tagged fish in small streams: Factors affecting detection probability. North American Journal of Fisheries Management, 30(2):323–336, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ogburn:2017:GJA

[OHSP17] Matthew B. Ogburn, Daniel J. Hasselman, Thomas F. Schultz, and Eric P. Palkovacs. Genetics and juvenile abundance dynamics show congruent patterns of population structure for depleted river herring populations in the Upper Chesapeake Bay. North American Journal of Fisheries Management, 37 (5):1083–1092, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ogle:2017:EAS

[OI17] Derek H. Ogle and Daniel A. Isermann. Estimating age at a specified length from the von Bertalanffy growth function. North American Journal of Fisheries Management, 37 (5):1176–1180, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

ODonnell:2017:IMP

[OL17] Matthew J. O’Donnell and Benjamin H. Letcher. Implanting 8-mm passive integrated transponder tags into small brook trout: Effects on growth and survival in the laboratory. North American Journal of Fisheries Management, 37(3):605–611, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Oele:2015:PBA

[OLM15] Daniel L. Oele, Zachary J. Lawson, and Peter B. McIntyre. Precision and bias in aging northern pike: Comparisons among four calcified structures. North American Journal of Fisheries Management, 35(6):1177–1184, December 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Odenkirk:2013:IAG

[OLOI13] John Odenkirk, Catherine Lim, Steve Owens, and Mike Isel. Insight into age and growth of northern snakehead in the Potomac River. North American Journal of Fisheries Management, 33(4):773–776, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olsen:2019:SSL

[Ols19] Zachary Olsen. The selection of stocking locations for fisheries enhancement in the estuarine environment: a decision support tool. North American Journal of Fisheries Management, 39 (6):1242–1259, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Osborn:2010:RBS

[OM10] Maury F. Osborn and Gary C. Matlock. Recall bias in a sportfishing mail survey. North American Journal of Fisheries Management, 30(3):665–670, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ohnuki:2015:NEC

[OMT^{+}15] Tsutomu Ohnuki, Kentaro Morita, Hiroshi Tokuda, Yasutaka Okamoto, and Kazumasa Ohkuma. Numerical and economic contributions of wild and hatchery pink salmon to commercial catches in Japan estimated from mass otolith markings. North American Journal of Fisheries Management, 35(3):598–604, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Olson:2012:MTG

[OOWW12] Justin A. Olson, Jenae M. Olson, Rachel E. Walsh, and Brian D. Wisenden. A method to train groups of predatornaive fish to recognize and respond to predators when released into the natural environment. North American Journal of Fisheries Management, 32(1):77–81, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Owensby:2017:MDH

[ORA17] Dylan P. Owensby, James A. Rice, and D. Derek Aday. Mortality, dispersal, and habitat use of stocked juvenile muskellunge in two Western North Carolina rivers. North American Journal of Fisheries Management, 37(1):108–121, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

ONeal:2016:CSR

[ORC^{+}16] Jennifer S. O’Neal, Phil Roni, Bruce Crawford, Anna Ritchie, and Alice Shelly. Comparing stream restoration project effectiveness using a programmatic evaluation of salmonid habitat and fish response. North American Journal of Fisheries Management, 36(3):681–703, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

ODonnell:2019:GPS

[ORD19] Timothy P. O’Donnell, Marcel J. M. Reichert, and Tanya L. Darden. Genetic population structure of white grunt in the Southeastern United States. North American Journal of Fisheries Management, 39(4):725–737, August 2019. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Oele:2016:DHS

[ORL^{+}16] Daniel L. Oele, Andrew L. Rypel, John Lyons, Paul Cunningham, and Tim Simonson. Do higher size and reduced bag limits improve northern pike size structure in Wisconsin lakes? North American Journal of Fisheries Management, 36 (5):982–994, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Powers:2016:ERE

[PA16] Sean P. Powers and Kevin Anson. Estimating recreational effort in the Gulf of Mexico red snapper fishery using boat ramp cameras: Reduction in Federal season length does not proportionally reduce catch. North American Journal of Fisheries Management, 36(5):1156–1166, October 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Petersen:2013:MPB

[PB13] Matthew J. Petersen and Phillip W. Bettoli. Mortality of palmetto bass following catch-and-release angling. North American Journal of Fisheries Management, 33(4):806–810, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perez:2017:CAF

[PBA^{+}17] Christina R. Perez, Scott A. Bonar, Jon J. Amberg, Bridget Ladell, Chris Rees, William T. Stewart, Curtis J. Gill, Chris Cantrell, and Anthony T. Robinson. Comparison of American fisheries society (AFS) standard fish sampling techniques and environmental DNA for characterizing fish communities in a large reservoir. North American Journal of Fisheries Management, 37(5):1010–1027, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perry:2012:CPB

[PC12] Robert C. Perry and John M. Casselman. Comparisons of precision and bias with two age interpretation techniques for opercular bones of longnose sucker, a long-lived northern fish. North American Journal of Fisheries Management, 32(4): 790–795, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Patrick:2013:EFF

[PCP^{+}13] Paul H. Patrick, Elaine Chen, Jason Parks, Jennifer Powell, J. S. Poulton, and Cherie-Lee Fietsch. Effects of fixed and fluctuating temperature on hatch of round whitefish and lake whitefish eggs. North American Journal of Fisheries Management, 33(6):1091–1099, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pritt:2019:SAR

[PCP^{+}19] Jeremy J. Pritt, Joseph D. Conroy, Kevin S. Page, Stephen M. Tyszko, and Richard D. Zweifel. Standard assessments reveal context-dependent responses of crappie populations to a length-based regulation in Ohio reservoirs. North American Journal of Fisheries Management, 39(5):999–1014, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phelps:2014:RBM

[PCT^{+}14] Nicholas B. D. Phelps, Meggan E. Craft, Dominic Travis, Katharine Pelican, and Sagar M. Goyal. Risk-based management of viral hemorrhagic septicemia virus in Minnesota. North American Journal of Fisheries Management, 34(2):373– 379, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Panek:2013:FST

[PD13] Frank M. Panek and Christine L. Densmore. Frequency and severity of trauma in fishes subjected to multiple-pass depletion electrofishing. North American Journal of Fisheries Management, 33(1):178–185, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pennock:2018:POR

[PDD^{+}18] Casey A. Pennock, Scott L. Durst, Bobby R. Duran, Brian A. Hines, C. Nathan Cathcart, Jason E. Davis, Benjamin J. Schleicher, and Nathan R. Franssen. Predicted and observed responses of a nonnative channel catfish population following managed removal to aid the recovery of endangered fishes. North American Journal of Fisheries Management, 38(3):565– 578, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pritt:2019:SGP

[PDH^{+}19] Jeremy J. Pritt, Michael A. Depew, Jason R. Herrala, Jana Hirst, Craig R. Jansen, and Katherine J. Zipfel. Spatial gradients in population characteristics of Ohio River sauger and implications for a length-based regulation. North American Journal of Fisheries Management, 39(6):1412–1425, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prichard:2018:CLR

[PDL^{+}18] Carson G. Prichard, James L. Dexter, Jr., Neil D. Ledet, Richard O’Neal, Martha Wolgamood, Jan P. VanAmberg, and Jory L. Jonas. Comparison of lake and river fishery performance between two strains of steelhead in Eastern Lake Michigan. North American Journal of Fisheries Management, 38(5):1138–1151, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Paradis:2012:CCE

[PDM12] Yves Paradis, Angélique Dupuch, and Pierre Magnan. Comparison of catch efficiencies between black and galvanized minnow traps. North American Journal of Fisheries Management, 32(3):539–543, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2017:EFA

[PFD17] Matthew L. Peterson, Andrea N. Fuller, and Doug Demko. Environmental factors associated with the upstream migration of fall-run Chinook salmon in a regulated river. North American Journal of Fisheries Management, 37(1):78–93, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pepino:2012:IHC

[PFRM12] Marc Pépino, Jan Franssen, Marco A. Rodríguez, and Pierre Magnan. Impacts of highway construction on redd counts of stream-dwelling brook trout. North American Journal of Fisheries Management, 32(6):1167–1174, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pennock:2016:STR

[PFWG16] Casey A. Pennock, Bryan D. Frenette, Micah J. Waters, and Keith B. Gido. Survival of and tag retention in southern redbelly dace injected with two sizes of PIT tags. North American Journal of Fisheries Management, 36(6):1386–1394, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Portley:2014:LRP

[PG14] Nicole Portley and Harold J. Geiger. Limit reference points for Pacific salmon fisheries. North American Journal of Fisheries Management, 34(2):401–410, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Powell:2010:PCF

[PHFK10] M. S. Powell, R. W. Hardy, T. A. Flagg, and P. A. Kline. Proximate composition and fatty acid differences in hatcheryreared and wild Snake River sockeye salmon overwintering in nursery lakes. North American Journal of Fisheries Management, 30(2):530–537, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pine:2013:IBM

[PHS^{+}13] William E. Pine III, Brian Healy, Emily Omana Smith, Melissa Trammell, Dave Speas, Rich Valdez, Mike Yard, Carl Walters, Rob Ahrens, Randy Vanhaverbeke, Dennis Stone, and Wade Wilson. An individual-based model for population viability analysis of humpback chub in Grand Canyon. North American Journal of Fisheries Management, 33(3):626–641, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pierce:2010:LTE

[Pie10] Rodney B. Pierce. Long-term evaluations of length limit regulations for northern pike in Minnesota. North American Journal of Fisheries Management, 30(2):412–432, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pate:2014:MCK

[PJLB14] William M. Pate, Brett M. Johnson, Jesse M. Lepak, and Daniel Brauch. Managing for coexistence of Kokanee and trophy lake trout in a montane reservoir. North American Journal of Fisheries Management, 34(5):908–922, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pacheco:2018:UTW

[PK18] James M. Pacheco and Jonathan H. Kohr. Update to the toe-width methodology: Determining preferred discharges for spawning and rearing salmonids. North American Journal of Fisheries Management, 38(6):1404–1414, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perret:2010:EHK

[PKKR10] Alexander J. Perret, Michael D. Kaller, William E. Kelso, and D. Allen Rutherford. Effects of Hurricanes Katrina and Rita on sport fish community abundance in the Eastern Atchafalaya River Basin, Louisiana. North American Journal of Fisheries Management, 30(2):511–517, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pedicillo:2013:ASD

[PLCG13] Giovanni Pedicillo, Massimo Lorenzoni, Antonella Carosi, and Lucia Ghetti. Application of stock density indices as a tool for broad-scale population assessment for four cyprinid species in Central Italy. North American Journal of Fisheries Management, 33(1):153–162, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2017:GCL

[PLCN17] Nick R. Peterson, Dijar Lutz-Carrillo, and J. Wesley Neal. Genetic composition of largemouth bass in Puerto Rico reservoirs. North American Journal of Fisheries Management, 37 (6):1199–1207, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pinter:2019:SBF

[PLUH19] Kurt Pinter, Erwin Lautsch, Günther Unfer, and Daniel S. Hayes. Snorkeling-based fish stock assessment by anglers — a valuable method for managing recreational fisheries. North American Journal of Fisheries Management, 39(1):82– 90, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pothoven:2013:IPL

[PM13] Steven A. Pothoven and Charles P. Madenjian. Increased piscivory by lake whitefish in Lake Huron. North American Journal of Fisheries Management, 33(6):1194–1202, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peer:2014:CCM

[PM14] A. C. Peer and T. J. Miller. Climate change, migration phenology, and fisheries management interact with unanticipated consequences. North American Journal of Fisheries Management, 34(1):94–110, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Patrick:2014:EEP

[PMP^{+}14] Paul H. Patrick, Elaine Mason, Jennifer Powell, Scott Milne, and J. S. Poulton. Evaluating the effectiveness of the Pickering Nuclear Generating Station Fish Diversion System Barrier Net. North American Journal of Fisheries Management, 34 (2):287–300, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2019:CMV

[PN19] Douglas P. Peterson and Helen M. Neville. Comparison of methods to verify upstream passage by trout at remediated culverts in four Rocky Mountain streams. North American Journal of Fisheries Management, 39(4):738–752, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pipal:2012:EEL

[PNHA12] Kerrie A. Pipal, Jeremy J. Notch, Sean A. Hayes, and Peter B. Adams. Estimating escapement for a low-abundance steelhead population using dual-frequency identification sonar (DIDSON). North American Journal of Fisheries Management, 32(5):880–893, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Page:2011:SEN

[PNWC11] Kevin S. Page, Mary T. Negus, Matthew C. Ward, and Tracy L. Close. Simulating effects of nonintrogressive hybridization with a stocked hatchery strain of rainbow trout on the sustainability and recovery of naturalized steelhead populations in Minnesota waters of Lake Superior. North American Journal of Fisheries Management, 31(6):1065–1076, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peirce:2011:RES

[POWF11] Joshua M. Peirce, Edward O. Otis, Mark S. Wipfli, and Erich H. Follmann. Radiotelemetry to estimate stream life of adult chum salmon in the McNeil River, Alaska. North American Journal of Fisheries Management, 31(2):315–322, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Price:2010:EME

[PP10] Alison L. Price and James T. Peterson. Estimation and modeling of electrofishing capture efficiency for fishes in wadeable warmwater streams. North American Journal of Fisheries Management, 30(2):481–498, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pritchett:2011:PFA

[PP11] Jennifer Pritchett and Mark Pyron. Predicting fish assemblages from substrate variation in a turbid river: Grab samples compared with pole probing. North American Journal of Fisheries Management, 31(3):574–579, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Purchase:2018:PIJ

[PPFC18] Craig F. Purchase, Brittany Palm-Flawd, and Louis Charron. Performance of instream Jordan–Scotty salmon egg incubators under different installation and sedimentation conditions. North American Journal of Fisheries Management, 38(2):462– 468, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Puchala:2018:SAS

[PPO18] Elizabeth A. Puchala, Donna L. Parrish, and Derek H. Ogle. Size and age of stonecats in Lake Champlain; estimating growth at the margin of their range to aid in population management. North American Journal of Fisheries Management, 38(6):1316–1323, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Parsley:2011:STR

[PPR11] Michael J. Parsley, Nicholas D. Popoff, and Jason G. Romine. Short-term response of subadult white sturgeon to Hopper dredge disposal operations. North American Journal of Fisheries Management, 31(1):1–11, February 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Price:2010:FPE

[PQB10] David M. Price, Timothy Quinn, and Robert J. Barnard. Fish passage effectiveness of recently constructed road crossing culverts in the Puget Sound region of Washington State. North American Journal of Fisheries Management, 30(5):1110–1125, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perales:2015:ELC

[PRM15] K. Martin Perales, Jay Rowan, and Peter B. Moyle. Evidence of landlocked Chinook salmon populations in California. North American Journal of Fisheries Management, 35(6):1101–1105, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pilgrim:2013:MGB

[PRR^{+}13] N. L. Pilgrim, R. Royer, L. Ripley, J. Underwood, J. B. Rasmussen, and A. Hontela. Multicompartment gravel bed flume to evaluate swim-up success of salmonid larvae. North American Journal of Fisheries Management, 33(1):71–78, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pritt:2015:ULF

[PRRD15] Jeremy J. Pritt, Edward F. Roseman, Jason E. Ross, and Robin L. DeBruyne. Using larval fish community structure to guide long-term monitoring of fish spawning activity. North American Journal of Fisheries Management, 35(2):241–252, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Price:2014:CPS

[PRTS14] Michael H. H. Price, Andrew G. J. Rosenberger, Greg G. Taylor, and Jack A. Stanford. Comment: Population structure and run timing of sockeye salmon in the Skeena River, British Columbia. North American Journal of Fisheries Management, 34(6):1167–1170, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Prager:2010:DAB

[PS10] Michael H. Prager and Kyle W. Shertzer. Deriving acceptable biological catch from the overfishing limit: Implications for assessment models. North American Journal of Fisheries Management, 30(1):289–294, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perkin:2010:MMA

[PSB^{+}10a] Joshuah S. Perkin, Zachary R. Shattuck, Preston T. Bean, Timothy H. Bonner, Ekaterina Saraeva, and Thomas B. Hardy. Movement and microhabitat associations of guadalupe bass in two Texas rivers. North American Journal of Fisheries Management, 30(1):33–46, February 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Perry:2010:ESM

[PSB^{+}10b] Russell W. Perry, John R. Skalski, Patricia L. Brandes, Philip T. Sandstrom, A. Peter Klimley, Arnold Ammann, and Bruce MacFarlane. Estimating survival and migration route probabilities of juvenile Chinook salmon in the Sacramento– San Joaquin River Delta. North American Journal of Fisheries Management, 30(1):142–156, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2015:EEM

[PSC15] James T. Peterson, Paul D. Scheerer, and Shaun Clements. An evaluation of the efficiency of minnow traps for estimating the abundance of minnows in desert spring systems. North American Journal of Fisheries Management, 35(3):491–502, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Peterson:2013:PMA

[PSCL13] N. Phil Peterson, Ryan K. Simmons, Tamre Cardoso, and Jeffrey T. Light. A probabilistic model for assessing passage performance of coastal cutthroat trout through corrugated metal culverts. North American Journal of Fisheries Management, 33(1):192–199, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pol:2011:FLT

[PSCS11] Michael V. Pol, Mark J. Szymanski, David M. Chosid, and Daniel Salerno. Fork length–total length conversions for haddock and pollock. North American Journal of Fisheries Management, 31(3):427–430, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Patterson:2016:PTL

[PSR16] Kristen A. Patterson, Jeffrey A. Stein, and Steven R. Robillard. Progress toward lake trout rehabilitation at a stocked and unstocked reef in Southern Lake Michigan. North American Journal of Fisheries Management, 36(6):1405–1415, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pratt:2011:PEL

[PT11] Thomas C. Pratt and Ron W. Threader. Preliminary evaluation of a large-scale American eel conservation stocking experiment. North American Journal of Fisheries Management, 31(4):619–628, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pacas:2015:NEI

[PT15] Charlie Pacas and Mark K. Taylor. Nonchemical eradication of an introduced trout from a headwater complex in Banff National Park, Canada. North American Journal of Fisheries Management, 35(4):748–754, August 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pouder:2010:CMD

[PTD10] William F. Pouder, Nicholas A. Trippel, and Jason R. Dotson. Comparison of mortality and diet composition of pelletreared advanced-fingerling and early-cohort age-0 wild largemouth bass through 90 days poststocking at Lake Seminole, Florida. North American Journal of Fisheries Management, 30(5):1270–1279, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phelps:2010:WTR

[PTH^{+}10] Quinton E. Phelps, Sara J. Tripp, William D. Hintz, James E. Garvey, David P. Herzog, David E. Ostendorf, Joseph W. Ridings, Jason W. Crites, and Robert A. Hrabik. Water temperature and river stage influence mortality and abundance of naturally occurring Mississippi River Scaphirhynchus sturgeon. North American Journal of Fisheries Management, 30 (3):767–775, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ptolemy:2013:PMD

[Pto13] Ronald A. Ptolemy. Predictive models for differentiating habitat use of coastal cutthroat trout and steelhead at the reach and landscape scale. North American Journal of Fisheries Management, 33(6):1210–1220, December 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Pulver:2017:DLC

[Pul17] J. R. Pulver. Does the lunar cycle affect reef fish catch rates? North American Journal of Fisheries Management, 37(3):536– 549, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phelps:2013:CME

[PVH^{+}13] Quinton E. Phelps, Ivan Vining, David P. Herzog, Ross Dames, Vince H. Travnichek, Sara J. Tripp, and Mark Boone. A comparison of methods to estimate shovelnose sturgeon mortality in the Mississippi River adjacent to Missouri and Illinois. North American Journal of Fisheries Management, 33(4):754–761, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Popowich:2011:VMC

[PVST11] Ryan C. Popowich, Paul A. Venturelli, Jim D. Stelfox, and Eric B. Taylor. Validation of morphological characteristics used for field identification of bull trout × brook trout hybrids. North American Journal of Fisheries Management, 31(3):548– 553, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Phelps:2013:DAC

[PW13] Quinton E. Phelps and David W. Willis. Development of an Asian carp size structure index and application through demonstration. North American Journal of Fisheries Management, 33(2):338–343, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Page:2012:DAK

[PZC^{+}12] Kevin S. Page, Richard D. Zweifel, George Carter, Nick Radabaugh, Michael Wilkerson, Matthew Wolfe, Michael Greenlee, and Kipp Brown. Do anglers know what they catch? Identification accuracy and its effect on angler survey-derived catch estimates. North American Journal of Fisheries Management, 32(6):1080–1089, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:2016:ARC

[QA16] Jeffrey W. Quinn and Diana Andrews. Angler returns of channel catfish stocked in Arkansas lakes and streams. North American Journal of Fisheries Management, 36(3):576–583, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:2011:MSB

[QK11] Jeffrey W. Quinn and Thomas J. Kwak. Movement and survival of brown trout and rainbow trout in an Ozark tailwater river. North American Journal of Fisheries Management, 31 (2):299–304, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Quinn:2010:SBT

[Qui10] Jeffrey W. Quinn. A survey of bowfishing tournaments in Arkansas. North American Journal of Fisheries Management, 30(6):1376–1384, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roni:2018:RTI

[RABK18] Philip Roni, Paul J. Anders, Timothy J. Beechie, and David J. Kaplowe. Review of tools for identifying, planning, and implementing habitat restoration for Pacific salmon and steelhead. North American Journal of Fisheries Management, 38(2):355– 376, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ranney:2018:QRE

[Ran18] Steven H. Ranney. Quantile regression estimates of body weight at length in walleye. North American Journal of Fisheries Management, 38(3):769–776, June 2018. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roni:2018:RAM

[RÅW18] Philip Roni, Ulrika Åberg, and Christine Weber. A review of approaches for monitoring the effectiveness of regional river habitat restoration programs. North American Journal of Fisheries Management, 38(5):1170–1186, October 2018. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Recsetar:2015:ETC

[RB15] Matthew S. Recsetar and Scott A. Bonar. Effectiveness of two commercial rotenone formulations in the eradication of virile crayfish Orconectes virilis. North American Journal of Fisheries Management, 35(3):616–620, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Radabaugh:2010:CSM

[RBB10a] Nicholas B. Radabaugh, William F. Bauer, and Michael L. Brown. A comparison of seasonal movement patterns of yellow perch in simple and complex lake basins. North American Journal of Fisheries Management, 30(1):179–190, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roper:2010:CPC

[RBB^{+}10b] Brett B. Roper, John M. Buffington, Stephen Bennett, Steven H. Lanigan, Eric Archer, Scott T. Downie, John Faustini, Tracy W. Hillman, Shannon Hubler, Kim Jones, Chris Jordan, Philip R. Kaufmann, Glenn Merritt, Chris Moyer, and Allen Pleus. A comparison of the performance and compatibility of protocols used by seven monitoring groups to measure stream habitat in the Pacific Northwest. North American Journal of Fisheries Management, 30(2):565–587, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rosenfeld:2016:DBB

[RBP16] Jordan Rosenfeld, Hal Beecher, and Ron Ptolemy. Developing bioenergetic-based habitat suitability curves for instream flow models. North American Journal of Fisheries Management, 36(5):1205–1219, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Radigan:2018:AUO

[RCGF18] William J. Radigan, Andrew K. Carlson, Brian D. S. Graeb, and Mark J. Fincel. Assessing the utility of otolith chemistry for management of six freshwater fishes from a river–reservoir system. North American Journal of Fisheries Management, 38 (2):316–326, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reeves:2018:SFP

[RCMB18] David B. Reeves, Edward J. Chesney, Ryan T. Munnelly, and Donald M. Baltz. Sheepshead foraging patterns at oil and gas platforms in the Northern Gulf of Mexico. North American Journal of Fisheries Management, 38(6):1258–1274, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rosenberger:2013:AEP

[RCP^{+}13] Stuart J. Rosenberger, William P. Connor, Christopher A. Peery, Deborah J. Milks, Mark L. Schuck, Jay A. Hesse, and Steven G. Smith. Acclimation enhances postrelease performance of hatchery fall Chinook salmon subyearlings while reducing the potential for interaction with natural fish. North American Journal of Fisheries Management, 33(3):519–528, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roell:2010:ECR

[RD10] Michael J. Roell and Robert J. DiStefano. Effects of a conservative rock bass length limit on angler participation, sport fish populations, and crayfish prey in a Missouri Ozark stream. North American Journal of Fisheries Management, 30(2):552– 564, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reynolds:2015:CES

[Rey15] James B. Reynolds. Comment: Experimental studies of electroshock effects on fish require in-water measurements and fish threshold observations to achieve electrofishing context. North American Journal of Fisheries Management, 35(2):205– 206, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roberts:2017:NTI

[RFHP17] James J. Roberts, Kurt D. Fausch, Mevin B. Hooten, and Douglas P. Peterson. Nonnative trout invasions combined with climate change threaten persistence of isolated cutthroat trout populations in the Southern Rocky Mountains. North American Journal of Fisheries Management, 37(2):314–325, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richer:2017:IGM

[RFKB17] Eric E. Richer, Eric R. Fetherman, Matt C. Kondratieff, and Taylor A. Barnes. Incorporating GPS and mobile radio frequency identification to detect PIT-Tagged fish and evaluate habitat utilization in streams. North American Journal of Fisheries Management, 37(6):1249–1264, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rhoten:2014:EAT

[RFLP14] Jason C. Rhoten, Shannon J. Fisher, Dale E. Logsdon, and Bruce J. Pittman. Effects of age at treatment on oxytetracycline mark formation in larval northern pike. North American Journal of Fisheries Management, 34(4):735–740, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rohde:2014:FAP

[RFQ14] Jessica Rohde, Kurt L. Fresh, and Thomas P. Quinn. Factors affecting partial migration in Puget Sound Coho salmon. North American Journal of Fisheries Management, 34(3):559– 570, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ranney:2010:ALR

[RFW^{+}10] Steven H. Ranney, Mark J. Fincel, Melissa R. Wuellner, Justin A. VanDeHey, and Michael L. Brown. Assessing lengthrelated bias and the need for data standardization in the development of standard weight equations. North American Journal of Fisheries Management, 30(3):655–664, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ranney:2011:ALR

[RFW^{+}11] Steven H. Ranney, Mark J. Fincel, Melissa R. Wuellner, Justin A. VanDeHey, and Michael L. Brown. Assessing lengthrelated biases in standard weight equations: Response to comment. North American Journal of Fisheries Management, 31 (4):661–665, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Raborn:2012:ETE

[RGC^{+}12] Scott W. Raborn, Benny J. Gallaway, John G. Cole, William J. Gazey, and Kate I. Andrews. Effects of turtle excluder devices (TEDs) on the bycatch of three small coastal sharks in the Gulf of Mexico penaeid shrimp fishery. North American Journal of Fisheries Management, 32(2):333–345, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Railsback:2013:CDR

[RGH^{+}13] Steven F. Railsback, Mark Gard, Bret C. Harvey, Jason L. White, and Julie K. H. Zimmerman. Contrast of degraded and restored stream habitat using an individual-based salmon model. North American Journal of Fisheries Management, 33 (2):384–399, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Romero:2010:WCF

[RGOR10] M. Alejandra Romero, Raúl A. González, and Matías Ocampo-Reinaldo. When conventional fisheries management fails to reduce the catch and discard of juvenile fish: a case study of the Argentine hake trawl fishery in San Matías Gulf. North American Journal of Fisheries Management, 30(3):702– 712, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Raabe:2014:ADM

[RH14] Joshua K. Raabe and Joseph E. Hightower. Assessing distribution of migratory fishes and connectivity following complete and partial dam removals in a North Carolina River. North American Journal of Fisheries Management, 34(5):955–969, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rudershausen:2016:RPV

[RH16] P. J. Rudershausen and Joseph E. Hightower. Retention probability varies with cull ring size in traps fished for blue crab. North American Journal of Fisheries Management, 36(1):122– 130, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rapp:2014:CAE

[RHC^{+}14] Tobias Rapp, Jan Hallermann, Steven J. Cooke, Stefan K. Hetz, Sven Wuertz, and Robert Arlinghaus. Consequences of air exposure on the physiology and behavior of caught-and-released common carp in the laboratory and under natural conditions. North American Journal of Fisheries Management, 34(2):232–246, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rook:2012:SSC

[RHG12] Benjamin J. Rook, Michael J. Hansen, and Owen T. Gorman. The spatial scale for cisco recruitment dynamics in Lake Superior during 1978–2007. North American Journal of Fisheries Management, 32(3):499–514, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rook:2013:BAF

[RHG13] Benjamin J. Rook, Michael J. Hansen, and Owen T. Gorman. Biotic and abiotic factors influencing cisco recruitment dynamics in Lake Superior during 1978–2007. North American Journal of Fisheries Management, 33(6):1243–1257, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ruetz:2015:RMR

[RHMH15] Carl R. Ruetz III, Brandon S. Harris, James N. McNair, and Jared J. Homola. Removal and mark–recapture methods for estimating abundance: Empirical and simulation results for mottled sculpin in streams. North American Journal of Fisheries Management, 35(1):62–74, February 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rugg:2014:VAF

[RHPH14] M. L. Rugg, M. J. Hamel, M. A. Pegg, and J. J. Hammen. Validation of annuli formation in pectoral fin rays from shovelnose sturgeon in the Lower Platte River, Nebraska. North American Journal of Fisheries Management, 34(5):1028–1032, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ross:2019:ERT

[RIMR19] Jared A. Ross, Dana M. Infante, Douglas J. Martin, and Mark Rey. Effects of riparian timber harvest on Southeast Alaska stream habitat after 30–40 years: Insights for management. North American Journal of Fisheries Management, 39(2):328– 342, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roman:2011:ARF

[RJC11] Sally Roman, Nikki Jacobson, and Steven X. Cadrin. Assessing the reliability of Fisher self-sampling programs. North American Journal of Fisheries Management, 31(1):165–175, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rotherham:2013:EGM

[RJMG13] Douglas Rotherham, Daniel D. Johnson, William G. Macbeth, and Charles A. Gray. Escape gaps as a management strategy for reducing bycatch in net-covered traps for the giant mud crab Scylla serrata. North American Journal of Fisheries Management, 33(2):307–317, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Romine:2015:RBC

[RJP^{+}15] Jason G. Romine, Nathan R. Jensen, Michael J. Parsley, Robert F. Gaugush, Todd J. Severson, Tyson W. Hatton, Ryan F. Adams, and Mark P. Gaikowski. Response of bighead carp and silver carp to repeated water gun operation in an enclosed shallow pond. North American Journal of Fisheries Management, 35(3):440–453, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rindone:2015:DAR

[RKB15] R. Ryan Rindone, G. Todd Kellison, and Stephen A. Bortone. Data availability for red snapper in Gulf of Mexico and Southeastern U.S. Atlantic Ocean waters. North American Journal of Fisheries Management, 35(2):191–204, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robert:2012:IHS

[RLC12] Justin J. Robert, Sarah M. Larocque, and Steven J. Cooke. Influence of hook size and style on short-term survival of deeply hooked bluegills. North American Journal of Fisheries Management, 32(5):869–874, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rose:2012:ECF

[RM12] Brien P. Rose and Matthew G. Mesa. Effectiveness of common fish screen materials to protect lamprey ammocoetes. North American Journal of Fisheries Management, 32(3):597–603, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Romer:2014:ALH

[RM14] Jeremy D. Romer and Fred R. Monzyk. Adfluvial life history in spring Chinook salmon from Quartzville Creek, Oregon. North American Journal of Fisheries Management, 34 (5):885–891, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ross:2016:CET

[RMTW16] J. E. Ross, C. M. Mayer, J. T. Tyson, and E. J. Weimer. Comparison of electrofishing techniques and effort allocation across diel time periods, seasons, sites, and habitat in the Ohio coastal waters of Western Lake Erie. North American Journal of Fisheries Management, 36(1):85–95, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roni:2010:ECC

[RPBM10] Philip Roni, George Pess, Tim Beechie, and Sarah Morley. Estimating changes in Coho salmon and steelhead abundance from watershed restoration: How much restoration is needed to measurably increase smolt production? North American Journal of Fisheries Management, 30(6):1469–1484, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robinson:2019:EDS

[RPC^{+}19] Anthony T. Robinson, Yvette M. Paroz, Matthew J. Clement, Thomas W. Franklin, Joseph C. Dysthe, Michael K. Young, Kevin S. McKelvey, and Kellie J. Carim. Environmental DNA sampling of small-bodied minnows: Performance relative to location, species, and traditional sampling. North American Journal of Fisheries Management, 39(5):1073–1085, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roop:2018:AAE

[RPJ18] Hunter J. Roop, Neelam C. Poudyal, and Cecil A. Jennings. Assessing angler effort, catch, and harvest on a spatially complex, multi-lake fishery in Middle Georgia. North American Journal of Fisheries Management, 38(4):833–841, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rudershausen:2019:EDM

[RPMB19] Paul J. Rudershausen, Stephen J. Poland, Wessley Merten, and Jeffrey A. Buckel. Estimating discard mortality for dolphinfish in a recreational hook-and-line fishery. North American Journal of Fisheries Management, 39(6):1143–1154, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ramirez-Rodriguez:2017:PAC

[RR17] Mauricio Ramírez-Rodríguez. A profitability analysis of catch quotas for the Pacific hake fishery in the Gulf of California. North American Journal of Fisheries Management, 37 (1):23–29, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Riley:2011:ITC

[RRH^{+}11] Stephen C. Riley, Jacques Rinchard, Dale C. Honeyfield, Allison N. Evans, and Linda Begnoche. Increasing thiamine concentrations in lake trout eggs from Lakes Huron and Michigan coincide with low alewife abundance. North American Journal of Fisheries Management, 31(6):1052–1064, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reeves:2011:RHM

[RS11] Keith A. Reeves and David F. Staples. Relative hooking mortality among walleyes caught on barbed and barbless octopus hooks and barbed jigs. North American Journal of Fisheries Management, 31(1):32–40, February 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reed:2017:ETD

[RS17] Jeffrey R. Reed and David F. Staples. Evaluation of two different stocking rates of small walleye fingerlings in Minnesota lakes. North American Journal of Fisheries Management, 37 (6):1243–1248, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richins:2018:SOF

[RS18] Shelby M. Richins and John R. Skalski. Steelhead overshoot and fallback rates in the Columbia–Snake River Basin and the influence of hatchery and hydrosystem operations. North American Journal of Fisheries Management, 38(5):1122–1137, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ratliff:2015:RYB

[RSHS15] Don Ratliff, Robert Spateholts, Megan Hill, and Eric Schulz. Recruitment of young bull trout into the Metolius River and Lake Billy Chinook, Oregon. North American Journal of Fisheries Management, 35(6):1077–1089, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Richter:2016:VSS

[RSI16] Jacob T. Richter, Brian L. Sloss, and Daniel A. Isermann. Validation of a side-scan sonar method for quantifying walleye spawning habitat availability in the littoral zone of Northern Wisconsin lakes. North American Journal of Fisheries Management, 36(4):942–950, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rach:2010:EWH

[RSLG10] Jeff J. Rach, Greg G. Sass, James A. Luoma, and Mark P. Gaikowski. Effects of water hardness on size and hatching success of silver carp eggs. North American Journal of Fisheries Management, 30(1):230–237, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roth:2019:EAE

[RSQ^{+}19] Curtis J. Roth, Daniel J. Schill, Michael C. Quist, Brett High, Matthew R. Campbell, and Ninh V. Vu. Effects of air exposure during simulated catch-and-release angling on survival and fitness of Yellowstone cutthroat trout. North American Journal of Fisheries Management, 39(1):191–204, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Roth:2018:EAE

[RSQH18] Curtis J. Roth, Daniel J. Schill, Michael C. Quist, and Brett High. Effects of air exposure in summer on the survival of caught-and-released salmonids. North American Journal of Fisheries Management, 38(4):886–895, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Reeves:2016:RIO

[RTJJ16] David B. Reeves, William B. Tate, Howard L. Jelks, and Frank Jordan. Response of imperiled okaloosa darters to stream restoration. North American Journal of Fisheries Management, 36(6):1375–1385, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Riley:2010:DTN

[RTMJ10] Jacob W. Riley, Neil F. Thompson, J. Ellen Marsden, and John Janssen. Development of two new sampling techniques for assessing lake trout reproduction in deep water. North American Journal of Fisheries Management, 30(6):1571–1581, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rennie:2012:ALR

[RV12] Michael D. Rennie and Richard Verdon. Assessing lengthrelated biases in standard weight equations: Response to comment. North American Journal of Fisheries Management, 32 (5):953–955, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robinson:2011:IBD

[RW11] Anthony T. Robinson and David L. Ward. Interactions between desert pupfish and Gila topminnow can affect reintroduction success. North American Journal of Fisheries Management, 31(6):1093–1099, December 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Robinson:2016:TBA

[RWAJ16] Jason M. Robinson, Michael J. Wilberg, Jean V. Adams, and Michael L. Jones. Tradeoff between assessment and control of aquatic invasive species: a case study of sea lamprey management in the St. Marys River. North American Journal of Fisheries Management, 36(1):11–20, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Radomski:2011:REP

[RWCP11] Paul Radomski, Kevin Woizeschke, Kristin Carlson, and Donna Perleberg. Reproducibility of emergent plant mapping on lakes. North American Journal of Fisheries Management, 31(1):144–150, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rude:2011:LTP

[RWPH11] Neil P. Rude, Gregory W. Whitledge, Quinton E. Phelps, and Shawn Hirst. Long-term PIT and T-bar anchor tag retention rates in adult muskellunge. North American Journal of Fisheries Management, 31(3):515–519, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rizzo:2017:PLP

[RWT17] Austin A. Rizzo, Stuart A. Welsh, and Patricia A. Thompson. A paired-laser photogrammetric method for in situ length measurement of benthic fishes. North American Journal of Fisheries Management, 37(1):16–22, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Rypel:2015:ERD

[Ryp15] Andrew L. Rypel. Effects of a reduced daily bag limit on bluegill size structure in Wisconsin lakes. North American Journal of Fisheries Management, 35(2):388–397, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Risley:2010:AEC

[RZ10] Casey A. L. Risley and Joseph Zydlewski. Assessing the effects of catch-and-release regulations on a brook trout population using an age-structured model. North American Journal of Fisheries Management, 30(6):1434–1444, December 2010.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stout:2016:ECA

[SAF16] J. Benjamin Stout, Brian W. Avila, and Eric R. Fetherman. Efficacy of commercially available quaternary ammonium compounds for controlling New Zealand mudsnails Potamopyrgus antipodarum. North American Journal of Fisheries Management, 36(2):277–284, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2016:CES

[Sam16] Steven M. Sammons. Catch and exploitation of shoal bass in the Flint River, Georgia, USA: Implications for harvest restrictions. North American Journal of Fisheries Management, 36 (3):606–620, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2019:ERA

[Sam19] Steven M. Sammons. The effects of regulated area location and size on the effectiveness of a catch-and-release season for a riverine black bass. North American Journal of Fisheries Management, 39(5):1103–1115, October 2019. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2017:EAT

[SAYJ17] Andrew A. Schultz, Virginia B. Afentoulis, Curtis J. Yip, and Michele N. Johnson. Efficacy of an acoustic tag with predation detection technology. North American Journal of Fisheries Management, 37(3):574–581, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schoenebeck:2011:GYS

[SB11] Casey W. Schoenebeck and Michael L. Brown. Gender- and year-specific mortality of yellow perch with evidence of compensatory mortality. North American Journal of Fisheries Management, 31(3):474–482, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sheppard:2015:USC

[SB15] John J. Sheppard and Michael S. Bednarski. Utility of singlechannel electronic resistivity counters for monitoring river herring populations. North American Journal of Fisheries Management, 35(6):1144–1151, December 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shoup:2018:EVD

[SB18] Daniel E. Shoup and Chance R. Broderius. Effects of vegetation density on the ontogeny to piscivory of juvenile largemouth bass. North American Journal of Fisheries Management, 38(3):630–638, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2019:DIA

[SBB19] Kevin M. Sullivan, Michael M. Bailey, and David L. Berlinsky. Digital image analysis as a technique for alewife fecundity estimation in a New Hampshire river. North American Journal of Fisheries Management, 39(2):353–361, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seekell:2011:LTC

[SBC^{+}11] David A. Seekell, Chase J. Brosseau, Timothy J. Cline, Raymond J. Winchcombe, and Lee J. Zinn. Long-term changes in recreational catch inequality in a trout stream. North American Journal of Fisheries Management, 31(6):1100–1105, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schaffler:2011:DAL

[SBD^{+}11] Jason J. Schaffler, Susan G. Barco, Linda R. D’Eri, W. Mark Swingle, and Cynthia M. Jones. Do alternative leaders affect fish catch in pound nets at the mouth of Chesapeake Bay? North American Journal of Fisheries Management, 31 (1):88–95, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schlechte:2016:SSM

[SBDB16] J. Warren Schlechte, Kristopher A. Bodine, Daniel J. Daugherty, and Greg R. Binion. Size selectivity of multifilament gill nets for sampling alligator gar: Modeling the effects on population metrics. North American Journal of Fisheries Management, 36(3):630–638, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stokesbury:2019:CDL

[SBG^{+}19] Kevin D. E. Stokesbury, N. David Bethoney, Daniel Georgianna, Susan Inglis, and Emily Keiley. Convergence of a disease and litigation leading to increased scallop discard mortality and economic loss in the Georges Bank fishery. North American Journal of Fisheries Management, 39(2):299–306, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sitar:2017:RPM

[SBHJ17] Shawn P. Sitar, Travis O. Brenden, Ji X. He, and James E. Johnson. Recreational postrelease mortality of lake trout in Lakes Superior and Huron. North American Journal of Fisheries Management, 37(4):789–808, August 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sloat:2011:EHS

[SBL11] Matthew R. Sloat, Peter F. Baker, and Franklin K. Ligon. Estimating habitat-specific abundances of PIT-Tagged juvenile salmonids using mobile antennas: a comparison with standard electrofishing techniques in a small stream. North American Journal of Fisheries Management, 31(5):986–993, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stamplecoskie:2012:RMS

[SBL^{+}12] Keith M. Stamplecoskie, Thomas R. Binder, Nicola Lower, Karl Cottenie, Robert L. McLaughlin, and D. Gordon McDonald. Response of migratory sea lampreys to artificial lighting in portable traps. North American Journal of Fisheries Management, 32(3):563–572, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schall:2018:QTT

[SBL^{+}18a] Megan K. Schall, Vicki S. Blazer, Robert M. Lorantas, Geoffrey D. Smith, John E. Mullican, Brandon J. Keplinger, and Tyler Wagner. Quantifying temporal trends in fisheries abundance using Bayesian dynamic linear models: a case study of riverine smallmouth bass populations. North American Journal of Fisheries Management, 38(2):493–501, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Struthers:2018:STP

[SBL^{+}18b] Daniel P. Struthers, Shannon D. Bower, Robert J. Lennox, Christine E. Gilroy, Elizabeth C. MacDonald, Steven J. Cooke, and Matthew K. Litvak. Short-term physiological disruption and reflex impairment in shortnose sturgeon exposed to catch-and-release angling. North American Journal of Fisheries Management, 38(5):1075–1084, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sakaris:2014:VDR

[SBS14] Peter C. Sakaris, David L. Buckmeier, and Nathan G. Smith. Validation of daily ring deposition in the otoliths of age-0 alligator gar. North American Journal of Fisheries Management, 34(6):1140–1144, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steinhart:2017:FPT

[SBTW17] Geoffrey B. Steinhart, Sierra Bode, Jennifer Thieme, and Christopher J. Winslow. Fish passage through three types of structures in diked coastal Lake Erie wetlands. North American Journal of Fisheries Management, 37(3):497–509, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sedinger:2012:EPS

[SBVB12] James S. Sedinger, Erik J. Blomberg, Amanda W. VanDellen, and Stephanie Byers. Environmental and population strain effects on survival of Lahontan cutthroat trout in Walker Lake, Nevada: a Bayesian approach. North American Journal of Fisheries Management, 32(3):515–522, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2016:EFM

[SBW^{+}16] Bradley J. Smith, Brian G. Blackwell, Melissa R. Wuellner, Brian D. S. Graeb, and David W. Willis. Escapement of fishes from modified fyke nets with differing throat configurations. North American Journal of Fisheries Management, 36(1):96–103, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schall:2017:EGP

[SBW^{+}17a] Megan K. Schall, Meredith L. Bartron, Timothy Wertz, Jonathan M. Niles, Cassidy H. Shaw, Vicki S. Blazer, and Tyler Wagner. Evaluation of genetic population structure of smallmouth bass in the Susquehanna River Basin, Pennsylvania. North American Journal of Fisheries Management, 37(4): 850–861, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2017:CSF

[SBW^{+}17b] Bradley J. Smith, Brian G. Blackwell, Melissa R. Wuellner, Brian D. S. Graeb, and David W. Willis. Contact selectivity for four fish species sampled with North American standard gill nets. North American Journal of Fisheries Management, 37(1):149–161, February 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Strong:2014:MME

[SC14] James W. Strong and Keith R. Criddle. A market model of Eastern Bering Sea Alaska pollock: Sensitivity to fluctuations in catch and some consequences of the American Fisheries Act. North American Journal of Fisheries Management, 34 (6):1078–1094, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sackett:2017:SHV

[SC17] Dana K. Sackett and Matthew Catalano. Spatial heterogeneity, variable rewards, tag loss, and tagging mortality affect the performance of mark–recapture designs to estimate exploitation: an example using red snapper in the Northern Gulf of Mexico. North American Journal of Fisheries Management, 37(3):558–573, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stuart:2018:CGI

[SC18] Ivor G. Stuart and Anthony J. Conallin. Control of globally invasive common carp: An 11-year commercial trial of the Williams’ Cage. North American Journal of Fisheries Management, 38(5):1160–1169, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steiner:2011:BBS

[SCA11] Erin M. Steiner, Keith R. Criddle, and Milo D. Adkison. Balancing biological sustainability with the economic needs of Alaska’s sockeye salmon fisheries. North American Journal of Fisheries Management, 31(3):431–444, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stout:2019:WAA

[SCB^{+}19] J. Benjamin Stout, Mary M. Conner, Phaedra Budy, Peter D. Mackinnon, and Mark C. McKinstry. We ain’t afraid of no ghosts: Tracking habitat interactions and movement dynamics of ghost tags under differing flow conditions in a sand-bed river. North American Journal of Fisheries Management, 39 (6):1337–1347, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smedley:2011:TSI

[SCG11a] Rosalyn A. Smedley, R. Allen Curry, and Michelle A. Gray. Testing the severity of ill effects model for predicting fish abundance and condition. North American Journal of Fisheries Management, 31(3):419–426, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sykes:2011:PEP

[SCG11b] Catherine L. Sykes, Colleen A. Caldwell, and William R. Gould. Physiological effects of potassium chloride, formalin, and handling stress on bonytail. North American Journal of Fisheries Management, 31(2):291–298, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schall:2017:UWF

[SCKZ17] Benjamin J. Schall, Timothy K. Cross, Eric Katzenmeyer, and Douglas L. Zentner. Use of wind fetch and shoreline relief to predict nearshore substrate composition in a North temperate lake. North American Journal of Fisheries Management, 37 (4):935–942, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2013:MPD

[SCMP13] Douglas W. Schultz, Andrew J. Carlson, Steve Mortensen, and Donald L. Pereira. Modeling population dynamics and fish consumption of a managed double-crested cormorant colony in Minnesota. North American Journal of Fisheries Management, 33(6):1283–1300, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stein:2014:CAB

[SCPC14] Jeffrey A. Stein, Julie E. Claussen, David P. Philipp, and Steven J. Cooke. Comment: Not all biases are created equal — a comment on the snorkel survey bias observed by hessenauer et al. (2014). North American Journal of Fisheries Management, 34(6):1204–1206, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sauls:2017:SME

[SCS17a] Beverly J. Sauls, Richard P. Cody, and Andrew J. Strelcheck. Survey methods for estimating red snapper landings in a higheffort recreational fishery managed with a small annual catch limit. North American Journal of Fisheries Management, 37 (2):302–313, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scharf:2017:FSS

[SCS17b] Frederick S. Scharf, J. Kevin Craig, and William E. Smith. Fine-scale spatial and temporal variation in fishing mortality of southern flounder: Management implications for a dynamic estuarine fishery. North American Journal of Fisheries Management, 37(5):1067–1074, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seekell:2013:CMR

[SCW13] David A. Seekell, Timothy J. Cline, and Raymond J. Winchcombe. Can management reduce harvest inequality in recreational fisheries? North American Journal of Fisheries Management, 33(1):148–152, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stearns:2017:DCT

[SCW^{+}17] George Stearns, Robert Conrad, David Winfrey, Nancy Shippentower-Games, and Deanna Finley. Dungeness crab trap catch efficiency related to escape ring location and size. North American Journal of Fisheries Management, 37(5): 1039–1044, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2017:IAV

[SCWP17] Christopher J. Sullivan, Carlos A. Camacho, Michael J. Weber, and Clay L. Pierce. Intra-annual variability of silver carp populations in the des Moines River, USA. North American Journal of Fisheries Management, 37(4):836–849, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stevens:2011:SUS

[SD11] Bryan S. Stevens and Joseph M. DuPont. Summer use of side-channel thermal refugia by salmonids in the North Fork Coeur d’Alene River, Idaho. North American Journal of Fisheries Management, 31(4):683–692, August 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scheibel:2016:INP

[SDDC16] Natalie C. Scheibel, Daniel J. Dembkowski, Jacob L. Davis, and Steven R. Chipps. Impacts of northern pike on stocked rainbow trout in Pactola Reservoir, South Dakota. North American Journal of Fisheries Management, 36(2):230–240, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sovic:2012:UFR

[SDF12] Michael G. Sovic, Jonathan C. Denlinger, and Paul A. Fuerst. The use of fluorescent randomly amplified polymorphic DNA markers to identify hybrids: a case study evaluating the origins of saugeye following the cessation of stocking in an Ohio reservoir. North American Journal of Fisheries Management, 32(4):671–678, August 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stich:2013:GPS

[SDF^{+}13] Daniel S. Stich, Victor Dicenzo, Emmanuel A. Frimpong, Yan Jiao, and Brian R. Murphy. Growth and population size of grass carp incrementally stocked for hydrilla control. North American Journal of Fisheries Management, 33(1):14– 25, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schuett:2016:EBM

[SDKS16] Michael A. Schuett, Chen Ding, Gerard Kyle, and J. Dale Shively. Examining the behavior, management preferences, and sociodemographics of artificial reef users in the Gulf of Mexico offshore from Texas. North American Journal of Fisheries Management, 36(2):321–328, April 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stewart:2015:AGA

[SDL^{+}15] Nathan D. Stewart, Michael J. Dadswell, Paul Leblanc, Rod G. Bradford, Cornel Ceapa, and Michael J. W. Stokesbury. Age and growth of Atlantic sturgeon from the Saint John River, New Brunswick, Canada. North American Journal of Fisheries Management, 35(2):364–371, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2017:GMA

[SDPB17] Kevin M. Sullivan, Kristin L. Duclos, Steven J. Parker, and David L. Berlinsky. Growth and maturation of Acadian redfish in the Gulf of Maine. North American Journal of Fisheries Management, 37(1):41–49, February 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stapanian:2012:PSA

[SE12] Martin A. Stapanian and William H. Edwards. A prototype splitter apparatus for dividing large catches of small fish. North American Journal of Fisheries Management, 32 (6):1033–1038, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmitt:2019:FED

[SEB^{+}19] Joseph D. Schmitt, Jason A. Emmel, Aaron J. Bunch, Corbin D. Hilling, and Donald J. Orth. Feeding ecology and distribution of an invasive apex predator: Flathead catfish in subestuaries of the Chesapeake Bay, Virginia. North American Journal of Fisheries Management, 39(2):390–402, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seekell:2011:RFA

[See11] David A. Seekell. Recreational freshwater angler success is not significantly different from a random catch model. North American Journal of Fisheries Management, 31(2):203–208, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sweka:2014:EPR

[SEM14] John A. Sweka, Sheila Eyler, and Michael J. Millard. An eggper-recruit model to evaluate the effects of upstream transport and downstream passage mortality of American eel in the Susquehanna River. North American Journal of Fisheries Management, 34(4):764–773, August 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steffensen:2011:ULP

[SEP11] Kirk D. Steffensen, Brandon L. Eder, and Mark A. Pegg. Using lindgren–Pitman hook timers to understand the dynamics of trotline catches in a large river. North American Journal of Fisheries Management, 31(5):980–985, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skalski:2014:ASC

[SEP^{+}14] John R. Skalski, M. Brad Eppard, Gene R. Ploskey, Mark A. Weiland, Thomas J. Carlson, and Richard L. Townsend. Assessment of subyearling Chinook salmon survival through the Federal hydropower projects in the main-stem Columbia River. North American Journal of Fisheries Management, 34 (4):741–752, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2019:VBC

[SEP^{+}19] Christopher J. Sullivan, Holly S. Embke, K. Martin Perales, Stephen R. Carpenter, M. Jake Vander Zanden, and Daniel A. Isermann. Variation in bluegill catch rates and total length distributions among four sampling gears used in two Wisconsin lakes dominated by small fish. North American Journal of Fisheries Management, 39(4):714–724, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schroeder:2014:FNP

[SF14] Susan A. Schroeder and David C. Fulton. Fishing for northern pike in Minnesota: Comparing anglers and dark house spearers. North American Journal of Fisheries Management, 34 (3):678–691, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Szymkowiak:2016:UDH

[SF16] Marysia Szymkowiak and Ron Felthoven. Understanding the determinants of hired skipper use in the Alaska halibut individual fishing quota fishery. North American Journal of Fisheries Management, 36(5):1139–1148, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schueller:2012:SEW

[SFH12] Amy M. Schueller, Andrew H. Fayram, and Michael J. Hansen. Simulated equilibrium walleye population density under static and dynamic recreational angling effort. North American Journal of Fisheries Management, 32(5):894–904, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scyphers:2013:VRF

[SFH^{+}13] Steven B. Scyphers, F. Joel Fodrie, Frank J. Hernandez, Jr., Sean P. Powers, and Robert L. Shipp. Venting and reef fish survival: Perceptions and participation rates among recreational anglers in the Northern Gulf of Mexico. North American Journal of Fisheries Management, 33(6):1071–1078, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Saunders:2011:AES

[SFW11] W. Carl Saunders, Kurt D. Fausch, and Gary C. White. Accurate estimation of salmonid abundance in small streams using nighttime removal electrofishing: an evaluation using marked fish. North American Journal of Fisheries Management, 31 (2):403–415, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sammons:2013:SHU

[SG13] Steven M. Sammons and David C. Glover. Summer habitat use of large adult striped bass and habitat availability in Lake Martin, Alabama. North American Journal of Fisheries Management, 33(4):762–772, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Syslo:2013:CHS

[SGC13] John M. Syslo, Christopher S. Guy, and Benjamin S. Cox. Comparison of harvest scenarios for the cost-effective suppression of lake trout in Swan Lake, Montana. North American Journal of Fisheries Management, 33(6):1079–1090, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sakaris:2019:OTD

[SGCB19] Peter C. Sakaris, John Galvez, Thomas P. Callier IV, and Allan Brown. Ontogenetic and temporal diet shifts of the invasive Asian swamp eel in South Florida. North American Journal of Fisheries Management, 39(6):1288–1300, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Somerton:2013:GPM

[SGF^{+}13] David Somerton, Scott Goodman, Robert Foy, Lou Rugolo, and Laura Slater. Growth per molt of snow crab in the Eastern Bering Sea. North American Journal of Fisheries Management, 33(1):140–147, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schueller:2011:IAR

[SH11] Amy M. Schueller and Daniel B. Hayes. Inbreeding and allele retention for lake sturgeon populations under different supplementation strategies. North American Journal of Fisheries Management, 31(6):1133–1145, December 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steffensen:2018:FRR

[SH18] Kirk D. Steffensen and Martin J. Hamel. Fin ray removal may be deleterious to pallid sturgeon. North American Journal of Fisheries Management, 38(2):439–445, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schultz:2017:BME

[SHK^{+}17] L. D. Schultz, Michael P. Heck, Brandon M. Kowalski, Collin A. Eagles-Smith, K. Coates, and Jason B. Dunham. Bioenergetics models to estimate numbers of larval lampreys consumed by smallmouth bass in Elk Creek, Oregon. North American Journal of Fisheries Management, 37(4):714–723, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shank:2016:ESU

[SHL16] Matthew K. Shank, Aaron M. Henning, and Andrew S. Leakey. Examination of a single-unit, multiple-pass electrofishing protocol to reliably estimate fish assemblage composition in wadeable streams of the mid-Atlantic region of the USA. North American Journal of Fisheries Management, 36 (3):497–505, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smolowitz:2010:UST

[SHM^{+}10] Ronald Smolowitz, Heather Haas, Henry O. Milliken, Matthew Weeks, and Eric Matzen. Using sea turtle carcasses to assess the conservation potential of a turtle excluder dredge. North American Journal of Fisheries Management, 30(4):993–1000, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sol:2019:CFA

[SHMJ19a] Sean Y. Sol, Amanda C. Hanson, Keith Marcoe, and Lyndal L. Johnson. Changes in fish assemblages at the Mirror Lake complex in the Lower Columbia River before and after a culvert modification. North American Journal of Fisheries Management, 39(6):1348–1359, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sol:2019:JSA

[SHMJ19b] Sean Y. Sol, Amanda C. Hanson, Keith Marcoe, and Lyndal L. Johnson. Juvenile salmonid assemblages at the mirror lake complex in the Lower Columbia River before and after a culvert modification. North American Journal of Fisheries Management, 39(1):91–103, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sabal:2016:HAN

[SHMS16] Megan Sabal, Sean Hayes, Joseph Merz, and Jose Setka. Habitat alterations and a nonnative predator, the striped bass, increase native Chinook salmon mortality in the Central Valley, California. North American Journal of Fisheries Management, 36(2):309–320, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2015:DLW

[SHS15] Cody Sullivan, Caleb Hasler, and Cory D. Suski. Do livewell temperatures differ from ambient water during black bass tournaments? North American Journal of Fisheries Management, 35(5):1064–1069, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sanchez-Hernandez:2016:IML

[SHSCA16] Javier Sánchez-Hernández, Stephanie L. Shaw, Fernando Cobo, and Mike S. Allen. Influence of a minimum-length limit regulation on wild brown trout: an example of recruitment and growth overfishing. North American Journal of Fisheries Management, 36(5):1024–1035, October 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sindt:2018:EUH

[Sin18] Anthony R. Sindt. Evaluation of unbaited hoop nets for simultaneously assessing channel catfish and flathead catfish populations in the Minnesota River. North American Journal of Fisheries Management, 38(3):538–548, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stich:2015:LDR

[SJM15] Daniel S. Stich, Yan Jiao, and Brian R. Murphy. Life, death, and resurrection: Accounting for state uncertainty in survival estimation from tagged grass carp. North American Journal of Fisheries Management, 35(2):321–330, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schofield:2018:LLT

[SK18a] Pamela J. Schofield and Jeffrey L. Kline. Lower lethal temperatures for nonnative freshwater fishes in Everglades National Park, Florida. North American Journal of Fisheries Management, 38(3):706–717, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stahr:2018:EEU

[SK18b] Kristopher J. Stahr and Rebecca L. Knudsen. Evaluating the efficacy of using time-lapse cameras to assess angling use: An example from a high-use metropolitan reservoir in Arizona. North American Journal of Fisheries Management, 38(2):327– 333, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snobl:2015:DTE

[SKBB15] Zachary R. Snobl, Ryan P. Koenigs, Ronald M. Bruch, and Fredrick P. Binkowski. Do tags exceeding 2% of total body weight impair lake sturgeon movement? North American Journal of Fisheries Management, 35(5):880–884, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Siepker:2012:ENP

[SKBK12] Michael J. Siepker, David S. Knuth, Elizabeth L. Ball, and Jeffrey B. Koppelman. Evaluating a novel passive integrated transponder tag in largemouth bass. North American Journal of Fisheries Management, 32(3):528–532, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skov:2011:DGD

[SKBSA11] Christian Skov, Anders Koed, Lars Baastrup-Spohr, and Robert Arlinghaus. Dispersal, growth, and diet of stocked and wild northern pike fry in a shallow natural lake, with implications for the management of stocking programs. North American Journal of Fisheries Management, 31(6):1177–1186, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sell:2016:CCH

[SKH^{+}16] Matthew T. Sell, David C. Kazyak, Robert H. Hilderbrand, Alan A. Heft, and Ryan M. Cooper. A comparison of circle hook size on hooking success, deep hooking rate, and postrelease mortality of hatchery-reared rainbow trout. North American Journal of Fisheries Management, 36(2):254–258, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Switzer:2015:IES

[SKS^{+}15] Theodore S. Switzer, Sean F. Keenan, Phillip W. Stevens, Robert H. McMichael, Jr., and Timothy C. MacDonald. Incorporating ecology into survey design: Monitoring the recruitment of age-0 gags in the Eastern Gulf of Mexico. North American Journal of Fisheries Management, 35(6):1132–1143, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stewart:2012:PCC

[SL12] David R. Stewart and James M. Long. Precision of channel catfish catch estimates using hoop nets in larger Oklahoma reservoirs. North American Journal of Fisheries Management, 32(6):1108–1112, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stewart:2014:RRR

[SL14] Christine T. Stewart and Marvin M. F. Lutnesky. Retardation of reproduction in the red shiner due to electroshock. North American Journal of Fisheries Management, 34(3):463–470, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shoup:2015:ETP

[SL15a] Daniel E. Shoup and W. Drew Lane. Effects of turbidity on prey selection and foraging return of adult largemouth bass in reservoirs. North American Journal of Fisheries Management, 35(5):913–924, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stewart:2015:ESE

[SL15b] Christine T. Stewart and Marvin M. F. Lutnesky. Experimental studies of electroshock effects on fish require in-water measurements and fish threshold observations to achieve electrofishing context: Response to comment. North American Journal of Fisheries Management, 35(2):207–209, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seung:2017:MAE

[SL17a] Chang K. Seung and Daniel K. Lew. A multiregional approach for estimating the economic impact of harvest restrictions on saltwater sportfishing. North American Journal of Fisheries Management, 37(5):1112–1129, October 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2017:OMJ

[SL17b] Richard A. Snow and James M. Long. Otolith marking of juvenile shortnose gar by immersion in oxytetracycline. North American Journal of Fisheries Management, 37(4):724–728, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skalski:2010:PTA

[SLB10] John R. Skalski, Shiquan Liao, and Rebecca A. Buchanan. Performance of three alternative estimators of stream residence time based on live and dead counts of salmonids. North American Journal of Fisheries Management, 30(1):191–208, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Spangenberg:2015:SDG

[SLG^{+}15] Dina K. Spangenberg, Donald A. Larsen, Ryan Gerstenberger, Chris Brun, Deborah L. Harstad, Shelly Nance, Larissa Rohrbach, and Brian R. Beckman. Stock differences in growth, smolting, and early male maturation in hatchery spring Chinook salmon: a common-garden experiment. North American Journal of Fisheries Management, 35(6):1090–1100, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schoenebeck:2015:UEC

[SLK15] Casey W. Schoenebeck, Seth A. Lundgren, and Keith D. Koupal. Using electrofishing catch rates to estimate the density of largemouth bass. North American Journal of Fisheries Management, 35(2):331–336, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schill:2010:SRF

[SLMM10] Daniel J. Schill, George W. LaBar, Elizabeth R. J. M. Mamer, and Kevin A. Meyer. Sex ratio, fecundity, and models predicting length at sexual maturity of redband trout in Idaho desert streams. North American Journal of Fisheries Management, 30(5):1352–1363, October 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stanfield:2013:OEI

[SLP13] L. W. Stanfield, N. P. Lester, and I. C. Petreman. Optimal effort intensity in backpack electrofishing surveys. North American Journal of Fisheries Management, 33(2):277–286, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Strobel:2017:EUE

[SLP17] Burke Strobel, Matthew B. Laramie, and David S. Pilliod. Exploring the use of environmental DNA to determine the species of salmon redds. North American Journal of Fisheries Management, 37(5):943–950, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smucker:2010:CBI

[SLR10] Byran J. Smucker, Robert M. Lorantas, and James L. Rosenberger. Correcting bias introduced by aerial counts in angler effort estimation. North American Journal of Fisheries Management, 30(4):1051–1061, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmalz:2016:VTO

[SLR^{+}16] Patrick J. Schmalz, Mark Luehring, Joe Dan Rose, John M. Hoenig, and Melissa K. Treml. Visualizing trade-offs between yield and spawners per recruit as an aid to decision making. North American Journal of Fisheries Management, 36 (1):1–10, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stewart:2016:SME

[SLS16] David R. Stewart, James M. Long, and Daniel E. Shoup. Simulation modeling to explore the effects of length-based harvest regulations for Ictalurus fisheries. North American Journal of Fisheries Management, 36(5):1190–1204, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmidt:2016:PPT

[SLSS16] Thomas Schmidt, Christian Löb, Benjamin Schreiber, and Ralf Schulz. A pitfall with PIT tags: Reduced detection efficiency of half-duplex passive integrated transponders in groups of marked fish. North American Journal of Fisheries Management, 36(4):951–957, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steffe:2010:IVD

[SM10] Aldo S. Steffe and Jeffrey J. Murphy. The importance of validating the duration of fishing trips: Are harvest rates from on-site recreational fishing surveys unbiased? North American Journal of Fisheries Management, 30(2):506–510, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stockton:2013:DTW

[SM13] Kelly A. Stockton and Christine M. Moffitt. Disinfection of three wading boot surfaces infested with New Zealand mudsnails. North American Journal of Fisheries Management, 33 (3):529–538, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Szedlmayer:2014:IAC

[SM14] Stephen T. Szedlmayer and Peter A. Mudrak. Influence of age-1 conspecifics, sediment type, dissolved oxygen, and the Deepwater Horizon oil spill on recruitment of age-0 red snapper in the Northeast Gulf of Mexico during 2010 and 2011. North American Journal of Fisheries Management, 34(2):443– 452, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2019:ISR

[SM19] Charles G. Snow and Ryan D. Mann. Identification of summer-run Chinook salmon spawning areas in the main-stem Columbia River upstream of Wells Dam, Washington. North American Journal of Fisheries Management, 39(2):307–314, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stolarski:2014:LMC

[SMCS14] J. T. Stolarski, F. J. Margraf, J. G. Carlson, and T. M. Sutton. Lipid and moisture content modeling of amphidromous Dolly Varden using bioelectrical impedance analysis. North American Journal of Fisheries Management, 34(3):471–481, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schill:2017:SEY

[SMH17] Daniel J. Schill, Kevin A. Meyer, and Michael J. Hansen. Simulated effects of YY-Male stocking and manual suppression for eradicating nonnative brook trout populations. North American Journal of Fisheries Management, 37(5):1054–1066, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2013:EDC

[SMK13] Charles G. Snow, Andrew R. Murdoch, and Thomas H. Kahler. Ecological and demographic costs of releasing nonmigratory juvenile hatchery steelhead in the Methow River, Washington. North American Journal of Fisheries Management, 33(6):1100–1112, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schwabe:2014:AHI

[SMP^{+}14] Martin Schwabe, Thomas Meinelt, Thuy M. Phan, Steven J. Cooke, and Robert Arlinghaus. Absence of handling-induced Saprolegnia infection in juvenile rainbow trout with implications for catch-and-release angling. North American Journal of Fisheries Management, 34(6):1221–1226, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2013:DHA

[SMS13] Christopher L. Sullivan, Kevin A. Meyer, and Daniel J. Schill. Deep hooking and angling success when passively and actively fishing for stream-dwelling trout with baited J and circle hooks. North American Journal of Fisheries Management, 33 (1):1–6, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smolowitz:2012:DEA

[SMW12] Ronald Smolowitz, Henry O. Milliken, and Matthew Weeks. Design, evolution, and assessment of a sea turtle deflector dredge for the U.S. Northwest Atlantic sea scallop fishery: Impacts on fish bycatch. North American Journal of Fisheries Management, 32(1):65–76, February 2012. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seung:2016:NEI

[SMW16] Chang K. Seung, Ben Muse, and Edward C. Waters. Net economic impacts of recent Alaska salmon fishery failures and Federal relief. North American Journal of Fisheries Management, 36(2):351–362, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sato:2016:RRP

[SNT^{+}16] Taku Sato, Atsushi Nanami, Takayuki Takebe, Masato Kobayashi, Takashi Iwasaki, Rihito Shinoda, and Masahiko Koiso. Retention rates of plastic-infusion-process passive integrated transponder tags and external tags in the whitestreaked grouper Epinephelus ongus. North American Journal of Fisheries Management, 36(1):116–121, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shepard:2014:FIS

[SNTZ14] Bradley B. Shepard, Lee M. Nelson, Mark L. Taper, and Alexander V. Zale. Factors influencing successful eradication of nonnative brook trout from four small Rocky Mountain streams using electrofishing. North American Journal of Fisheries Management, 34(5):988–997, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Salierno:2018:OHS

[SOB18] J. D. Salierno, A. S. Overton, and C. L. Benson. Optimization of hook size to reduce discards and injury in the New Jersey summer flounder hook-and-line fishery. North American Journal of Fisheries Management, 38(2):393–399, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Siepker:2010:EPL

[SOBW10] Michael J. Siepker, Kenneth G. Ostrand, Elizabeth L. Ball, and David H. Wahl. Electrofishing practices have limited effects on largemouth bass feeding behavior. North American Journal of Fisheries Management, 30(6):1395–1400, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Small:2015:CSG

[SOS^{+}15] Maureen P. Small, Serena D. Rogers Olive, Lisa W. Seeb, James E. Seeb, Carita E. Pascal, Kenneth I. Warheit, and William Templin. Chum salmon genetic diversity in the Northeastern Pacific Ocean assessed with single nucleotide polymorphisms (SNPs): Applications to fishery management. North American Journal of Fisheries Management, 35(5):974–987, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stewart:2010:BCI

[SOSL10] Thomas J. Stewart, Robert O’Gorman, W. Gary Sprules, and B. F. Lantry. The bioenergetic consequences of invasiveinduced food Web disruption to Lake Ontario alewives. North American Journal of Fisheries Management, 30(6):1485–1504, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seibert:2013:EAS

[SP13] Justin R. Seibert and Quinton E. Phelps. Evaluation of aging structures for silver carp from Midwestern U.S. rivers. North American Journal of Fisheries Management, 33(4):839–844, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Siddons:2016:AGM

[SPHK16] Stephen F. Siddons, Mark A. Pegg, Nick P. Hogberg, and Geoff M. Klein. Age, growth, and mortality of a trophy channel catfish population in Manitoba, Canada. North American Journal of Fisheries Management, 36(6):1368–1374, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steffensen:2010:AHR

[SPK10] Kirk D. Steffensen, Larkin A. Powell, and Jeff D. Koch. Assessment of hatchery-reared pallid sturgeon survival in the Lower Missouri River. North American Journal of Fisheries Management, 30(3):671–678, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2018:PFO

[SPL18] Richard A. Snow, Michael J. Porta, and James M. Long. Precision of four otolith techniques for estimating age of white perch from a thermally altered reservoir. North American Journal of Fisheries Management, 38(3):725–733, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schrandt:2015:HUF

[SPM15] Meagan N. Schrandt, Sean P. Powers, and John F. Mareska. Habitat use and fishery dynamics of a heavily exploited coastal migrant, Spanish mackerel. North American Journal of Fisheries Management, 35(2):352–363, April 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steffensen:2012:PSH

[SPP12] Kirk D. Steffensen, Larkin A. Powell, and Mark A. Pegg. Population size of hatchery-reared and wild pallid sturgeon in the Lower Missouri River. North American Journal of Fisheries Management, 32(1):159–166, February 2012. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sanft:2018:EIS

[SPP^{+}18] Eric J. Sanft, Anthony P. Porreca, Joseph J. Parkos III, Thomas M. Detmer, and David H. Wahl. Effects of ingestion of soft plastic fishing lures on largemouth bass. North American Journal of Fisheries Management, 38(3):718–724, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sindt:2012:FSG

[SPQ12] Anthony R. Sindt, Clay L. Pierce, and Michael C. Quist. Fish species of greatest conservation need in wadeable Iowa streams: Current status and effectiveness of aquatic gap program distribution models. North American Journal of Fisheries Management, 32(1):135–146, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Simpson:2016:EWV

[SPS16] William G. Simpson, Douglas P. Peterson, and Kurt Steinke. Effect of waveform and voltage gradient on the survival of electroshocked steelhead embryos and larvae. North American Journal of Fisheries Management, 36(5):1149–1155, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sechler:2012:HAS

[SPT^{+}12] Dawn R. Sechler, Quinton E. Phelps, Sara J. Tripp, James E. Garvey, David P. Herzog, David E. Ostendorf, Joseph W. Ridings, Jason W. Crites, and Robert A. Hrabik. Habitat for age-0 shovelnose sturgeon and pallid sturgeon in a large river: Interactions among abiotic factors, food, and energy intake. North American Journal of Fisheries Management, 32 (1):24–31, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smallwood:2012:EAR

[SPW^{+}12] C. B. Smallwood, K. H. Pollock, B. S. Wise, N. G. Hall, and D. J. Gaughan. Expanding aerial–roving surveys to include counts of shore-based recreational fishers from remotely operated cameras: Benefits, limitations, and cost effectiveness. North American Journal of Fisheries Management, 32 (6):1265–1276, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stuck:2015:CPD

[SPWC15] Jason G. Stuck, Anthony P. Porreca, David H. Wahl, and Robert E. Colombo. Contrasting population demographics of invasive silver carp between an impounded and free-flowing river. North American Journal of Fisheries Management, 35 (1):114–122, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stott:2010:GSD

[SQGK10] Wendylee Stott, Henry R. Quinlan, Owen T. Gorman, and Timothy L. King. Genetic structure and diversity among brook trout from Isle Royale, Lake Nipigon, and three Minnesota tributaries of Lake Superior. North American Journal of Fisheries Management, 30(2):400–411, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sindt:2012:HAF

[SQP12] Anthony R. Sindt, Michael C. Quist, and Clay L. Pierce. Habitat associations of fish species of greatest conservation need at multiple spatial scales in wadeable Iowa streams. North American Journal of Fisheries Management, 32(6):1046–1061, December 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skaggs:2017:ECA

[SQS^{+}17] Jordan Skaggs, Yasmín Quintana, Stephanie L. Shaw, Micheal S. Allen, Nicholas A. Trippel, and Michael Matthews. Effects of common angler handling techniques on Florida largemouth bass behavior, feeding, and survival. North American Journal of Fisheries Management, 37(2):263–270, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shoup:2016:LSS

[SR16] Daniel E. Shoup and Ryan G. Ryswyk. Length selectivity and size-bias correction for the North American standard gill net. North American Journal of Fisheries Management, 36 (3):485–496, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sagarese:2018:WTF

[SRCC^{+}18] Skyler R. Sagarese, Adyan B. Rios, Shannon L. Cass-Calay, Nancie J. Cummings, Meaghan D. Bryan, Molly H. Stevens, William J. Harford, Kevin J. McCarthy, and Vivian M. Matter. Working towards a framework for stock evaluations in data-limited fisheries. North American Journal of Fisheries Management, 38(3):507–537, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scoppettone:2015:PDC

[SRFS15] G. Gary Scoppettone, Peter H. Rissler, Mark C. Fabes, and Sean P. Shea. Population dynamics of the cui-ui of Pyramid Lake, Nevada: a potamodromous catostomid subject to failed reproduction. North American Journal of Fisheries Management, 35(5):853–864, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scoppettone:2014:AWP

[SRFW14] G. Gary Scoppettone, Peter H. Rissler, Mark C. Fabes, and Donna Withers. American white pelican predation on cui-ui in Pyramid Lake, Nevada. North American Journal of Fisheries Management, 34(1):57–67, February 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Scoppettone:2012:EBT

[SRSS12] G. Gary Scoppettone, Peter H. Rissler, Sean P. Shea, and William Somer. Effect of brook trout removal from a spawning stream on an adfluvial population of Lahontan cutthroat trout. North American Journal of Fisheries Management, 32 (3):586–596, June 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmalz:2011:FAW

[SS11a] Patrick J. Schmalz and David F. Staples. Factors affecting walleye catch in short-term gill-net sets in a large Minnesota lake. North American Journal of Fisheries Management, 31 (1):12–22, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2011:PSS

[SS11b] William E. Smith and Frederick S. Scharf. Postrelease survival of sublegal southern flounder captured in a commercial gill-net fishery. North American Journal of Fisheries Management, 31 (3):445–454, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schmitt:2013:DHM

[SS13a] Joseph D. Schmitt and Daniel E. Shoup. Delayed hooking mortality of blue catfish caught on juglines. North American Journal of Fisheries Management, 33(2):245–252, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stolarski:2013:PAT

[SS13b] Jason T. Stolarski and Trent M. Sutton. Precision analysis of three aging structures for amphidromous Dolly Varden from Alaskan Arctic rivers. North American Journal of Fisheries Management, 33(4):732–740, August 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stahr:2016:EMS

[SS16] Kristopher J. Stahr and Daniel E. Shoup. The effects of macrophyte stem density and structural complexity on foraging return of invertivorous juvenile largemouth bass. North American Journal of Fisheries Management, 36(4):788–792, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2017:CTP

[SS17] Bradley J. Smith and Darin G. Simpkins. A comparison of three paired modified fyke nets for characterizing fish assemblages in the nearshore zone of Lake Michigan. North American Journal of Fisheries Management, 37(5):962–969, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Shaw:2019:EAH

[SSE19] S. L. Shaw, G. G. Sass, and L. D. Eslinger. Effects of angler harvest on adult muskellunge growth and survival in Escanaba Lake, Wisconsin, 1956–2016. North American Journal of Fisheries Management, 39(1):124–134, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stepien:2018:GDN

[SSK18] Carol A. Stepien, Matthew R. Snyder, and Carey T. Knight. Genetic divergence of nearby walleye spawning groups in Central Lake Erie: Implications for management. North American Journal of Fisheries Management, 38(4):783–793, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schall:2019:STV

[SSK19] Benjamin J. Schall, Casey W. Schoenebeck, and Keith D. Koupal. Spatial and temporal variability in a large-reservoir fish assessment and application of a stratified random sampling approach. North American Journal of Fisheries Management, 39(5):1086–1102, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2018:ETP

[SSP18] Richard A. Snow, Daniel E. Shoup, and Michael J. Porta. Effects of turbidity on prey selection and foraging rate of hatchery-reared juvenile tiger muskellunge. North American Journal of Fisheries Management, 38(2):487–492, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sterling:2019:SBP

[SSP^{+}19] Edward M. Sterling, Kylie B. Sterling, Quinton E. Phelps, Tyler D. Ham, and Sara J. Tripp. Smallmouth bass population demographics in Missouri Ozark streams and an evaluation of a statewide 305-mm minimum length limit. North American Journal of Fisheries Management, 39(6):1175–1191, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Snow:2019:EWS

[SSPP19] Richard A. Snow, Daniel E. Shoup, Michael J. Porta, and Chas P. Patterson. Effects of wintertime stocking of rainbow trout on the forage community of an Oklahoma impoundment. North American Journal of Fisheries Management, 39(2):289– 298, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2015:SMP

[SSS15] N. J. Smith, T. M. Sutton, and J. W. Savereide. Seasonal movement patterns of inconnu in an Arctic Estuary delta complex. North American Journal of Fisheries Management, 35 (4):698–707, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stark:2012:CMG

[Sta12] James W. Stark. Contrasting maturation and growth of northern rock sole in the Eastern Bering Sea and Gulf of Alaska for the purpose of stock management. North American Journal of Fisheries Management, 32(1):93–99, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schwinghamer:2019:UUT

[STP19] Christopher W. Schwinghamer, Sara Tripp, and Quinton E. Phelps. Using ultrasonic telemetry to evaluate paddlefish spawning behavior in Harry S. Truman Reservoir, Missouri. North American Journal of Fisheries Management, 39(2):231– 239, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skalski:2010:CTA

[STSH10] John R. Skalski, Richard L. Townsend, Tracey W. Steig, and Steven Hemstrom. Comparison of two alternative approaches for estimating dam passage survival of salmon smolts. North American Journal of Fisheries Management, 30(3):831–839, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Smith:2012:RCC

[STW^{+}12] Matthew W. Smith, Amy Y. Then, Catarina Wor, Gina Ralph, Kenneth H. Pollock, and John M. Hoenig. Recommendations for catch-curve analysis. North American Journal of Fisheries Management, 32(5):956–967, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Soetaert:2018:NIE

[SVD^{+}18] M. Soetaert, B. Verschueren, A. Decostere, J. Saunders, H. Polet, and K. Chiers. No injuries in European sea bass tetanized by pulse stimulation used in electrotrawling. North American Journal of Fisheries Management, 38(1):247–252, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schramm:2010:FAM

[SVFG10] Harold L. Schramm, Jr., Bruce Vondracek, William E. French, and Patrick D. Gerard. Factors associated with mortality of walleyes and saugers caught in live-release tournaments. North American Journal of Fisheries Management, 30(1):238–253, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Seegert:2013:LVO

[SVPD13] Greg Seegert, Joe Vondruska, Elgin Perry, and Doug Dixon. Longitudinal variation in the Ohio River fish community. North American Journal of Fisheries Management, 33(3):539– 548, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Schueller:2017:DDG

[SW17] Amy M. Schueller and Erik H. Williams. Density-dependent growth in Atlantic menhaden: Impacts on current management. North American Journal of Fisheries Management, 37 (2):294–301, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sullivan:2016:BRW

[SWG^{+}16] Brittany G. Sullivan, Alexander D. M. Wilson, Lee F. G. Gutowsky, Paul H. Patrick, Michael Sills, and Steven J. Cooke. The behavioral responses of a warmwater teleost to different spectra of light-emitting diodes. North American Journal of Fisheries Management, 36(5):1000–1005, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Skalski:2016:SAY

[SWH^{+}16] John R. Skalski, Mark A. Weiland, Kenneth D. Ham, Gene R. Ploskey, Geoffrey A. McMichael, Alison H. Colotelo, Thomas J. Carlson, Christa M. Woodley, M. Brad Eppard, and Eric E. Hockersmith. Status after 5 years of survival compliance testing in the Federal Columbia River power system (FCRPS). North American Journal of Fisheries Management, 36(4):720–730, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steffensen:2015:CCP

[SWHA15] Kirk D. Steffensen, Josh J. Wilhelm, Justin D. Haas, and J. David Adams. Conditional capture probability of pallid sturgeon in benthic trawls. North American Journal of Fisheries Management, 35(4):626–631, August 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Stewart:2014:LIU

[SWW14] Heather A. Stewart, Max H. Wolter, and David H. Wahl. Laboratory investigations on the use of strobe lights and bubble curtains to deter dam escapes of age-0 muskellunge. North American Journal of Fisheries Management, 34(3):571–575, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Sigourney:2015:TDP

[SZHC15] Douglas B. Sigourney, Joseph D. Zydlewski, Edward Hughes, and Oliver Cox. Transport, dam passage, and size selection of adult Atlantic salmon in the Penobscot River, Maine. North American Journal of Fisheries Management, 35(6):1164–1176, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Steed:2010:PVA

[SZKK10] Amber C. Steed, Alexander V. Zale, Todd M. Koel, and Steven T. Kalinowski. Population viability of Arctic grayling in the Gibbon River, Yellowstone National Park. North American Journal of Fisheries Management, 30(6):1582–1590, December 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Teel:2014:GIC

[TBH^{+}14] David J. Teel, Daniel L. Bottom, Susan A. Hinton, David R. Kuligowski, George T. McCabe, Regan McNatt, G. Curtis Roegner, Lia A. Stamatiou, and Charles A. Simenstad. Genetic identification of Chinook salmon in the Columbia River Estuary: Stock-specific distributions of juveniles in shallow tidal freshwater habitats. North American Journal of Fisheries Management, 34(3):621–641, June 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

TenHarmsel:2019:ISA

[TBIJ19] Hailey J. TenHarmsel, B. Bynum Boley, Brian J. Irwin, and Cecil A. Jennings. An importance–satisfaction analysis of trout license holders in Georgia. North American Journal of Fisheries Management, 39(6):1227–1241, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Turner:2018:EED

[TCB18] Sara M. Turner, Bradford C. Chase, and Michael S. Bednarski. Evaluating the effect of dam removals on yellow-phase American eel abundance in a Northeastern U.S. watershed. North American Journal of Fisheries Management, 38(2):424–431, April 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tatara:2017:AMR

[TCG^{+}17] Christopher P. Tatara, Matt R. Cooper, William Gale, Benjamen M. Kennedy, Chris R. Pasley, and Barry A. Berejikian. Age and method of release affect migratory performance of hatchery steelhead. North American Journal of Fisheries Management, 37(4):700–713, August 2017. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2019:INA

[TCM19] Timothy N. Taylor, Benjamin K. Cross, and Barry C. Moore. Investigating nonlethal age estimation in a lentic brook trout population. North American Journal of Fisheries Management, 39(6):1385–1394, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Truesdale:2019:FKP

[TDM19] Corinne L. Truesdale, Tracey M. Dalton, and M. Conor Mc-Manus. Fishers’ knowledge and perceptions of the emerging Southern New England Jonah crab fishery. North American Journal of Fisheries Management, 39(5):951–963, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Twardek:2019:CFG

[TEB^{+}19] W. M. Twardek, L. K. Elmer, M. C. Beere, S. J. Cooke, and A. J. Danylchuk. Consequences of fishery gear type and handling practices on capture and release of wild steelhead on the Bulkley River. North American Journal of Fisheries Management, 39(2):254–269, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trammell:2016:ESK

[TFH^{+}16] Jeffrey L. J. Trammell, David E. Fast, Douglas R. Hatch, William J. Bosch, Ryan Branstetter, Andrew L. Pierce, Joseph W. Blodgett, and Chris R. Frederiksen. Evaluating steelhead kelt treatments to increase iteroparous spawners in the Yakima River basin. North American Journal of Fisheries Management, 36(4):876–887, August 2016. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tabor:2011:HUJ

[TFP^{+}11] Roger A. Tabor, Kurt L. Fresh, Richard M. Piaskowski, Howard A. Gearns, and Daniel B. Hayes. Habitat use by juvenile Chinook salmon in the nearshore areas of Lake Washington: Effects of depth, lakeshore development, substrate, and vegetation. North American Journal of Fisheries Management, 31(4):700–713, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:2015:ETR

[TGA15] Brandon C. Thompson, Daniel C. Gwinn, and Micheal S. Allen. Evacuation times of radio transmitters consumed by largemouth bass. North American Journal of Fisheries Management, 35(4):621–625, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trested:2010:SSS

[TGKI10] Drew G. Trested, Reuben Goforth, James P. Kirk, and J. Jeffery Isely. Survival of shovelnose sturgeon after abdominally invasive endoscopic evaluation. North American Journal of Fisheries Management, 30(1):121–125, February 2010. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thomas:2019:SEB

[TGKZ19] Nathan A. Thomas, Christopher S. Guy, Todd M. Koel, and Alexander V. Zale. In-situ evaluation of benthic suffocation methods for suppression of invasive lake trout embryos in Yellowstone Lake. North American Journal of Fisheries Management, 39(1):104–111, February 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Teuscher:2015:PAW

[TGS^{+}15] David M. Teuscher, Matthew T. Green, Daniel J. Schill, Arnold F. Brimmer, and Ryan W. Hillyard. Predation by American white pelicans on Yellowstone cutthroat trout in the Blackfoot River Drainage, Idaho. North American Journal of Fisheries Management, 35(3):454–463, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tingley:2019:CAP

[THI^{+}19] Ralph W. Tingley III, Jonathan F. Hansen, Daniel A. Isermann, David C. Fulton, Andrea Musch, and Craig P. Paukert. Characterizing angler preferences for largemouth bass, bluegill, and walleye fisheries in Wisconsin. North American Journal of Fisheries Management, 39(4):676–692, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thornton:2015:EDL

[THQ15] Emily J. Thornton, Ronald W. Hardy, and Thomas P. Quinn. Experimental determination of the limits of using stable isotopes to distinguish steelhead and rainbow trout offspring. North American Journal of Fisheries Management, 35(4): 810–817, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tyszko:2017:ARL

[THZ^{+}17] Stephen M. Tyszko, Matt A. Hangsleben, Richard D. Zweifel, Jeremy J. Pritt, and Joseph D. Conroy. Assessing reservoir largemouth bass standardized boat electrofishing: Effect of catchability on density and size structure indices. North American Journal of Fisheries Management, 37(3):587–598, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trumbo:2016:ECV

[TKH^{+}16] Bradly A. Trumbo, Michael D. Kaller, A. Raynie Harlan, Tiffany Pasco, William E. Kelso, and D. Allen Rutherford. Effectiveness of continuous versus point electrofishing for fish assemblage assessment in shallow, turbid aquatic habitats. North American Journal of Fisheries Management, 36(2):398– 406, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2015:QMT

[TL15] Andrew D. Taylor and Matthew K. Litvak. Quantifying a manual triangulation technique for aquatic ultrasonic telemetry. North American Journal of Fisheries Management, 35(5): 865–870, October 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tatara:2019:ARS

[TLC^{+}19] Christopher P. Tatara, Donald A. Larsen, Matt R. Cooper, Penny Swanson, Mollie A. Middleton, Jon T. Dickey, Deborah Harstad, Michael Humling, Chris R. Pasley, and Barry Berejikian. Age at release, size, and maturation status influence residualism in hatchery steelhead. North American Journal of Fisheries Management, 39(3):468–484, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Twardek:2018:PSW

[TLL^{+}18] W. M. Twardek, R. J. Lennox, M. J. Lawrence, J. M. Logan, P. Szekeres, S. J. Cooke, K. Tremblay, G. E. Morgan, and A. J. Danylchuk. The postrelease survival of walleyes following iceangling on Lake Nipissing, Ontario. North American Journal of Fisheries Management, 38(1):159–169, February 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2018:HGS

[TLSB18] Andrew T. Taylor, James M. Long, Michael R. Schwemm, and Shannon K. Brewer. Hybridization and genetic structure of neosho smallmouth bass in the Ozark highlands. North American Journal of Fisheries Management, 38(6):1226–1240, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Titzler:2010:AAR

[TMC10] P. Scott Titzler, Geoffrey A. McMichael, and Jessica A. Carter. Autonomous acoustic receiver deployment and mooring techniques for use in large rivers and estuaries. North American Journal of Fisheries Management, 30(4):853–859, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tuckey:2010:MSF

[TO10] Troy D. Tuckey and John E. Olney. Maturity schedules of female American shad vary at small spatial scales in Chesapeake Bay. North American Journal of Fisheries Management, 30 (4):1020–1031, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tyszko:2017:COS

[TP17] Stephen M. Tyszko and Jeremy J. Pritt. Comparing otoliths and scales as structures used to estimate ages of largemouth bass: Consequences of biased age estimates. North American Journal of Fisheries Management, 37(5):1075–1082, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thompson:2016:ECI

[TPB^{+}16] Brandon C. Thompson, Wesley F. Porak, Kimberly I. Bonvechio, Brandon L. Barthel, Micheal S. Allen, and Charles M. Adams. Economic and conservation impacts of stocking wild Florida bass into large Florida lakes. North American Journal of Fisheries Management, 36(3):452–464, June 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Throckmorton:2015:DIE

[TPBF15] Elizabeth Throckmorton, Amber Peters, Travis Brenden, and Mohamed Faisal. Direct and indirect evidence suggests continuous presence of viral hemorrhagic septicemia virus (Genotype IVb) in Budd Lake, Michigan: Management implications. North American Journal of Fisheries Management, 35(3):503– 511, June 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tiffan:2015:SGT

[TPC^{+}15] Kenneth F. Tiffan, Russell W. Perry, William P. Connor, Frank L. Mullins, Craig D. Rabe, and Doug D. Nelson. Survival, growth, and tag retention in age-0 Chinook salmon implanted with 8-, 9-, and 12-mm PIT tags. North American Journal of Fisheries Management, 35(4):845–852, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Thomas:2013:SSR

[TPF^{+}13] Michael J. Thomas, Matthew L. Peterson, Nick Friedenberg, Joel P. Van Eenennaam, Joseph R. Johnson, Jan Jeffrey Hoover, and A. Peter Klimley. Stranding of spawning run green sturgeon in the Sacramento River: Post-rescue movements and potential population-level effects. North American Journal of Fisheries Management, 33(2):287–297, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Trippel:2018:PSC

[TPL18] Nicholas A. Trippel, Wesley F. Porak, and Erin H. Leone. Poststocking survival of conditioned and pond-reared compared to indoor pellet-reared advanced fingerling Florida bass. North American Journal of Fisheries Management, 38(5): 1039–1049, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tatara:2011:EHF

[TRB11] Christopher P. Tatara, Stephen C. Riley, and Barry A. Berejikian. Effects of hatchery fish density on emigration, growth, survival, and predation risk of natural steelhead parr in an experimental stream channel. North American Journal of Fisheries Management, 31(2):224–235, April 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tattam:2013:IRS

[TRB^{+}13] Ian A. Tattam, James R. Ruzycki, Peter B. Bayley, Hiram W. Li, and Guillermo R. Giannico. The influence of release strategy and migration history on capture rate of Oncorhynchus mykiss in a rotary screw trap. North American Journal of Fisheries Management, 33(2):237–244, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Talmage:2011:MWA

[TS11] Philip J. Talmage and David F. Staples. Mortality of walleyes angled from the deep waters of Rainy Lake, Minnesota. North American Journal of Fisheries Management, 31(5):826–831, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

TenBrink:2013:RBP

[TS13] Todd T. TenBrink and Paul D. Spencer. Reproductive biology of Pacific Ocean perch and northern rockfish in the Aleutian Islands. North American Journal of Fisheries Management, 33 (2):373–383, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Topping:2019:CPR

[TSFS19] Tara S. Topping, Matthew K. Streich, Mark R. Fisher, and Gregory W. Stunz. A comparison of private recreational fishing harvest and effort for Gulf of Mexico red snapper during derby and extended Federal seasons and implications for future management. North American Journal of Fisheries Management, 39(6):1311–1320, December 2019. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tuten:2010:CHS

[TSNH10] Travis Tuten, Will A. Strong, Eric J. Nagid, and Marty M. Hale. Comparisons of haul seine and otter trawl catches of black crappie and the evaluation of haul seines for future research. North American Journal of Fisheries Management, 30 (4):964–975, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Taylor:2018:GPS

[TTS^{+}18] Andrew T. Taylor, Michael D. Tringali, Steven M. Sammons, Travis R. Ingram, Patrick M. O’Rouke, Douglas L. Peterson, and James M. Long. Genetic population structure of shoal bass within their native range. North American Journal of Fisheries Management, 38(3):549–564, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Tabor:2017:DDS

[TWL^{+}17] Roger A. Tabor, Frithiof T. Waterstrat, Daniel W. Lantz, Hans B. Berge, and Martin C. Liermann. Distribution, density, and size of migratory and fluvial sculpins in relation to barriers in Puget Sound lowland streams. North American Journal of Fisheries Management, 37(4):729–742, August 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Uthe:2016:LHC

[UACZ^{+}16] Patrick Uthe, Robert Al-Chokhachy, Alexander V. Zale, Bradley B. Shepard, Thomas E. McMahon, and Tracy Stephens. Life history characteristics and vital rates of Yellowstone cutthroat trout in two headwater basins. North American Journal of Fisheries Management, 36(6):1240–1253, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ulibarri:2017:CEA

[UBR^{+}17] Roy M. Ulibarri, Scott A. Bonar, Chris Rees, Jon Amberg, Bridget Ladell, and Craig Jackson. Comparing efficiency of American fisheries society standard snorkeling techniques to environmental DNA sampling techniques. North American Journal of Fisheries Management, 37(3):644–651, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Underwood:2014:CSP

[UMC14] Zachary E. Underwood, Christopher A. Myrick, and Robert I. Compton. Comparative swimming performance of five Catostomus species and roundtail chub. North American Journal of Fisheries Management, 34(4):753–763, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Uphoff:2011:ISS

[UML^{+}11] James H. Uphoff, Jr., Margaret McGinty, Rudolph Lukacovic, James Mowrer, and Bruce Pyle. Impervious surface, summer dissolved oxygen, and fish distribution in Chesapeake Bay subestuaries: Linking watershed development, habitat conditions, and fisheries management. North American Journal of Fisheries Management, 31(3):554–566, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Usvyatsov:2013:DBS

[UWL13] Sima Usvyatsov, James Watmough, and Matthew K. Litvak. Differentiating between sampling- and environmentrelated mortality in shortnose sturgeon larvae collected using anchored D-frame nets. North American Journal of Fisheries Management, 33(3):595–605, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vadas:2016:CCT

[VBBK16] Robert L. Vadas, Jr., Hal A. Beecher, Steven N. Boessow, and Jonathan H. Kohr. Coastal cutthroat trout redd counts impacted by natural water supply variations. North American Journal of Fisheries Management, 36(4):900–912, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vandergoot:2012:ETS

[VBT^{+}12] Christopher S. Vandergoot, Travis O. Brenden, Michael V. Thomas, Donald W. Einhouse, H. Andrew Cook, and Mark W. Turner. Estimation of tag shedding and reporting rates for Lake Erie jaw-tagged walleyes. North American Journal of Fisheries Management, 32(2):211–223, April 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Velez-Espino:2016:RDM

[VEIW^{+}16] L. Antonio Vélez-Espino, James R. Irvine, Ivan Winther, Roger Dunlop, Geoff Mullins, Kristin Singer, and Nicole Trouton. Robust and defensible mark–recapture methodologies for salmonid escapement: Modernizing the use of data and resources. North American Journal of Fisheries Management, 36(1):183–206, February 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDoornik:2013:GMT

[VEW^{+}13] Donald M. Van Doornik, Debra L. Eddy, Robin S. Waples, Stephen J. Boe, Timothy L. Hoffnagle, Ewann A. Berntson, and Paul Moran. Genetic monitoring of threatened Chinook salmon populations: Estimating introgression of nonnative hatchery stocks and temporal genetic changes. North American Journal of Fisheries Management, 33(4):693–706, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vincent:2014:CAF

[Vin14] Matthew T. Vincent. Comment: Accounting for fishing mortality when comparing density-dependent with densityindependent mortality in Gulf of Mexico red snapper. North American Journal of Fisheries Management, 34(2):349–351, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Venditti:2013:RBS

[VJK13] David A. Venditti, Chris A. James, and Paul Kline. Reproductive behavior and success of captive-reared Chinook salmon spawning under natural conditions. North American Journal of Fisheries Management, 33(1):97–107, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vandergoot:2011:SET

[VKBL11] Christopher S. Vandergoot, Patrick M. Kocovsky, Travis O. Brenden, and Weihai Liu. Selectivity evaluation for two experimental gill-net configurations used to sample Lake Erie walleyes. North American Journal of Fisheries Management, 31(5):832–842, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vine:2019:USS

[VKH^{+}19] Joshua R. Vine, Yoichiro Kanno, S. Chad Holbrook, William C. Post, and Brandon K. Peoples. Using side-scan sonar and N-mixture modeling to estimate Atlantic sturgeon spawning migration abundance. North American Journal of Fisheries Management, 39(5):939–950, October 2019. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Varkey:2016:MCD

[VMA^{+}16] Divya A. Varkey, Murdoch K. McAllister, Paul J. Askey, Eric Parkinson, Adrian Clarke, and Theresa Godin. Multicriteria decision analysis for recreational trout fisheries in British Columbia, Canada: a Bayesian network implementation. North American Journal of Fisheries Management, 36(6):1457–1472, December 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Vandergoot:2011:ETF

[VMC^{+}11] Christopher S. Vandergoot, Karen J. Murchie, Steven J. Cooke, John M. Dettmers, Roger A. Bergstedt, and David G. Fielder. Evaluation of two forms of electroanesthesia and carbon dioxide for short-term anesthesia in walleye. North American Journal of Fisheries Management, 31(5):914–922, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Veinott:2018:RAL

[VPV18] Geoff Veinott, Levi Pike, and Asokan Mulayath Variyath. Response of anglers to less-restrictive harvest controls in a recreational Atlantic salmon fishery. North American Journal of Fisheries Management, 38(1):210–222, February 2018. CO-DEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

VanDoornik:2011:GMR

[VWB^{+}11] Donald M. Van Doornik, Robin S. Waples, Melissa C. Baird, Paul Moran, and Ewann A. Berntson. Genetic monitoring reveals genetic stability within and among threatened Chinook salmon populations in the Salmon River, Idaho. North American Journal of Fisheries Management, 31(1):96–105, February 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Welsh:2014:SAM

[WA14] Stuart A. Welsh and Joni L. Aldinger. A semi-automated method for monitoring dam passage of upstream migrant yellow-phase American eels. North American Journal of Fisheries Management, 34(4):702–709, August 2014. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walker:2016:MLS

[WA16] Daniel J. Walker and J. Brian Alford. Mapping lake sturgeon spawning habitat in the Upper Tennessee River using side-scan sonar. North American Journal of Fisheries Management, 36 (5):1097–1105, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wang:2017:SPR

[WALL17] Yuanming Wang, Ruidong An, Yong Li, and Kefeng Li. Swimming performance of rock carp Procypris rabaudi and Prenant’s schizothoracin Schizothorax prenanti acclimated to total dissolved gas supersaturated water. North American Journal of Fisheries Management, 37(6):1183–1190, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:2019:GSI

[WAW^{+}19] Hillary G. M. Ward, Paul J. Askey, Tyler Weir, Karen K. Frazer, and Michael A. Russello. Genetic stock identification reveals that angler harvest is representative of cryptic stock proportions in a high-profile Kokanee fishery. North American Journal of Fisheries Management, 39(3):415–425, June 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winans:2010:GPB

[WBB10] Gary A. Winans, Melissa C. Baird, and Jon Baker. A genetic and phenetic baseline before the recolonization of steelhead above Howard Hanson Dam, Green River, Washington. North American Journal of Fisheries Management, 30(3):742–756, June 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Worthington:2013:STV

[WBF13] Thomas A. Worthington, Shannon K. Brewer, and Nicole Farless. Spatial and temporal variation in efficiency of the Moore egg collector. North American Journal of Fisheries Management, 33(6):1113–1118, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wirgin:2015:OAS

[WBF^{+}15] Isaac Wirgin, Matthew W. Breece, Dewayne A. Fox, Lorraine Maceda, Kevin W. Wark, and Tim King. Origin of Atlantic sturgeon collected off the Delaware Coast during spring months. North American Journal of Fisheries Management, 35(1):20–30, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:2017:EFH

[WBG^{+}17] Taylor D. Ward, Jacob W. Brownscombe, Lee F. G. Gutowsky, Robert Ballagh, Nicholas Sakich, Derek McLean, Geneviève Quesnel, Sahil Gambhir, Constance M. O’Connor, and Steven J. Cooke. Electric fish handling gloves provide effective immobilization and do not impede reflex recovery of adult largemouth bass. North American Journal of Fisheries Management, 37(3):652–659, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Worthington:2013:SEM

[WBGM13] Thomas A. Worthington, Shannon K. Brewer, Timothy B. Grabowski, and Julia Mueller. Sampling efficiency of the Moore egg collector. North American Journal of Fisheries Management, 33(1):79–88, February 2013. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Webber:2013:TSE

[WBH13] P. Aaron Webber, Kevin R. Bestgen, and G. Bruce Haines. Tributary spawning by endangered Colorado River Basin fishes in the White River. North American Journal of Fisheries Management, 33(6):1166–1171, December 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Widmer:2010:DPE

[WBK^{+}10] Ann M. Widmer, Laura L. Burckhardt, Jon W. Kehmeier, Eric J. Gonzales, C. Nicolas Medley, and Richard A. Valdez. Detection and population estimation for small-bodied fishes in a sand-bed river. North American Journal of Fisheries Management, 30(6):1553–1570, December 2010. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Widmer:2013:DPE

[WBK^{+}13] Ann M. Widmer, Laura L. Burckhardt, Jon W. Kehmeier, Eric J. Gonzales, C. Nicolas Medley, and Richard A. Valdez. Detection and population estimation for small-bodied fishes in a sand-bed river: Response to comment. North American Journal of Fisheries Management, 33(2):453–458, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:2013:EIT

[WBP13] Jonathan P. Wagner, Reginald B. Blaylock, and Mark S. Peterson. Evaluation of internal tag performance in hatcheryreared juvenile spotted seatrout. North American Journal of Fisheries Management, 33(4):783–789, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winters:2017:ETS

[WBT17] Lisa K. Winters, Phaedra Budy, and Gary P. Thiede. Earning their stripes: The potential of tiger trout and other salmonids as biological controls of forage fishes in a Western reservoir. North American Journal of Fisheries Management, 37(2):380– 394, April 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weaver:2019:OAS

[WBZ19] Daniel M. Weaver, Michael Brown, and Joseph D. Zydlewski. Observations of American shad Alosa sapidissima approaching and using a vertical slot fishway at the head-of-tide Brunswick Dam on the Androscoggin River, Maine. North American Journal of Fisheries Management, 39(5):989–998, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wuellner:2010:IBW

[WCWA10] Melissa R. Wuellner, Steven R. Chipps, David W. Willis, and Wells E. Adams, Jr. Interactions between walleyes and smallmouth bass in a Missouri River reservoir with consideration of the influence of temperature and prey. North American Journal of Fisheries Management, 30(2):445–463, April 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilcox:2018:CIS

[WCY^{+}18] T. M. Wilcox, K. J. Carim, M. K. Young, K. S. McKelvey, T. W. Franklin, and M. K. Schwartz. Comment: The importance of sound methodology in environmental DNA sampling. North American Journal of Fisheries Management, 38(3):592– 596, June 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williamson:2010:CTU

[WD10] Craig W. Williamson and Ryan R. Dirnberger. A comparison of techniques using dorsal spines to estimate sauger age. North American Journal of Fisheries Management, 30(4):1016–1019, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:2014:STD

[WDD^{+}14] Tyler Wagner, Jefferson T. Deweber, Jason Detar, David Kristine, and John A. Sweka. Spatial and temporal dynamics in brook trout density: Implications for population monitoring. North American Journal of Fisheries Management, 34(2):258– 269, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wolter:2013:FLE

[WDW13] Max H. Wolter, Corey S. DeBoom, and David H. Wahl. Field and laboratory evaluation of dam escapement of muskellunge. North American Journal of Fisheries Management, 33 (4):829–838, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Withers:2019:IAT

[WEC^{+}19] Jonah L. Withers, Donald Einhouse, Michael Clancy, Lori Davis, Rachel Neuenhoff, and John Sweka. Integrating acoustic telemetry into a mark–recapture model to improve catchability parameters and abundance estimates of lake sturgeon in Eastern Lake Erie. North American Journal of Fisheries Management, 39(5):913–920, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weber:2017:EPI

[WF17] Michael J. Weber and Mark Flammang. Effects of passive integrated transponder tag size and implantation site on age-0 walleye and muskellunge tag retention, growth, and survival. North American Journal of Fisheries Management, 37(3):480– 488, June 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weber:2013:EEM

[WFS13] Michael J. Weber, Mark Flammang, and Randall Schultz. Estimating and evaluating mechanisms related to walleye escapement from Rathbun Lake, Iowa. North American Journal of Fisheries Management, 33(3):642–651, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Winner:2014:MER

[WFWSV14] Brent L. Winner, Kerry E. Flaherty-Walia, Theodore S. Switzer, and Julie L. Vecchio. Multidecadal evidence of recovery of nearshore red drum stocks off West-Central Florida and connectivity with inshore nurseries. North American Journal of Fisheries Management, 34(4):780–794, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weeber:2010:ERS

[WGJ10] Matthew A. Weeber, Guillermo R. Giannico, and Steven E. Jacobs. Effects of redd superimposition by introduced Kokanee on the spawning success of native bull trout. North American Journal of Fisheries Management, 30(1):47–54, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Woodcock:2011:DMS

[WGM^{+}11] Skye H. Woodcock, Bronwyn M. Gillanders, Andrew R. Munro, David A. Crook, and Andrew C. Sanger. Determining mark success of 15 combinations of enriched stable isotopes for the batch marking of larval otoliths. North American Journal of Fisheries Management, 31(5):843–851, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Williams-Grove:2016:MER

[WGS16] Laura Jay Williams-Grove and Stephen T. Szedlmayer. Mortality estimates for red snapper based on ultrasonic telemetry in the Northern Gulf of Mexico. North American Journal of Fisheries Management, 36(5):1036–1044, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waterhouse:2011:IRT

[WH11] Lynn Waterhouse and John M. Hoenig. Instantaneous-rates tagging models allowing for delayed mixing of newly tagged cohorts: Partial year tabulation of recaptures. North American Journal of Fisheries Management, 31(6):995–1004, December 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waterhouse:2012:TME

[WH12] Lynn Waterhouse and John M. Hoenig. Tagging models for estimating survival rates when tag visibility changes over time: Partial-year tabulation of recaptures. North American Journal of Fisheries Management, 32(1):147–158, February 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weber:2011:SPR

[WHB11] Michael J. Weber, Matthew J. Hennen, and Michael L. Brown. Simulated population responses of common carp to commercial exploitation. North American Journal of Fisheries Management, 31(2):269–279, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Waldman:2014:GMS

[WHB^{+}14] John Waldman, Daniel Hasselman, Paul Bentzen, Michael Dadswell, Lorraine Maceda, and Isaac Wirgin. Genetic mixedstock analysis of American shad in two Atlantic Coast fisheries: Delaware Bay, USA, and Inner Bay of Fundy, Canada. North American Journal of Fisheries Management, 34(6): 1190–1198, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilber:2012:RCA

[WHC12] Dara H. Wilber, Nancy H. Hadley, and Douglas G. Clarke. Resident crab associations with sedimentation on restored intertidal oyster reefs in South Carolina and the implications for secondary consumers. North American Journal of Fisheries Management, 32(5):838–847, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walter:2014:RBFa

[WHC14a] John F. Walter, John M. Hoenig, and Mary C. Christman. Reducing bias and filling in spatial gaps in fisherydependent catch-per-unit-effort data by geostatistical prediction, I. Methodology and simulation. North American Journal of Fisheries Management, 34(6):1095–1107, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walter:2014:RBFb

[WHC14b] John F. Walter, John M. Hoenig, and Mary C. Christman. Reducing bias and filling in spatial gaps in fishery-dependent catch-per-unit-effort data by geostatistical prediction, II. Application to a scallop fishery. North American Journal of Fisheries Management, 34(6):1108–1118, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilder:2016:FPI

[WHG^{+}16] Richard M. Wilder, Jason L. Hassrick, Lenny F. Grimaldo, Marin F. D. Greenwood, Shawn Acuña, Jillian M. Burns, Donna M. Maniscalco, Patrick K. Crain, and Tien-Chieh Hung. Feasibility of passive integrated transponder and acoustic tagging for endangered adult delta smelt. North American Journal of Fisheries Management, 36(5):1167–1177, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Warpinski:2016:ASF

[WHGC16] Stephanie Warpinski, Mark Herrmann, Joshua A. Greenberg, and Keith R. Criddle. Alaska’s sablefish fishery after individual fishing quota (IFQ) program implementation: an international economic market model. North American Journal of Fisheries Management, 36(4):864–875, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wegener:2017:MHU

[WHKB17] Matthew G. Wegener, Kathryn M. Harriger, John R. Knight, and Mark A. Barrett. Movement and habitat use of alligator gars in the Escambia River, Florida. North American Journal of Fisheries Management, 37(5):1028–1038, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Warren:2014:UEB

[WHMS14] D. R. Warren, C. J. Harvey, M. M. McClure, and B. L. Sanderson. Use of an ecosystem-based model to evaluate alternative conservation strategies for juvenile Chinook salmon in a headwater stream. North American Journal of Fisheries Management, 34(4):839–852, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wellenkamp:2015:UME

[WHV15] William Wellenkamp, Ji X. He, and Darren Vercnocke. Using maxillae to estimate ages of lake trout. North American Journal of Fisheries Management, 35(2):296–301, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wegleitner:2017:CTV

[WI17] Eric J. Wegleitner and Daniel A. Isermann. Comparison of two viewing methods for estimating largemouth bass and walleye ages from sectioned otoliths and dorsal spines. North American Journal of Fisheries Management, 37(6):1304–1310, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weaver:2014:SCC

[WKP14] Daniel M. Weaver, Thomas J. Kwak, and Kenneth H. Pollock. Sampling characteristics and calibration of snorkel counts to estimate stream fish populations. North American Journal of Fisheries Management, 34(6):1159–1166, December 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wanner:2010:CGW

[WKS^{+}10] Greg A. Wanner, Robert A. Klumb, Dane A. Shuman, Kirk Steffensen, Sam Stukel, and Nicholas J. Utrup. Comparison of green and white mesh trammel nets and gill nets to assess the fish community in a large river. North American Journal of Fisheries Management, 30(1):12–25, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Watkins:2015:ISL

[WKTQ15] Carson J. Watkins, Zachary B. Klein, Marc M. Terrazas, and Michael C. Quist. Influence of sectioning location on age estimates from common carp dorsal spines. North American Journal of Fisheries Management, 35(4):690–697, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wallen:2016:MER

[WLK^{+}16] Kenneth E. Wallen, Adam C. Landon, Gerard T. Kyle, Michael A. Schuett, Jeremy Leitz, and Ken Kurzawski. Mode effect and response rate issues in mixed-mode survey research: Implications for recreational fisheries management. North American Journal of Fisheries Management, 36(4):852–863, August 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wang:2015:TAC

[WLL^{+}15] Yuanming Wang, Kefeng Li, Jia Li, Ran Li, and Yun Deng. Tolerance and avoidance characteristics of Prenant’s schizothoracin Schizothorax prenanti to total dissolved gas supersaturated water. North American Journal of Fisheries Management, 35(4):827–834, August 2015. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wilson:2015:CSM

[WMB^{+}15] Kyle L. Wilson, Bryan G. Matthias, Andrew B. Barbour, Robert N. M. Ahrens, Travis Tuten, and Micheal S. Allen. Combining samples from multiple gears helps to avoid fishy growth curves. North American Journal of Fisheries Management, 35(6):1121–1131, December 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wolf:2019:CPW

[WMB19] Skylar L. Wolf, Robert Mollenhauer, and Shannon K. Brewer. Coldwater periods in warmwater streams: Microhabitat shifts from autumn to winter by smallmouth bass. North American Journal of Fisheries Management, 39(6):1360–1372, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Withers:2019:LCS

[WMD^{+}19] Jonah L. Withers, Nicholas Markley, Lori Davis, Thomas Kehler, and John Sweka. Labeling calcified structures in young lake sturgeon: a comparison of oxytetracycline hydrochloride, alizarin red s, and calcein. North American Journal of Fisheries Management, 39(6):1217–1226, December 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Whitney:2019:EFR

[WMMB19] Don W. Whitney, Kevin A. Meyer, Joshua L. McCormick, and Brett J. Bowersox. Effects of fishery-related fight time and air exposure on prespawn survival and reproductive success of adult hatchery steelhead. North American Journal of Fisheries Management, 39(2):372–378, April 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:2013:ELA

[WMSH13] David L. Ward, R. Morton-Starner, and Shaula J. Hedwall. An evaluation of liquid ammonia (ammonium hydroxide) as a candidate piscicide. North American Journal of Fisheries Management, 33(2):400–405, April 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:2018:HRR

[WMSV18] David L. Ward, Rylan Morton-Starner, and Ben Vaage. Are hatchery-reared rainbow trout and brown trout effective predators on juvenile native fish? North American Journal of Fisheries Management, 38(5):1105–1113, October 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weber:2011:SFD

[WP11] Christine Weber and Armin Peter. Success or failure? Do indicator selection and reference setting influence river rehabilitation outcome? North American Journal of Fisheries Management, 31(3):535–547, June 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Watson:2018:OCB

[WPJ^{+}18] Nicole M. Watson, Carson G. Prichard, Jory L. Jonas, James J. Student, and Kevin L. Pangle. Otolith-chemistry-based discrimination of wild- and hatchery-origin steelhead across the Lake Michigan Basin. North American Journal of Fisheries Management, 38(4):820–832, August 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Watkins:2019:DWM

[WPS^{+}19] Owen B. Watkins, Andrew J. Paul, Stephen C. Spencer, Michael G. Sullivan, and Lee Foote. Dude, Where’s my transmitter? Probability of radio transmitter detections and locational errors for tracking river fish. North American Journal of Fisheries Management, 39(4):753–761, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:2015:LET

[WPY^{+}15] David L. Ward, William R. Persons, Kirk L. Young, Dennis M. Stone, David R. Vanhaverbeke, and William K. Knight. A laboratory evaluation of tagging-related mortality and tag loss in juvenile humpback chub. North American Journal of Fisheries Management, 35(1):135–140, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walrath:2015:TEN

[WQF15] John D. Walrath, Michael C. Quist, and Jon A. Firehammer. Trophic ecology of nonnative northern pike and their effect on conservation of native westslope cutthroat trout. North American Journal of Fisheries Management, 35(1):158–177, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:2013:ACM

[WQP13] Hillary G. M. Ward, Michael S. Quinn, and John R. Post. Angler characteristics and management implications in a large, multistock, spatially structured recreational fishery. North American Journal of Fisheries Management, 33(3):576–584, June 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walton-Rabideau:2019:ETR

[WRNJ^{+}19] Sarah E. Walton-Rabideau, Mark Newell, Amanda L. Jeanson, Elodie J. I. Lédée, John M. Farrell, and Steven J. Cooke. Evaluation of tag retention, healing, growth, and behavior in age-0 muskellunge following acoustic transmitter implantation. North American Journal of Fisheries Management, 39 (4):652–663, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:2011:EHD

[WS11a] Tyler Wagner and John A. Sweka. Evaluation of hypotheses for describing temporal trends in Atlantic salmon parr densities in Northeast U.S. rivers. North American Journal of Fisheries Management, 31(2):340–351, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Weis:2011:LBN

[WS11b] Jerome J. Weis and Greg G. Sass. Largemouth bass nest site selection in small, North temperate lakes varying in littoral coarse woody habitat abundances. North American Journal of Fisheries Management, 31(5):943–951, October 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Westhoff:2013:MGC

[WS13] Jacob T. Westhoff and Nicholas A. Sievert. Mortality and growth of crayfish internally tagged with PIT tags. North American Journal of Fisheries Management, 33(5):878–881, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walker:2016:GIF

[WS16] Brian M. Walker and Trent M. Sutton. Growth-increment formation using otoliths and scales for age-0 Chinook salmon. North American Journal of Fisheries Management, 36(5):995– 999, October 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wellenkamp:2018:EPW

[WSA18] William Wellenkamp, Shawn P. Sitar, and James Aho. Evaluation of partial and whole maxillary bone excision as a nonlethal age estimation structure in lake trout. North American Journal of Fisheries Management, 38(6):1375–1380, December 2018. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Walleser:2014:STV

[WSH^{+}14] Liza R. Walleser, Mark B. Sandheinrich, David R. Howard, Mark P. Gaikowski, and Jon J. Amberg. Spatial and temporal variation of the gill rakers of gizzard shad and silver carp in three Midwestern rivers. North American Journal of Fisheries Management, 34(5):875–884, October 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Ward:2013:EAS

[WSS13] Matthew C. Ward, Donald R. Schreiner, and David F. Staples. An evaluation of age-1 steelhead stocking locations on a Minnesota tributary to Lake Superior. North American Journal of Fisheries Management, 33(5):1063–1070, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Welty:2015:MAR

[WTB^{+}15] Ethan Z. Welty, Christian E. Torgersen, Samuel J. Brenkman, Jeffrey J. Duda, and Jonathan B. Armstrong. Multiscale analysis of river networks using the R package linbin. North American Journal of Fisheries Management, 35(4):802–809, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wathen:2017:AGC

[WWB^{+}17] Gus Wathen, Nicholas Weber, Stephen Bennett, Nicolaas Bouwes, and Chris E. Jordan. An assessment of the geographic closure assumption in mark–recapture abundance estimates of anadromous steelhead populations. North American Journal of Fisheries Management, 37(5):951–961, October 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Woodard:2013:GSL

[WWD13] Stephen R. Woodard, Russell A. Wright, and Dennis R. De-Vries. Growth and survival of largemouth bass following supplemental feeding of bluegills in small impoundments. North American Journal of Fisheries Management, 33(1):170–177, February 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wiedenmann:2013:EHC

[WWM13] John Wiedenmann, Michael J. Wilberg, and Thomas J. Miller. An evaluation of harvest control rules for data-poor fisheries. North American Journal of Fisheries Management, 33 (4):845–860, August 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wood:2018:GAB

[WWP18] Darren M. Wood, Amy B. Welsh, and J. Todd Petty. Genetic assignment of brook trout reveals rapid success of culvert restoration in headwater streams. North American Journal of Fisheries Management, 38(5):991–1003, October 2018.

CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:2014:PSR

[WWS^{+}14] Katie A. Wagner, Christa M. Woodley, Adam G. Seaburg, John R. Skalski, and M. Brad Eppard. Physiological stress responses to prolonged exposure to MS-222 and surgical implantation in juvenile Chinook salmon. North American Journal of Fisheries Management, 34(4):863–873, August 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Wagner:2015:SCI

[WWW15] Curtis P. Wagner, Michael J. Weber, and David H. Wahl. Structural complexity influences littoral coarse woody habitat selection by juvenile muskellunge. North American Journal of Fisheries Management, 35(1):14–19, February 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Xie:2014:EAE

[XM14] Yunbo Xie and Fiona J. Martens. An empirical approach for estimating the precision of hydroacoustic fish counts by systematic hourly sampling. North American Journal of Fisheries Management, 34(3):535–545, June 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Yuen:2011:BEI

[YC11] Henry Yuen and Robert Conrad. Bias in the estimation of impacts of simultaneous mark-selective and nonselective fisheries on ocean salmon. North American Journal of Fisheries Management, 31(6):1043–1051, December 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Younk:2010:SLT

[YHP10] Jerry A. Younk, Brian R. Herwig, and Bruce J. Pittman. Short- and long-term evaluation of passive integrated transponder and visible implant elastomer tag performance in muskellunge. North American Journal of Fisheries Management, 30(1):281–288, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Young:2012:MCE

[YS12] Michael K. Young and David A. Schmetterling. Movement and capture efficiency of radio-tagged salmonids sampled by electrofishing. North American Journal of Fisheries Management, 32(5):823–831, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zeller:2010:EVI

[ZC10] Megan M. Zeller and Stefan H. Cairns. Evaluation of visible implant elastomer (VIE) for marking fingerling blue catfish. North American Journal of Fisheries Management, 30 (1):254–258, February 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhang:2012:EHC

[ZC12] Yuying Zhang and Yong Chen. Effectiveness of harvest control rules in managing American lobster fishery in the Gulf of Maine. North American Journal of Fisheries Management, 32 (5):984–999, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhang:2015:DAI

[ZC15] Chongliang Zhang and Yong Chen. Development of abundance indices for Atlantic cod and cusk in the coastal Gulf of Maine from their bycatch in the lobster fishery. North American Journal of Fisheries Management, 35(4):708–719, August 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zemeckis:2014:SDA

[ZDC14] Douglas R. Zemeckis, Micah J. Dean, and Steven X. Cadrin. Spawning dynamics and associated management implications for Atlantic cod. North American Journal of Fisheries Management, 34(2):424–442, April 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhao:2011:RSC

[ZEM11] Yingming Zhao, Donald W. Einhouse, and Thomas M. Mac-Dougall. Resolving some of the complexity of a mixed-origin walleye population in the East Basin of Lake Erie using a mark–recapture study. North American Journal of Fisheries Management, 31(2):379–389, April 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zajicek:2011:TGC

[ZGW11] Paul Zajicek, Andrew E. Goodwin, and Tina Weier. Triploid grass carp: Triploid induction, sterility, reversion, and certification. North American Journal of Fisheries Management, 31(4):614–618, August 2011. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhao:2013:DSR

[ZKM13] Yingming Zhao, Patrick M. Kocovsky, and Charles P. Madenjian. Development of a stock–recruitment model and assessment of biological reference points for the Lake Erie walleye fishery. North American Journal of Fisheries Management, 33 (5):956–964, October 2013. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zielinski:2010:DMC

[ZLB10] C. M. Zielinski, H. V. Lorz, and J. L. Bartholomew. Detection of Myxobolus cerebralis in the Lower Deschutes River Basin, Oregon. North American Journal of Fisheries Management, 30(4):1032–1040, August 2010. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhao:2017:ESS

[ZM17] Yingming Zhao and Yolanda E. Morbey. Estimating the size selectivity of trap nets using a gill-net selectivity experiment: Method development and application to lake whitefish in Lake Huron. North American Journal of Fisheries Management, 37(6):1341–1349, December 2017. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zydlewski:2014:GSL

[ZOC^{+}14] Joseph Zydlewski, Andrew O’Malley, Oliver Cox, Peter Ruksznis, and Joan G. Trial. Growth and smolting in lower-mode Atlantic salmon stocked into the Penobscot River, Maine. North American Journal of Fisheries Management, 34(1):147–158, February 2014. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zeigler:2019:PPR

[ZRR^{+}19] Matthew P. Zeigler, Kevin B. Rogers, James J. Roberts, Andrew S. Todd, and Kurt D. Fausch. Predicting persistence of Rio Grande cutthroat trout populations in an uncertain future. North American Journal of Fisheries Management, 39 (5):819–848, October 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zielinski:2016:BCD

[ZS16] D. P. Zielinski and P. W. Sorensen. Bubble curtain deflection screen diverts the movement of both Asian and common carp. North American Journal of Fisheries Management, 36(2):267– 276, April 2016. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zorn:2011:DUF

[ZSW11] Troy G. Zorn, Paul W. Seelbach, and Michael J. Wiley. Developing user-friendly habitat suitability tools from regional stream fish survey data. North American Journal of Fisheries Management, 31(1):41–55, February 2011. CODEN NA-JMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zorn:2012:RSE

[ZW12] Troy G. Zorn and Todd C. Wills. A reconnaissance survey of the effects of sediment traps on Michigan streams. North American Journal of Fisheries Management, 32(5):1005–1016, October 2012. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zhu:2015:CTA

[ZWH^{+}15] Xinhua Zhu, Rick J. Wastle, Kimberly L. Howland, Deanna J. Leonard, Susan Mann, Theresa J. Carmichael, and Ross F. Tallman. A comparison of three anatomical structures for estimating age in a slow-growing subarctic population of lake whitefish. North American Journal of Fisheries Management, 35(2):262–270, April 2015. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).

Zeug:2019:AJC

[ZWS^{+}19] Steven C. Zeug, Jesse Wiesenfeld, Kirsten Sellheim, Annie Brodsky, and Joseph E. Merz. Assessment of juvenile Chinook salmon rearing habitat potential prior to species reintroduction. North American Journal of Fisheries Management, 39(4):762–777, August 2019. CODEN NAJMDP. ISSN 0275-5947 (print), 1548-8675 (electronic).